A Bibliography of Publications in Datamation Magazine

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

23 May 2013 Version 1.14

Title word cross-reference

#1 [Ano92-118].

n [Dar96h].

-tier [Dar96h].

000 [Ano94-71].

1-800 [Woo91]. 1.44MB [Ano96o]. 10 [Ano92-71, Liv95]. 100 [Ano93-303, Ano93-305, Ano93-286, Ano95-35, Ano95-36, Ano95-37, Ano95-38, Ano97-27]. 100EX [Ano97-42]. 1988 [Ano93-104]. 1990s [Ano90a, Fra90a, Tom88]. 1992 [Ano92a]. 1994 [Moa95f, Moa95g]. 1TB [Ano96p].

2 [Ano93-119, Ano93-151, Ano93-240, Ano94-243, Sim96j]. 2.0 [Ano93-81]. 2.9V [Ano95-47]. 2000 [Ano96-35, Ano97m, Ano97u, App97b, DB97, FM97, Hay97t, dJ96a, Jag96b, Jon97a, War97, de 97d, de 97e, de 97f, de 97a, de 97g, de 97b, de 97c, de 97h, de 97i, de 98a, de 98b]. 21st [Ano92-80, Joy89b, Jon97c]. 2Q [Ham90a]. 2s [Ano93-214].

3 [Ano96v, Ano96l, Ano96w, Ano96f, Ano97n]. 3-D [Ano96f]. 3.0 [Moa90i]. 32 [Ano93-191, Coh95, Var95b]. 32-bit [Fol96b, Ano93-191, Coh95, Var95b]. 390 [Ano97-48, Sim96r, Var96j].

4 [Sim96j, Ano93-87]. 4.0 [Ric92b]. 4.8-pound [Ano95-47]. 40 [Ano93-326]. 400 [Ama95b, Ano92-60, Ano92-82, Ano93-84, Ano93j, Ano93-172, Ano93-296, Ano93-345, Ano94d, Ano96b, Ano96n, Ano97m, Gib97a, Hok96, Kre93a, McC93e, Nee95, Sim96p, Sim96q, Sim96s, Tay96]. 400-based [Sim96q]. 400s [Ano93-230, Sim96c, Sim96l]. 486s [Ano93-83]. 4GL [Ano92b]. 4GLs [Bau94b, Ano93a, Ano93-294]. 4th [Ano93b].

5 [Ano96u]. 5.25 [Ano96p]. 5.25-Inch [Ano96p]. 50 [Ano92-152]. 500 [McC97e].

64-bit [Var96a]. 64Bit [Ano96-51].

’70s [Hen95].

85 [Ano93c]. ’88 [Ano88c].

9 [Ano96u]. 90MHz [Ano95-47]. 911 [Mik97c]. ’92 [Ano93-301]. ’94 [Ano93-326, Ano94-27]. 95 [SF96k]. ’96 [Ano96u]. ’97 [Ano97s, Ano96z, Ano97n].

abound [Ano94-278, McC94c]. abroad [Kor93]. academia [Lei90d]. Access [Ano93-319, App96f, Dar95g, Dar96s, Lin94c, Dar95a, Dar95b, Dar96n, Dav89f, Fis96b, Kay97, Lei90c, Sta96, Ama95b, Ano94b, Ano94-245, Ano94-270]. accolades [Ano96-33]. accommodate [Dar96s]. according [vSS96]. accountants [Sne92g]. Accounting [Ano92g, Ano92-88, Man92a, Sne92g]. accounts [Ano97x]. AcerPAC [Ano92-147]. AcerPAC-kage [Ano92-147]. achievable [Joy89c]. achieve [Ano93-350, Inm97a, Vog96c]. acquired [Ano95-113]. acquiring [Dar96d]. acquisition [Dar95a, Mae96, McC96o]. Across [Ano93-284, Ano94-183, Ano94-244, Cor96, Dar96l, Lin94b, Tuc97k, Var96b]. act [FM97, Tuc97g]. action [de 97f]. Active [Dav97c, Dav97j]. ActiveX [Ano96-30, Ano97a, Lin96c]. ActiveXecutables [Ano97m]. actually [Ova96]. Actuate [App96h]. ad [Ano96-44, Car90c, Fos90a, MB90, Sch90c]. AD/Cycle [Fos90a, MB90, Sch90c, Car90c].

Adams [Ano96y]. adapt [Ric94h]. Adaptation [Mar89a]. adaptive [Lin96b]. Add [Ano94-70, Ano94-277, Dav92a, Rya94, Fly96b, Hur94a]. add-ons [Fly96b]. added [Kno97d]. Adding [Ano93-314, Hay96e, McC96d]. Address [McC96g]. Adds [Ano93-234]. adherence [Hay96d]. administer [Val96]. administration [Hir96]. administrator [Ril96h]. Administrators [Ham94c, Ano96b, Ril95e]. adopt [Ano94-278]. adopters [Kay96f]. adopting [Ano95-56]. adults [Ril95h]. Advanced [Ano94-182, Car88a]. Advantage [Ano96c, Ano96d, Ano96g, Ano96v, Ano97c, Ano97d, Ano97e, Ano97j, Ano97l, Ano97i, Ano97k, Ano97n, Ano97o, Ano98a, Inm97a, MCBH88, Sim96d]. Advantages [Sim96r, Sim96p]. Advertiser [Ano97b]. Advertisers [Ano95a]. advice [Rad96a]. Aetna [Bra89]. Affect [Ano93-112, Mac95, Moa93b]. affordable [Dri96, McC96m]. African [Mad90]. After [Ano93-170, Ano94-272, Str95e, SF96l, Moa94a]. aftershock [Ano97-28, Ruh88]. AG [Car90e]. Again [Ano93-294, SF98, Dar96f, Moa89e, SF96a, Sim97a, Vog96b]. Against [Ber94]. age [Bus88, Hay96c, McM90f, McM90i, Rot88a, Kay95j]. agency [App96d]. Agenda [Ano92-112, Moa90e]. agent [Ano97k]. Agents [Dav94, Ano95-126]. ago [Hur96, Str96a]. agony [Kay96a, Var96g]. Agriculture [Sch89c]. Ahead [Ano93-189, Var94b, Car89a, Coo92b, Hay96a, Ryp88, Sem96e, Sim97h, Ano93-173, Ano93-174, Ano93-175, Ano93-176, Ano93-177, Ano93-178, Ano93-179, Ano93-180, Ano93-181, Ano93-182, Ano93-183, Ano93-184, Ano93-185, Ano93-186, Ano93-187, Ano93-188, Ano94-153, Ano94-154, Ano94-155, Ano94-156, Ano94-157, Ano94-158, Ano94-159, Ano94-160, Ano94-161]. AI [Ker90b]. Aid [Moa94e]. aided [Gib89]. Aim [Kee97a, Wil97a]. Aiming [LR89]. aims [Ano97l, Sna90a]. ain’t [Ano95-30, Ril95e, Sar95, SF96m, Sem94e]. Airine [Dar95o]. airline [Dar95c]. airport [Ano97e]. Al [The96a]. alienation [Ano95-135]. Aligned [Sta95c]. aligning [Ova96]. all-stars [Esk97c]. alliances [Lam90]. allies [McM90g, MK90a, Mod90]. allowing [May96]. Almost [Ano92-115, App97b, Kau96, Sne92f, Wat96a]. alone [Dav89h, Dri96, Kay96b]. Alpha [Sha95, Ano94c]. Alphabetical [Ano95-32]. alphabetically [Ano97-27]. AlphaServer [Ama95a]. already [OHE94, The96a, Var96f]. Also [Ano93-343, Ril95h]. Alternative [Sha95, Var94a, Ano95-69, Ker89c, Var96b, Ano94-181]. Alternatives [Sim96s]. Always [Ano92-71, Sim96b, The94b]. am [Esk97g]. Amalgamated [Ano92h]. Amdahl [Ano96p]. Amencan [Ano95-32]. America [Ano92h, Hay96a]. American [Ano93-305, Ano93-232, Ano95-37, Ano95-38, Ano96r]. Americans [Sta96]. AMEX [Ano92-72, Dit92]. amnesia [Car89c]. Amoco [Ano97x]. among [Ril95h]. amount [Ano96b]. Analog [Mik97a]. analyses [Sem96a]. Analysis [Ano94-170, Ano96m, Ano96l, Ano96w, Ano96-38, Ano97m, Ano97n, Ano97p, Ano97o, Ano97r, Ano97q, Jag96a, Moa95h, Ric94b]. Analytical [Ano92v]. Analyze [Ano97v]. anchored [Ano97-32]. and-most [Dar96n]. and-reporting [App96h]. Andersen [Ano95-76, Ano95-148]. Andrew [Ano92i]. anesthesia [Hay97g]. anew [Kim90a]. angry [Moa91b]. angst [Ano98a]. animation [Ano96f, SF96q]. Anna [Ano96d]. announced [Kno96a]. annual [Ano97-27, Str96b]. anoint [Ano93-350]. ANSI [Ano93-350]. ANSI-supported [Ano93-350]. Answer [Ano93-147, Rym94, Ano95-135, Fry96, McC96d, The96a].

Answers [Ano93-280, The95o]. ante [Kno97k, Ric92b, Sed90a]. any [Fry96]. AnyLAN [Ano93-120]. anymore [Fra89h, Sta93, Ano93-233]. Anyone [Ano94-107, Ril96m, SF96a, SF96e]. Anything [Ano92-115, Bla96, Fre96b, Sne92f]. Anyware [Ano98a]. anyway [Ril96k]. Anywhere [Ano92-115, Sne92f, Ano96-50]. App [Ano92-31, Dar95m, Hab95c, Lin96c, The95j, The95k, Vin95b, Ano96a, Ano96g, Ano96z, Ano97t, Ano97-32, Ano97-44, Dar96e, Fol96b, Fre97d, Rad96b, Ril96m, SF95c, Var96i, Ano92d, Dar96c, McC96k]. App-Ware [McC96k]. appears [Lin96a]. appetite [Kay90]. Apple [Ano92j, Fra89h, Ric90a, Ril96k, SF95g, Sch88b]. Applet [Mee96, Ano96x]. AppleTalk [Jen89a]. appliances [Ano95-85]. Application [Ano92o, Ano92-58, Ano93-295, Ano96a, Moa88b, Ano94-278, Ano95-77, Dar96f, Dar96o, FC96, Fre97d, Hay96c, Hod90]. Applications [Ano92-49, Ano93-70, Ber95a, Ber95b, Ano98g, Bry88, Chi92, Dar96h, Dav90d, Esk97n, Fol96e, Fre97f, Ger90, Gil96a, Gos88, Hay96e, Jon97a, Ker90a, McC96k, McM89a, Moa89i, Ril96i, Sco96, Sim96b, Sim96q, Sim97i, Wat96c, Ano93-66, Ano93-253]. Applix [Ano98a]. APPN [Ano92-165]. Apprentices [Ano92-136]. Approach [Ric94g, Car89e, Ham90b, McM91b, Moa95h, Ano94-242, Ano97n]. Approachable [Ano93-337, Ano93-193]. approaches [Mar95d, Rea95]. approval [Dav89e, Kee97a]. Apps [Ano92-151, Ano92-50, Ano92-67, Ano92-105, Ano93v, Ano93-54, Ano93-144, Ano93-237, Ano93-255, Ano93-294, Ano94-38, Ano94-68, Ano94-284, Ano94-278, Ano96r, Dar95e, Fol96e, Kay95i, Mar95c, McC96k, Ric94d, Sch95i, Sch95j, The95e, Var96h, Ano95-131, Ano96-50, AR96, Dar96h, Fly96a, Gar95a, Kno96a, Mar95b, Ril95i, SF96c, Sha93b, Sim96i, Tay96, The95d, The95q, Var96b, Dar95i, Lin95c, Qui95, Str94n, Var95b]. AppWare [Rym94]. Architecture [Moa88b, AR96, Inm97b, Jen89b, Kno96c, Moa94b, Ste96a, Sch95h, Var95a]. architectures [Hay96d, Sim97n, Str95j]. archival [Ger96a]. ARCnet [Siv89c]. area [TB90]. arena [Ale96c, Ano96k, Sim96k, Sim97n]. aren’t [Ano95y, Kay96b]. Arm [Var96e]. array [Sim96j]. arrays [Sim96k]. arrive [Lei90a, Sim96k]. Arrives [Ano92-32, Pin90, Ano93-49]. art [Cod90a, Cod90b, Mar88]. article [Dar96d]. article-there [Dar96d]. arts [Ano94-278]. AS/ [Ano92j, Ano92-29, Ano92-82, Ano92-92, Ano92-148, Sta89a, Ano92o, Ano92p]. AS/400 [Ama95b, Ano92-60, Ano93-84, Ano93j, Ano93-172, Ano93-296, Ano93-345, Ano94d, Ano96b, Ano96n, Ano97m, Gib97a, Kre93a, McC93e, Nee95, Sim96p, Sim96q, Sim96s, Tay96, Hok96]. AS/400-based [Sim96q]. AS/400s [Ano93-230, Sim96c, Sim96l]. Ashton [McM90a]. Ashton-Tate [McM90a]. Asia [Ano92n, Joh90c]. Asian [Ano92m]. Ask [Fry96, Hen94, Kad97, Dar96r]. asked [Hir96]. Asking [Moa89a, Tay96]. aspects [Sch88a]. assemble [Dar96m]. Asset [Ano93m, Bla96, Fra93a]. assets [Dit92, Ano92-72]. Associated [The95p]. Associates [Ano95r, Car89b, Car90a, Ano92-144]. assorted [Tuc97c]. assurance [Joh90a]. AST [The92a]. ATA [Dar95c]. Atlanta [Sim96l]. ATM [Ano93-293, Ano93-341, Ano95x, Ano95-56, Ano96d, Dar96i, Dar96l, McC96a, McC96d, Ano93k, Ano93-254, Ano93-335, Ano94e, Dar95t, McC95b, Str94d]. ATMs [SF96r]. attack [Ale96b]. attacks [Ano95-131]. attorneys [App97b]. attract [Bel90]. attraction [Esk97d]. attractive [Ano93-311]. Audio [Ano93l, Hal96]. Audit [Ril97b]. Austin [Ano92q]. Australian [Hea89]. authenticate [McC96j]. Authentication [McC96j]. Automate [Ano93m, Fra93a, The95n, Dav92b, McC95j]. Automated [Kel90b, Asb97b, Hay97r]. Automates [The96a, Ano96d, Dav90a]. automatically [Ale96b, Ano97v]. Automation [Ano93b, Ano96-31, Ano95-69, Asb97g, Asb97b, FT88, FR89, Hod89c, Moa88d, The94c, Tra90b]. Availability [Str94j, Ano97z, Fol96f, Sim96c]. available [McC96h, OHE94, Sim96p]. Average [Kno97i, Ano97-45]. AViiON [Ano93n]. Avoid [Ano92r, Ano93o, Str96c, Jam97b, Kay96b]. avoiding [Gre96, Sta96, Ano93p]. await [Ril96j, Sch88b]. award [The96a]. Awards [Str96b]. Aware [Kay96d]. awash [SF95j]. away [Ano96-54, Gil96a, Ril95e, Var96f].

B [Ano92-42]. Back

[Ano93-46, Bla95a, Fol96a, McW89a, Pan96a, SF97a, Str94a, Ano95-113, Fra88, McM90a, Pan96c, SF96a, Sem96i, Ano95-81]. Back-Office [Fol96a]. back-up [Fra88]. Backaches [Ano93r]. Backbone [Ano93-297, Dar95t, Dar96i, McC96e, Ano96-48, Dar96l, McC96g]. Backbones [McC96a, Str94d]. Backing [Bau94a, Lin95b]. BackOffice [Fol96f, Sem95a, Str95f]. Backup [Ano93r, Ril95d, Sim96k]. backups [The95p]. bad [Ano95-149, Jag96a, Ril95j]. Bag [Win95]. Bagging [The95b]. balance [McC97e]. Balancing [Dar90, Hab89]. balloon [Dar96r]. ballot [Ano95-138]. Banco [Dar96r]. Bancom [Ano92s]. Bandwidth [SF97b, Ano96i, SF95e, Sch97d, Ano92t]. Bank [Ste90, Dar95h, SF96r, Ano93-304, Ano96g, Ano97j]. BankAmerica [Ano93-170]. Banker [Wil97b]. bankers [Bel90]. banking [Ano97-48, Eth89]. bankrupt [Hur94a, Ano94-277]. Banks [McW88a, Moa89j]. banner [dJ96a]. bar [Mee96]. bar-chart [Mee96]. Barbarians [Cas95b, Cas95c, Cas95a]. barbecue [Str95e]. Bare [Ano92-160, Sne92a]. Bargain [Gut94, Rub95b]. bargains [McC94c]. barrier [Bun90]. Barriers [May96, Moa89c]. bars [Inm97b]. base [McC89b]. based [Ano95-47, Ano96f, Ano97-41, Ano97-48, Fra91c, Fre97d, Gil96a, Kno96a, Lin96a, MF90, Man92b, Sha96, Sim96q]. Basement [Gut94]. Basic [Ano93-336, SF95c, Fly96b, Sch89b, Ano94-106]. Basics [Str94a, Ano92-168]. Bass [Ano94f]. bathwater [Mik97f]. Battle [Ano93-300, Ano93-272, Fly96a, Ker90c, Ano93-297, Ale95c]. Battling [Ano93-278, Ano93q]. Baxter [Dar95d]. Bay [Sch95d, Sch95e]. Be [Ano93-327, Ano93-352, Ano94-172, Ano94-283, Dai95, Moa95b, Sim95g, Ale96c, Ano93-350, Ano94-215, Ano95t, Ano95w, Ano95-56, Ano95-113, Ano96q, Ano96t, Ano96-52, AR96, Dar96a, Dar96b, Dav97d, FGR88, Fry96, Ger96a, Gil96a, Hal96, Hen95, Hir96, Inm97b, Lin96a, Lin96b, McC96d, McC96e, McC96j, Mik97h, Mod90, Ril95g, Ril95l, Ril96g, Ril96h, Ril97e, SF95j, Sco96, Sem96b, Sim96n, Sim96t, Sta96, Str95j, Tay96, The96b, Tuc97c, Var96b, Var96c, Wes90]. BEA [Ano97r]. beachhead [Wil97b]. beast [Dal98]. became [Bea96]. because [Ano95-82]. Become [Mae95, Ger96b, Gut96, Jam97c, OHE94, Tuc96a]. becomes [Asb97d]. Becoming [She95, Ano95-126, Bal96b, Joy89a, McC96m, Scr95b, Sim96p]. bed [Siv89d]. bedfellows [Tuc98b]. Beef [Dar96n, Sch89c]. Beefs [Ano93-235]. been [Ano95-128, Fre97d, Hay96c, Hay96e, Lin96c, SF95d, SF96d, Ste96b, Str95j]. Before [App96b, Hur94b, Mar95c, Ril96b, ZNZA95, Joy89a, Mar95b, Mar95d, McC96b, SF96a]. begin [Sim96l]. begins [Hod90, Jag96b, SF95d]. Behind [Ano92-138, Ril96b, Ker90a, Sim96l, Lee95]. Being [Ano95-56, Ano95-128, Gil96a]. Believe [Ano95r, Tuc97c]. Bells [Ano94-280]. bench [Ano96x]. Benchmark [Ano96n, Can88, Ser89]. Benchmarking [Ano94-242, Sch98, Mik98a]. Benchmarks [Ano92u, Ano94g, Ano94h, Ano94u, Ano94v, Ano94i, Ano94j, Ano94k, Ano94l, Ano94m, Ano94n, Ano94o, Ano94p, Ano94q, Ano94r, Ano94s, Ano94t, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g]. benefit [Moa95a, RK89, Var96a, Kee98a]. Benefits [Str95m, Inm97b, Kay96a, Kee97d, Sem94a, Var96g, de 97d, Ano94-109]. Best [Ano92-149, Ano92-90, Ano93-301, Ano94-250, Jon98, Str95g, Ano95-138, Ano95-162, Ano96-54, Ano97n, Bro89, FR89, Ric94a, Ril95i, Ril96p, Sha96, Sim96m, Sne92c, Str95i, Ano95-148]. best-laid [Ril96p]. Best-selling [Str95g]. bet [Ano96-54, Fra89e, Sim96e, Ano92-139, Sem95a]. beta [Sim96g, The95c]. Better [Ano93t, Ano93-50, Ano95-72, Hir96, Liv95, SF96q, Win90, Ano96-51, Ano97-36, Ano98g, BF90, Dar96f, Dar96o, Dav96a, Fre97d, Joy89a, Kee98b, Rad96c, Ril96d, Str96a, The94b, Var96d, Wra98, Ano95-98]. Betting [Joh90b, Sim96e]. between [Ano95-162, Cas96, McC97e, SF95d, Str97a, Tuc97g]. bevy [Kno96a]. beware [Inm97b]. Beyond [Ano92-108, Ano92-172, Ano93-68, Ano96u, Dav97b, Esh89, Gol95, Liv92, Str94e]. Bids [Ano92-102]. Big [Ano92-42, Ano92-79, Ano92-152, Ano92-135, Ano92-136, Ano93s, Ano93-286, Ano97r, App96a, Hab96, Mae95, Moa89l, Pan96a, SF95g, Sha93a, Sim97c, The94b, Ano95-56, App92a, Dav90f, Hay98b, HE90, Hir96, Jag96a, Kre93b, Lin96a, SF95c, SF96f, Sim96d, Sim97o, Str97a, Tuc97c, War97, Ano93-121, Fol96d, Sem94b, Sim97b, Var96j, Ano93r]. big-bucks-generating [App96a]. big-deal [App96a]. Big-LAN [Ano93r]. big-league [App96a]. big-time [SF95c]. big-volume [App96a]. Bigger [Ano93-50, Ano97-36, Mae96, Ano96-51, Kno96f, Pan96a]. biggest [Fry96, Kee98a, Sim96t]. Bill [Esk97h]. billion [Dav89a, Jon97a, Kno97b, Mar92]. binary [Gil91]. Birth [Fre97a]. bit [Fol96b, Fol96d, Ril96f, Var96a, Ano93-191, Coh95, Var95b]. Bits [Var96a, Hal96]. Biz [Ano92v]. Black [Dar96d, de 98a]. Bleeding [Hal96]. bless [Ano93-350]. blessing [DB97]. blips [Sim96u]. BLOBs [Ano93q]. bloc [Pur89]. Blood [Con97]. blossom [Fre98b]. blown [McC96d]. Blue [Moa89l, Tat89b, War97, Ale96a, Ano92w, Ano96t, Hur96, War97]. Blueprints [Bal96a]. blues [Ano97-38]. blurry [Sim96u]. board [McW89a, Dar96b]. Boards [Ano93x, Ano95-127, Woo93]. boat [Ano95-70]. Bob [Ril97g]. body [Tra95]. bogus [Dar96c]. Bomb [McC96g]. Bonding [Ano93t]. Bonkers [Ric94f]. book [Str97a]. books [Ano95t, Sar95, Str95g]. Boom [Ano93-153]. boomed [Tuc96a]. booming [Ano95-29]. Boomtown [Sem95h]. boon [Ste96b]. Boost [Ano92-122, Hay97k, Sne92e, Tuc96a, Ano92-113]. Boosts [Ama95a]. borders [Car90b, Esh89]. Borland [Ano92x, McM91a, Vin95a]. Born [Ano93-294, SF95k]. Born-Again [Ano93-294]. boss [Hay97j, Ano95-102]. Boston [Dar96j]. Both [Hok96, Ale96c, Sim96k, The96d, Ano95-162]. bottleneck [Sim96o, Sim97d, Eld95]. Bottlenecks [Ano93-70]. bottom [Ano95-132, McC97c]. bottom-line [Ano95-132]. boundaries [Cas96]. Box [Ano95-70, Ano95-71, Ano95-82, Ano95-73, Ano95-72, Ano95-74, Ano95-75, Ano95-76, Ano95-69, Ano95-77, Ano95-78, Ano95-79, Ano95-81, Ano95-80, Ano96-31, Ano96-32, Ano96-33, Ano96-34, Dar96d, Dar96f, Ano96w, Ano94-125, Ano94-126, Ano94-113, Ano94-114, Ano94-115, Ano94-116, Ano94-117, Ano94-118, Ano94-119, Ano94-120, Ano94-121, Ano94-122, Ano94-123, Ano94-124, Ano95-61, Ano95-62, Ano95-63, Ano95-64, Ano95-65, Ano95-66, Ano95-67, Ano95-68]. BPR [The95n]. Brace [Ano96u]. Branch [Ano93-271]. Branches [Ano94-274, Ano94-277, Hur94a]. Branding [Lin95d]. Bravado [Ano93-206]. brave [MF96]. Bravo [Ano93u]. Brazilian [Moa88c]. Break [Ano93-249, May96, Sne91a]. breakdown [Ano97-27]. Breaking [Gut95a, Gut95b]. breakneck [Ano96-52]. Breakout [Ano96-28]. breaks [Bun90, Moa90g]. breed [Ano96j, McM90h, Sim96u]. breeze [Ano95-79]. Bridge [Ano95-75, Tuc97b, Sta89b]. BridgeWare [Ano95-47]. Bridging [Sno98]. Brillant [Ric94f]. bring [Ano97l, TMS96]. Bringing [McC97a, FC96]. brings [Ano97v]. broader [McC90a]. browser [Kno97m]. bucks [App96a, Sim96d, Sim97o, War97]. budget [Hod89a, Vog96b]. budgeting [Fre97h]. Budgets [Ano93-318, Ker88c, Gut94]. Bug [Ano96-45, Mar95c, Mar95b]. Bug-Free [Ano96-45]. Bugs [Ano96-46]. Build [Ano92-67, Ano93q, Ano93v, Ano96r, Ano97-42, Dar96c, Dar96a, Fis96b, Ham96, Kno96c, McC94a, Sem95d, Sim96a, Vog96a, Ano96m, Kig96, Kno96a, Kno97j, MF96, Mee96, Sta95c, Ano94w]. Builder [Ano93e]. Builders [Ano93-212, Fol96b]. Building [Ben98, Bes90, Dar95e, Dav93a, Hab97, McC96b, The95d, Ano96f, App96e, McC96h, Pan96b, SF96e, SF96r, Sha96, Ano93w]. Builds [Dar96b, McC96e, Ano95-102, Ano96-50]. built [Dar96a]. Bull [Ano96r]. bulldoze [AR96]. bullet [de 97g]. Bulletin [Ano93x, Woo93]. Bulletproof [Abr91]. bundle [Mun97, Sim97o]. bundled [Fol96a, Sim96r]. Bundles [Ama95c, Ama95d]. bungee [Kno96d]. Buoy [Ano92-102]. Burden [Ano93-306, Hur95b, McC93c, Sta90b, Ano93-75]. buried [McM90c]. Burnout [Bau95c]. Bus [Ano93-279, Ano94-200, Eld95, Fra89i, Fis96a]. Business [Ano92y, Ano92v, Ano93x, Ano93-142, Ano93-270, Ano93-287, Ano96c, Ano96d, Ano96g, Ano96v, Ano96-53, Ano97c, Ano97d, Ano97e, Ano97j, Ano97l, Ano97i, Ano97k, Ano98a, Bau95d, Fis96b, Fol96e, Kay95b, Kay95e, Lic94, SF96i, Sco96, Sem96b, Sem96g, SF96r, Sha94b, The95o, The96e, Woo93, Ano94-278, Ano95-76, Ano95w, Ano96-54, Asb97a, Dar95d, Dar96b, Dug96b, ES97, Esk97h, Fra90k, Gut96, Hay97l, Hen90, Jag96a, Kay95a, Kee97b, Kno96c, Kno97c, Kor93, Lei90d, LR89, Maj90, Moa89h, Ova96, Sem96h, SF95l, Var96b, Vog96a, Ano93-52, Ano97n, Ano97o, Ano97x, Moa94a, SF96r, Str95m, Ano95-29]. business-drive [Dar96b]. business/academia [Lei90d]. businesses [FC96]. Bust [Sim97d, Car90b]. busters [Sne94a, Sne94b]. busts [Sim96o]. busy [Ano95s, Lin96c]. Button [Lin94a, Ano94x]. Buttoned [Ano92z]. Buttoned-Up [Ano92z]. Buy [Bow94, Cor95c, Kno97j, Sem95d, Ger90, The95u]. buyer [Ano97-45, Der92, SF95j, Ano93y, Ano93z]. Buyers [Sim95h, Ano95-136]. Buying [Hur94b, Moa90b, Dar95d, Var96f]. Buys [Ano93-246, Fra93d, The91]. buzzwords [Ano96-32]. bytecode [LP96]. Bytes [Sem96f, Hal96].

C [Ano93-193, Ano93-337, Ano94-284, Ano95-133, Ano96r, Ano96f, Car97, FC96, Sem95d, Vin95c, Ano93-337, Ano94-38]. C/S [Ano94-284, Ano95-133, Ano96r, Ano96f, Car97, FC96, Sem95d, Vin95c, Ano94-38]. C10 [Dai95, Mae95]. CA [Ano93-28, Ano96-37, Car90a, McC96h, Ano93-27, Ano93-44]. CA-Unicenter [Ano96-37, McC96h]. cable [Ano96i]. CAD [GFR89, Lin92a]. CAD/ [GFR89]. CAE [GFR89]. Cairo [SF96k]. calculate [Sim97f]. Calculating [Moa95a]. calculation [Ano95-80]. Caldera [CS95]. Calendar [Ano92-44]. Calendars [Ano92-95, Pal92]. California [Ano95-78, Dar96b]. Call [App96b, MF96, Asb97b, SF96d, Sim96v, Siv90a]. Called [vSS96, Hir96]. Calling [Mik97b]. calls [Var96d]. Calm [Var96c]. CAM [GFR89]. CAM/ [GFR89]. came [Ano96-47, SF96a]. camera [McC97a]. Campbell [Siv90h]. Can [Ano92-104, Ano93-117, Ano95-85, App96c, BS94, Dar95p, Dav89b, Dav89h, Dri96, Fol96a, Ham90a, Ker88a, Kno97a, McC95c, Moa88a, Moa91a, Pin92, Rad94b, Rad96a, Sha94c, Sim96b, Tab92a, The95n, Tra89a, Tuc97c, Ano95-79, Ano95y, Ano95-98, Ano96-54, Ano97p, AR96, Cor96, Dar96b, Dar96r, Fis96b, Hal96, Hin93, Inm97a, LP96, McC96a, McC96e, McC96j, Rad93, Rad94a, Ril95e, Ril96g, SF96a, Scr95b, Sem96b, Sim96d, Sim97o, Ste92b, Str95i, The96d, Var96a, Var96b, Var96d, Vog95, Ano92-27, Ano92-28, Ano92-29, Ano93-28, Ano93-29, Ano94y, Moa95b, Sem94c, The95e]. candidate [Ano98f, Sim96c]. Can’t [Sem94h, Cor96, Rad96a, The96c, Sim95d]. Capacity [Ano93d, Bry89a, McC96l, Sim96c]. capacity-planning [McC96l]. Card [Ano93-63, Ano98c, McC93a]. Cards [Ano92-83, Wil97d]. Care [App96g, Moa94a, Inm97b]. Career [Ano97f, Moa95i, Dav89b, Fra89a, Kno97b, Ano93-30, Ano93-31, Ano93-32, Ano93-33, Ano93-34, Ano93-35, Ano93-36, Ano93-37, Ano93-38, Ano93-39, Ano93-40, Ano93-41, Ano93-42, Ano93-43]. careers [ZSZ88]. carefully [Wil97a]. Case [Ano96e, Kay95b, Kay95e, Sem94d, Kay95a, Ril96j, Wil97e, Ano92f, Ano93-151, Ano93-280, Ano95-162, Car88c, Car90c, Gri89, HC90, LN89, Moa89l, Sch90c, Sem90, Sta89g, Web89, Ano93z, Ano93-51, Sem94b]. cases [Kee97b]. cash [Fre96a, Hin93]. Cast [Ano95-138, Roc97a]. Casts [Ano92-103]. catalyst [Ras89]. Catch [Kno96a, SF96h, Ano93-208]. Catch-Up [Kno96a, Ano93-208]. Categories [Ano93-103]. Cavalry [App96b]. caveats [Dri96]. CAX [GFR89]. CCITT [Emm89a]. CD [Ano94-163, Fra93b, Pin91, Ste95c, Ste95a, Ano94z]. CD-ROM [Pin91, Ano94z]. CD-ROMs [Ano94-163, Fra93b]. CeBIT [Ano97s, Ano94-27]. Cellateral [de 97d]. Center [Ano93-317, Ano93-343, Asb97b, Bun90, Dav90a, FP88, Fre97h, Moa88d, Pet88, Ste89, Tab92b, Wal89b]. centers [Ham96, Hir96, Ano93-320]. centralized [Ano95-30]. Centre [Jua90]. centres [Hea89]. Centric [FC96, SF96q]. century [Ano95-80, Jag96b, Joy89b, Ano92-80, Jon97c]. CEO [Moa94c]. certain [Inm97a, SF95f, Inm97a]. certainly [App96c]. certification [de 98b]. Certify [Ano93-113, Moa93c]. chain [Asb97f, Fre97g]. Challenge [Wil94, Hur96, McW88b, Meh88, Tay90, Ano94-127, Kay95c, Kay95d, Ano93-350]. Challenges [Ano93-287, Str95h]. chance [Fol96g, Fol96i, McC96b]. Chances [Kay96a, Var96g]. Change [Dar95k, Cus93, Dar96g, GH88, Ker89j, MBGS88, MB94, Rea95, Ros89, Sco96, Ano93-312, Ano95-136]. changed [Sco96]. Changes [Ano92-79, Ano93-192, Ano97j, App92a, Esk97h, Rob96, Ano92-66, Ano93-316]. Changing [Ano92-63, May96]. Channeling [Ano92-30, Par92]. channels [Van90]. Chaos [Sne94b, ES97, Hay97f, Pal93, Sne94a, Ano93f, Lin94c, Rym94]. charge [Sta88]. Chargeback [Bus88]. charging [OHE94]. Charm [Ano94-267]. Chart [Ano92-31, Mee96]. Chase [Ano93-65, Ano93-45, App96h]. chat [Ril96o]. Cheap [Ano93-322, Ano94-251, Ric94e, Val96, Dav97f, Ril97c]. cheapboxes [Ano96-42]. Cheaper [Ano93-342, Ano94-73, McC95e, Str94b, Fra93g, Hen96, Kre93c, Str94f]. Cheapest [Dur94e]. Cheaply [Ano93-274, Str93]. Check [Ano94-220, Ano95-78, Cha95, Hur94b, The95f]. CheckFree [Gib98]. checks [Fra90i]. cheer [Dar96c]. chemical [Ost90]. Chevy [Hen95, Ril95i]. Chicken [The94c, SF95i]. Chiefs [Ano93-341]. chip [Ano95-47, Str95e]. Chips [Ric94e, Var94b]. choice [Bal96b, Dav90e, Hin93, Inm97a, OHE94, Rad96c, Sim96n, Sim96p, VM96a]. choices [Ril96g, Sem96a]. chokes [Sch97a]. Choking [Ano94-69]. Choose [Ano97g, Hec97, Kay96c, Ale96c, Dav97h, Roc97c, ZSZ88, KK95]. choosing [Pan96a, Wat98]. Choreography [Cha95]. chose [Ano95-162]. churn [Roc97f]. churning [Var96f]. CICS [Ano93-49, Ano93-212, Ano94-107, Ker89e]. cigar [Cor95a, Sch95a]. CIM [Bou90, Heb90, Siv90h, Wes90]. Cinch [Kol94]. Cinches [The94h]. CIO [Kno97b, Ano92-172, Ano93-267, Mik97h, Ano94-279]. CIOs [Ano93-232, Ano93-46, Str94a, WP94]. Circle [Mik97a]. circuiting [Blo89b]. circus [Dar96c]. Cisco [Ano96n, Ano96l]. Cisco/StrataCom [Ano96n]. City [SF95g]. claim [Hal96]. claiming [Tra95]. claims [Dar96c, Sed90a]. clarifies [Ano95-71]. Clark [Fra90c]. Class [Ano93-282, Ano95-77]. Clean [Ano93-114, Ano94-28]. cleaned [Str95i]. Cleans [Ste96b]. Clear [Ano96n, McM90b, Sim97f]. CLEO [Ano93-47]. Client [Ano92b, Ano92-32, Ano92-93, Ano93-27, Ano93j, Ano93-272, Ano93-313, Ano93-294, Ano93-328, Ano94-250, Ano94d, Ano94-272, Ano96f, Ano96-52, Bel90, Fra90a, Gar95a, Hab95a, HS94, Ham94a, Ham94b, Hen94, Kay95l, Mar95c, Nee95, Qui95, Rad94b, Sch95f, Sem94c, Sim95e, Sne94a, Ste96a, Str94f, Str94m, Ano90a, Ano95-160, App95a, Bal96a, Bal96b, Dar96p, Dar96r, Dri96, Fra93f, Lin96c, Mar95b, MB94, Pan96b, Rad94a, Ric92c, Ric94a, Ric94c, Sco96, Sim95b, Sim96l, Siv89a, Str94b, Ano92-42, Ano92-163, Ano93-48, Ano93-49, Ano93-50, Ano93-51, Ano93-317, Ano93-65, Ano93-153, Ano94-29, Ano94-30, Ano94-217, Ano97q, App95d, App95e, Bau94a, Gar95b, Hab95b, Sem94d, Sem95b, Sem95c, Sne94b]. Client-Server [Str94m, Siv89a]. client/ [Sco96]. Client/Server [Ano92-32, Ano92-93, Ano93-27, Ano93j, Ano93-272, Ano93-313, Ano93-294, Ano93-328, Ano94-250, Ano94d, Ano94-272, Ano96f, HS94, Ham94a, Ham94b, Hen94, Mar95c, Nee95, Qui95, Rad94b, Sch95f, Sem94c, Sim95e, Ste96a, Str94f, Ano96-52, Bel90, Fra90a, Gar95a, Hab95a, Sne94a, Ano95-160, App95a, Bal96a, Bal96b, Dar96p, Dar96r, Dri96, Fra93f, Lin96c, Mar95b, MB94, Pan96b, Rad94a, Ric92c, Ric94a, Ric94c, Sim95b, Sim96l, Str94b, Ano92-42, Ano92-163, Ano93-48, Ano93-49, Ano93-50, Ano93-51, Ano93-317, Ano93-65, Ano93-153, Ano94-29, Ano94-30, Ano94-217, App95d, App95e, Bau94a, Gar95b, Hab95b, Sem94d, Sem95b, Sem95c, Sne94b, Ano97q]. clients [Ano95-129, SF95f, Sco96]. Climbing [Fra89a]. cloaked [App97b]. clogged [McC96g]. Clogs [The95g]. Cloned [Moa95b]. Close [Dug96a, Als88, Cor95a, Dar96m]. close-up [Als88]. Closer [McC97c]. Closes [Ano93-27]. cloud [Str95e]. clouded [Bal96b]. Cluster [Sim95d, Sim96c, Ano96z]. clustered [Ano97v]. clustering [Sim96c]. Clusters [Ano96n, Ano96p, Sim97h, Ano93-330, Ste94b]. CMOS [Sim95f, Sim96d]. Co [Sch94]. coalition [Dar96q, Dar96q]. COBOL [Ano92l, Ano93-338, Ano93-52, Ano92-62]. cockroach [Ano95-130]. CODA [Ano92-33]. coddle [McC96c]. Code [Ano93-344, Ste96b, Win95, Kau96, SF96c, Var96c, Ano94-240]. Coding [Bau95d, Ano93-255]. collaborate [Rad96c]. collar [Siv90i]. collect [Ano97t].

Collection [Ano92z]. college [ZSZ88]. Collide [Hur95c]. colonize [Mus88]. Color [Ano92-126, Ano93s, Lin91a, Rub95a, Dav92a, Sha93a, Ano92-34]. Column [Ano93-122]. columnist [SF96b, SF96g]. columnists [SF95g]. COM [SF97g]. combine [SF96j]. combining [Cas96]. Combo [Var96b]. Combos [Ano94-112]. COMDEX [Ano93-301, SF95f, Dat94]. Come [Ano93s, Pat96, Bus88, Hay96c, Kno97l, OHE94, Ric92c, Sha93a, Tuc97c]. Comeback [Ano92-75, Sne91c]. Comes [Ano93-48, Ano93-67, Ano93-89, Rad96c, Ano96w, Fol88a, Fra91b, McM90i, Mik97a, Ril95k, Sim96s, Tat89c, Wat98]. Coming [Ano92-169, Fol96e, Var96g, Ano96t, Bar90c, Fis96a, Fol96a, Hen96, Ker89a, Sim96k, The96d, Ano92-35]. Command [Ano92-172]. Commentary [Ano94-31]. Commerce [Ano96p, Ano97n, Ano97q, Hay96a, Sem96i, Tuc97n, VM96b, Var96j, VM96c, Wil97d]. Commercial [Ano93-349, Ano95-77, Gil96b]. committed [App96e]. committee [Ano97-31]. committees [RR89]. Communicate [Sch95i]. Communications [Ano93-47, Dav90c, Hab89, Rub95c]. Communicator [Ano94a]. compact [McC90d]. Companies [Ano93-110, Ano96-36, Ano89c, Ano93-311, Ano95-28, FM97, Hay96e, Mae96, McC97c, McC97e, Var96d, Ano93-111]. Company [Ano95-32, Ano95-31, Ano96-28, App96a, Fis96b, Fry96, Hab96, Ham96, McC93e, Ost90, Ril95c, Sem96b, Sem96e, The96e]. Compaq [Ano92-36, Ano92-37, Ano93-53, Ric93a, Sim96e]. Compatibility [Ano96q]. compelling [Sim96n]. Compete [App96i]. Competency [Sta95c]. Competing [Kno96c, Kno96e, Kno96d, Kno96f, Kno97i, Kno97j, Kno97k, Kno97l, Kno97m]. Competition [Ano96-44, Ano97k, Asb97b, MK89, Sem96e, Siv90b]. Competitive [Ano93-256, Ano94-170, App96i, Inm97a, Joy89a, MCBH88, Moa90c, Ric94b, Sem96b, Var96f]. Competitiveness [Ano94-127, Fra89b]. competitors [Fol96c]. compiled [LP96]. complete [Kno97g]. Complex [Kay95i, Str94d, Sna90b]. Complexity [Ano96-44, Kay96f]. compliant [Jon97a]. Component [OH95, Sem94g, Fre97d, OHE94]. component-based [Fre97d]. Components [SF95l]. Compound [OHE94]. Comprehensive [Ano95-34]. Compression [Ano92-27, Ano92-55]. comprised [Fol96e]. Computer [Ano92y, Ano93-149, Ano95r, Car89b, Car90a, Hab96, Pic94a, Ano96t, Bar90a, Blo89b, FM89a, Fol93, Gib89, Hod88, Jag96b, Kay96c, Lin96a, Sar95, Sch89b, Str95j, Ano94-32, Ril96k, Sim95i, Sim96u, Str97a]. computer-aided [Gib89]. Computers [Ano93-325, Ano95-127, Cor95b, Lin93, Ril96d, Sim97o, Str97a, Sim96u]. Computing [Ano93-250, Ano93-328, Bla95b, FC96, Mar88, McC97f, Pan96b, Sem94c, Sim95e, Var96b, Ano95-59, Ano95-125, Dav90c, Kay96f, Lib88, McC94e, McW89b, MB94, Pin90, Ric94a, SF95f, Sco96, SF96q, Siv90a, Sta88, Ano94-250, Mel95, Str94m, Ano95-127]. Compuware [Ano96-49]. concentrate [Scr95b]. concerns [Hay96a]. concert [Ker89g]. concurrent [Ril95j]. Conferencing [Ano93-142]. confesses [Mit96]. conformance [Ano97t]. Confused [Rad96c]. confusion [Coo92a]. Connect [Jef97a, Lic94, Ric92a, Dav90b, Tit96]. connected [Ano97-48, Sim96l]. Connection [Dar96k, Dar96t, Sch90b, Ano93-235]. Connections [Hok96, Ano95-30]. Connectivity [Ano93-328, Dur94c, SF96j, Sim96s, Str95k, Ano93-47]. conquer [Cor95b]. Consider [Ste96a, Ano97-49, Sim96i, VM96a]. Considerations [Ano93-268]. considered [Ano95-128]. Considering [Tay96]. consistently [Dar96d]. Consolidate [Cor95b, Dur94c]. consolidation [Car90h, Cas96, McM89b]. constraints [Sch97a]. Consultant [The96b, Wat96a]. consultants [Bry89a, Kay96b, Kay96c]. Consulting [Ano95v, Ano95-148, Ano96-36, The96b]. contacts [The96b]. contender [McM90e, McM91a, Ric90d]. contenders [Ale96c]. Content [Ano96n, Ano95-30]. contest [Dav96b, Ano92a, Ano94-218, Ano95-138]. continue [Ano93-311, Fol96e, Sim97n]. continues [Hay96e]. contract [McC94c, Ano95-159]. Contracting [Moa89b]. contracts [Abr91, Str95b]. Control [Ano92-143, Ano93-101, Ano93-295, Ano93-296, Ano94-245, Ano94-246, Ano96-53, Ano97-33, App95f, Dar95f, Dur94d, Kay95g, Sta89d, Str94l, Bla96, Dar96d, Dav90g, Fra90d, FM97, Kay96f, Kel90d, McC96a, Rad92b, Ril96h, Ril96i, Ano92-138, Ano92-172, Ano93-54, Kay95h, Str94g, Vin95a, Ano92-38]. Controls [Ano96r, Ano95-130, Ano92-39]. conundrum [CC89, Sim96t]. conversion [Ano97u]. Conversions [Ano92-81]. converted [Var96c]. cookin [Siv90h]. Cooking [The94c, Tuc98a]. cool [Ano95-57, Fol96b, Sim97a]. cooled [Pan96b]. cooperate [Dar96m]. Cooperative [Siv89a]. coordinating [Wil90]. cop [Ril97f]. Cope [Bau95c]. copier [Bla88]. Copy [Kno97b]. copying [Mar89c]. Corp [Ano93-256, Sta89d, Ano92i, Ano92w, Ano92-76, Ano92-114, Ano92-119, Sem94f]. Corporate [Ano92-30, Ano92-40, Ano93-171, Kau96, Sim96n, Ano95-29, Ano96k, App93a, Car89c, Esk97k, Fly96a, Fra90b, Fry96, Hay96a, Ker89f, Ker89g, Moa90i, Par92, Ano93-55, Ano93-56, Ano93-57, Ano93-58, Ano93-59, Ano93-60, Ano93-62, Ano93-61, Ano94-33, Ano94-34, Ano94-35, Ano94-36, Ano94-37]. corporate.com [Ano94-269]. corporation [Tuc97c]. corporations [McC96h, Sim97f]. Corral [Sim97e]. Cos. [Dar96o]. Cost [Ano92-161, Ano96n, Fre97d, Gib98, Gre88, Hay97p, Jon97a, Kay95g, Kay96f, Kel90d, McC96c, McC92, Ric90a, Tay96, Ano92-87, Ano93-80, Ano94-163, Hay96a, Kay95h, Str94j]. Cost-Management [Ano96n]. Costs [Ano94-29, Ano96u, McC96d, Sim95e, Ano96g, Asb97f, Bla96, Cor95b, McC96j, Rad96a, Sch95a, Sim96r, Sim96s, Sim97f, Sim97e, Sim97m, The96d, Var96f, Ano93-306]. Could [Ano93-352, SF96r, FM97, Ger96a, Kno97b, McC96o, OHE94, Sem96f, Sta96, The96a]. couldn’t [Mae96]. count [Ano96-54]. Countdown [Sch97c]. counted [FGR88]. Counting [Esk97e]. course [App96a, Inm97b]. Cover [Ano97h, Jam97a, Sim96f, Sch95b]. coverage [Str95k]. covered [Ano95-70]. cow [Ril96f]. coworkers [Ano96s]. Cows [Ril96f]. CPU [Ano96-51]. Cracks [Ano93-252]. craft [Ano94-278]. Crazy [Dur94c]. create [McM89a, Str95i]. created [Ano93-189, Ano95r, Can88]. creates [Bar90a]. creating [Ril96g, SF95g]. Creation [Kno97e]. creator [McC96f]. Credit [Ano93-63, McC93a, Ano97d]. creepy [Ril96h]. crime [Blo89b]. Crisis [App96d, Moa94i, Dam89, Hay97t]. Critical [Ano93-237, The95j, The95k, App96c, Dar96h, Fol96f]. crooks [Ano95s]. Cross [Ale96c, Ano93-101, Ano97u, Sim96i, War97, Moa89f]. cross-functional [Moa89f]. Cross-Platform [Ano93-101, Ano97u]. Cross/Blue [War97]. crossroads [Car89b]. Crown [Ano92x]. Cruisers [Ano96r]. Crumbles [Moa95i]. Crunch [The95h]. CSP [Moa88b]. CTI [McC96c]. Cultivating [Ano92-41]. Cultural [Moa89c]. culture [Tom88]. cultures [Ker89f]. Cup [Ano97e]. cupboards [Dar96n]. cure [McA89, SF95i]. Cured [Ano93f, Pal93]. Cures [Moa94h, Wal89a]. currency [Kno97e]. current [McC96k, Ril96e]. Curriculum [Rid88]. curve [Fre97d]. Custom [Ano93v, Ano95-69]. Customer [Ano92-134, Hab96, The96d, Ano97-48, Ano98a, Asb97b, Dug97b, Joy89a, McC97c, Moa93e, Siv90d, Var96f, Var96d, Ano93-145, Sch95h]. Customers [Ano93t, Ano93-64, Ano93-87, Moa95c, Ano95-29, Ano97-30, Asb97b, Dar96c, Dar96t, Fol96g, Ger96b, Kno97i, Kno97f, McC96c, McC97c, MK90b, Moa90h, The96d, Var96f]. Cut [Ano93-318, Jef97b, McC96d, Sch95a, Sim97f, Cas90a, Var96d, McC95d, Sim95e]. cuts [Jag96b]. Cutting [Ano92l, Ano95v, Ano96g, Ano96v, Ano96y, Ano96x, Ano96h, Ano96i, Ano96j, Ano96k, Ano96s, Ano96n, Ano96m, Ano96l, Ano96w, Ano96o, Ano96p, Ano96q, Ano96z, Ano96r, Ano96t, Ano96u, Ano97j, Ano97l, Ano97i, Ano97k, Ano97m, Ano97n, Ano97p, Ano97o, Ano97r, Ano97q, Ano97s, Ano97t, Ano97u, Ano97v, Ano97z, Ano97y, Ano97w, Ano97x, Ano98c, Ano98b, Coo92a, Gri89, Mit96, Bla96, The96d, Val96, Ano94-39, Ano94-40, Ano94-41, Ano94-42, Ano94-43, Ano94-44, Ano94-45, Ano94-46, Ano94-47, Ano94-48, Ano94-49, Ano94-50, Ano94-51, Ano94-52, Ano95r, Ano95s, Ano95t, Ano95u, Ano95w, Ano95x, Ano95y, Ano95z, Ano95-27, Ano95-28, Ano95-29, Ano95-30, Ano95h, Ano95i, Ano95j, Ano95k, Ano95l, Ano95m, Ano95n, Ano95o, Ano95p, Ano95q]. Cyber [Ano95s]. CyberMall [Mar95a]. cyberpunks [LP96]. Cycle [Ano93z, Fos90a, MB90, Sch90c, Car90c].

D [Ano92-84, Ano96f, Lei90d, Ano92-42]. DAL [Ano92j]. Daly [Ano92-43]. Dancing [Cha95]. dangerous [Hay97h, Kay96c]. darned [Ano95-82]. DASD [Dar95n, Moa89e]. Dash [Laz98]. Data [Ama95b, Ano92z, Ano92-30, Ano92-54, Ano92-55, Ano92-61, Ano92-66, Ano92-146, Ano93-317, Ano93-79, Ano93-171, Ano93-268, Ano93-284, Ano93-320, Ano93-343, Ano94b, Ano94-69, Ano94-219, Ano96-27, Ano97-45, Ano97-46, App95b, App96i, Bla95a, Cha95, Dar95g, Dar96m, Ger96a, Gre96, McC96d, Pet88, Rad95, Sim97g, Sta89d, Str96c, Tuc97a, Var96h, Ano95-74, Ano95-80, Ano96-42, Ano96-50, Ano97j, Ano97-43, Ano97-47, App93a, App96e, App96f, App96h, Asb97a, Bun90, Car90c, Dar95h, Dar96f, Dar96n, Dav90a, Fis96b, FP88, Fol96d, Fre97a, Fre97b, Hal96, Ham96, Hea89, Inm97a, Jef97e, Kay96a, Kay96f, Ker89d, Kig96, Kno96e, Kno97d, MF96, Mar95d, McC96b, McC96h, McC89b, Moa88d, Moa90a, Par92, Pin91, Ric93d, Ril95e, Ros96, Rot97, Rot91, Sem96e, Sim96q, Sim97j, Ste89, Str95i, Tab92b, Var96g, Wal89b, Wat96b]. data [Wat96c, Ano92-45, Ano92-138, Ano93-114, Ano93-195, Ano93-319, Ano94-280, Ano97-43, App96e, CM95, Dar95q, Hac95, Lin94c]. Database [Ano92-56, Ano92-74, Ano92-97, Ano93e, Ano93-302, Ano93-336, Ano97y, Fol96d, Gre88, Ham94c, Sem96h, Var96f, Wat96a, Ano95-160, Bal96b, Cod90a, Cod90b, Dar96e, Hal96, Ham96, Jen91, Kim90b, Kno97c, Lin91b, McC93b, Moa89d, Moa90h, Sch92, The95p, Tuc97i, Wat96b, Wat96c, Wat98, Ano96v]. Databases [Ano92-65, Ano93-239, Ano96u, Hal96, Ano96-50, Ham96, Hod89d, Kim90a, Mar89b, McC94a, McM90h, The94b, Tra90a, Wat96a, Ano94w]. datacenter [Gib98]. Datacomp [Ano92-173]. datamart [Ano96w, Inm97b]. Datamarts [Var96g, Dar97b, ES97, Kay96a]. Datamation [Ano93-303, Ano93-304, Ano93-305, Ano88a, Ano95-86, Ano95-138, Ano96-28, Ano96-41, Ano96-42, Car89g, Mae96, McC96f, Sem96b, Sim96m, VM96a, Wat96c, Ano92-44, Ano94-66, Ano94-53, Ano94-54, Ano94-55, Ano94-56, Ano94-60, Ano94-62, Ano94-63, Ano94-64, Ano94-65, Ano95-32, Ano95-31, Ano95-34, Ano95-33, Ano95-39, Ano95-40, Ano95-41, Ano95-42, Ano95-44, Ano95-45, Ano95-46, Ano96a, Ano96w, Ano97-27, Ano97-34, Ano98e, Ano92-149, Ano94-57, Ano94-58, Ano94-59, Ano94-61, Ano95-35, Ano95-36, Ano95-37, Ano95-38, Ano95-47, Ano95-43]. Datamining [Ano96r, Dar97a, Fre97b, Ger96a, Asb97a, McC97c, Mik97g]. datawarehousing [Ano96h]. Dataweb [Ano96-29]. Datawebs [Ano96-39, Var96h]. date [Lin96a]. dates [Jag96a]. dawn [Dav90b]. day [Bea96, SF95g, Str95l, Esk97j, Ano95-73]. Daytona [Ano94-169]. DB2 [Ano92-66, Ano93i, FGR88, GRF90, Mar89b, McC89b, Sta90a, Ano92-150]. DBase [Ano92x, Rad94a, Rad94b]. dBASEics [McM90a]. DBMS [Ano92c, Ano92-89, Bro88, Bry89b, O’C88, Wil94, Ano92-117, Ano93-65]. DBMSs [Ano93-71, Sem94g]. DCE [Ric94d]. deadline [App96d, dJ96b]. deadlines [dJ96b]. Deal [Ano92-90, Ano96n, Ano96y, App96a, Bla96, Kno97b, Moa93d, Sne92c, The95v, Tuc98a]. Dealing [Ano92-63, Ano95-159, Ano97-28]. deals [Ano95-28]. Dear [de 97e]. death [Ano95-162, Jag96b, Mar95b, SF96l, You97, Mar95c]. debate [Sch88a]. DebitCredit [Ser89]. debugger [Ano96-44]. debut [Fol96e]. DEC [Ano92-157, App91, Sch88b, Siv90c, Sta89e, Ano92-46]. DEC-IBM [Sta89e]. DEC/ [Sch88b]. decade [Hod90]. Decentralizing [Gra91]. Decide [Sim96t, SF96h]. Decision [KS94, Rad95, Dug96d, WN90, Ano92-45, Dar96c]. Decisions [The96e, HE90, Sim96t]. deck [Fra89j]. decline [Ano97-45]. decode [McC90b]. Deep [Ano96t, Dar95g]. Deeper [Ano96r]. defenders [Ano95-81]. Defining [Kim90a]. delayed [McC96e]. deliver [Ano97n]. delivered [Str96a]. delivering [Fol96g]. Delivers [Ano93-106, Ano97n, Ano97q, Jud90a]. Dell [Ano92-47]. Delphi [Vin95a]. deluge [Rot91]. delving [Dar96s]. Demand [Ano93-105, Ano93-268, Ger95, Arn90, Joy89a, Wil97e]. Demands [Ano92t, Sch95h, Dav89b, FM89b, Sem96h]. demise [Lin96b]. Demystified [Sch95g]. Department [Sch89c, Sta88, Ano95-132, Vog96b, Ova96]. departmental [Ano95-129]. depend [FM97]. deploy [Kno96a, Sim96b]. depopulation [Pet88]. depth [Sem96a]. descending [Sar95]. describe [SF95l]. Design [Ano92y, Mae96, Key89, Wal89b]. desire [Bob89]. Desk [Kay96c, Kay96d, Liv95, Moa94e, The96a, The96d, Var96d, Dar96d, Gib97b, Hec96, Mik97c, Ano93-107, The94a]. desk-top [Dar96d]. Desks [Ano93-147, The96d]. Desktop [Ano90a, Ano92-48, Ano92-154, Ano93-54, Ano93-142, Ano96u, CS95, Cor96, Dar96d, Fol96b, Fra89d, Kay96e, Sim95k, Sim96n, Ste94a, Var96b, Woo92, Ale92, App96g, Cor95b, Fra89c, Fra91b, Fre97c, Laz98, Lin92a, Maj90, McM90e, Pan96b, Ril95a, SF95e, Sim95a, Sim97f, Siv90b, Ano92-167, Ano94-67, McC95d]. Desktops [Coh95, Cor95c, App96c, Hok96, Kay96e]. Desperately [Kay96b, Kor94]. Despite [Ano92n, Gut94, Kay96f, Lin96b]. determine [Fly96a]. Dev [Dar96c, Ano96a, Ano96g, Ano96z, Ano97-44, Fre97d, Lin96c, Rad96b, Vin95b]. DevDays [SF95c]. Develop [Ano92-49, Str95n, Dar96h, Dar96a, Fly96a, SF96c]. developed [Hay96d, McC97e, Var96b]. developer [Dri96, Ste96b]. Developers [Ano92-50, Dav97c, Ano96f, Fly96a, Lin96c, Ste96b]. developing [Fol96d, Gol95, Jag96a, Ker89l, SF95h, Ano93-66, Ano94-68]. Development [Ano92-60, Ano92-100, Ano92-133, Ano93o, Ano93-231, Ano96-29, Bau95e, Dri96, Fol96b, Rad96b, Sem94b, Ano94-278, Ano95-77, Ano95-133, Ano96a, Bal96a, Dar96f, Dar96n, Dar96o, Dev97, FC96, Fre97d, Ger96b, Hay96c, Hum89, Joy88a, Joy89b, Ray95, Rob96, SF96d, Sch89c, Sem96d, Tuc98a, Vog96b, Ano92-39, Ano93-101, Sch95k, Vin95c]. Devices [App96f, Dar96s]. Diagnostic [Ano93-67]. Dial [Sch95d, Sch95e, Woo91, Pin91]. Dial-Up [Sch95d, Sch95e, Pin91]. Dialing [Joy89a, Ril96c]. Dialup [Ril96c, Dar96s]. Diary [Ano96w]. did [Ova96]. dies [Mik97j]. differ [Siv90e]. Difference [Ano93-340, Fre97g]. Different [Ano92-51, dJ96b, Mar95d, Pan96c, SF96p]. difficult [Fre97d]. Dig [Ano96r, Dar95g]. Digging [Fis89]. Digital [Lei90b, McW89a, Moa90h, Siv90b, Dar96s, Ama95a, Ano93-68, Ano93-69, Var96a]. digitize [Hal96, McC88]. dilemma [Sch89b]. dime [McC96l]. Dimension [Ano93b, Fra89c, Moa89d]. dinosaur [Bea96]. Direct [App96f]. direction [Ano95-70]. directly [Ano97u]. Directories [Ano94-108, Ano93-291]. Directory [Fol96h, Hur95a, Ano88a, Ano95-30, Dav96a, Ker89a, Rot89, Ano92p, Ano93-329, Dav97j]. Dirk [Ano96-33]. Dirty [Ano93-114, CM95, Ano95-162, Kno97d, Var96h]. disarray [Hec96]. Disaster [Moa90a, SM94, Car97, Rot88b, Str95i, MS94]. Disaster-proof [Moa90a]. Disasters [Ano92r, Fra93e, Ano93-266]. Disc [Rya95]. discontinuity [GH88]. disguise [DB97]. Disk [Ano96p, Rya94]. diskless [Fra90g]. disks [McC90d]. dismiss [Dar96q]. Dispersed [Bau95e, Moa94j]. disrepair [Hec96]. dissecting [Ger96a]. distinguish [Hab96]. Distribute [Str95a, Dav90c, Sco96, Ano94-69]. Distributed [Ano92r, Ano93o, Ano93-266, Sch95i, Str95m, Sud95, Fra93e, Hay96e, Rea95, Siv90a, Ano93-70, Ano93-71]. Distribution [Ano94z]. distributors [Ano96g, Ano97-32]. district [Dar95u]. Diversity [Hea90a, Moa94d]. Divide [Ale96c, Sim96i]. DLT [Sim96k]. DMI [Ano96u]. DNA [Ano95-125]. Do [Ano93-63, Ano93-351, Ano94-128, Ano94-284, Ano94-278, Esk97f, McC96j, Ril96d, Sim97h, Ano95-133, Ano96-54, Ano97-46, App96a, App96d, Dar96h, Fol96c, Hok96, McC93a, Ril96b, SF95b, SF96a, SF96c, SF96j, SF96r, Str96c, Str96a, The96b, The96d, Vog96a, Wor90, Ano97-46, Fre97d, Sch94]. do-it-yourself [Wor90]. doc [Ano96w]. docs [Ano96q]. doctor [Mik98b]. Document [Ano93-197, App95f, Ker88b, App92b, Dav92d, Fly96b, Man91, McW88a]. Documentation [Ano96-40, McC88, Ste96b]. Documents [Ano94-183, Ano97u, App93a, McC96h, OHE94, Ano93-171]. DOD [Ker88b, Ano92-52]. Does [Ano92-161, Ano93-317, Ano93-73, Cor95c, Dar95n, Dar96e, Inm97b, Moa93a, Ril95c, SF95b, Sem96c, Sim95g, The95h, Ril96e, The96e, Var96f, Wat96b, Ano92-53, Ano93-74, Ano93-72, Ano94-70, Moa94c, Sim95f]. doesn’t [Kno97i, Sim96b, Sim96h, The96b, Jon97a]. Dog [Ano94f]. doing [Esk97g, Scr95b]. Dollars [Ano92-125, Fre97b, Ray95, Sch97b, Sch98, Dar95p]. done [Var96c, Ano93-327]. donned [Fol96f]. Don’t [Ano94-29, Ano94-112, AR96, Bow94, Dar96q, Dug96b, Fol96h, Gre96, Hay97l, MS94, Mik97f, SM94, Var95d, Win95, Ano96k, App96h, Bla96, Dar96m, FM97, Hen95, Inm97a, SF95e, Sim96s, CM95, Ham94a]. Doors [Kau96, Ano92-141]. dope [Ano95-75]. Double [Ano92-54, Ano92-55, Ano92-69, Wat96b]. Doubt [Vog96c]. Dough [Ano97-45, Ano97-46, Tuc97a]. Down [Ano92-88, Ano92-168, Ano93-52, May96, Ano95u, Ell88, Fra90k, Fra91b, FM97, Ker88c, McC96b, Moa90g, Sta89g, Tra95, Var96h, Var96d]. down-and-dirty [Var96h]. downpour [Sim97j]. downright [Ril96h]. Downsize [Ano92-56]. Downsized [Ano92-133]. Downsizing [Ano92-58, Ano92-57, Ano92-164, Ano93-342, Ano94-70, Ano96u, Sim95b, Kre93c, Sha92, Ano92-52]. Downtime [Sim95d]. Dozen [Ano96-36, vSS96]. dozens [App96c]. DP [Ano89b, Ano89a, Kel89, Kel90d, Kir89, Tra89a]. dragon [Tat89b]. Draino [The95g]. drawing [McW89a]. draws [Kel90b]. dream [Ril96h]. dreck [Dar96k]. Drive [Ano93-64, Rub95b, Ano93-311, Dar96b, Fra93b]. Drives [Ano92-55, Ano92-167, Pan96b, Als88, Arn90, Eth90a, Mar89a, Ano93-106]. driving [Sim96q, Siv90f]. drug [McC88]. DSM [Tuc96b, Tuc97d]. DSS [Dar96c, Dar96f, Fra89e, The95q, The95u]. dual [Cor95b, Emm89a]. dual-protocol [Emm89a]. Duck [Ano94f]. Ducking [Siv90b]. due [Sim96d]. duel [Emm89a]. DUET [Lin93]. Duking [Siv90c]. Dumb [Ano92-129, SF95f]. dummies [Ano96l]. Dump [Kno97d]. duo [SF96b]. duplicate [Ril96i]. during [SF95g]. Duty [Var95a]. dynamic [Ano97-43, Lei90c].

e-cash [Fre96a]. e-commerce

[Ano97q, Hay96a, Sem96i, Tuc97n, VM96b, Var96j, VM96c]. E-Forms [The94d]. e-frontier [Ano98e]. e-mail [Ano90c, Fol96h, MF90, McC93b, Ril96e, The93e, The93f, Ano94-269, Ano92-96, Ano93-274, Ano93-273, Ano93-291, Ano93-324, Chi91, Ker89a, Ste92a, Str93, Ano93-302]. E-payments [Ano97i]. each [Ano96-28, Kay96c, Mae96]. Eager [Sim96b]. Eagle [Fol96a]. Early [Fos90b, Fol96e, Gib97b, Kay96f, OHE94, Sim96g]. Earn [Mae95]. Ease [Ano93-316, Ano93-230, Bar90c, Dar97b, Ano92-59]. Eases [Ano94-72, Hur95a, Ano98a]. Easier [McC95e, Ano97-44, Ano98g, Con92, Fol96h, Gre96, Hab94, Ano92-101]. easily [McC96h]. Easing [Ano92-60, McC93c, Ano93-75]. Easy [Ano93-319, Ano94-251, Dur94f, Val96, Ano93-350, Dar96h, Dar96r, Fis96b, Hay97o, McC96m, McC89a, McM91b, The96e]. eats [Vog96b]. ECCS [Ano93-76]. EcoCLlENT [Ano96-49]. Economic [Ano93-77, Ano93-78]. EDA [Ano96r]. EDA/SQL [Ano96r]. Edge [Ano92l, Ano95r, Ano95s, Ano95t, Ano95u, Ano95v, Ano95w, Ano95x, Ano95y, Ano95z, Ano95-27, Ano95-28, Ano95-29, Ano95-30, Ano96g, Ano96v, Ano96y, Ano96x, Ano96h, Ano96i, Ano96j, Ano96k, Ano96s, Ano96n, Ano96m, Ano96l, Ano96w, Ano96o, Ano96p, Ano96q, Ano96z, Ano96r, Ano96t, Ano96u, Ano96-40, Ano96-45, Ano96-46, Ano96-41, Ano96-42, Ano96-43, Ano96-44, Ano96-47, Ano97j, Ano97l, Ano97i, Ano97k, Ano97m, Ano97n, Ano97p, Ano97o, Ano97r, Ano97q, Ano97s, Ano97t, Ano97u, Ano97v, Ano97z, Ano97y, Ano97w, Ano97x, Ano97-34, Ano97-35, Ano98c, Ano98b, Hal96, Mik97a, Mik97b, Mik97c, Mik97d, Mik97f, Mik97g, Mik97e, Mik97i, Mik97j, Mik97h, Mik98a, Mik98b, Mit96, Moa90c, Val96, Ano92-166, Ano94-39, Ano94-40, Ano94-41, Ano94-42, Ano94-43, Ano94-44]. Edge [Ano94-45, Ano94-46, Ano94-47, Ano94-48, Ano94-49, Ano94-50, Ano94-51, Ano94-52, Ano95h, Ano95i, Ano95j, Ano95k, Ano95l, Ano95m, Ano95n, Ano95o, Ano95p, Ano95q]. EDI [Ano93-256, Dav89g, Eth88, Lei90b, McC93d, McC94b, McW88a, Sna90a, Tuc97e, Ano94-111, Dav94]. EDI/ [Ano93-256]. Edition [Ano95-86]. Editorial [Ano97-29]. educated [Ano97n]. education [Dav89b]. educational [Bra89]. Edwards [Ano92-84]. effective [Bal96a, Dar96b, Mik97c, Str96b]. Effectively [Kno96c, Kno96e, Kno96d, Kno96f, Kno97i, Kno97j, Kno97k, Kno97l, Kno97m]. efficient [Dar96m, Gib98]. effort [Dar96o]. Egads [App96a]. Eight [Ano96p, Gos88, Sha96]. Eight-Way [Ano96p]. EIS [Arm90b, Arm90a]. EISs [Dav89h]. either [Rad96c, Ril95k]. either/or [Rad96c]. elbow [App96f]. Electronic [Ano96p, Dys90a, Ker89d, McW88a, Pur89, SF95g, Sch89e, The93c]. element [Ano95-125]. Eli [Ano95-70]. eliminate [Vog96c]. eliminating [Ril95j]. Ellison [Ano92-106, Pat96]. Elusive [Fol96h]. EMASS [Ano93-79]. Embattled [Ano95-71]. Embrace [Dar96g, Hay97f]. Embraces [Sem96c, Lin96c]. emergency [Ano95v]. emerges [Str95e]. Emerging [Sim96q]. emphasizing [Mad90]. Employ [Wil97c]. employees [Ano94-278, Gib97b, McC96h]. Emulex [Ano93-80]. enable [Rad96c]. enabled [Chi92, Fol96b, Var96h]. enables [Ano96q]. enabling [Sim96p]. Encapsulation [Ano95-112]. Encrypt [Ril96e, Rot97]. encryption [Ker89h, Ril96e, Str93, Ano93-274]. End [Ano92-109, Ano93-307, Ano93-210, App96f, App96h, Bla95b, McM89a, Roc97f, Ano95-96, Dav97f, ES97, Fol96b, Gil96a, Hea90a, McC93f, Mee96, Nor88, Ril96h, Sha93c, Sim96d, Siv90c, Ano93-236]. End-User [Ano92-109, Bla95b, Fol96b]. end-users [Ano95-96]. Enemy [Foc95, Ale95a, Ano95-97]. Energizes [Hac95]. Enfin [Ano93-81]. Engage [Ano97t]. engine [Ril96c, Siv90f, Ano93-336]. engineer [Ano95u]. Engineering [Ano93-280, Kni90, Gib89, Gos88]. EngineerSpeak [Ano95u]. enhanced [Sim96m]. enhancements [Ano96f]. Enough [Ano94-110, Kay95i, Dar96m, Kay96b, Mae96, McC97e]. ensure [Rot88b]. Entering [de 98a]. Enterprise [Ano92-64, Ano93b, Ano93-207, Ano93-209, Ano93-253, App96c, Dar96h, Dar96k, Dev97, Fol96a, Fol96f, Hur94c, Mel95, Pan96b, SF95k, Sem94e, SF96q, Tuc98a, Ale96c, Ano95-30, Ano95-59, Ano96a, Ano97s, Cor96, Dar95f, Dar96o, Dar96p, Dar96s, Dav97a, Dav97i, Fol96c, Gil96a, Inm97b, Lin94b, McC96e, McC93g, Ric92a, Ric93a, Sem93, Sim96b, Sim96e, Sim96t, The93e, TMS96, Tuc97c, Var96b, Ano93-48, Ano93-53, Ano93-82, Ano93-273, Ano93-290, Ano93-348, Ano94-183, Ano94-244, Ano96-51, Ano98a, Cor96, Dar96h, Hac95, McC96e, Rad96b]. enterprise-computing [Ano95-59]. enterprise-level [Sim96b, Sim96t]. enterprise-quality [Fol96f]. Enterprise-wide [Ano93b]. Enterprisewide [Sch95j, Dar96r]. enterprising [Dav96c]. Enters [Kay95j]. entirely [FC96]. entry [Ric90c]. Environment [Dar96n, Ano96z, Bal96b, Car90c, Dar96f, Dar96o, TMS96, Ano92-148]. environments [Ano96a, Bal96a, Dev97, McW89b]. EPA [Dav92b]. equal [Can88]. equals [Ano94-28]. equitable [Ser89]. equivalent [Ril95i]. era [Bro88, Nor88]. ERP [App97a, Hec97]. error [Joy89c]. error-free [Joy89c]. Escape [Ano93-261, Bla90]. ESS [Ano97t]. Essential [Sem94d]. estate [Ano97z]. Ethernet [Ano92-71, Ano94-71, Ano96h, Dar96i, Dar96j, Dar96l, McC95c, McC96a, Siv89c]. Euro [Kno97e]. Europe [Eth88, Jud90b, Esh89, Kay90, Lam90, Pur89, Tat89a, Tat89c, Tat89b]. European [Ano95-86, Eth89, Mar89a, Ros89, Tra89a, Ano97l, Bar90a, Tom88]. Evaluating [SF96b, Dar96s]. Evaluation [The95l, Ger95, The95d, VM96a]. Evaluations [Dar95s]. Even [Str95i, Ano95-30, Ano95-98, Ano95-126, Ano95-136, Dar96k, Hir96, Kay96c, McC96h]. Ever [Ano93-311, Ano96r, Ano95s, Ano97-37, Joy89a, Tuc96b]. Ever-more-attractive [Ano93-311]. Every [Ano92-49, Ano92c, Ano92d, Fol96d, Hen96, Ril96j]. Everybody [Ova96]. Everyone [Dar96k, Ano95-76, Ano95-148, App96f, May96, McC96b, Ril95i]. Everything [Mee96, Fre96b, Gib97b, Roc97d, Ano95-161]. Everywhere [Ano96u, Fly96a]. evidence [Str97a]. Evolution [Ano93-244]. evolve [Ker89i]. evolved [Ham96]. evolving [Sim96m]. exact [Ano95-136]. exactly [Val96]. example [Roc98a]. Excavate [Ger96a]. Excel [Sta93, Ric92b]. Excellence [Ano93-45]. excess [Asb97f]. Exchange [McC96e]. exclusive [Ano96w, McC96f]. execs [HE90, Tat89a]. execution [Dar96d]. executive [Fry96, Mik98a]. Existing [Bec94, Str95f]. expand [Ano95-131]. Expands [Gut94]. expect [Dar96c, Fol96h]. expectations [Jag97]. expenses [Ano97i]. Expensive [Ric94e, Lin96a]. experience [Bra89, Ger96b, Kay96b, The96b]. experimental [Ano95-130]. experimentation [OHE94]. Expert [Ano92-164, McC90e, New90, PC89, Sha92, Wat96b]. expertise [The96b]. Experts [New90, Ano93-124]. explain [Sem96b]. explained [Wat96a]. explode [McC96g]. exploit [Ano97j, Dar96o]. Explore [Kno96e]. Exploring [Ano93-326]. Explosion [App95f, Ano93-197]. Expo [Ano96-47]. exponential [Pan96a]. Extension [Mik97c]. extensions [Ano93-350]. extinguish [Sch95a]. extra [Sim96s]. extranet [Str97c]. extremely [Ano95-128]. Extremes [Ano93-123]. Eye [Fre97c, App91, Dar96m, McW88a, Sta90c].

F&D [Ano97d]. FAA [Moa89j]. Face [Ano92-117, Ano92-120, Dug97a, Hen96, SF96e, Ano97r, May96, McC96g, Sta96, Str95i]. faced [App96d]. Facelift [Ano93j]. faces [Hay97q, Kee97b, Str95h]. Facing [Sch95c]. fact [Lon90, Sim96g]. facto [Tuc96a]. factor [Arm90a]. factories [Web89]. factors [VM96a]. Factory [Ano96-37, Asb98a]. facts [Ale96a, Ano96-34]. Fails [KS94]. Failure [Ano96-45]. faint [Gre96]. faire [Ker89d]. faith [Eth90b]. Fall [Ano93-301, Ano97-45, Sim96k, Ano93-264]. falling [Kay96c, Sim96k]. Fallow [Kee98b]. falls [Car88c]. family [Hay96c]. fan [Ano96-31]. fans [Ano96-33]. far [Ano95w, Sta89b, Var96a]. Farewell [Ano98d]. Fargo [Ano97-48]. farms [Esk97b, Fre98b]. Fascinates [Ano93-341]. fashion [SF96d]. fashionable [Scr95b]. Fast [Ano92-61, Ano93-203, Ano93-319, Ano94-200, Ano96h, Dar96s, Sne91a, Sne93, Bun91, Dav91, McC96p, Sim96k, Ano93-83, Dur94a]. Faster [Ano93-84, Ano94-71, Ano96-51, Hen96, Joy89a, Kre93a, SF96k, Sim96h, Wal90b, Ano92-67]. Fastest [Ano93-85, Dur94e, Vog96c]. fat [SF95f]. Fatal [Cod88a, Cod88b]. fate [Bal96b]. father [Ano97k]. Fault [Ano93-86, Sim96m]. fault-tolerant [Sim96m]. favorite [SF96g]. Fax [Ano92-160, Ano93-63, Ano93-75, Ano93-149, Lev95, McC94d, Bla88, Coo92b, Fol93, Hin93, McC92, McC93a, McC93c, Sne92a, Ste92b, Ano92-104]. FDDI [Ano92-77, Ano93-85]. feature [Ano95-86, Ano96a, Ano96s]. Features [Ano93-83, Ano95v, Dar96s, Sim96m, Sim97h, Sne93, vSS96]. fee [Lei90d]. Feeding [Kay90]. Feel [Ano93-232, SF96n, HE90, Var96b]. feeling [Joh89]. feels [Ril96h]. fences [FS89]. Ferraris [Ril95i]. Fever [Gil96b]. few [Ano95z, Hug90b, Ril96b, Sim97h]. fiascos [Jam97b]. fiber [Bun90, Jef97c]. fiction [Lon90]. field [Joy89a]. fields [The92a]. fifth [Ano95-127, Meh89]. fight [Ano95-81]. Fighting [Car89c, Fly96a]. figment [Pat96]. figure [Jon97a, Man92b]. file [App96c, Ril95e, Tuc96a]. file-sharing [Tuc96a]. files [McC96h, Ril95c, Kol94]. Fill [Ano94y, Fol96c]. filled [Ker89k]. Filling [Jef97c]. Final [Ano98d]. Finally [Ano93-49, Ano93-89, Gil91, Sem95b, Sem95c, McC96m, Sim96p, Var96b]. finance [Bus88]. Financial [Ano93-144, Car90b, Sha93b, App96i, The95h]. Financials [Ano94-70, Hab95b, Hab95a]. Find [Ano92-164, Ano93-239, Ano98f, App95d, App95e, Bec94, Dar96k, Kno97f, Roc97b, Sem96e, App95a, Dar96s, Dri96, Fra91c, McC97e, Sha92, Var96f]. Finding [Lin96a, Sta90a, App96i, Sim96i, Ano93-86]. finds [Dar95d, Ets98, Gib98]. fingertips [Ano96b]. Fininvest [Mar89d]. Finland [Ano95x]. Fire [Ril96n, Sta89d]. Firewall [McC96b, Ano95-102, Ril96b]. Firewalls [Ano96p]. firm [Dar96a, Roc97b]. firms [Ano96l, Fra90e, McC88, Sha96, The96b]. First [Ano93-250, Ano96p, Ano96u, Ano96-45, Moa94e, Wat96c, Dav97c, Ger96b, McC96b, Mee96, Str95i]. fistful [App96c]. Fit [Dar95n, Ano93-189, MK90a, Ano93-85]. Fits [Ano93y, Ano93-344, Nee95, Kay95f]. Five [Ano93-148, Ano95t, Ano95y, Var96d, Vog96a, Ano92-62]. Five-Platform [Ano92-62]. Five-Year [Ano93-148]. Fix [KS94, Ano96s, Jag96a]. fix-it [Ano96s]. fixed [Ker89b]. fixes [App97b, Sta90a]. Fixing [Cel96, Lin96a, Ril96d]. Flags [Bec94]. Flat [Ano93-321]. flawless [Dar96d]. flaws [Cod88a, Cod88b]. fleeing [Asb97b]. Fleer [Ano98c]. Fleer/Skybox [Ano98c]. flex [Sim97c]. Flexed [Kay98]. Flexes [Ano92-36]. Flexibility [App96g, TMS96]. flexible [Dav94]. flies [Mar95d]. flight [Tuc97j]. Flood [Ano92-30, Par92]. floppies [Ano96o]. Flow [Ano93-283, Dar95h]. Fly [Ano93-192]. Focus [Ano93-87, Ano95-47]. focuses [Ano95-86]. foe [Ano97m]. folks [App96e, SF95b]. follow [Ano95-86]. follow-on [Ano95-86]. followed [OHE94]. following [OHE94]. Folly [Moa94e]. Food [Ano98b]. foothold [Sim96e]. force [Jon97c, Ker89f]. Forges [Ano93-209, Hug90a]. Forget [Hay97m, McC95e, McC96k, MS94, Gib97b, SM94, Asb97a]. form [Ano95-128]. Former [Mit96]. forms [The94c, Ano93-283, The94d]. formula [Gib97b, Joh90a, Str97a]. Fortune [McC97e]. Forty [Str97b]. Fossilitis [Moa94h]. found [App96g, Var96b]. Foundation [Ano93-124, Kno96c, Ano96-48]. Four [Fra90i, Jon97b, Sem95e]. FourGen [Ano93-87, Ano93-87]. Fox [Ano96-35]. fractional [Siv90g]. Frame [Ano92-161, McC96d, Ano94-72, Ano94-73]. Framework [Ale96a, Fol96c, Lin96a, Sem96a, Sem96d, Sem96g, Str96a, Var96i, Ano96m, Tuc97c, Tuc97d, Sem96b, Sem96c, Sem96f, Sem96e, Ano93-88, Ano94-74, Ano94-75, Ano94-92, Ano94-93, Ano94-94, Ano94-95, Ano94-96, Ano94-76, Ano94-77, Ano94-78, Ano94-79, Ano94-80, Ano94-81, Ano94-82, Ano94-83, Ano94-84, Ano94-85, Ano94-86, Ano94-87, Ano94-88, Ano94-89, Ano94-90, Ano94-91, Ano95-57, Ano95-58, Ano95-59, Ano95-48, Ano95-49, Ano95-50, Ano95-51, Ano95-52, Ano95-53, Ano95-54, Ano95-55, Dar96c, Sch95a, Str95d, Str95l, Str95e, Str95f, Str95g, Str95h, Str95j, Str95i, Str95k, Ano95-56]. Framework-Another [Str95l]. Frameworks [The95e, Ano96o, Ano95-77]. Framwork [Ano95-60]. Franchise [Dar96b]. Free [Ano96-45, Joy89c, Kau96, SF97b, The93c, Ano93-247]. frenzy [Ano95-27]. fresh [Moa95h]. Fried [The94c]. friend [Ano97m]. Friendly [Ano92-29, Ano93-260, The96c]. Frito [Fra89e]. frontier [Ano98e]. Fruit [Ano96g]. fruitful [Sch88b]. frustration [Ril96c]. FSIOP [Ama95b]. FTAM [Ano92-153]. fuel [Sim97i]. fugitive [Mit96]. Fujitsu [Moa89g, Ano92-63]. Full [Ano93-206, Mik97a, Hay96a, McC96d, Ano93-89]. full-blown [McC96d]. full-steam [Hay96a]. Full-Strength [Ano93-89]. Fun [Ano93-212]. function [Jag96b]. functional [Moa89f]. functionality [Hay96e, TMS96]. Functions [The95m, Ste96a]. fundamental [Rob96]. furiously [Hay96e]. fusion [Cod90a, Cod90b]. Future [Ano92-64, Ano94-129, Ano94-272, Sch95c, Ste94a, Var96a, Var96b, Ano96j, Ano96-51, Bee89a, Dar96k, Fly96a, Fra90g, Tat89a, Var96f, Was90, Ano92-120, Ano93-90, Ano93-91, Ano93-92, Ano93-93, Ano93-94, Ano93-95, Ano93-96, Ano93-97, Ano93-98, Ano93-99, Ano93-100, Ano93-275, Ano94-97, Ano94-98, Ano94-99, Ano94-100, Ano94-101, Ano94-102, Ano94-103, Ano94-104, Ano94-105]. fuzzy [Joh89].

Gain [Ano93-204, Dav97k, Sim96e]. gaining [Dys90b]. Gains [Ano96r]. gambles [Mik97d]. game [Fre97e, Ger90, Hur96, Ric94b, Ano94-170]. games [Sim96l]. gap [Dug96a, Ric90c, Sch90a]. garbage [Ano96x]. garnering [Hay96c, Sim96g]. Gartner [Mik97d]. gate [OHE94]. Gates [Esk97h, de 97e]. Gateway [Ano92-64, Cas95a, Cas95b, Cas95c]. gave [Dar96t]. gazillions [Var96a]. gear [Ale92, Bau94b]. gee [Ano96f]. gee-whiz [Ano96f]. General [Ano93-213, dJ96a, The93a, The96c, Var96b, Dav94]. general-purpose [The96c, Var96b]. generating [App96a]. generation [Ano95-134, AR96, Hab96, Meh89, Rad96c, Sim96d, Str95j]. Geographic [Ano92-65]. George [SF95d, SF95c]. Georgia [Siv89b]. Georgia-Pacific [Siv89b]. Get [Ano92u, Ano92-27, Ano92-105, Ano92-121, Ano92-146, Ano93-52, Ano93-115, Ano93-251, Ano93-269, Ano94-29, Ano94-111, Ano94-276, Ano95-127, Ano96i, Bau94b, Dar96h, Fis96a, Fre98a, Gil96b, Hab94, Kno97g, Moa94c, Rad94b, Ray95, Ril95a, Str94a, Str95b, The94d, Ano95-133, Ano98g, Dar96a, Fra90k, Fry96, Gut96, Hal88, Jag96a, Kee97a, Mae96, McC96a, McC96b, McC93d, McC94b, McM90d, Mik97f, OHE94, Pic94d, Rad94a, Ril95e, SF96j, Sch90b, The93d, The96e, Var96a, Ano92-66, Ano93-101, Ano93-102, Ano94-106, Ano94-107, Bla95a, FH94, Ham94b, Ril97d, Str94g, Sud95]. Getcha [Gre96, Ham94a]. Gets [Ano92-113, Ano92-116, Ano92-168, Ano93j, Ano93-154, Ano93-247, Ano93-336, Ano94-71, Ano94-273, Ano95-134, Ano97i, Ano98c, Dav89e, Ker89k, Lin92c, McC96b, McM90a, Ric92d, Sne92e, The93c]. Getting [Ano95-72, Str94o, Ano93-350, Fol96d, Pen97, Var96c, Sch95i, The94e]. giant [Gad89, Sem96e]. giants [Dav89a, Sim96k]. gigabytes [Var96a]. girds [Rot91]. Give [Ano93-338, Sta95b, Sta95a, Ano95y, Fis96b, Fol96f, Fol96i, Inm97a, Var96a, Ham94c]. Glass [Gut95a, Gut95b, Moa89h]. glitch [Lin96a]. Global [Ano93w, Ano93-315, Ano95-35, Ano95-36, Ger96b, Hur95a, Run89a, Ano88b, Dav93a, Ker89a, Rec89, Sna90a, Sne92d, Ano93-103, Ano94-108, McC95e]. globe [Mar90]. glory [Sim96n]. Glue [Ano93-242, Sch95j]. Go [Ano92-65, Ano94-284, Ano96u, Dar95h, Dar96s, Mae95, Man92b, Moa95e, Sim96h, The95u, WP94, Dav89g, Inm97a, Kay96b, Rad96b, SF96a, Sim97n, Str95h, Str96a, Ano92-67, Ano93-347, Ano95w, Ano95-57]. Goal [Ano92-86]. goals [Ova96]. gobbling [Ano96l]. gods [Esk97l]. Goes [Ano92-48, Ano93-123, Ano94-217, Meh89, Moa89k, Ril95k, Sne92d, Woo92]. Going [Ano92-79, Ano93-327, Ano94-277, Ano95u, Ale96c, Ano97-49, App92a, App96c, Bla96, Fis96a, Fol96g, Hal96, Ham96, Hen95, Hur94a, Rec89, Sco96, McC95f]. Gold [Ale94, Dar95g]. Good [Ano92-117, Ano93-205, Lee95, Ano95x, Ano95-28, Ano95-149, App96e, Dar96a, Inm97a, Kau96, Kay96a, Ril95d, Ril97e, Sim96k, The95b, The96e, Var96b, Var96g]. Goodbye [Sim96r, Ano96o]. goof [Ril95h]. Gosling [McC96f]. got [Ano95-114, App96d, Bea96, FM97, Sim96n, Str96c]. Gotchas [Gre96, Ham94a, Ano98c]. gotta [App96d]. Government [de 97f, Tra89a]. grade [Dar95u, The94i]. graduating [App96a]. grand [Lam90]. Graphical [Bau94b]. graphics [Ano94-278, Bar90c, McM90c]. great [Fol96g, Gil96a, Rad96b, Ril95h, Wat96c]. Greater [OH95]. gremlins [McC96l]. GRiD [Ano90b]. Grip [Ano92-66]. Grocers [The95p]. grooming [Sta89c]. ground [Dys90b, Ker89k, Sne91a, Ano95-134]. Group [Ano92-44, Ano92-68, Ano93-314, Ano88a, Ano93-350, Ano97z, Car90a, McC97e, Rot89, Tat89b, Vog96b, Ano92-145, Mik97d]. Groups [Ano92-159, Car88a, Rot88a]. Groupware [Ano94y, Bau95a, Rad96c, The95m, Dys90a, Dys90b, Hur96, The93f, Var96k, Ano93-324, Bau95b, The95c]. Grow [Ano92-159, Ste94b, Dav90d, Str96c]. growing [Ano96b, Fol96c]. grown [Var96c]. Grows [Ano92-110, Ano93-330, Fol96f, Fra92b, McC90c]. Growth [Ano93-111, Als88, App96f, Hod89a, Hod89b, Sna90a, Ano93-104]. GTE [Siv90d]. guarantee [Str97a]. Guaranteed [Dug97b]. guarantees [Hug90b]. guardedly [Dar96q]. Guarding [Mar95a]. Guest [Ano96n]. guests [Ano97e]. GUI [Ano92-140, Ano93y, Ano93e, Ano93-212, SF96k, SF95l]. Guide [Ano93y, Ano93z, Str95b, Bar90b, Der92, Esk97k, McC96h, Mik98a, Wor90, Ano94-27, McC95i, Sch95b, Str94c, Ano95u]. GUIs [Ano92-35, Ano92b, Ano92-59]. guise [Ril96l]. gums [Ano95t]. gun [App96d]. Guns [Ano96r]. Gupta [Ano92-69]. gushing [Jef97e].

Haas [Ost90]. Habit [Rya94]. hacker [Ano95v]. Hackers [Ale95c]. hail [McC90e]. hall [Ano95u]. Hallmark [Joh90a]. Halls [SF95g]. hammer [Gib97b]. Hancock [Ano92-80]. handful [Pat96]. Handle [BS94, Ham94b, The96d, Tuc97c]. Hands [The95i]. Hands-On [The95i]. handy [Ano95u]. Hang [Ano97-30]. Hanging [SF97c]. happen [App96i, Fis96a, Str95g]. happened [Ril96k]. happens [Ano95-58]. happy [The96d]. Hard [Hab95b, Rya94, Ale96a, Ano96-34, Dar96h, Hab95a, Kee97d, Kir89, Rub95b, Ano93-105]. Hardware [Ano93-236, Ano93-251, Ham94d, Sem95b, Sem95c, Ano96-51, Bla96, Fis96a, Fol96g, Hay96d, McC96l, The93d]. Harness [Ano94-251]. Harrods [Sed90b]. Harvest [Esk97b, MF96]. harvests [LN89]. Hash [Rad92a]. Hassle [Hur96]. hates [Ano95-148]. haven’t [Hay97b]. having [Ril95d]. HDS [Ano93-106]. Head [Ano93-282, Ano93-287, Ano95-27, Ano95-58, Ell88, Kno97c]. Headed [Ano94-285]. heading [Ano95-70]. headlines [Asb97a, dJ96a]. Heads [Str94e]. headway [McC90f]. health [Sch88a]. Healthcare [Dar95d]. healthier [McC97c]. hear [Ril95j]. heart [Gre96]. hearts [Fly96a]. heat [Ano96-49]. Heavy [Ano93-306, Ano93-191, Ano94-38, Var95a]. Heavy-Duty [Var95a]. Hel [Ano96n]. held [Ano96-47]. Helical [Ano92-125]. Hello [Sim96r, Sch95d, Sch95e]. Help [Ano92-70, Ano92-164, Ano93-107, Ano93-147, Ano94-110, Ano95-96, Ano97-38, Bry89a, Dar96j, Hab96, Kay96c, Kay96d, Liv95, Moa94e, The94a, The96a, The96d, Var96d, Ano94-268, Ano95-85, Ano96v, Ano96-54, App96d, Fis96a, Fol96d, Fre98b, Gib97b, Hec96, Joy89a, Kay96b, Kno96a, McC96a, McC96h, McC97e, Mik97c, MK90b, Rad96a, Ril97c, Sha92, Ano93-107]. Help/Systems [Ano92-70]. helpful [Sta96]. helping [Kay96f]. Helps [Ano96r, Ano95-132, Ano96g, Ano98c, Jam97c]. Here [Ano92-151, Ano93-331, Ano94-278, Ano95u, Ano95-159, Ano96w, App94, App96d, Bau95c, Bla96, CS95, Dar96h, Dar96k, Gut96, Hur96, Kay96b, LP96, McC96h, Pat96, Rad96a, Ric92c, SF95h, Sem94a, Sim95e, Sim96d, Sim96i, Sim96n, Spe96, Str96c, The96c, The96e, Vog96a, vSS96, Kno97l, Roc97f, SF96b, Str96a, Vog95, Ano92-71, Ano94-109, Asb97b, Ova96, Sim97f]. Hermes [Sem94e]. hero [Dal98]. Hershey [Inm97b]. heterogeneity [Dav97g]. Hewlett [Ano96r]. hiccups [Kay96f]. hidden [Rad96a]. Hiding [Bec94]. Hierarchical [Rya94, Ano95-74, Ano93-108]. High [Ano93-236, Ano93-210, Ano93-292, Ano96-44, App96f, App96h, Gil96a, Kay97, Moa95b, Ano96-48, Bry88, Fol89, Hay97p, Kay96b, Kee97a, Lin92c, McC93f, Siv90c, Sna90a]. High-End [Ano93-210, App96f, McC93f]. high-speed [Ano96-48]. High-tech [Ano96-44]. high-volume [Fol89]. Higher [Str94j, Sim96c, Vog96c]. Highest [Ano93-104]. highlight [Str96b]. highlights [VM96a]. highs [CF91]. Hike [Ano96u]. Hilton [Ano97e]. Hire [The96b, Gut96]. history [Str97b]. hit [Ril95f, The96b]. hit-or-miss [The96b]. Hitchhiker [McC95i]. Hits [Gar95b, Ano89c, Ano95w, Gar95a, Kee98a]. Hitters [Ano93-191]. hobgoblins [Ano95-112]. hokey [Asb97a]. Hold [Ano93-349, Ano95x, Ano97r, Fly96b]. holding [Sem96i]. holdingback [Ano95-130]. holds [Hug90b]. Holdup [Ano94-241]. Hole [Ano93-27]. holes [Kno96f]. Holography [Ano95-128]. home [Ano96s, App96a, McC96m, Owe89, Roc98b, Var96c]. home-grown [Var96c]. Honey [Ano93-109]. honor [Ano96l, de 97a]. Hood [Dar95l, The95f]. Hook [Dar96l, McC96e, McC96m]. Hope [Ano96r]. hopes [Ano93-350]. hopped [Ano95-70, Ano96b]. horse [Joh90b]. Horsepower [Ano93-293]. horses [Sim96l]. Hospital [Ano96d]. host [Var96j]. hosts [SF95c]. Hot [Ano94-277, Ano96-49, Fol96d, Run89b, Var94b, Moa90b, SF96f, The96b, Ano93-110, Ano93-111]. hottest [Joh90c]. Hour [Mee96]. hours [Ano95-132, Mee96]. house [Joy88a, Moa89h, Gut95a, Gut95b]. households [Ano96k]. How’d [McC96l]. however [Hay96e]. How’re [Dug96c]. HP [Ano96v, Bro90, Moa91b, Sta95b, Sta95a, Ano92-75, Ano93-118, Ano93-120]. HR [Ano97t, Dav91, Gra91, Jam97c, Ric92c, The95j, The95k]. HSM [Sim95g]. HTML [Mee96]. HTTPD [McC96q]. Hub [Ano93-293, Ano94-129, Ano94-112]. Hub/Router [Ano94-112]. Hubs [Ano93-316, Ano93-282, Ano93-102]. huge [McC96a, Ste94b]. Human [Sta90b, The95r, The95s, Hir96, The95a]. humor [Ano96-41, Ano96-42, Ano97-34]. Hundreds [McC96h]. hungry [Inm97b]. Hunt [Roc97a, Roc97b, Roc97c, Roc97f, Roc97d, Roc97e, Roc98a, Roc98b, Tuc97f]. hurts [Str95i]. hybrid [Sim96u]. Hydra [Dal98]. Hype [Ano93k, Dar96b, Gri89, Sim96n]. hypertext [The92b].

I-commerce [Wil97d]. I-Muxes [Ano92t]. I-Net [SF96b, Wil97a]. I-Nets [McC96h, Pan97]. I/O [Ano93-45, Sim97d, Tuc96a]. I/T [Ano95v]. IBI [Ano96j]. IBM [Ano92-28, Ano92-73, Ano92-157, Ano93-47, Ano93-69, Ano93-119, Ano93-235, Ano95v, Ano95-160, Ano96r, Ano96t, Ano97q, Ano97v, App92a, Bun90, Car89d, CM89, Car90c, Fol96a, Fol96g, Fol96h, Fol96i, FS89, Ger96b, Hay96d, Hay96c, Hok96, Jen89b, Jua90, Ker88a, Ker89b, Ker89c, Ker89e, Lei90c, McC95g, McC96o, MK89, MK90a, MK90b, MK90c, Moa91a, Pin92, Ric90c, Ric94f, Ryp88, SF95i, Sim96j, Sim96l, Sim96s, Siv89d, Tat89b, Var96j, Wat96b, Ano92-79, Ano92-76, Ano92-80, Ano92-77, Ano92-78, Ano93-120, Sta89e]. IBM-developed [Hay96d]. IBM/DEC [Ano92-157]. IBM/HP [Ano93-120]. IBM/Lotus [McC95g]. IBS [Ano92-81]. ice [de 97h]. iceberg [Ano96-35]. idea [Ril95j, Tay96]. Ideal [Hab95c]. idiot [Ano95-82, Esk97k, dJ96a]. idiot-proof [Ano95-82]. idiots [Ano95-82]. If [Dar96d, Dar96f, Fis96b, Hab96, Hal96, Hay97b, Hen95, Kay96e, McC96d, SF95e, SF96b, Sim96c, Vog96b, Ano96b, Ano96-47, App96c, Bla96, OHE94, Ril95d, Sim96s, Sta96, Str95i, Val96, FM97, Inm97a, Tuc97c]. Ignore [Dar96c, Var95d, Sim96s]. ignored [SF96h]. II [Cod90b, Fol88a, Siv89b]. III [Bro90, Sta88]. IM [Ano92-82]. Image [Ano93q, Bla90, Dys90a, Fis89, Hal96, Kni90, Lei90e, Man91, May90, Wal90b, Ano92-105]. imagination [Pat96]. Imaging [Ano94-251, Dur94a, Fly96b, Kre93b, Dys90a, Ano93-121]. imminent [Lin96b]. Immortalize [Dar95i]. Impact [Ano92-69, Eth89, FC96, Jag96a, SF96i]. implementation [Ano97-28, Dar97b, Ruh88]. implementations [Fol96e]. Implementing [Gib89]. implements [Ano96v]. important [McC96h, Sco96, Sim96p, vSS96, Str95f]. important-it [Sco96]. imposes [Sch95l]. impossible [Ano95-82, The96a]. Impressive [Hea89, Sim96j]. imprisonment [Sta96]. Improve [Ano93-281, Ova96, Ker90b]. Improved [Ano92-150, Ano93-106, Ber95a, Hab96, Ber95b]. improves [Ano97x]. Improving [Hum89, Ano96-52, The96d]. impulse [Ano95-136]. in-depth [Sem96a]. in-house [Joy88a]. Inch [Ano96p]. include [Ano96f]. included [Ano95-86]. Includes [Ano96-28, Ano95-86]. including [Hay96c]. increase [McC96l]. increases [Pan96a]. increasingly [Bal96b]. indecent [Sta96]. indeed [Sim97o]. Independence [Sem95b, Sem95c]. Independent [Ano92-159, Dav90f]. Index [Ano95a, Ano95-32, Ano95-34, Ano97b, Ano95-31, Ano95-33]. Index/Sales [Ano95a]. Indiana [Ano96d]. Indigo [Ano93-123]. Industry [Ano93-309, Ano93-287, Ano89b, Ano89a, Ano96-49, Bar90a, Car88b, Cas96, Dav89a, Kel89, Sch93]. industry-by-industry [Dav89a]. inept [Ril96l]. Inexpensive [Ano93u]. infections [Ano95w]. Inferno [Ano96u]. Infiniti [Siv90f]. influence [Rot88a]. info [Ano97t, Kay90]. InfoBahn [FH94]. Information [Ano93-304, Kay95j, Kel90c, Ano96b, Ano97d, Ano97-27, Ano97-32, CM89, Car90g, Fol96d, Ger96a, LS89, McC97c, MCBH88, Ros96, The96b, The96e, Van90, Ano93-124, McC96p, Sem94f]. Informix [Ano96-50]. InfoVista [Ano97t]. infrastructure [Fly96a, Moa95h, Pan96b, Sim96l]. ingenious [Ano95-82]. inheritance [Ano95-112]. inhouse [Ger96b]. Initiating [Dav89d]. ink [Tuc97f]. Innovative [Joy88a]. Innovators [Sim96m]. input [May96]. Ins [Dar95r, MB90]. insider [Sta90c]. insiders [Fis96b]. Inso [Ano96p]. Insoft [Ano93-142]. Insource [Ano93-144, Sha93b]. install [McC96m]. installing [Dar96d]. Insulate [Ber94]. insurance [Ano97k]. integral [Gut96]. Integrate [Ano93-302, Dar96m, Hur94d, The95e, AR96, McC93b, Dar95j]. Integrated [Gil95, McW89b, McC96o, Sim96j]. integrates [TMS96]. Integrating [Tra90a]. Integration [Ano93-76, SF97d, Bro88, Dar96m, Dav91, Eth90a, Hab96, Mar89a, Moa89f, Sim96i, Tuc98b, Ano94-72]. Integrator [Dar96a, Dar96b, Ger96b, Hen94]. Integrators [Ano93-50, Kay96c, KJH88, Moa89b]. Integrity [Ano93-80]. Intel [Ano93-150, Var95a, Hen96, Kay96e, Sim96h]. Intellectual [Ano97z]. intelligent [Ano95-126, Dar96n, McC96o]. intelligently [Scr95b]. intension [Ano95v]. Inter [Val96]. interactive [Jen91, Var96j, Ano97n, Ano97-35]. interactivity [Ano96f]. InterBase [Ano92x]. interchange [Ker88b, Ker89d, McW88a]. InterDev [Ano97-44]. Interest [Eth89, Kel90b]. interested [Dar96q]. interface [SF96e]. Interfaces [SF96e]. Intermezzo [Ano96n]. Internacional [Dar96r]. internal [Dar95c, McC96h]. International [Ano92w, Ano92-44, Ano95-86, Rot89, Ano88a, Str95k, Ano92-85, Ano93-143, Ano95-85, Ano95-83, Ano95-84, Ano95-86]. Internet [Ano94-269, Ano95-134, Ber94, McC96m, Ril95h, Sem95h, Fly96a, Ano96-32, Ano96k, Ano96e, Ano96-42, Ano96-30, Ano96-48, Ano97s, App95c, Cas96, CS95, Dar96k, Dar96s, Dar96t, Dav97b, Dri96, Fis96b, Fol96b, Fre96a, Fry96, Hay96e, Hur96, Kay97, Kno96a, Kno97k, LP96, McC96b, McC96e, McC96g, McC96h, McC96n, Rad96c, Ril97f, SF96b, SF96p, SF96m, SF96h, SF96j, Sco96, Sem96d, SF96q, TMS96, VM96c, Wat96c]. Internet-enabled [Fol96b]. Internetworking [Ano93-314]. interns [Sta89c]. InterOffice [Fol96a]. interoperability [TMS96]. interview [McC96f]. intraenterprise [Kay96f]. intranet [Ano96-50, Ano97g, Fly96a, Lin96c, McC97b, Rad96c, SF96d, Sno98, Tit96, Ano96f, Kay96f]. Intranets [McC96h, SF96d, Fry96, Var96k, Ano96w]. intrepid [SF95g]. introduces [Ano95-47]. Intruders [Ber94]. Inventory [Sem94f, Jag96a, SF96r]. invest [FR89]. Investment [Ano94-109, Hay97p, Moa95h, Sem94a, Wil98]. Invisible [SF96e]. inviting [Gut96]. IP [Ale96b, Ano92-166, Ano93-203, McC95i, McC96g, Str94h, TMS96, Ano93-297]. IPO [Ano95-27]. Ireland [Ano97j]. Iron [Pan96a, PC89, Sim97c, Sim97b, Var96j]. Irvine [Ano95-78]. IS-department [Ova96]. IS-Friendly [The96c]. IS-user [Dug96a]. ISDN [Ano96u, Dar96s, Ker90a, McC96m, Mus88, Str94c]. islands [Mus88]. isn’t [Gre96, Joh90c, Roc97d, SF96k, Sim96n, The94b]. ISO [Ano93-112, Moa93b]. ISPs [Dar96t, McC96g]. ISSC [Ano95-149]. issue [App97b, Sim97k, Ano95-34]. issues [Ano95-86, Dar96c, Hay96a, Str95k]. IT-savvy [Ano95-78]. itself [Hab96, Ril96e, Rob96, Vog96a].

J. [Ano92-84]. Jackpot [Fre97e]. Jaffray [Dar96o]. Jaggies [Ano92-130]. James [McC96f]. January [Ano95-113]. Japan [Als88, Dam89, Joh90c]. Jaques [Ano96d]. Java [Ano96u, Dav97d, FC96, Kno96b, LP96, Lin96c, McC96f, Mee96, SF96f, SF97k, Sco96, Sem96d, SF96q, vSS96]. Java-compiled [LP96]. JavaMan [Ano97m]. JBA [Ano92-85]. JFK [Ano94-215]. Job [Ano92-118, Sem95a, Ano96-54, Gre96, Rad92a, SF95g, Ano93-150]. Jobs [BS94, Sem94b, McC93f, Ano93-210]. Johnson [Pan90a]. Join [Ano92-153, Hir96]. Jones [Esk97m]. Jonesy [Ano95-102]. judge [Moa89a]. juggle [Moa88a]. juggling [Tuc97g]. Jump [McC96h, Kno96d, Sim96h]. jumps [Var96j]. jungle [Tuc97h]. junkies [Ano96k]. Just [Ano93-233, Dar96r, Gar95c, LP96, McC96b, Ste96a, The94f, The95h, Ale96a, Ano94-278, Ano95-70, Ano96-42, Ano98g, App96d, AR96, Dar96b, Dar96o, Dar96s, Esk97i, FT88, Fis96b, Fra89h, Fry96, Kir89, Mae96, McC96a, Pat96, Sar95, Sem96d, SF96q, Sta93, Str96c, The96d, Str94h]. just-in-time [Ano94-278]. justify [Fre97d]. Justifying [May90]. JWorks [Ano96-50].

kage [Ano92-147]. Keep [Ano92-134, Ano93-320, Cas95a, Dar96m, McC96i, Spe96, LP96, Sch91, Var96f, Wat96b, Cas95b, Cas95c, The95l]. Keeps [Ano92-169]. Kentucky [The94c]. Kerberos [Ano93-29]. kernel [Ras89]. Kevin [Mit96]. Key [Ano93-312, Sch95f, Boo90, Cas96, Cus93, Fos90b, MBGS88]. Kick [Rya94]. Kidding [Ano96-41, Ano96-42]. kids [Ano96k, Fol96f].

Kill [Ano93-116, McC96l]. killed [Kno97b]. killer [Bla96, Fol96g, Ano94-182]. Killing [Ano96-51, Asb97b, Vog96b]. Kimberly [Fra90c]. Kimberly-Clark [Fra90c]. Kind [Ano92-51, Tuc97c]. kinks [Ano97-50]. kitchen [Ano95-85]. knew [App96i]. knight [Tat89b]. Knock [Ano92-141]. Know [Ano92-162, Ano94-128, Vog96a, Ano95-133, Ano96b, Hay96e, Hin92, Hir96, McC96j, Mee96, Ril96d, SF95f, SF95h, Sem96f, The96b, Liv95, SF95b]. knowledge [MF96]. knowledgebase [Vog96a]. KnowledgeWare [Ano93-151]. Konica [The96a]. Korean [Gad89].

l [Dav89d]. lab [Siv90h]. LabVIEW [Ano98g]. lack [Hay96a, SF95i]. Ladder [Moa95i, Fra89a]. LAGO [Ano93-152]. laid [Ril96p]. laissez [Ker89d]. laissez-faire [Ker89d]. Lan [Ano93-306, Ano92-95, Ano92-101, Ano93-75, Ano93-316, Ano93-172, Ano93-211, Ano93-352, Ano94x, Ano94-108, App93b, Cas96, Cas90b, Con92, Dar95b, Dar96l, Der92, Hod89b, Hok96, Lin94a, MF90, McC92, McC90c, McC93c, McC93e, McC94d, Pal92, Ric93d, Sch91, Sch94, Sch95d, Sch95e, Ano93r, Ano93-108, Ano93-153, Ano93-154, Ano93-155, Ano93i, Ano94-129]. LAN-based [MF90]. LAN-to-As [Hok96]. LAN-to-As/400 [Hok96]. Land [Ano97x, Kay96a]. landings [McW89c]. Landmark [Ano93-156]. landscape [Cas96]. lane [McC96p, Ano94-200]. LANF [Dar96l]. language [Ano93-350, Sem96d]. languages [Car89a, Jua90]. LANs [Ano93-269, Ano93-335, Dav89c, Dav90b, Ker89g, McC95a, McC90d, Rad93, Ano92-87, Hur94d]. LANSCAPE [Ril96b, Ril96c, Ril96e, Ril96f, Ril96d, Ril96l, Ril96g, Ril96h, Ril96i, Ril96j, Ril96a, Ril96k, Ril97a, Ril97b, Ril97c, Ril97d, Ril97e, Ril97g, Ril97f, Ano94-139, Ano94-140, Ano94-141, Ano94-142, Ano94-143, Ano94-132, Ano94-133, Ano94-137, Ano94-138, Ano95-97, Ano95-96, Ano95-87, Ano95-88, Ano95-89, Ano95-90, Ano95-91, Ano95-92, Ano95-93, Ano95-94, Ano95-95, Ril95b, Ril95c, Ril95e, Ril95d, Ril95f, Ril95h, Ril95g, Ril95i, Ril95k, Ril95j, Ril95l, Ano92-123, Ano94-130, Ano94-131, Ano94-134, Ano94-135, Ano94-136, Ano95-98]. laptop [Lin91a, O’C91]. laptops [FT88]. Large [Moa90b, Ano95z, Ano96-28, Ker89l]. large-scale [Ano96-28]. large-volume [Ano95z]. Larry [Pat96]. Laser [Ano93-210, McC93f]. LaserJet [Bro90]. lasers [Sne91a]. LaserVend [Ano95-136]. Last [Str95a, Var95a, Ano93-311, Mae96, Moa94b, SF96a, Sim96j, Ano92-151]. late [Ano95-56, Gib97b, Ril95f]. Lately [SF96d]. later [Ril96a]. Latest [Ano92v, Ano93-118, Ano96i, Ano96s, Ano96-42, Fol96b, McM90e]. launched [Arm90b]. Launches [Sim96j, Ano97n]. law [Sch88a]. lay [Var96k]. lays [Fra89e]. Lead [The94a, Ano94-28, Kee98b, Sem96c]. Leaders [Ano93-103, Hal96]. Leadership [Mey97, Esk97k, Sch90a]. Leading [Ano93-79, Ano96-36, Roc98a, Ano93-311, Ano95-28, Ano96a]. league [App96a]. Leaner [Ano92-47]. Leap [Ano93-118]. Leaping [Car89a, Dar95a]. leaps [Car90e]. Learn [Kno96d, Ano94-278]. learned [McC96h]. learning [Fre97d, Lin96c]. Leased [Ano94-73]. Leasing [Sne91b, Hal90, Ham90a]. least [Ale96c, Ano95w, Ano95-135, Gre88, McC96l]. least-cost [Gre88]. leave [Ano95y, Ano96-50]. ledgers [The93a, Ano93-213]. Lee [Ano97-32]. left [Ano96-52, HE90]. Legacy [Ano93-319, AR96, Dar95e, Dar95q, Var96h, Dar96l, Gil96a, Tuc97b, Var96i, Win95]. Legal [Ker89d, Ril95a]. Legent [Ano92-86]. Legent/Goal [Ano92-86]. Legitimizes [Ano92-77]. Legos [OHE94]. lend [Ril96e]. Less [Ano93-86, Ril97b, Ano93n, Ano93-190, Ano93-351]. lesson [MS95]. Let [Gre96, Ham94a, Ril96o, Sch95i, Sim96l, Tit96, Ano96s, App96e, Dar96f, Hay97l, McC96k, Moa95c]. Lets [Wil94, Ano96-50, McC96c, Var96f]. Letter [Dal98, Dav96c, Dav96a, Dav96d, Dav96b, Dav96e, Dav97b, Dav97c, Dav97f, Dav97d, Dav97g, Dav97e, Dav97a, Dav97h, Dav97i, Dav97k, Dav97j, Ano95-71, Sch89e]. Letters [Ano93-157, Ano93-158, Ano93-161, Ano93-159, Ano93-160, Ano93-162, Ano93-163, Ano93-164, Ano93-165, Ano93-166, Ano93-167, Ano93-168, Ano93-169, Ano94-144, Ano94-145, Ano94-146, Ano94-147, Ano94-148, Ano94-149, Ano94-150, Ano94-151, Ano94-152]. level [Kee97a, Sar95, Sim96b, Sim96t, The96a]. levels [Kay96b]. Leverage [Ham94c]. Leverages [Ano92-72, Dit92]. Levi [Ano93-189]. liberally [The96b]. libraries [Ano95-77]. licensing [McM90j, Ril95a, Ril95j, Sim96r]. lie [Ano95-98, Hay98b]. Life [Ano93z, Ano93-213, Ano93-214, SF97e, Bee89b, Fol96h, Gib97a, Gol95, SF96l, SF97g, Sco96, Str95l, The93a, Ano93-170]. Lifting [Ano94-38]. Lighter [Ano92-47]. lights [Cas90b]. like [Fol96f, Hok96, McC96d, McC96l]. likelihood [Ano95-59]. likely [FC96]. Lilly [Ano95-70]. Limb [Moa95e, SF96o, SF98]. Limelight [Ano93a]. limit [Ste96a]. limits [Ano97-44]. Line [Ano92-48, Ano93-245, Ano96p, Ano95-132, Ano96g, Ano96-51, Ano97d, Ano97-32, Ano97-48, Ano98f, Dav90d, Hay96c, McC96d, McC97a, McC97c, Mey97, MK90c, Pan96a, Ryp88, Siv90d, Woo92]. Lines [Ano94-73, Var96c]. lining [de 97i]. Link [Ano93-172, Ano93-195, Ano96u, App93a, Dur94e, Fre96b, McC93e, Var96h, Var96d, Ano96d, Dar96k, Dar96t, Hok96, Moa88b, Tab92b, Var96i, Ano93-171]. Linking [Var96d, Fol96h]. Links [Ano94w, Dar95e, Fis96b, Hug90a, McC94a, Ano93-209]. Linux [Kau96]. listed [Ano97z]. listing [Ano95-32, Ano95-34, Ano97-27]. lists [Ano97-27]. literally [Ano96b]. Litlenet [Ano97n]. litter [Ril95l]. Little [Ano96r, The94f, Kay96e, Ril95f, Ril96f]. Live [McC95f, Ano96l]. lives [App96f, Ril96j]. Living [McC95h]. Lloyd [Sed90a]. load [Ste96a, Str95a]. loafers [Fol96f]. Local [Ano93-279, Dar95t, TB90]. Lock [Ano93-252, dJ96b, Ano90b, Dav96d]. logical [Dav90c]. Logics [Ano93-231]. Long [Sem94g, Fol96g, Hur96, Mar89b, Rad96b, Ril96g, SF95d]. long-term [Fol96g]. longer [Ano94-278, Jen89a, McC96a, Pat96, Rad96c, Sim97n]. Look [Ano92-108, Ano92-129, Hab95b, Pan96c, Sem96b, Sem96a, Wat96c, Ano96b, Dar96h, Hab95a, Hab96, Hal90, Liv92, McC96b, McC96h, Sta89f, The96c, Var96b, Var96h, Wal90a, Ano93-189, Ano93-173, Ano93-174, Ano93-175, Ano93-176, Ano93-177, Ano93-178, Ano93-179, Ano93-180, Ano93-181, Ano93-182, Ano93-183, Ano93-184, Ano93-185, Ano93-186, Ano93-187, Ano93-188, Ano93-338, Ano94-153, Ano94-154, Ano94-155, Ano94-156, Ano94-157, Ano94-158, Ano94-159, Ano94-160, Ano94-161, Ano94-162, Dar95l, The94f]. looked [Dar96a]. Looking [Ano95x, Ano96-34, App96g, Val96, Var96b]. looks [Gil96a]. Loom [Ano96g]. looming [Ano95-129, SF95e]. loose [Sim96j]. Lords [Tuc97h]. lose [Hay96d]. Loses [Ano96r]. Losing [Ril95e]. Loss [Ano93-122]. Lost [Ano94e, Ano95-102, Ano95-99, Ano95-100, Ano95-101]. Lot [Ano96r, Fol96c, Kay96a, Pan96c, Ril95e, Ril96h, Sar95, Sem96d, SF96q, Str95f, Tay96, Var96g, vSS96]. lotta [App96f]. Lotus [Ano93-62, Ano95-113, Ano97n, Hur96, McC95g]. Lotus-1-2-3 [Ano93-62]. loves [Ano95-148]. Low [Ano93-80, Ano93-307, McC92, Gil96a, Hay97p, Hea90a, Kay96b, Ric90a, Sha93c, Ano92-87, Ano94-163]. Low-cost [McC92, Ric90a, Ano92-87, Ano94-163]. Low-End [Ano93-307, Gil96a, Sha93c]. Lower [Dug97d, Str94j, McC96c, Sim96r, Sim96s, Tay96, Hur95b]. lowly [Dri96]. lows [CF91]. Loyalties [Sta89d]. loyalty [Var96f]. LTV [Chi92]. LU [Dav90e]. Lufthansa [Gut92, Ano92-155]. lurking [Lin96a]. Luster [Ano96n].

Ma [Ano94-162]. Mac [McM90h, Ric90a]. Mach [Ras89]. machine [Sim96u, The92a, LP96, Tra95]. machines [Ano95-136, Gre88, Var96a]. Macintosh [Fra89h]. MacMaverick [Fra89h]. Macros [Ano93-204]. Macs [Sch90b]. made [Ano97-44, McC89a, Str96b, Hen95]. Madness [Moa95d]. Maestro [Ano98c]. magazine [Str96a]. Magic [Ano92-135, Dav94, SF96c, Kir89, Str95g]. Mail [Ano92-96, Ano93-246, Ano93-274, Ano93-273, Ano93-291, Ano93-324, Cas90a, Ano90c, Ano96s, Chi91, Chi92, Dav89e, Fol96h, Fra93d, Ker89a, MF90, McC93b, Ril96e, Sch89e, Ste92a, Str93, The91, The93e, The93f, Ano93-302, Ano94-269]. mail-enabled [Chi92]. main [Var96a]. Mainframe [Ano92-87, Ano92-88, Ano92-90, Ano93-281, Ano96-51, Ric94f, Ric94e, Sim95h, Ano97-32, Sim95b, Sne92c, Sta89d, Ste96a, Ano92-89, Ano93-190, Ano93-265, Ano93-352, Bau94c, Dar95m, Sim95c, Ano95y]. Mainframe-Killing [Ano96-51]. Mainframes [Ano92-155, Ano93-342, App95d, App95e, Lin96b, Ano97-48, App95a, Dar96r, Gil96a, Gut92, Kre93c, Sim96d, Sim97g, Sim97a, Str94b, Var96j, Gut95a, Gut95b, Sha95, Str94f, Wil94]. mainstream [Ano96-54]. maintain [Str96c]. Maintaining [Moa90c, Ano95-29]. Maintenance [McC94c, Ano94-273, Asb97c, Ker89b, Vog96b]. major [Fol96d, Hay96a]. Make [Ano92-29, Ano92-74, Ano92-101, Ano93-29, Ano93-212, Ano95-149, Ano97-39, Fra91a, Kay96d, Kno96f, McC95k, Ano96k, App96i, App97b, Con92, Dar96c, Esk97c, Fis96a, Fly96b, Fol96h, Ger90, Gre96, Ham96, Inm97a, Jon97a, MF96, McC96h, McC96l, Ril96a, SF96a, Sch92, Sim96h, Sim96v, Str96c, The96d, Vog95, Vog96a, Ano93-191, Ano93-192, Ano94-163]. maker [Dug96d]. Makers [Rad95, Ric94h]. Makes [Ano93-340, FC96, Str94e, The94i, Ano97y, Car90h, Dar95u, Dar96l, Fre97g, Kay96c, Str95g, Var96j]. Making [Ano92-90, Dur94c, Lei90c, Ric93b, Sne92c, VM96c, You97, Ano95-96, McC90f, Mod90, Rob96, VM96a, Ano93-193, Ano93-194, Ano93-195, Ano93-196]. Mall [Var96g]. Malpractice [Blo89a]. man [Ano95-135, Ano98c, Dav90b]. Manage [Ano93-28, Ano93-117, Ano96-37, Dar96n, Gut96, Jag97, MB94, Sch97d, Cor96, FC96, Kno96a, McC96h, McC97d, The96c, McC94d, Str94i]. Manageable [Ano94-172]. Managed [App96g, Tuc97c]. Management [Ano92-32, Ano92-102, Ano92-101, Ano92-143, Ano92-137, Ano93m, Ano93-102, Ano93-156, Ano93-197, Ano93-201, Ano93-209, Ano93-243, Ano93-260, Ano93-331, Ano94-217, Ano94-244, Ano94-241, Ano96n, Bla96, Cor96, Dur95, Hur95b, Ril96m, Ril96p, Ril96n, Ril96o, Sem96g, SF96r, Sim95k, The95e, The96c, Ano96w, Bus88, Car88a, Car89c, Car91, Con92, Dar95f, Dar90, Dav92d, Esk97l, FT88, Fos90b, Fra93a, Gil95, Kee97a, Ker88d, Kir89, Lin94b, Man91, MS95, McC96o, Moa89i, Rea95, Rec89, Sim95a, Sim96m, Sim97f, Sim97e, Sim97k, Ste92a, Str95g, The93g, Tuc97c, Win90, Ano92-96, Ano92-91, Ano92-157, Ano93-27, Ano93-69, Ano93-259, Hur94c, SF96r, Ano95-74, Ano95-132]. Manager [Cor96, Sem94e, Ano97n, McC90c, Sim96u].

Managers [Ano92-163, Ano93-248, Str95a, Ano96y, Ano96-54, App96i, Dug96b, Gil96a, Hay96a, Hen96, dJ96a, McC96m, MK90c, Pan96a, Pan96b, Rad92b, Rad96a, Sem96a, Var96b, Var96c]. manages [App96c]. Managing [Ano92-93, Dav89d, Dug96a, Dug96b, Dug96c, Dug96d, Dug97a, Dug97b, Dug97c, Dug97d, Dug97e, FM89a, MBGS88, Tuc97i, Ano92-92, Ano93-197, Dur94b, McC95i]. Maneuver [Ano93-256]. Manhattan [App96h]. Mania [Fry96]. manipulate [App96e]. Manufacturing [Ano92-94, Ano93-64, Fre97f, Hod89c, Lin92b, May96, Sem95d, Asb98b, Ost90, Ric92a, Sta90d, The95b, Ano92-38]. many [Hay96a, Hay96e, Ril96i, Wat96a]. map [McC97e, Str97c, Ell88]. MAP/ [Ell88]. mapping [OHE94]. march [You97]. margins [Ric94h]. Mark [Pal92, McC96o, Ano92-95]. Market [Sim95h, Als88, Ano95-27, Ano95-86, Ano97l, Car90g, Eth89, Gil96a, Gil96b, Hea90a, Hec96, Her90, Mar89a, McM90c, Pan90b, Pan96b, Ros89, SF95j, Sim96e, Siv90b, Tat89c, Tra89a, Tuc96a, Ano95-160]. Marketing [Ano92-74, Ano96-46, Ano96v, Dar96b, Sch92, Siv90c, Var96f]. markets [Ano89b, Kel89, Moa89g]. marks [Lin92c]. Marriage [Ano93-278]. marrying [Hal96]. Mass [Ano93-76, Ano93-79]. masses [Asb97a, Dar97a, Ano92-158, Ano94-171]. massive [Ano95-128]. massively [Ano95-125]. master [Fre97d, Ger96a]. Mastering [Cod90a, Cod90b, Ste92a, Ano92-96]. Masterpiece [Ano92g, Man92a]. masters [Emm89b]. Masthead [Ano97-29, Ano95a]. Match [Ano92b, Dar95m, Kel90a]. Matches [Ano93h]. math [Esk97f]. MathWorks [Dar96t]. matter [O’C88, Sem96c, Str95h, de 97a]. maturation [GRF90]. mature [Hay96a]. maturity [Hod88]. may [Ano94-215, Ano95-56, Ano95-134, Ano96k, Ano96t, AR96, Dar96q, Hir96, McC96d, McC96g, Ril95l, Ril96h, Sar95, Sim96n, Sim96p, Sim96t, Var96b]. Maybe [Ano95y, Ril95c, Str97a, Tay96, Str97a]. MBAs [Ano95-78]. McCool [McC96q]. MCI [Ano96-48]. meal [Inm97b]. Mean [Cor95c, Sem94h, Bla96, Sim96h]. Meaning [Ano96r]. Means [Ano92-86, Lin95d, Sem96b, Ano97-49, Ril96i]. Measure [Ano92-53, Ano97-40, Rad96a]. Measuring [Hen90]. Media [Ano93-198, Ano93-199, Ano93-200, Ano94-164, Ano94-165, Ano94-166, Ano94-167, Ano94-168]. meet [Ano96w, Dav89b, Dri96, Meh88, Sem96h, Tay90]. Meeting [Ano93-268, Hur96]. Meets [Ano93-202, Ano93-287, The95j, The95k, Ano95-58, Ano97e, Fra93c, Kno97h]. megadeal [Ano97-50]. melting [Cas96]. memo [Pic94a, Ano94-32]. memory [Sem96f, Var96a, Ano93-201]. mends [FS89]. mentality [Hay97d]. mention [Str95g]. mentioned [Ano95-34]. Menu [Ano93-267]. mercenary [Hay97d]. mere [Ham96]. Merger [Ano92-86, Ano93-170]. Mergers [Mae96]. Merle [Var96c]. Merlin [Ano96n]. mesmerized [SF96d]. Mess [Str94i, Jag96a]. messages [Ano96s, Cas90a]. messaging [Dav89e, McC96e, McC92, Str94i]. messy [Lin96a, OHE94]. Metadata [Dar96q]. Method [Lei90c]. methodologies [Dar96d]. methodology [Ano95t, Car88c]. Methods [Lee95, Moa95d, Ano94-278]. Mickey [SF95d]. Micro [Ano95-47]. microprocessor [Sha93d]. microprocessors [Coo92a, Hen96]. Microscope [Ano96n]. Microsoft [Ano92-98, Ano95r, Ano95z, Ano95-59, Ano95-137, OHE94, Ril95j, Sem94g, Ano92-97, Ano92-98, Ano94-169, Ano95-137, Ano96m, Ano96z, Ano96-30, Ano97-44, App96c, Cor96, Dav96c, Dav96d, Fly96a, Fol96a, Fol96c, Fol96f, Fol96h, Hok96, Kno96a, McC96e, McC96n, McM90f, Rad96b, SF96p, SF96h, SF97e, Sem96c, Sha96, Sim96g, Ste96b, VM96a, Wat96c]. Microsoft-based [Sha96]. Microsystem [McC96f]. middle [Ker89k]. Middleware [Fre96b, Ano96j, Ano97r, Fol96g, Gar95a, HS94, Sch95i, Sch95g]. Midrange [Ano92-57, Ano92-93, Sim96m, Eth90b, Ric90c, Sim96e]. Midsized [Tra90b]. might [Inm97b]. migraines [Asb97c]. Migration [Ano92-58, Ano93-124, Sch95b, Dar96r, Fra90f, Fre97e, Mar89b, Sim95b, Sim96i]. millennium [Ano97w, Roc97e, McC96i]. million [Var96c]. Millions [Ano92-74, Sch92]. Mills [Ano96-40]. mind [Key89]. minds [Fly96a, Ano95-112]. Mine [Ale94]. Mini [Ano93-261, Kel90a]. minister [Ano97w]. minisupers [App91]. minus [Esk97j]. Mips [Sim97n]. mis [Ano94-269, Ano92-162, Fra93c, Hin92, Moa94f, The91, Ano93-202]. MISfortune [Ano96-40]. miss [The96b]. missed [Ano95-70]. missile [Ano95-130]. missing [SF95e, Sim97h]. Mission [Ano97-33, App96c, Dar96h, Dar96o, Fol96f]. mission-critical [Dar96h, Fol96f]. mistake [Ano95-56]. mistakes [Ano95-28]. Mitnick [Mit96]. Mix [Ano92b, Sim96i]. Mixing [Ano93-203]. Mobil [Ano97k]. Mobile [Kol94, Ano97-43]. Model [Vin95b, FC96, Fra90a]. Modeling [Fre97g]. Models [Ano92-156]. Modem [Ano92-137, Ano96i, Emm89a]. Modems [Ano93-63, BF90, Bry88, Emm89b, McC93a]. Moderate [Ric94e]. modern [Asb98b, MF96, Ano93-204]. modular [Fra91a]. modularity [Bal96a]. Moguls [Ano93-284]. moment [Ril95l]. monetary [Ano95-132]. Money [Bla96, Mae95, Roc97d, Sim95f, Ale96c, Ano97y, App97b, Esk97c, Ham90a, Hir96, Ril96l, Tuc97k, Str94e]. Monitor [Del94, SF96r, Del94]. Month [Ano96-46, Sim96j]. months [Ano93-350, Hen96]. Moonlight [Fol96b].

Morph [The96d]. Most [App96h, Dar96k, Kau96, Str94o, Dar96n, Gut96, Kay96a, McM90d, Ray95, Spe96, Str96b, Str97d, Tuc97c, Var96g, vSS96]. Motherboard [Ano92-99]. moths [Ano96-42]. Motif [Ano93-231]. Motivating [Cou88]. Mouse [SF95d]. Move [Ano92-93, Ano93-71, Ano93-230, Kay95k, Dar96r, Fre96a, Gar98, Gil96b, Ste96a, The96a]. moved [Ano97-34]. Moves [Ano92-88, Ano92-99, Ano93-68, Gil96b]. Moving [Kor93, SF95f]. MP [Bow94]. MPP [Sim96a]. MQSeries [Tuc97j]. Mr [Sch89e]. Mr. [Esk97m, de 97e]. MRP [Fol88a, Siv89b]. MS [Kno96a, Ano93-205]. MS-DOS [Ano93-205]. MSDN [Ste96b]. Much [Ano94-110, Ano95-125, Ano97p, Asb97b, Hur96, Inm97a, McC96e, McC97e, OHE94, Sim96h, Sim97o, Var96c, Vog96b]. much-delayed [McC96e]. Muddy [Fra89g]. multi [Ano96-51]. multi-CPU [Ano96-51]. Multidimensional [Ric94b, Ano94-170]. Multimedia [Ano92-99, Ano92-100, Ano93u, Ano93-202, Ano93-194, Ano94-278, Kay95c, Sem93, Sne92d, Ano95-47, Fra90k, Fra93c, Kay95a, Kay95d, McC96a, Ric93b, Ano93-206, Ano93-207, Ano94-171, Kay95b, Kay95e]. Multiple [Ano94w, Inm97a, McC94a]. Multiprocessor [Ano93-208]. Multithreaded [Ano94-279]. Multitier [Bal96a]. multitudes [Ano95-77]. multiuser [Fra90f]. multivendor [Ger96b]. Muscle [Ano92-36, Pau91, The95f]. muscular [Dar96h]. Museum [Ano96-46]. Must [Dar95m, Dar96a, Rea95, Tuc97c, Wes90, Ano93-205]. muster [Sim97h]. Mutant [Ano93-116]. muxes [Tay90, Ano92t]. MVS [Sim96r, Ano93-252]. My [Dav96a]. myth [Dav97g]. myths [MCBH88].

Name [Ril95f, Ano95r]. narrows [Joy89a]. NASA [Joy89b]. National [Ano96-48]. nations [Mad90]. NatWest [Ano96g]. NaviGate [Ric91]. Navigating [Moa89f]. NCR [Moa90d]. NCs [Ano97p, Dav96e, Sim97o]. Near [Ano93-245, Var96g, Ano96t]. nearly [The96a]. necessary [Ril96j]. Need [Ano95z, Ano96p, Coh95, Dar96a, Kay96b, Sem94b, Ale96a, Ano95-136, Ano95-159, Dar96e, Dar96h, Dar96k, Gut96, Mee96, SF95e, SF95h, Sem96f, SF96r, Sim96c, Sim96s, The96c, Var96a, Var96b, The95m]. Needed [Gil95, App96d, App96g, App96h, App96i, Bal90, Dar96t, Ste96a, The96a]. Needs [Ano94y, Rub95a, Ano96-54, App96f, Dri96, dJ96a, Sim96u, The96d, Dar95m, The95j, The95k]. Negotiate [Wra98, Ano95z]. Net [Ano92-157, Ano93-69, Ano96m, Ano97n, Ano98b, Hur95b, Kno97l, McC96i, Pat96, SF95h, SF96b, Str95a, Ker90c, McC96l, Roc97a, Wil97a, Ano92-103, Ano93-209, Ano94-69, Ano96r, Ano98c, Moa94j, Tuc97e]. Net2000 [McC96k]. Nets [Ano92-37, Ano93-346, McC95i, Arn90, Pan97, McC96h]. Netscape [Ano95-27, Fly96a, McC96q, Sem96c, VM96a]. NetView [Ano93-109, Ano94-267]. NetWare [Ano94-182, Ano95-131, Ano95-129, Hur94d, Sim96g, Str94g, Ano93-243, Ano94-172, Hur95a, Hur95c, Str94j]. Network [Ano93k, Ano93w, Ano93-308, Ano93-284, Ano93-341, Ano96e, App93b, Cas96, CS95, FC96, Kay96d, McC95f, McC96k, McC93f, McC94f, Ril95e, Ril96m, Ril96p, Ril96n, Ril96o, Sch97e, Str94k, Ano94-181, Ano95-127, Ano96-48, Ano97n, Ano97v, Cas95a, Dav90g, Dav93a, Fis96b, Fra93c, Gil95, Hab89, Jen89b, Joh90b, Kay96f, Ker88d, McC96b, McC96l, McC97d, McM90j, McW88a, Ric90b, Ril95c, Ril96h, Rub95b, Sch97a, SF96q, Sim96j, Sna90b, Sne91a, Sta96, Ste90, Ano93-202, Ano93-210, Ano93-211, Kay96f, Sim96u, Sim97o, Str94l].

Network-Aware [Kay96d]. Network-Centric [FC96, SF96q]. network-ready [McC96k]. networked [Cas96, SF96k, Kay95c, Kay95d]. Networking [McC95e, Str94f, Ano96l, Ano96-49, Fol96h, McC90a, Sta89b, Str94b, Tay90, TMS96]. Networks [Ale96b, Ano92-80, Sch95d, Sch95e, Ano89c, Ano97-42, Fra90h, McC96g, McM90f, Nor88, Ril96i, TB90, Tuc96a, Wal90a, Ano93q, Ano93-117, Ano93-315, Ano93-278, Ano94-252, Dur94a, Dur94b]. Never [Ano93-327, Dar96r, Str95g]. News [Ano96m, Ano96l, Ano96w, Ano96-38, Ano97m, Ano97n, Ano97p, Ano97o, Ano97r, Ano97q, Ano95-149]. Next [Ano93-352, Ano94-272, Bul95, Moa89g, Sim96d, Sim96n, Ano95-134, AR96, Bal90, Fol88a, Hab96, OHE94, Rad96c, Sim96t, The96a]. Next-generation [Sim96d, Ano95-134, AR96, Rad96c]. NeXTSTEP [Ano93-150]. NFS [Ker89c, Tuc96a]. ng [Dar96a]. niche [Fra91c, Ano93-239]. Nine [Ano93-103]. Nixdorf [Ano96m, Ano97l]. NLM [Ano92-142]. No [Ale96a, Ano93-120, Ano94-162, Dav96b, Dri96, Gar95c, Ger96a, Ger96b, Hay97j, Hay97o, Kir89, McC96a, Ril96f, Sim95c, Sim95k, Sim97k, Str95h, Ano94-278, Ano95-78, Ano95-149, App97b, Cor95a, Ker89h, Pat96, Rad96c, Ril95d, SF96m, Sim95a, Sim97n, Str97a, Str95f, VM96a]. nobody [Ano96-47]. Node [Dur94d]. Nodes [Dur94f, Ano95-30]. noise [Wal89a]. non [Rad92b, Ano93-231]. non-IS [Rad92b]. none [Ril95k]. Norman [Var96c]. North [Ano92h, Ano93-305, Ano95-32, Ano95-37, Ano95-38, Ano93-232]. NOS [Ano97g, Dar95k]. noses [Kau96]. notch [Fol96f]. Notebook [Ano92-47, Ano95-47, Bea96]. Notebooks [Ano92k, Ano93s, Sha94b, Fra91c, Sha93a, Sne92b]. Notes [Ano95-72, Bau95b, Bau95a, Ano93-234, Ano94-68, Ano94-287, Hur96, Kay96f, McC95g, Tab92a]. Noteworthy [Ano92k, Sne92b]. Nothing [Bau95b, Bau95a, Dar96d]. notion [Ale96b]. Novell [Ano95-131, Ano92-102, Ano92-103, Ano93-235, Fol96h, McC96k, The95v]. NT [Ale96c, Ano93-68, Ano94-169, Ano94-286, Ano95-129, Ano96n, Ano96p, Ano97v, Asb98a, Bal96b, Dav96b, Fol96f, Gil96a, Gil96b, Hur95c, Ril95l, SF96k, Sem94h, Sim95j, Sim96b, Sim96i, Sim96n, Sim96t, Sim96v, Sim97h, Sim97n, Tit96, TMS96, Tuc97c, Tuc98b, Var96b, Ano92-98, Ano93-236, Ano93-237, Ano93-238, Ano93-248, Ano94-182, Ano94-238]. NT/RISC [Var96b]. NTs [Ano95-131]. nuggets [Asb97a]. NUMA [Ano96-33, Sim96o]. number [Ano93-350, Ano96b, Ril96g, Sim96e, Var96d]. numbers [Pan96a, Tuc97c, Bow94, The95h]. nurture [Str96c]. Nuts [Moa94f]. NYNEX [Ano96v].

O [Ano93-45, Sim97d, Tuc96a]. OA [Siv90i]. Object [Ano92-120, Ano92-156, Ano93-240, Ano96o, Ano96u, Rym94, Vin95a, Ano93-350, Ano96m, AR96, Bob89, Dav96d, Fly96b, Fre97d, Kim90a, MS95, McC96k, SF96g, SF96j, Sta89f, Wat98, Ano93-239, Ano96l, OH95]. object-oriented [Ano93-350, Bob89, Fre97d, MS95, McC96k, SF96g, Sta89f]. object-wrapping [AR96]. objective [Sta89f]. OBJECTS [Sem95g, Ano94-280, Fol96e, Sem95f, Vin95c, Dav97h, SF95j, Bau94c, Str95m, Sud95, Ano95t]. obscure [Ano94-278]. Obsolescence [Str96c]. Obsydian [Vin95b]. obvious [Sim96p]. OCR [Bal90, Fra91b]. of’em [Ril95e]. Off [Ano94-243, App96f, Bal96b, Dar96q, Ano95u, Ano95-59, Ano95-102, Ano95-134, Dug96b, Gib98, LP96, MF90, McC96i, Ril95h, SF97h, Sch91, Sch97e, Ste96a, Tab92a, Wil98, You97, Ano93-308]. off-load [Ste96a]. Offer [Ano96p, Kay96c, Kay96a, Var96g]. offered [Ano96-32]. Offering [Ril96g, Ano96b]. Offers [App96g, Ano97d, Str97a]. Office [Fol96a, Ano96s, Ano98g, Ber95c, FR89, Fol96b, Fre98a, Pic94c, Ric90d, Ano93-271, Ano94-216, Ano96z, Fol96b]. Offices [Dur94e]. OfficeVision [Tab92a]. Official [Sha94a]. Offshore [Ano93-339]. Oil [Fra90e, WN90]. Oject [Moa95d]. OK [Bla96]. OK. [Sim96p]. Okay [Ril95f, Val96, Ril96g]. O’Lakes [Ano97x]. OLAP [Ano96h, Ano96n, Ano96-39, Ano97x, App96e, Dar96f, Pen97, Wat96b, The95o, Var95c]. Old [Ano93-213, Ano93-214, Ano94-215, PC89, The93a, Tuc97g, Ano92-105]. OLE [Ano95-114, Ano96-31, OHE94, SF95d, SF95j, SF96k, Str95n, Vog95, Ano93-242]. OLTP [Ano93-249, Ano94d, Del94, McW89a, Str95f, Str95i, Ano92-107, Ano93-334, Sch95f]. olympian [McW88b]. Olympics [Sim96l]. On-Line [Ano96p, Dav90d, Ano97d, Ano97-32, Ano97-48, Ryp88]. On-Site [App96h]. once [Hen96]. One [Ano93y, Ano93-244, Ano93-344, Ano94-278, Dar95s, Dar96t, Mee96, Ste96a, The96d, Wat96c, Ano95-135, Ano96h, App96d, App97b, Bla88, CS95, Dar96o, Hir96, Kel89, Lin93, Ova96, Ril95e, Ril95f, SF95h, Sim96b, Sim96t, Sim97f, The96b, The96e, Var96b, Ano93-325, Dar95n, Moa94d]. One-Size-Fits-All [Ano93y]. one-upmanship [Ano96h]. ones [Var96f]. online [Dav89f, Ano97n]. Only [Ano96-41, SF95f, Ano95-86, Ano96-35, Ano96-54, Hur96, SF96f, Sim96n, Sim96u, Var96a]. ons [Fly96b]. Onsite [The94h]. Onto [Ano93-150]. OO [Ano95-112, Dar96e, Fre97d]. ools [Fol96g]. OOP [Ano92-167, Ano93-115]. Open [Ano92-94, Ano92-148, Ano93-28, Ano93-124, Ano93-326, Ano93-334, Ano96r, Dar96d, Her90, Lin92b, Mar89e, McC96g, McC90f, Moa88a, Moa90e, Ras89, Run89c, Sim97g, Siv90e, Ste89, Ste90, Ano93-243, Ano94-183, Del94, The95l]. OpenDoc [Ano96l, McC96k, OHE94]. Opening [Ano93-328, Heb90, Asb97a]. Opens [Eld95, Kau96, Ano97n]. Operating [Ano93-323, Moa94d, Ale96c, Ano94-181, App96c, Bal96b, Dav93b, Eth90b, Fol96g, Ano93-244]. operating-system [Ale96c]. operation [DD89, Fre97h, Mik97j]. operations [Ste89, Tuc96a]. Operator [Ano92-41]. Opinions [Ano96-32]. Opportunities [Ano97f, Heb90, Ano92o, Ano93-30, Ano93-31, Ano93-32, Ano93-33, Ano93-34, Ano93-35, Ano93-36, Ano93-37, Ano93-38, Ano93-39, Ano93-40, Ano93-41, Ano93-42, Ano93-43, Ano93-328]. Opportunity [Kay96a, Run89c]. optical [Dav92c, Fra88, Wal90a, Ano93-245]. optimistic [Dar96q]. option [Fol89, Tuc98b, Ano93-269]. Options [Sim96k, Hab89, Sim96i]. Oracle [Ano92-106, Ano92-107, Ano96z, Fol96a, McW88b]. Orchestrate [Sem94c]. Order [Ano93-246, Ano95-73, Fra93d, Pan90a, Sch95l, The91]. Ordering [Ano98g]. ordinary [Tuc97l]. organization [Car90f, Inm97a, MF96, Ova96, Rec89, Ano94-281, Sta95c]. Organizational [Pin90]. organizations [Jon97a]. Organizing [Was90]. oriented [Ano93-350, Bob89, Fre97d, MS95, McC96k, SF96g, Sta89f]. Original [Rho95b]. originally [Sim97o]. OS/ [Ano92-151, Sem90]. OS/2 [Ano93-119, Ano93-151, Ano93-240, Ano94-243, Sim96j, Fra89d, Fra89j, Fra90f, Sim96j]. OS/390 [Sim96r]. OS/400 [Tay96]. other [Ano95-70, FM97, Hal96, Kay96c, Kay96f, Mae96, SF95f, Sim96p]. Our [Ano93-311, Ano95-27, Ano95-113, Esk97c, Esk97i, Fra93d, SF95g, SF96g, Str96b, Wat96a, Ano95-86, Ril96j, SF96b, SF97a, Str95k, Wat96b, Ano93-246]. ours [SF96a]. outdated [Ril96l]. Outfitting [Pic94b, Ano94-184]. outlandish [Ale96a]. Outlook [Ano93-287]. Output [Ano92-126]. outrageous [Dar96c]. Outs [Dar95r, MB90]. Outsource [App96a, Asb97c, Bla95b, Kay95k, Kay96f, Dug97c, Rad93, Scr95b]. Outsourcer [Gut96, Kay96c, Ano95-159, App96g, Dar96a, Ger95]. Outsourcers [Kay96c, Rub95b]. Outsourcing [Ano92-108, Ano93-269, App96d, App96g, Asb97d, Dar96d, Kel90d, Liv92, Ric94c, Ano94-28, Ano95-28, Ano96-32, App96a, Dug97d, Gar98, Ger95, Kay96c, Kno97b, Moa93d, Ano94-185, Mik97f, Moa94e, Moa95e]. overall [The96d]. Overnight [Jon97c, Jon97b, Jon98]. overseeing [Ano96b]. Overwhelmed [MB94]. Overwork [Ano94-112]. Own [Str94n, Gol95, Jag97, McC96h, Ril95c, The96a]. ownership [Gib98]. Oxymoron [Sem94d]. Ozone [Ano95-102, Ano95-99, Ano95-100, Ano95-101].

p [McC96h, Ale95f, Ale95g]. P6 [Var95a]. Pace [Ano93-320, Ano96-52]. Pacific [Ano94-273, Bau96b, Ker89j, Ost90, Siv89b]. PacifiCare [App96e]. pack [Ril97c]. Package [The95u, Rad96c, The96d, Ano93-62]. Packaged [Ano93v, Dar97b]. packages [Dav90c, Sim96r, Spe96]. Packard [Ano96r]. page [Ano97-34, App96a, Mee96]. pages [Ano96f, Sem96a]. Pain [Ano94-108, Kay96a, Tra95, Var96g]. Paint [Ano92-34]. pair [SF96g]. Palmtops [Ano93-310, The93b]. panacea [Sim95a, Sim95k]. Paper [The93c, Bla90, Vog95, Ano93-247]. Paper-free [The93c, Ano93-247]. paperless [Ano97-49]. papermate [Esk98b]. Paradigm [Dar96b, SF96a, SF96i, SF96b, SF96n, SF96c, SF96p, SF96d, SF96e, SF96m, SF96f, SF96g, SF96h, SF96o, SF96l, SF96j, SF96k, SF97a, SF97b, SF97c, SF97d, SF97e, SF97h, SF97f, SF97g, SF97j, SF97i, SF97k, SF98, Mik97d, SF95e, Ano93-44, Ano94-186, Ano94-187, Ano94-188, Ano94-189, Ano94-190, Ano94-191, Ano94-192, Ano94-193, Ano94-194, Ano94-195, Ano94-196, Ano94-197, Ano94-198, Ano94-199, Ano95-112, Ano95-113, Ano95-114, Ano95-103, Ano95-104, Ano95-105, Ano95-106, Ano95-107, Ano95-108, Ano95-109, Ano95-110, Ano95-111, SF95a, SF95f, SF95b, SF95d, SF95c, SF95e, SF95g, SF95i, SF95h, SF95j, SF95l]. Paradox [Ano92x, The95w]. Parallel [Ale94, Ano92-107, Wat97, Sch89a, Ric94e, Wil94, Ano95-125]. Parallelism [Ano92-106]. Part [Dar96f, Gut96, Ril95j, Ril96j, SF95i, SF96f, SF95l]. part-time [Ril95j]. parties [Hok96, Sim96s]. Partitions [Dar96p]. partner [Fre97i, Sta90b, Wes90]. Partnering [Dar96b, Tuc97e]. Partners [Ano93-351, Cla90, Dar95d, FM97]. partnership [Dar96b]. parts [Ano96l, OH95]. party [Fol88b, FS89]. pass [Sim97h]. passage [Joy88b]. passion [SF96n]. Password [Spe96]. passwords [Ano95-96]. Past [SF97f, Ano97k]. paste [Jef97b]. Path [Ano96n, Ano95-136, Ano92-34]. paths [Fra89a]. patient [Mik97j]. Patricia [Kno97b]. Patrolling [The95p]. patterns [Var96f]. Pave [Ano93-335]. Paves [Ano93-153]. Paving [Ano92-109]. Pay [Ano94-220, App97b, Esk97a, Sim95g, You97, App94]. payable [Ano97x]. payback [Tuc97d]. payments [Ano97i]. payoff [Kay97, Sim97n, Ano94-108]. Payoffs [Sta89g]. pays [Gib98, SF97h, Wil98]. PBX [Str94m]. PC [Ale95f, Ale95g, Ale95b, Ano92-39, Ano92-104, Ano92-152, Ano92-163, Ano93-300, Ano93-82, Ano93-311, Ano93h, Arn90, Bar90c, Bea96, Bla96, Bry89b, Coo92b, Cor95b, Dar95d, Dar95f, DD89, Fra88, Fra89a, Fra89g, FM89b, Fra89i, Fra92a, Fra92b, Gil91, Gil96b, Hab94, Ham90b, Hay96d, Hea90a, Hur94c, Joy89a, Ker88a, McM89a, Meh88, Moa89i, Ric93c, Ric94g, Ric94h, Sim97k, Siv90b, Ste92b, Tuc96a, Woo91, Wor90, Ano92-110, Ano93-248]. PC/ [Coo92b]. PC/workstation [Cor95b, Hea90a]. PC50 [Ano93-311]. PCI [Ano94-200]. PCMCIA [Ano94-283]. PCs [Ano92-28, Ano93-311, Ano94-128, Ano96m, Dar96d, Dav96e, Fra91a, Fra93d, Kay96e, McM90c, Pat96, Pau91, Pin92, Ril96h, Sne93, The92b, Ano92-75, Ano93-89, Ano93-150, Ano93-246, Ano93-249, Ano95-56]. peak [Ano89a]. peek [App96e]. Peering [Dav90e]. Pen [Ano93-149, Ano93-323, Fra91c, McC94e, The93d, Dav93b, Fol93, Man92b, Ano93-250, Ano93-251]. Pen-based [Fra91c, Man92b]. Pentagon [Ano95-130]. Pentium [Ale95f, Ale95g, Ale95b, Ano93h, Ano95-47, Kay96e, Sim96h, Ano93-257, Cor95b, Hur94b, Kay96e, Sha93d, Sim96h]. Pentiums [Ano94-239]. People [Ano93-312, Ano95-133, Arm90a, Cus93, Dug97e, vSS96]. percentage [Ano96-28]. perfect [Dar96d, Ril95l]. PerfectOffice [The95v]. perform [Kno96b, MF96]. Performance [Ano92-31, Ano93-80, Ano93-87, Ano93-156, Ano93-292, Ham94a, Ham94d, McC96l, Moa95b, Ano94-278, BF90, Ham96, Hea89, McC97e, Tuc96a, Tuc97i, Ama95a, Ano93-295, Ano93-333, Ham94b]. peripheral [Bla88]. peripherals [Ano96-28]. perish [Hay97n]. Perle [Ano92-111]. person [Ano95-132]. person-hours [Ano95-132]. personnel [Cou88, Rid88, Roc97f]. Perspective [Bea96, Siv89a, Ano93-148]. Peugeot [Eth90a]. Phone [Ano93-149, Dav89e, Sch94, Fol93, Wil97f, Wra98]. physique [Dar96n]. Pick [Ano93-250, Fre97i, HS94, Ril95l, SF96g, Ano94-201, Sha94b, Sha94c, The94g]. picked [Dar96o]. picking [Ano96x]. Picks [Ano95-47, Sim96m, Ano94-53, Ano94-54, Ano94-55, Ano94-56, Ano94-57, Ano94-58, Ano94-59, Ano94-60, Ano94-61, Ano94-63, Ano94-64, Ano94-65, Ano95-39, Ano95-40, Ano95-41, Ano95-42, Ano95-43, Ano95-44, Ano95-45, Ano95-46, Ano94-62]. Picture [Ano94-67, Ano94-282, Kno97g, SF96f]. Pictures [Ano93-288, Ano94b]. pike [Sta89g]. Pilot [Ano96r]. pioneer [Wat96b, Ano96v]. pioneers [Gad89]. Pipe [McC96d]. pipeline [Bou90]. Piper [Dar96o]. pirates [Mar89c]. pit [Bla90]. pitch [Ker89b]. Pitfalls [Ano93o, Ano93p, LS89, Gre96, Ril96j]. Pitt [Ano96-33]. place [Ano95x, Kay96c, Ril95h, Ril95d]. placing [Ano95-136]. Plan [Ano93f, Ano93-261, App96g, Ham96, Jag96a, McC95j, McC96g, Pal93, McC95j, Ano93-272, Ano96u, Str94k]. Planned [Str96c]. Planning [Ano93d, Ros89, Wal89b, Bry89a, LS89, McC96l, McC89a, Sim96g]. plans [McC96n, Ril96p]. Platform [Ano93-101, SF96q, Ano97u, Ano97g, Ano92-62]. Platforms [Ano93-243, Ril96e, Sim96l, Sim96p, Tuc97c, Var94a]. Platt [Ano92-112]. Play [Ano92-42, Ano94-283, Ano96p, Cor95c, Fis96a, Fol96a, Dar96f, Hab94, Hea90b, Maj90, Ril97f, The95q, TB90, Ano93-346]. players [Cas96, Ano93-270]. Playing [Ker89g, McC90d]. Plays [Ano93-208, Kno96a, Mik97b, Ric90a]. Plenty [Pau95a, Pau95b]. ’plex [Sim97c]. plight [Moa88c]. Plore [Ano92-171]. plowing [Hay96a]. Ploy [Ano96-46]. Plug [Ano94-283, Ano96p, Ano97-42, Cor95c, Fis96a, The95q, Hab94, Sim95b]. Plug-and-Play [Ano96p]. Plugin [Ano97-36, Ano97-38, Ano97-39, Ano97-40, Ano97-41, Ano97-37, Ano98e, Ano98f]. Plumbing [Rad96c]. plummeting [Ano93-311]. plunge [SF96b]. Plus [Ano93-83, Ano93-254, Cas96, Sne93]. PM [Ano94-128]. PnP [Fis96a]. Point [Ano92q, Dav89f, Esk97d, Esk97e, Esk97j, Esk97f, Esk97g, Esk97k, Esk97h, Esk97i, Esk97l, Esk97m, Esk98a, Esk98b, Fol96i, Hay96b, Hay97b, Sem96h, Sem96i, Ano95-80, Fra97]. policy [Ril96l]. Politics [Ano92-52, Hal97, Ker88d]. pollution [Tra89b]. polymorphism [Ano95-112]. pop [Kay96a, Var96g]. popular [Fol88b]. popularized [Ale96b]. Port [Tay96]. portability [Mar89e, Lin95b]. portable [Fra92a, Sne91c]. Portables [Hal88]. portfolio [Fre97i]. ports [Jud90b]. position [Ano95-71]. POSIX [Ano93-252]. possible [Fly96b, McC96l]. possibly [Sem96f]. Post [Fol88a, MF96, Ruh88]. post-implementation [Ruh88]. post-modern [MF96]. Post-MRP [Fol88a]. postings [Sta96]. Postman [Sch89e]. PostScript [Sne92e, Ano92-113]. potato [Str95e]. pound [Ano95-47]. pour [Tuc97k]. powder [Str95e]. Power [Ale95b, Ano92-89, Ano92-97, Ano92-155, Ano93-190, Ano93-238, Ano93-330, Ano94-271, Kay95i, Arn90, Bro90, Dar96p, Fra89f, Gut92, Hea90b, Lin91a, SF97i, Sem90, Tuc97m, Ull94, Wal89a, Ale95f, Ale95g, Ano93-204, Ano94-251]. PowerBuilder [Kay95i]. PowerBuilding [Ano93-253]. powered [Kay96e]. Powerful [Has95, Jef97b, OHE94]. PowerPC [Ano93-72, Ano93-73, Cor95a]. PowerPCs [Sim97n]. pp [Ano93-87, Ano93-326]. PR [Dug96c]. practical [Dar96l, Ril95i, The93d, Ano93-251]. practically [Dar96l]. practice [Tra89a, Ano95v]. practiced [Ano94-278]. Praise [Ano93-333]. Pre [Ano93-254]. Pre-ATM [Ano93-254]. preach [Tra89a]. precise [Sta96]. predecessor [Sim96h]. predict [Tat89a, Var96f]. predicted [Ano95-113]. predictions [Lin96b, SF96a]. predicts [Var96f]. Prefab [Ano93-255]. Premenos [Ano92-114, Ano93-256].

PREMIERE [Ano96-41]. prepared [Hay97f]. prepares [Dam89]. Preparing [Sha93d, Ano93-257]. preps [Ano97s]. presence [Ano95-129]. Press [Ano96-48, dJ96a, Ano94-215, Ano94-202, Ano94-203, Ano94-204, Ano94-205, Ano94-206, Ano94-207, Ano94-208, Ano94-209, Ano94-210, Ano94-211, Ano94-212, Ano94-213, Ano94-214, Ano95-125, Ano95-127, Ano95-128, Ano95-134, Ano95-126, Ano95-131, Ano95-130, Ano95-132, Ano95-133, Ano95-135, Ano95-137, Ano95-136, Ano95-129, Ano95-115, Ano95-116, Ano95-117, Ano95-118, Ano95-119, Ano95-120, Ano95-121, Ano95-122, Ano95-123, Ano95-124]. Pressure [Ano93-232]. pretty [Hur96, Ova96]. prevent [Var96f]. Prevents [Ano96-30]. Price [Ano92q, Ano92-108, Ano96u, BF90, Gib98, Hay96d, Ker89b, Liv92, Ano93-333]. Price/Performance [Ano93-333]. Prices [Ano93v, Ano93-264, Ano93-311, O’C91, Sim96k, Ano93-82, Ano93-258, Ric94g]. Pricing [Ano93-44, Ano95-28, Dar96t, Kay98, The95v]. pride [Ano97x]. prime [Ano89c, Sim96c]. prime-time [Ano89c]. Primer [Ano93d]. Print [Ano92-115, Sne92f, Pau95a, Pau95b]. Printer [Ano92-37, Ano97n, Bla88, Fol89, Dar95s, McC94f]. printer/copier/scanner/fax [Bla88]. Printers [Ano92-34, Ano93-210, McC93f]. Printing [Ano92-113, Dar95j, Rub95a, Dav92a, Sne92e]. ’prise [Sim96e]. prison [Sta96]. Private [Nor88, Fis96b, Kay96f]. privileges [Ano95-30]. Pro [McM91a, Fra90h, Kay96e, Sim96h]. probably [App96c, Hab96, Ril95l, Str95f]. probe [Ale96b]. Problem [Var96c, Ano95-80, Ano96-35, Ano97m, Ano97r, App97b, Dri96, Hod89c, Jag96a, dJ96a, Lin96a, Sch89a, Sim96b, Sim97f]. Problems [Cel96, Lin96a, Siv90g, Wal89a, Rea95]. Process [Hay97n, Sem96g, SF96r, Hum89, MBGS88, May96, McC96h, Moa90g, Rob96, VM96a]. Processes [Ano96-53, Hay97l]. processing [Bla90, Can88, CF91, Kni90, Lei90e, May90, Sch89a, Sch95l, Sed90a, Siv89a]. processor [Sim96h]. Procurement [Dar96b]. Product [Ano92p, Ano93-103, Ano93-148, Ano93-329, Ano95-34, Ano95-33, Ano90b, Ano95r, Ano95-138, Ano97-32, Dar96k, Hay96c, The96d]. Production [Ano96-37, Ano98c, Vog96c]. productive [DD89, Vog95]. Productivity [Ano92-122, The95w, Car89a, Dar90, Esk97l, Joy88b, Siv90i, Ano93-281]. Products [Ano92a, Ano93-45, Ano93c, Ano93-228, Ano94-66, Ano94-181, Ano95-138, Ano96q, Ano96z, Ano96-49, Ano96-50, Ano96-54, Ano97s, Ano97t, Ano97u, Ano97v, Ano97-42, Ano97-43, Ano97-44, Ano98g, Gre97, Lee95, The94a, Ano95-34, Cor96, Dar96m, Fol96b, Hok96, Lin96c, McC96b, OHE94, Pat96, Ste96b, The96a, The96b, Ano93-143, Ano93-156, Ano93-215, Ano93-216, Ano93-217, Ano93-218, Ano93-219, Ano93-220, Ano93-221, Ano93-222, Ano93-223, Ano93-224, Ano93-225, Ano93-226, Ano93-227, Ano94-173, Ano94-174, Ano94-175, Ano94-176, Ano94-177, Ano94-178, Ano94-179, Ano94-180, Ano95-138, Ano97o, Dar96c, Lin95a, The95m, Ano94-278]. Professionals [Ano96u, Moa89h]. profile [Ano96f]. Profiles [Ano92m, Ano95-32, Ano95-86]. Profit [FH94, Ano97-49, Fre97h, Ham96, Sno98, Tuc97d]. profitability [Str97a]. profitable [Ben98, Tuc97e]. Profits [Ano92n, Ano93x, Esk97b, Fra90c, Ger96a, Hay97a, Kno97d, Var96f, Wil97a, Woo93]. profound [FC96, SF96i]. programmer [Bro89, Fol88b]. Programmers [Ano96r, Fol88b, SF96c, SF96j]. Programming [Ano93-204, Ano93-212, Ano93-255, Ano93-339, Sch89a, Bob89, Car89c, Car89a, Moa89c, OHE94, SF96g, Sta89f]. Programs [Ano93-112, Moa93b, Ano92-140]. progress [Ano95-97]. Progresses [Ano94d]. Project [Ano96-45, Fol96a, McC89a, Pic94c, The93g, The96c, Ano96w, Car91, Dar96r, FM97, Hay97b, War97, Win90, Ano93-259, Ano93-260, Ano94-217, Ano94-216]. projects [Ano96a, Dav89d, Hal97, dJ96b, Rad96a, The96c, You97]. Prolog [SF97f]. prom [Esk98a]. Promise [Ano92-106, Sem96f, Gib89, Hug90b, Sch90c, Siv90g]. promised [Sim97o, Str96a]. promises [Ale96a, Sim96h]. promising [Ano94-278, Dav90b]. Proof [Esk98a, Ano95-82, Ano95-114, Moa90a, Ric93d]. Proofing [MS94, SM94]. proposition [The96b]. proprietary [Eth90b, Ano93-261]. Pros [Ano92-100, Ker89i, Kay96e]. Protecting [Ano95-74]. Protection [Ano93-284, Ano96p, Ril96a, Wal89a, Coh95]. Protocol [Ano95-134, Emm89a, Moa89d]. protocols [McC90b, TMS96, Ano93-241]. prototype [Jef97b]. Prototyping [Til89, CC89]. provide [Ano95-77, Bal96a, Var96d]. Provider [Sha96]. providers [McW88a]. provides [Ano97-32]. providing [Sim96m]. proving [Lin96b]. Pry [Dar96d]. PS [Ano93-214, Ale95b]. P’s-Pentium [Ale95b]. PS/2s [Ano93-214]. Psych [Moa94f]. public [Ano95-127]. Publishing [Ano92-48, Ano92-116, Ano93-247, Ric92d, Ano95t, Fra90b, Lin91b, The93c, Woo92]. pull [Sim95b]. pulling [Ano95-59]. pulls [Kel90d]. Pump [Esk97n, Ano97-35]. Punch [Ano92-107]. Pundits [Dar96q]. purchase [Ano95z]. purchases [Dar95d]. purchasing [Ham90b]. purists [Ano95-112]. purpose [The96c, Var96b]. Pursuing [Bla88, Mar89c]. push [Fra89f, Moa88d, Siv90i]. pushers [Vog95]. Pushes [Ano92q, Ano97-44]. pushing [Sim96c]. Put [Ano93-46, Ano96n, Sim96p, Ano95r, Ano95-159, Ril95c, Ril96b, Sem96e, Ano93-262]. Puts [Ano94z]. Putting [Ano95-132, Kay96e, Ano92-117]. puzzle [Jen89b, KJH88]. puzzles [Ano97x]. Pyramid [Ano97-44].

Qbytes [Sem96f]. QIC [Sim96k]. QMS [Ano97n]. qualified [Kay96b]. Quality [Ano92-118, Ano93-112, Ano93-152, Hay97k, Hay97l, Hay97m, Hay97p, Hay97n, Hay97s, Hay97r, Hay97o, Hay97q, Hay98b, Rad92a, Bar90b, Dar90, Fol96f, Fra90d, Joh90a, Moa93b, RK89]. Quantify [Ano94-109, Sem94a]. Quattro [McM91a]. Query [Ano96p, App96h]. query- [App96h]. Quest [Hay97k, Hay97l, Hay97m, Hay97p, Hay97n, Hay97s, Hay97r, Hay97o, Hay97q, Hay98b, Ano94-218, Str96b]. question [SF95i]. questioning [Hur96]. Questions [Hen94, Kad97, The95o]. quests [Fol96e]. quick [Ano95-162]. quick-and-dirty [Ano95-162]. quickly [Ano97-42, Sim96h]. Quintillions [Sem96f]. quite [Hay96a]. Quixx [Ano92-119].

R [Ano96v, Ano96l, Ano96w, Ano97n, Lei90d]. R/3 [Ano96v, Ano96l, Ano96w, Ano97n]. R13 [Ano97-38]. race [Sch89d, Sem96h, Sha93c, Ano93-307]. racing [Joh90b]. Rack [Ste95b]. racked [Tra95]. RAD [Ano92-67, Ano95-162, Rob96]. radar [Sim96u]. RAID [Ano92d, Ano92e, Ano94-201, Ano96u, Dar95m, Rub94, Sim96m, Sha94c, Ano93-263, Sim95h]. RAID-5 [Ano96u]. RAID-The [Sim96m]. Rainbow [Ano92w]. Raises [Ano93-263]. Raising [Ano97-45, Ano97-46, Kno97k, Tuc97a]. ranked [Ano97-27]. ranking [Ano93-311]. Ranks [Ano93-321, Ano93-249]. rapidly [Ano94-278, May96]. rare [Rho95b].

Rashomon [Hay97e]. Rate [Ano93-104, Eth89, Sne92g]. rates [Kay96b, Wra98]. Rating [Sta95a]. Ratings [Ano93-229, Sta95b]. rats [Ril97c]. rave [Hay96c, Sim96g]. raves [Hal88]. RDA [Moa89d]. RDBMS [BS94, Ano92-120, Ano93-264]. RE [Ano94-269]. Reach [Dar95b, Jen89a]. Reaches [Ano93-108]. Read [Ano92-104, Ano95-79, Ano96-42, Var96d, Ste92b]. Reader [Ano96m, Str95l]. Readers [Ano95-75]. reading [SF96e]. Ready [Ano92l, Ano93-237, Fis96a, Fol96f, FH94, Kno97l, Ste94a, Sud95, Ano95-127, Ano96i, Asb97a, Car90c, Hab94, Hay96a, McC96d, McC96k, Sem93, Ano93-207]. readying [Pat96]. Real [Ano92u, Ano92-121, Ano93-280, FC96, Foc95, Ale95a, Ano95-74, Ano95-75, Ano95-114, Ano96f, Ano97z, Asb97a, Jon97a, Kno97b, McC96n, Moa95a, Pen97, SF95c, The93c, Ano93-247, Ano94-219]. Real-Time [Ano92-121, Ano96f, Ano94-219]. real-world [SF95c]. realities [MCBH88]. Reality [Ano94-252, Ano95-58, McC95a, SF96i, Ano95-85, Ano96f, Dar96l, McC94e, Sch90c, Ano94-220]. realize [Moa91a]. Really [Ano92-29, Ano92-161, Ano93-29, Ano93-74, Cor95c, Sem94h, Ano95-133, Ano97p, Kno97f, McC96j, Moa93a, Ril95j, Ril97e, Tuc96b, Vog95, Ano95x]. Reap [Mik97g]. Reaping [LN89]. reasons [Ril96e, Sim96n]. Rebuilding [Ano92-73]. recasting [Lin96c, Ano93-265]. Recently [Ano96b, Sha96]. reception [Fra89i]. recipe [Fre97f, Lin91b]. Reclaim [Ano93a]. recognized [Sha96]. recommendations [SF96b, The96e]. record [App97b]. Recover [Fra93e, Ano93-266]. Recovery [Car90d, Rot88b]. Recruit [Bul95, Cla90]. Recruiting [App95c]. Red [Bec94, Cas90a, Tuc97f]. redefines [MK89]. Redmond [Dal98, Dav96c, Dav96a, Dav96d, Dav96b, Dav96e, Dav97b, Dav97c, Dav97f, Dav97d, Dav97g, Dav97e, Dav97a, Dav97h, Dav97i, Dav97k, Dav97j, Fol96f]. Reduce [Ano96u]. reduced [Ano94-28, Ril96g, Sim96r]. Redundant [Rub94]. Reenergize [ZNZA95]. Reengineer [Ano92-122, Sem95d, Ano94-268, ZNZA95]. Reengineering [Ano93-74, Ano93-189, Moa94a, Moa95d, Ova95, Ano95-161, Hay96b, MF96, Moa93a, Ano93-267, Moa94g]. reengineers [Ano94f, Ano93-229]. Reevaluate [Sim96k]. refill [SF96r]. Refiner [Ano96p]. refueling [Ano94-268]. Refuseniks [Gar95c]. regional [Dar95h]. regress [Hay97l]. regular [Vog95]. Regulations [Mac95]. regulatory [Ano97j]. reigns [Ker88b]. Reinventing [Dar95c, Dar95o]. reinvigorate [Sch88a]. Relational [Ano92-65, Ano92-156, Ano94-280, Ano96u, Hod89d, Mar89b, Ano93-268]. relations [Ova96]. Relationship [Dar96a, Gut96, Str97a]. relationships [Lei90d, McC97c]. relatively [Ano93-350, Fis96b]. Relay [Ano92-161, Ano94-72, Ano94-73, McC96d]. release [Ano95-130]. released [Kno96a]. releases [Ano97t, Fol96b]. Reliability [Ano93-314, Dar96t, Fol96f, Ano93-263]. Reliable [Dar96t, Hen95]. ReliaStar [App96i]. relies [Dar96t]. Remain [Ano93-321]. remake [Cas96]. remedy [Lei90b]. Remember [Ano94-215, McC96b]. Remote [Ano92-143, Ano94-245, Ano94-270, Dur94e, Dur94d, Dur94f, Ril96h, SF97g, Vog95, Dar95a, Dar95b, Fis96b, Gar95a, Kay96f, McC96j, Rad93, Ano93-269, Dur94c, Dur94d, Dur95]. renaming [SF96q]. renewed [Sim96q]. rent [Ril96m]. rent-an-app [Ril96m]. repair [Wor90]. Replaces [Sem94f]. Replacing [Ano93p, Rob96]. replenish [SF96r]. Replicate [Sim96q]. replication [Sim96q, Wat96a, Ama95c, Ama95d].

Report [Ano93-287, Moa94g, Dar96n, Ano93-198, Ano93-199, Ano93-200, Ano93c, Ano93-228, Ano94-164, Ano94-165, Ano94-166, Ano94-167, Ano94-168, Ano94-236, Ano94-237]. reporter [Ano95-27]. Reporting [App96h, Dar96n, Fre97c]. reports [App96h, Dav92b, Ano97d]. repository [Car90c, Ste96b]. required [SF96e]. Requirements [Ano92-63, Sim96q]. rescue [McC96o, Ano93-67]. Research [The96e, Mik97d, Sch89b, The96b, Ano93-45]. resemblance [SF95d]. Reservations [Ano92-155, Gut92]. resource [Dar95f, Hir96, Ril95h]. resources [Ano97v, Dar96d, Sta90b, The95a, The95r, The95s]. response [Ano95v, Hab96, Jen91]. response-computer [Hab96]. rest [The93g]. result [McC97c, The96d]. resulted [Ano95-28]. results [Fre97d, Til89]. Resurge [Ano92-28, Pin92]. retail [Bel90, New90]. Retailers [Dav89g]. Rethink [Ano93z, Hay96b]. Retool [The95a, The95r, The95s]. retooling [Dar96o]. retrieval [App92b]. retrieve [Hal96]. returns [Sim96g]. reusability [Moa89c]. Reusable [Joy88b]. Reveals [The95i, Ano93-311]. Revenue [Ano92n, Dar96b, Ano96-28, Ano97-27]. Revenues [Ano93-148, Ano97-27]. review [The95v]. reviews [Ano90b, Hay96c, Sim96g]. Revisiting [de 98b]. Revive [Ano92t, Ano93-51, Ker88a]. Revolution [Cas96, Moa90f]. Revolutionizing [Ano93-142]. Revs [Ano93-151, Fra90h]. rewards [Cox90, Mik97g]. Rich [Dar95q, McC97c]. riddled [Rad96a]. Ride [Ano93-279, Dar96b, Ano94-29]. Right [Ano94-201, App95b, Bau95d, KK95, Ano97g, Ano98f, Car90g, Esk97a, Fol96c, Fre97f, Gib97b, Hec97, Inm97a, Kad97, Roc97b, Sta90a, Til89, Tuc97c, Ano95-133]. Ring [Str95c, Bun91, Dar96c, Siv89c]. RIP [Esk98b]. RISC [Ano92-139, Car88b, Fra90j, Hen96, McW89c, Sim97n, Var96a, Var96b]. RISCy [Ano93-270]. Rise [Ano93-343, Ano96-40, CM89, Coo91, Eth89, Pan90b, Ano92n]. Rises [Ano94-127]. Rising [Bul95, Asb97e]. Risk [Ova95, Rub95c, Dug97d, Fra90j, Jag96a]. risks [Cox90]. rival [Siv89c]. Road [Ano94-184, Dar96i, Gil96a, Lin95c, Str97c, Ano97i, Ell88, Gar95a, Hal88, Kno97h, Mar89b, McC97e, Mik97b, Pic94b, Ano94-275, Gar95b, Sha94b]. road-warrior [Ano97i]. roadblock [GH88]. roadside [Esk97d]. Roaming [Ano92-123]. roars [Sim97b]. Rocket [Ano94e, Dar96b]. Rockwell [Arm90b]. Rohm [Ost90]. ROI [Ano95-72, Con97, Fra97, Kee97a, Kee97b, Kee97d, Kee97c, Kee98a, Kee98b, Moa95h, Rad96a]. Role [App95d, App95e, Ano90a, Ano95-129, App95a, Sta89a]. Roles [Ano92j]. Roll [Ano92-98, Var96h]. Rolling [McC96m]. Rollout [Ano93-312, Cus93]. ROM [Ano94z, Pin91]. ROMs [Ano94-163, Fra93b]. room [Ano95-134, App96f]. Roomier [Ano93-84, Kre93a]. rooms [Ano97e]. roost [Sim96s]. roots [SF97a]. rose [Ano93-311]. Rough [McC95d]. Route [Ano92-142, Gre88]. Router [Ano93-345, Ano94-112]. Routers [Ano93-271, Dar95p]. Routes [Var95c, Ano93-261]. RPC [Siv90a]. RT [Ker88a]. rubber [Kno97h]. Rugged [Lin95c]. Rule [McC95k]. Rules [Ano93-229, Ano96r, McC95g, Sha94a, Esk97h, Sim96s, Sta96, Bau95d]. Run [Ano96n, Sch94, Sem94h, SF96r, Ano95-76, Ano95-134, Dav89a, Jef98, LP96, MK90c]. Running [Gre96, Bea96, Rot88b]. Runway [SF96d]. RUSH [Sem95g, Sem95f, McC96b, McC96o]. Russia [Heb89]. rusting [Dav96d]. Rx [Wal90b]. Ryder [Ano94-268].

S [Ano94-284, Ano95-133, Ano96r, Ano96f, Ano97-48, Car97, Esk97j, FC96, Moa91a, Sem95d, Var96j, Vin95c, Ano93-87, Ano93-326, Ano94-38]. S&L [Hay97t]. S-4 [Ano93-87]. S-40 [Ano93-326]. S-Day [Esk97j]. S/390 [Ano97-48, Var96j, Moa91a]. SAA [Maj89, Moa88b, Moa89j, Sim96g, TB90]. Sabre [Ano97i]. Safe [McC97f, Ril95d, VM96c, Str94h]. Safeguard [Ste95b]. Safety [Moa94j]. Sagister [Ano96r]. said [SF96q]. sake [Mar88]. Salaries [Ano92-124, Ano93-321, Pan90b]. Salary [Ano96-43, Ano88c, Car89g]. sale [Fra97]. Salem [Ano96x]. Sales [Ano95a, Ano97b, App95a, Ano93-311, Asb97g, FT88, Kel90c, Mar92, Sim97i, App95d, App95e]. salesforce [Ano95-69, Var96e]. Same [Dat94, Pan96c, Ril96j, Sch95a, Str95h, The96d]. SAP [Ano96v, Ano96l, Ano96w, Ano97n, Ano97-28, Kay96b, Ano93-272]. SATAN [Ale96b]. satisfaction [Dug97b]. Satisfied [Ano92-134]. Satisfies [Ano93-145, Moa93e]. satisfy [Pan96a]. Save [Ano92-125, Ano96-33, Bla96, Dar95p, Sim95f, Ano97p, Car97, Ham90a, Hir96, Ril96a, Sim97o]. saved [Bla96]. saver [Gib97a]. savers [Ull94, Ano94-271]. saves [Lin91a, Sim96d, War97]. saving [Ril96l]. savings [Kee97d, Ker90c, Ano93-254]. savvy [Ano95-78]. Say [Ano92-163, Gar95c, Sim96r, Ano96o, Dar96h, Hal96, Kay96f, Sar95, SF96h]. says [Ale96c, Dri96, Hay97j, Mar95d, Wat96b]. Scalability [Sch97c]. Scalable [Asb94]. Scale [Ano92-124, Sim96b, Ano96-28, App96h]. Scales [Ano94-221, App96h, Esk97a]. Scaling [The93e, Ano93-273]. Scan [Ano92-125, Ano97u]. scanner [Bla88]. scanning [Bar90b]. Schedule [Hay97s, Ril95f]. scheduling [Ano97i, SF96r]. scheme [Ano95-59]. schemes [Wat96a]. School [Ano96u, Moa94i, Dar95u]. Science [Ano94e, Ano96-48, Hod88]. scope [Gol95]. Score [Pan96c]. Scorecard [Ano92-40, Ano93-62, Ano93-55, Ano93-56, Ano93-57, Ano93-58, Ano93-59, Ano93-60, Ano93-61, Ano94-33, Ano94-34, Ano94-35, Ano94-36, Ano94-37]. Scorecards [Pan96c]. Scott [Ano96y]. Scramble [Ano93-334]. scrambling [SF96h]. scratching [Ano95-27]. screaming [McC96m]. Screens [Ano93s, Sha93a, Sim96u]. seals [The95v]. seamless [Dar96n, Lib88]. Search [Ano93-44, Ano93-313, Hab95c, Ale96b, Fra93f, Mik97e, Roc97b]. Searches [Ano96n]. Seasons [Fra90i]. Second [Fol96g, Kno96b, Fol96i, Moa89i, Str96b]. Second-wave [Kno96b]. secrecy [App97b]. Secret [Ano93-246, Fra93d, Hay97h, Ker89h, SF97g]. Section [Ano96-28, Ano96-41, Ano96-42, Ano97-27, Ano97-34]. section-from [Ano96-28]. Sector [Moa95f, Moa95g]. Secure [Ano93-29, Ber95c, LP96, Str93, Fis96b, Fra90g, Fra92a, McC96b, Val96, Ano93-274, Mel95]. Securicor [Jud90a]. Securing [Ano93-275]. Security [Ale96b, Ano92-54, Ano92-110, Ano93-211, Ano93-146, Ano95v, Foc95, Sch95k, Ale95a, Ano95-73, App93b, Cas95a, Der92, FM89a, Fra92b, Fra93f, Hab89, Hay96a, Ker89h, Ker90b, Lin94a, Lon90, Mad90, McC96b, McC96j, McC97e, Ric93d, Spe96, Ano93-313, Ano94x, Vac94]. See [Ano92-50, Ano98g, Pan96c, Rad96a, SF95d, SF96a]. Seeing [Kno97m, Str95j]. seek [Fra90e]. Seeking [Kay96b, Kor94]. Seeks [Ano93-146]. seers [Ano95-113]. sees [Fra90c, SF95c]. Seize [Str94l]. Sekizawa [Ano92-63]. self [Ano97t]. self-service [Ano97t]. sell [Ano95-85, Ano96g, Dar96k, Inm97b, Sar95, The96b]. Selling [Dug96d, Jef98, Ano93-62, Ano95r, Car89f, Str95g, Var96j, Kay95j]. Sells [The95t, Car90a]. Semantics [Jua90, O’C88]. Send [Roc98b, Asb97b]. sense [Ano95-149, Ano96k, Ano97y, Car90h, VM96a, Var96j]. Sequel [Ano96r]. Sequent [Ano94-221]. Serial [Fis96a]. series [Ano96f]. Serious [Ano92-116, Ano94-68, Ano97i, Ric92d, Var96j, Ano93-154]. seriously [dJ96a]. Server [Ama95c, Ama95d, Ano92-36, Ano92b, Ano92-32, Ano92-51, Ano92-93, Ano93-27, Ano93j, Ano93-308, Ano93-272, Ano93-313, Ano93-289, Ano93-294, Ano93-328, Ano93-330, Ano94-250, Ano94-182, Ano94-200, Ano94d, Ano94-238, Ano94-272, Ano96q, Ano96f, App95b, Fol96a, Gil96b, HS94, Ham94a, Ham94b, Hen94, Mar95c, Nee95, Qui95, Rad94b, Ric94f, Sch95f, Sem94c, Sha94a, Sim95e, Sim95d, Sim95c, Sim96j, Ste96a, Str94f, Str94m, VM96a, Wat96c, Ano95-160, Ano96o, Ano96-52, Ano97q, Ano97v, App95a, Bal96a, Bal96b, Bel90, Car90e, Dar95a, Dar96p, Dar96r, Dri96, Fra90a, Fra93f, Fre98b, Gar95a, Gil96a, Hab95a, Lin96c, Mar95b, MB94, Pan96b, Rad94a, Ric90b, Ric92c, Ric94a, Ric94c, Ril95e, Sco96, Sim95b, Sim96e, Sim96l, Sim96p, Sim97i, Sim97l, Siv89a, Sne94a, Str94b, TMS96, Ano92-42, Ano92-163, Ano93-48, Ano93-49]. Server [Ano93-50, Ano93-51, Ano93-317, Ano93-65, Ano93-153, Ano93-276, Ano93g, Ano94-29, Ano94-30, Ano94-217, App95d, App95e, Bau94a, Cor96, Gar95b, Hab95b, Hab95c, Kay95l, Sem94d, Sem95b, Sem95c, Sim96g, Sne94b, Wat96c]. Servers [Ano96n, Ano96-51, Bow94, Hur94d, Hur95c, Lev95, Ric94g, Sem95e, Ste95b, The95u, Ano95-129, Ano97v, Ano97-48, App96c, CS95, Dar95b, Gil96a, Mun97, SF95f, Sim96t, The95q, Tuc96a, Tuc97n, VM96a, Asb94, Sch95d, Sch95e]. Serves [Ano92-78, Ano93n]. Service [Ano93t, Ano96u, Fol96h, Sch95h, Str94o, Ano96s, Ano96-48, Ano97t, App96g, Asb97b, Asb98a, Hab96, Hir96, Joy89a, McM90g, Siv90d, The96e, Var96d, Ano93-276, The94h, The95t]. serviceability [Fol96f]. Services [Ano93-285, Moa95f, Moa95g, Ano96-32, Dav89f, Esk97i, Ger96b, Hab95a, Hir96, Kay96c, Kay96f, Sim96r, Str96b, TMS96, Ano93-300, Ano94-222, Ano95-147, App96g, Ger96b, Hab95b, Ano94-223, Ano94-224, Ano94-225, Ano94-226, Ano94-227, Ano94-228, Ano94-229, Ano94-230, Ano94-231, Ano94-232, Ano94-233, Ano94-234, Ano94-235, Ano95-148, Ano95-149, Ano95-139, Ano95-140, Ano95-141, Ano95-142, Ano95-143, Ano95-144, Ano95-145, Ano95-146, Scr95a, Scr95b]. Serving [Emm89b]. Sessions [Ano92-128]. set [Fis96b, Ril96a]. setbacks [Kee97c]. sets [Ano95-160]. Setting [Ano95-29]. settled [Fly96b]. setups [Str95f]. Seven [Chi91, McC96m, Inm97b]. seven-course [Inm97b]. Several [Ano96n, Hok96, Ano93-244]. SGI [Var96a]. Shape [Ano96n, Kel89, Sch93, Ano93-309]. Shaping [FP88]. Shared [Ano92-126, Hir96]. shareware [Kau96]. sharing [Sim97g, Tuc96a]. Shearson [Dys90a]. shells [PC89]. Sherlock [Kee98b]. Shield [War97]. Shift [Ano95-113, Ano95-114, SF95f, SF95b, SF95d, SF95c, SF95e, SF95g, SF95i, SF95h, SF95j, SF96a, SF96i, SF96b, SF96n, SF96c, SF96p, SF96d, SF96e, SF96m, SF96f, SF96g, SF96h, SF96o, SF96l, SF96j, SF96k, SF97a, SF97b, SF97c, SF97d, SF97e, SF97h, SF97f, SF97g, SF97j, SF97i, SF97k, SF98, SF95l, Ano95-112, Ano94-186, Ano94-187, Ano94-188, Ano94-189, Ano94-190, Ano94-191, Ano94-192, Ano94-193, Ano94-194, Ano94-195, Ano94-196, Ano94-197, Ano94-198, Ano94-199, Ano95-103, Ano95-104, Ano95-105, Ano95-106, Ano95-107, Ano95-108, Ano95-109, Ano95-110, Ano95-111, SF95a, SF95e]. shifted [Fol96d]. Shifts [Ano88b, Ano93-300]. Shines [Ano93-245]. ship [Pat96]. shipped [Ano96q]. shipping [Sim96u]. ships [Ano96z]. shock [Rec89, Moa94h]. shoes [Fol96c]. shoestring [Dav93a, Ano93w]. Shoot [Ano93-332, Sim97l]. Shoot-Out [Ano93-332]. Shootout [VM96a, The94h]. Shop [Ano92-53, App96c, Fre97d, Kay96c]. Shopper [Ano93-246, Fra93d]. Shopping [Ano96e, BM90, Bry89b, Dar95f, Dav89g, Ano93-277]. shops [Ano97m, Kau96, Sim96i, Sim97c, Str96b, Ano95-28]. Shores [Ano92-82]. Short [Blo89b, Car88c]. Short-circuiting [Blo89b]. shortfall [Rid88]. shot [Ano94-215]. Shotgun [Ano93-278]. Should [Ano92-162, Ano95-76, Fol96i, Kay96f, Moa95b, Ril95g, Sim96e, Str94a, Ale96c, Ano93-350, Hab96, Hay96b, Hin92, Hur96, Mod90, SF95j, SF95l, Sim96h, Sim96i, Tay96, Ano93-279, Kay95k, Ova95]. shout [dJ96a]. Show [Ano93c, Ano93-228, Ano94-27, Fol88b, Sim96u, Ano94-236, Ano94-237]. shows [Ano96-51]. shrink [Ano95-69, Dav97a, OHE94]. shrink-wrapped [Ano95-69, Dav97a, OHE94]. Shrunk [Ano93-109]. Sidestepping [Ano92-127]. Siemens [Ano96m, Ano97l]. Sieve [Ano93-211, App93b]. Sight [Hur94c]. sights [Ano95-160]. Sigma [Ano93-280]. Sign [Vac94]. significance [Dar96r]. Signs [Ano96-45, Car88b, Jef97d]. silicon [Ano95-125, de 97b]. silver [de 97g, de 97i]. SIM [Ano97n]. simple [App96a, McC96h, Str97a, Wil97e]. Simplify [Bla95b]. simply [Rob96]. Simulate [Ale96b]. simulators [Ale96b]. Simultaneous [Ano92-128]. Since [Ano93-104]. single [App96g, Fry96, Mar89a, Ros89, Tra89a, Var96b]. Sinking [Kay95l]. siren [SF96d]. sit [Tra95]. Site [App96h, Dur95, Ano97t, Ano97-35, App96a, Kay96a, Var96g]. sites [Ano95-29, Ano96b, Hay96e, Kay96f, McC96a, McC96h, Sim96g, Var96j]. situation [Sim96k]. six [Kee97b, Kno97j, Ano93-280, Ano93-281, Ano97r, Fol96d]. Size [Ano93y, Dar95n]. Sizzles [Ano94c]. skeptical [Sim95g]. skills [Dug96c, Mik97e, Ano93-105]. skip [Dar96d]. sky [Ale96a, Ril96n]. Skybox [Ano98c]. skyscrapers [Bes90]. slash [O’C91]. Slashed [Ano93-258]. Slashing [Ano96g]. slate [Ano94-28]. slaves [SF96d]. SLC [Ric90d]. sleeves [Var96h]. slides [Hod89a]. slim [Ric94h]. sloppy [SF96j]. Slow [Ano93-342, Dur94a, McC96b, Sim96h, Kre93c, Moa89c]. slowdowns [Ric91]. slower [Ano95-125, Ano95-162]. slower-than-death [Ano95-162]. slowing [Bun91]. Slowly [Fol96e, App92a, Sim95b, Ano92-79]. Slows [Ano96u]. Small [Ano93-121, Ano95-112, Kre93b]. Smalltalk [Has95, Vin95c]. Smart [Ano92-129, Ano93-102, Ano93-147, App92b, Ale96c, Kig96, Sne91b, Ano93-282]. Smarten [Lev95]. Smartens [Ano93-201]. smarter [Pic94d, Ano94-276, The94d, Ano95-126]. smarts [The96a, Ano92-35]. Smartsourcing [Str94o]. SmartSuite [Fol96b]. smiling [McC97a]. Smooth [Ano92-41, Dar96m]. Smoothing [Ano92-130, Ano93-283]. SMP [Ano93-333, Ano97-44, Sim96o, Var95d, Ano94-239, Ano94-238, McC95k, The95u]. SMS [Cor96]. SNA [Ano93-203, Ano94-72, Sim96g, Ano92-73, Ano93p, McC95h]. snake [de 97b]. Snap [ES97, Hal96, OHE94, Ano94-240]. Snap-on [ES97, Ano94-240]. Soar [Ano93-321, Kel90c, Ano93-285]. Soared [Moa95f, Moa95g]. Soars [Ano94-239]. society [Ano95-135]. Soft [Ano93-105, Hab95b, McW89c, Hab95a, Kee97d, Meh89, Mik97e, Pan90b]. SoftLanding [Ano92-131]. Software [Ano92h, Ano92w, Ano92-40, Ano92-68, Ano92-116, Ano92-122, Ano92-132, Ano92-135, Ano92-136, Ano92-133, Ano92-134, Ano92-144, Ano93-64, Ano93-55, Ano93-56, Ano93-57, Ano93-58, Ano93-59, Ano93-60, Ano93-62, Ano93-61, Ano93-317, Ano93-67, Ano93b, Ano93-87, Ano93-208, Ano93-260, Ano93-314, Ano93-274, Ano93-147, Ano93-280, Ano94z, Ano94-33, Ano94-34, Ano94-35, Ano94-36, Ano94-37, Ano96-45, Ano97t, Car90e, Dav90f, Mar90, The95h, Abr91, Ano95-126, Ano95-136, Ano96d, Ano96z, Bes90, Car89a, Car90a, Dam89, Dav97f, Der92, Fis96a, Fry96, Ger90, Gib89, Gra91, Ham90a, Hay96a, Hay96d, Heb89, Hec97, Hen95, Hum89, Joh90a, Joy88b, Joy89a, Joy89c, Mar89c, McM90i, Moa90f, Pal92, Ric92a, Ric92d, Ril95a, Ril96l, Ril96f, Sim96p, Sim96r, Sne92g, Spe96, Sta90d, Str93, The95b, The96c, Tuc97l, Tuc98a, Web89, Ano92-95, Ano93-286, Ano93-284]. Software [Ano93-285, McC94d, VM96a]. software-vending [Ano95-136]. SolarNet [Ano94-181]. Sold [Dav97e]. solid [Ano95-77, Mae96, Wat96c]. Solution [Sha96, Fre97d, Kni90, McC96g, Ste96a, Tra89b, de 97g, Ric94e]. Solutions [Ano93-79, Asb97f, Tuc97l, Ano93-152, Kor94]. Solves [Var96c]. Solving [KJH88, Sim96t]. Some [Ano94-239, Ano95-133, Ano96n, SF96h, SF96r, Dar96k, Dri96, Hur96, Kay96f, McC96d, Rad96a, SF95c, SF96b, SF95l, Sim96g, Sim96i, Sim96n, Sim96p, Sim97f, Sim97n, Spe96, Dat94, Ste96a, Tuc97c, Tuc97c]. something [Str95f]. Sometimes [App96d, Str95i, Dar96r, Ril95i]. Son [Moa89j]. Soon [Ano92-35, Ano95-134, Ano95-136, Hir96, Ril95k, Str95j]. sooner [Ano96t, Fis96a]. sophisticated [Fly96b]. Sorcerer [Ano92-136]. sort [Dar96k, Rad96a]. Sorting [Ano93-292]. Sound [Ano94-280, Ano97y]. soup [Meh88, SF95i, Tuc96a]. Souped [Ano93g]. Source [Ano92-126, Ger96a]. Space [Ano94e]. span [Dar96h]. Spanish [Jua90]. Spanning [Sta89e]. spans [Mar90, Ric90c]. SPARCstation [Ric90d]. Spare [Ano93-236]. Speak [Ano93-248]. speaks [Ano96-31]. special [Ano95-86, Ano93-287]. specialists [Ano94-278]. Specialization [Ano95-97]. Speed [Dar96l, Sch95k, Ano95-30, Ano95-161, Ano96-48, Ano97l, Ano97k, Ano97-49, Bry88, Kay97, Sch97e, Sim96s, Sne93, Ano92-78, Ano93-83]. speeding [Tuc96a]. SPEEDS [Sem95g, Ama95b, Ano93-119, Sem95f, Sim96h]. Spell [Ale95b, Ale95f, Ale95g]. spend [Ano96-34, Jon97a]. Spending [Pan96b, Car90d, Dav89a, Ker89l, McC96l, Pan97, Str97a]. Spin [Ano94z, Mik98b]. Spinning [Ost90]. sport [SF96k]. spot [Ano90b, Joh90c]. spread [Var96b]. spreading [Var96c]. Spreads [Ano92-139, Ano97x]. Spreadsheet [Ano93-233, McM89b, Ano93-62, Man92b, McM91a, Ric92b, Sta93]. spreadsheets [Fra89g]. spree [Dar95f]. sprinkled [The96b]. spur [Pan97]. Spy [Ric93d]. SQL [Wat96c, Ano92-109, Ano93-195, Ano96r, Ano96-50, Bee89a, Bee89b, Cod88a, Cod88b, Dar95l, Wat96c, Ama95c, Ama95d, Ano93-288]. Squandered [Str97a]. Square [Bal96b]. Squawks [Ano92-137]. Squeeze [Ano92-27, Asb97f]. Squeezing [Hay97a]. Staff [Ano97b, dJ96b, Moa94f, Mik97i, Roc98b]. Staffing [Roc97e, App96d, App94]. stages [HC90]. stagger [Mar92]. Stalking [Ano97-47, Fol96h]. stamp [Dav89e]. stand [App97b, Dav89h, Dri96, FGR88]. stand-alone [Dav89h, Dri96]. Standard [Ano93-231, Ano93-350, Cas90b, Fra89b, Gil91, Moa90i, Moa94b, Ril96g, Sim96n, Tuc96a, Ano94-241]. standardize [Ano97z]. Standards [Ano92-121, Ano93-241, Moa94b, Fly96b, Fra90e, Fra93b, Mar89e, Moa90e, Moa90g]. Star [Ano96r, Ric90b, Bul95]. StarGROUP [Ric90b]. stars [Esk97c]. Start [Ano93-121, Ano93-155, Ano96-52, Cel96, McC96h, She95, Ell88, Kre93b, OHE94, Sim96u]. Started [Ano93-115, Ano95r, Jag96a]. starters [Kay96a, Var96g]. Starting [Ano94-172, Esk97d, Esk97e, Esk97j, Esk97f, Esk97g, Esk97k, Esk97h, Esk97i, Esk97l, Esk97m, Esk98a, Esk98b, Fol96i, Hay96b, Hay97b, Sem96h, Sem96i, Fol96e, Sim96u]. startup [Ano89c]. state [Ril96e]. stations [Ano96k]. statistics [Ano95-98]. Status [Tra95]. Stay [Ste94b, Ano97-43, App96i, SF95h, Sem96b]. Staying [Dav90g]. Stays [Bau96b]. Steak [Ano94c]. steam [Dav97k, Hay96a]. Steel [Gad89]. steep [Fre97d, Sim97f]. Steering [RR89]. Stellar [Ale95d, Ale95e]. Step [Ano94-272, Ano96p, Ano96i, Dar96o, Sem96e, Asb94]. stepping [Str95k]. Steps [Ano93-290, McC96m, Sem95e, Ano96j, Car97, Chi91, Fos90a, Gos88, Sim97h, Tra90b, Dur94f]. stick [Hur96, Moa94b]. Stickier [Ano93-242]. Still [Ano92-166, Ano93-85, Ano93-238, Hal96, Kay95c, Kay95d, McC95d, McC95a, SF96o, Wil97f, Ano95-27, Ano95-148, Fol96c, McC96m, Rad96b, Sim96p, Sim97h, Hay96a]. stock [Ano95-27, Sim96k]. Stop [Ano92-137, Moa89g, Ham94d]. Stopgap [Ano93-350]. stops [Hab97, Roc97f]. Stor [Ano97-42]. Storage [Ano92-87, Ano92-102, Ano92-125, Ano93-79, Ano93-108, Ano93-154, Ano93-152, Ano93g, Ano93-331, Ano93-352, Ano94-172, App96f, Kor94, Sim96s, Str94g, Ano95-74, Dav92c, Kel90b, Mik97a, Sim97e, Sim97j, Sim97m, Ano93-76, Ano93-79, Ano93-245, Ano93-289, Rya94, Ano95-128]. Store [Pau95a, Pau95b, Ano96-51, Sed90b]. Stored [Ano92-30, Par92]. stores [Ham96]. Stories [Ano93-340]. Story [Ano97h, Jam97a, Sch95b, App96i, GFR89, Var96j, Sim96f]. straight [RR89]. Strange [Tuc98b, Tay96]. Strassmann [Ano92-52, Str97a]. StrataCom [Ano96n]. Strategic [Gar98, Ano90a, Asb97d, Dar96o, Ger96a, GH88, Jam97c, LS89, Sta90b]. strategic-retooling [Dar96o]. strategically [Ano95-136, Dug97c]. strategies [Run89a, SF96p, Siv90c]. Strategy [Ano92-138, Var96c, Moa90d, SF96p, Siv89d, Dar95j, KK95, OHE94]. Stratus [Ano92-139]. Strauss [Ano93-189]. Streamlined [Vac94]. streets [Ano95w, Ril95f]. Strength [Ano93-89]. strengthens [Ano97r]. Stress [Ano92-140]. Stretch [Ano93-289]. stretched [Ste96a]. Strike [Dar95g, McC97c, Ets98]. Striking [WN90]. strings [Kel90d]. Strong [Ano92-169, Sim96k]. Stronger [Ano92-135, Ano92f, Ano93-209, Ano93-242]. Structure [Ano92-66]. Struggle [Ano93-315]. students [ZSZ88]. studies [Sch88a]. Study [Ano96e, Sta89g]. stuff [SF95c]. stupid [Ano95-96, Ano96y]. subcontractor [Gut96]. subsystems [Sim96m]. succeeded [Var96d]. succeeding [Asb97g]. Success [Ale95f, Ale95g, Ano93-340, Cas95d, Cas95e, Jon97c, Jon97b, Jon98, Ale95b, Arm90a, Fos90b, Gil96a, GFR89, Lei90e, Mik97j]. successes [Kee97c]. successful [Chi91, Ano94-242]. successor [Hod89d, Roc97c]. succumb [Ano95-131]. such [Ano95-78, Ano96f, Lin96c]. suffer [Ril97b]. Suffolk [Sch88a]. suit [Ril95g]. Suite [Fol96b, The95v, Ano97t, SF97j]. Suites [Ano94y, Fol96a, Fol96b, OH95, Ano96o]. suiteware [The95d]. summary [Ano96a]. Summer [Ano96u]. Sun [Ano93-343, App91, SF95d, Ano96r, Ano96-51, Eld95, McC96f, Ric90d, Sem96d, SF96q]. SunSelect [Ano94-181]. Super [Tuc97l]. Supercomputer [Ano92-61, Sch89d]. supercomputers [BM90, Her90, Ano92-141]. superfloppy [Coo91]. Superglue [McC96k]. Superhighway [McC96p]. Superior [Ano93-277]. SuperLAN [Ano93-120]. superpowers [Bar90a]. Superserver [Ano93-332, Ano93-277]. supplement [Ano95-86]. Suppliers [Ano90c, Ano95-86, FM97, McM90g, Siv90e]. supplies [Ano98g]. Supply [Fre97g, Asb97f, Wil97e]. supply-chain [Asb97f]. Support [Ano92-127, Ano92-165, Ano93-334, Ano96u, App96c, Dar96c, Hab96, Kay96b, KS94, Sna90b, The96d, Ano94-278, Ano95-131, Dav97e, Dug97e, Fly96b, Fol96f, FM89b, Hab94, Hir96, Joy89a, MF96, McC96m, The96a, Var96d, Wat96c, WN90, Woo91, Ano94-246, Rad95, The94f]. supported [Ano93-350]. Supporting [Dug97e, Dar96d, Ger96b]. Sure [Ano95-98, Ano95-148, Fol96f, McC95e, Ril95h, Ril95i, Ham96, Vog96a, SF95f]. Surf [Ano96f, Ano95-127]. Surfing [SF95h]. Surge [App95d, App95e, App95a, Hay96e, Sim96q]. surges [Hod89b]. Surprise [Ano94-243, Sta89a, Kay95l, Lev95]. survey [Ano88c, Ano96m, Car89g, Dav89a, Ker89l]. surveys [Ano95-98]. Survival [Ano94-185, Sch95b, Sta90d, Str94c, Dav90a, Ric94c]. survive [App97a, Ruh88]. susceptible [Ano97q]. sustain [Str96c]. sweet [Tuc98a]. swift [Ano95-128]. Switch [Dur94b, McC97b, Sim96g, Str95j, Sim95i]. Switched [Ano94-252]. Switches [Dar96i, Dar96j, Ano96-49, Dar95r]. Switching [Dar96t, Sch97b, Ano94-129]. Sybase [Ano93-65, Ano92-142, Ano93-290]. symbiosis [Inm97a]. Synch [Kol94]. Synchronize [Ano93-291]. Syncsort [Ano93-292]. Synergy [Mar89d]. SynOptics [Ano93-293]. System [Ano92-143, Ano93-76, Ano93-79, Ano93-244, Dar96c, Fra90h, Lin94b, Moa94d, The95e, Ale96c, Ano94-181, Ano95-82, Ano96l, Ano97z, Ano97x, Ano97-48, App96i, Car97, Dar96a, FR89, Fos90b, Gil95, Jef97e, Joy89b, Kay96c, Rea95, Ric91, Ril96a, Rob96, SF96r, Sim96j, Siv90e, Ano92-144, Ano94-244]. Systematizing [Jua90]. Systems [Ale94, Ano92s, Ano92-57, Ano92-70, Ano92-94, Ano93-28, Ano93-117, Ano93-124, Ano93-156, Ano93-152, Ano93-326, Ano96f, AR96, Cor96, GH88, Key89, McC95k, Moa88b, Nee95, Sch89c, Sed90b, Tra89b, Ano90a, Ano96b, Ano96-28, App96c, Bel90, Bus88, Car90b, Car90e, Car90g, Cor95b, Dar95c, Dar96a, Dar96b, Dav89d, Dav93b, Eth90b, Fol96f, Fol96g, Ger96b, Hay96a, Jag96a, Jag96b, Jen89b, Jen91, KJH88, Ker89l, Kno96a, LS89, Lib88, Lin92b, Lin96a, MF96, Mar89e, McC96o, McC90e, McC90f, McW89c, Moa88a, Moa90b, Moa90e, New90, Ost90, Pan96a, PC89, Ras89, Ric92c, Run89c, Ryp88, SF96r, Sha96, Sim96m, Sim97g, Ste96a, Tuc97b, Var96b, Ano92-92, Ano92-111, Ano92-131, Ano93-296, Ano93-323, Ano97r, Bec94, McC96o, SF96r, Ano95-69].

T [Ano93-46, Ano95v, Ric90b]. T1 [SF95e, Siv90g, Tay90]. tab [Dav89a]. Tackle [Ano93-210, Ano93-243, Ano97m, McC93f, Ano93-294]. Tackles [The95w, Lin96c]. tackling [Sim97f]. tactic [Ric94c, Ano94-185]. Tactical [Ano96-36]. Tag [Ano92-108, Liv92]. tailorable [Fol96e]. Take [Ano93-81, Ano93-271, Ano93-296, Ano94-32, Ano95-162, App95f, Hab95b, Jag96b, Kor93, Ova95, Ril96i, Sch89e, Sha94c, Ano95-57, Ano96b, FM97, Hab95a, dJ96a, Pic94a, Rad92b, Sim96d, Sim96k, Tab92a, The96a, Tuc97j, Var96h, Ano93-295, Ano94-245, Ano94-246, Rub95c]. Taken [Ano94-29, Hay96d]. Takes [Ano92j, Ano92-142, Ano93-53, Ano93-309, Ano93-308, Ano93-348, Ano93-349, Ano94-275, Ano96n, App96f, Dai95, Gil96a, Ano98c, Ger96a, MF90, McC93g, Pin91, Ric93a, Sch93, Sch97e, Sta95c]. Taking [Ano94-243, Moa94a, Sta89f, Fra90j, McC96h, SF96b, Sim97n]. talent [Roc97a]. Talk [Ano97z, Ano97y, Ano97w, Ano97x, Ano98c, Ano98b, Sch95i, App97b, Kno97a]. Talking [Str94e, Ano95s, Ano95-132]. talks [Cor96, Ova96]. Tame [McC96a, Moa95d]. Tandem [Ano97v]. Tangent [Ano92-145]. tap [Rad96c, Dar95q]. Tape [Sim96k, Cas90a, Sim97m]. Tardy [Hin93]. targets [Jon97b, LR89]. task [Dar96r]. tasks [Vog96c]. tassel [Fol96f]. tassel-topped [Fol96f]. taste [de 97c]. Tate [McM90a]. taverns [Ano94f]. Tax [Dar96b]. TCO [Ets98]. TCP [Ale96b, Ano92-166, Ano93-203, TMS96, Ano93-297]. TCP/IP [Ale96b, Ano93-203, TMS96, Ano93-297]. Team [Ano92-153, Ano94-242, Vin95c, Ano95-133, Hay97k, Kno97b, Ray95]. Teams [Bau95e]. teamwork [Sta95c]. tech [Ano96-44]. techies [Car89d]. technical [Lin96a]. techniques [AR96]. technological [MBGS88, Jua90]. Technologies [Fol96d, Ano96t, App96i, Cas96, Lin96c, May96, McC96d, McC96k, Run89b, Sem96b, Sem96e, Sem96h, Sim96e, The96b, Was90]. Technology [Ano92-72, Ano93-82, Bau95c, Cox90, Gol95, Lic94, Ric94d, Sem96e, Ano96-52, Car88a, Dar96o, Dar96s, Dit92, Esk97e, Fis89, Hab94, Kel90c, Lin96b, MS95, MCBH88, Ril96e, Sem96a, Sim96m, Wat96a, Wat96b, Wil90, Ano93-298, Ano93-299, Ano94-247, Ano94-248, Ano94-249]. tee [Ano95-102]. Telecom [Ano96-44, Ker89i, Mac95, Str95b]. Telecommunications [Lei90c, Ker89j]. telecommuters [Ano96s, Jef97a]. telephone [Hab96, Jen91, McC93e]. Telephony [Hab96]. Tell [Ano95-138, Jag96a, SF95g, SF96r, VM96a]. tempted [AR96]. Ten [Car97, Fos90a]. Tensions [Hur95a]. terabyte [Fre97a]. Terabytes [Ano92-146]. term [Fol96g]. terminal [Lei90c]. Terminals [Ano92-129, Ano92-171, Ano96-32, Dav97k, McM90e, McM90c, SF95f]. Terminate [Str95c]. terms [Ano97z]. Territory [Ano92-171]. Test [Mar95b, Con97, Hay97i, Hay97k, Hay97r, Jag96b, Maj89, Mar89e, Siv89b, Siv89d, Ano93-262, Jag96b, Mar95c, Qui95]. test-bed [Siv89d]. Testing [Ano93-288, Bro89, Dar95r, Hay96e, Hay97o, The95i, Ano96d, Con97, Hay97q, Sim96r, de 97c]. testosterone [Hay97i]. Tests [Ano92-140, Moa94f]. Texas [App96d]. text [Tra90a]. Thanks [Ano96f, Sim96m, May96, Rad96c, McC97c]. that’ve [Str96b]. Their [Ano93-239, Moa95c, Ano95-76, Ano95s, Ano95-30, Ano96b, Ano96s, Ano97e, Ano97-27, Dar96c, Dar96s, Fol96e, Fol96f, Fol96h, Fra91c, FM97, Hay96e, Jon97a, Kau96, Lin96c, McC96g, McC96h, McC97c, SF95g, SF96g, Sha96]. Them [Ano92-146, Ano94-163, App96e, AR96, Dar96c, Dar96k, FM97, Gib97b, Jag96a, Jam97b, McC96g, Sco96, Sim96d, Sim96e, Ano94-112]. themselves [Sim96i]. Therapy [Moa94h]. There [Ano93-238, Ano95-77, Ano96b, Ril96e, Sim96b, Sim96s, Wat96a, Ano95-78, Ano95y, Dar96c, Dar96d, Dar96s, Kay96b, MF96, OHE94, Rad94a, Ril96b, SF96l, VM96a, Var96c, Rad94b]. thereby [Ril96g]. these [Gre96, Ano93-316]. they’ll [Ril96a]. They’re [Sim97f, Fol96f, SF95g, Sim96l, The96d]. they’ve [McC96h]. thin [de 97h]. Thing [vSS96, Ano95-78, Inm97a, SF95h, Sim97f, The96d, Rho95b]. Things [Liv95, Fol96f, OHE94, Pan96c, Ril95e, Ril96b, Vog96a, vSS96]. Think [Ano93-121, Ano95-78, Ano95-79, Dar96f, Sco96, Tay96, The95m, Ano96t, App96c, Fis96a, Fol96f, Kay96e, Kre93b, Mae96, SF95e, Vog96b]. Thinking [Hay97i, Hay97h, Hay97f, Hay97j, Hay97g, Hay97d, Hay97e, Hay97c, Hay98a, App96c, Kay96e]. THINKWRAP [Dar96q, dJ96b, Pan96c, Vog96c, Ano96-52, dJ96a, Lin96b, McC96n, McC96q, Rob96, Sta96, Vog96b, Ale95c, Ano94-253, Ano94-254, Ano94-255, Ano94-256, Ano94-257, Ano94-258, Ano94-259, Ano94-260, Ano94-261, Ano94-262, Ano94-263, Ano94-264, Ano94-265, Ano94-266, Ano95-159, Ano95-160, Ano95-161, Ano95-162, Ano95-150, Ano95-151, Ano95-152, Ano95-153, Ano95-154, Ano95-155, Ano95-156, Ano95-157, Ano95-158, Gol95, Hen95, Rea95, Rho95a, Rho95b, Sar95, Sta95c, Tra95, Vog95]. third [Fol88b, FS89, Hok96, Sim96s, Ano94-267]. third-parties [Hok96]. third-party [FS89]. Thoroughly [Asb98b, Dar96a]. those [Ano96l, Jag96a]. Thought [Str94h, Rho95b]. thousand [Jag96b, Ril96e]. Thousands [Ano93-54, App96c, Tuc97c]. Threat [Foc95, Ale95a, Sta96, Str95f]. threaten [Dav96e]. Three [Ano94-268, Kel89, Ste96a, App96g, Dar96c, Hay97q, Hen96, Sim96t, Sim97h, Vog96a, Var95c]. three-ring [Dar96c]. Three-Tier [Ste96a]. three-tiered [Ste96a]. Throwing [Ham94d]. thrown [Mik97f, Str95j]. ticketing [SF96r]. tide [Ano93-311, Asb97e, Fol96d, Ker89j]. Tier [Ste96a, Dar96h]. tiered [Ste96a]. tight [Ano93-189, Ano93-318]. tighten [McC96j]. Till [Mar95a]. Time [Ano92-49, Ano92-121, Ano93-144, Bal96a, Dar95k, Dav97h, McC96g, Moa95h, Sim96k, Sim96t, Ano89c, Ano94-278, Ano96-34, Ano96f, App96a, Dar96s, Ger96a, Moa89e, OHE94, Ril95j, Ril96a, SF95c, Sha93b, Var96c, Var96h, Wat96b, Ano92-153, Ano94-219, Str95c, Sch95a]. Timeplex [Rot91]. Times [Ano94-71, Ano94-28, Car88b, Kim90b, Win90]. Tinker [OHE94]. Tinkering [Mik97i]. tip [Ano96-35]. tips [Ano95z]. Tivoli [McC96o]. Tl [Ker89k]. Today [Ano93-270, Dar96a, LP96, Sim96m].

Together [Ano94-106, McC96d, FM97, Inm97a, The96d, Ano93-155, Moa95e, Sch95j]. Token [Str95c, Arn90, Bun91, Siv89c]. tolerant [Sim96m]. Tolerate [Sim95d]. toll [Hay96d]. Tomorrow [Ric94d, Car90f]. Too [Ano93-63, Ril96i, Ano94-215, Fol96b, Gib97b, Inm97a, McC96e, McC93a, Moa88a, Ril95k, SF95d, Sim96e, Sta89b, Ste96a, Vog96b, Asb97b]. Tool [Ano92v, Bau96a, SF95k, SF96q, Ano97y, App96e, Dar96f, Dar96h, Dar96o, Lin96c, McC96m, Sch90c, SF96r, The92b]. ToolBook [McM91b]. Tooling [Dav97i]. toolkit [Dar96m, SF96b]. Toolkits [SF96b, OHE94]. Tools [Ale96c, Ano92-100, Ano93y, Ano93-51, Ano93-155, Ano93-207, Ano93-238, Ano93-310, Ano93-231, Ano93-230, Ano93-240, Ano93-338, Ano94-278, Ano94-286, Ano96-29, Ano96-27, Ano96-53, Bau95d, Dri96, Fol96b, McM90i, Rad96b, The94g, Ano96f, App96h, Con97, Dar96g, Fre97c, Fre98b, Hay97r, Kno96a, Kno96b, McC96j, McC96l, McC97c, OHE94, Rob96, SF96g, Sem93, SF96r, The93b, Tuc96b, Tuc97j, Var96h, Ano92-109, Ano93-151, Ano93-319, Ano93-320, Ano94-270, SF96r]. tools-and [Ano96f]. Top [Ano96-45, Ano96-54, Gre97, McC88, Sne92g, Sta95b, Sta95a, Ale96c, Ano89a, Ano90b, Ano93-62, Ano95-32, Ano95-86, Ano97-27, Dar96d, Dav90f, Dav90g, Fol96f, Ell88, Ano93-321]. top-selling [Ano93-62]. Topology [Ano96r]. topped [Fol96f]. Tops [Ano93-267]. Toshiba [Ano95-47]. Toss [Moa94j]. total [Ano96-28, Gib98]. Totally [Ano92-67, Lin95c, Ric94f, Sco96]. touch [Ano97-43, Dar95b]. Tough [Ano94a, Str95b, The95o, Dar96h, Rad96a, Sim97f, Hen94]. Toughest [BS94]. Town [Ano97z, Ano97y, Ano97w, Ano97x, Ano98c, Ano98b, Hur96]. Toy [SF96q, Sem96d]. Toys [Ano93-310, OHE94, The93b]. TP1s [Can88]. TQM [Bec94]. Track [Ano92-61, Bau96b, Ano94-273, Dav91, Fol96c]. Tracking [Car91, McC89b, Tab92b]. tracks [Ano97t, Wat97]. trade [Ano88b, McM90i]. trading [Ano98c, Cla90, Pur89, Sta90c]. trading-card [Ano98c]. traditional [Hal96, Ova96, VM96a]. Traffic [Dar96j, McC96a, Dar96l, McC96d]. Trail [Ano93-115]. Train [Gib97b, Gib97b]. Training [Ano93-113, Boo90, KK95, Moa94i, Ano94-278, Ano97-41, Gib97b, Moa89h, Moa93c, Moa95a, Ray95]. transaction [Can88, Joh90b, Pan96a, Sch95l]. transactions [Kno97l, LP96, Tuc97k]. transfer [Ano96s]. transition [Boo90]. transitional [Car88b]. translating [Ger96b]. translation [Ano95u]. transmission [Wal90b]. Transportable [Lin95c]. Traps [Ano92-127]. trash [Dar96k]. Traveling [Ano92-154, Fol88b]. travelling [Ano97-30]. Trenches [Ano93-71, Moa94g]. trend [Fry96, Hir96]. trends [Moa90b, Ano93-77, Ano93-78, Ano93-298, Ano93-299, Ano94-247, Ano94-248, Ano94-249]. Trial [McC95b]. trials [Ano95x, Ano96x]. tribes [Ano97y]. Tricks [Dur95]. tricky [Dar96m]. tries [Sim96e]. Trilogy [Dar96f]. trims [Ker88c]. Trio [Hok96]. Triple [Ano92x, Rot97]. Triple-DES [Rot97]. trivial [Lin96a]. trouble [SF96r, SF95h]. trouble-ticketing [SF96r]. truck [Hab97]. True [Moa94d, SF97i, Fre97d]. Truly [Ano96-45]. Trusted [Str95c]. Truth [Var94a, Hay97r, Str95i, VM96a, The95i]. Try [Ano95x, Mar95d, OHE94, Rot97]. trying [Hab96, Inm97b]. tug [Fra89d]. Tumble [Ano92n]. Tune [Fra93g, Ano93-322]. tuned [Kno96b, Tuc97n]. Tunes [Ano94-169]. Tuning [Dur94f]. Tunneling [Fis96b]. turks [Ano93-311]. Turmoil [Dav93b, Kay96a, Ano93-323]. Turn [Fre97h, Kee97d, Kee97c, The93f, The96e, Ull94, Ano95-80, Gil96a, Kau96, Ros96, Ano93-324, Ano94-271]. turn-of-the-century [Ano95-80]. turned [App96h, Hen96, Sim96j]. turnover [Var96f]. turns [Dar96r]. tutorial [Mee96]. TV [Ano96i, Ano97y, Tra90a]. Tweaking [Ano92-170]. twice [App96c, Sco96]. Twin [Ano92-155, Gut92]. twists [Dar96r]. Two [Ale95f, Ale95g, Ano93h, Lin93, Str96a, Ale95b, Ano95-57, Emm89b, Mee96, Sem96c, Sta96, Ano93-325, Dar95s]. Two-Pentium [Ano93h]. two-wheelers [Ano95-57]. two-year [Sem96c]. type [Ano96-54]. types [Hal96, Inm97a]. typical [Ano94-278, Gol95]. tyranny [Hay97s].

U. [Ano96d]. U.S. [Ano97l, Dav89a, Hen95, Jon97a]. U.S.-made [Hen95]. Ultimate [Ano93k, Ano94-184, Pic94b, Ril96c]. ultra [Ano96-48]. UltraSPARC [Ano96-51]. unauthorized [Mar89c]. uncertain [Win90, Ano94-272]. Unchained [Sem95b, Sem95c]. unclog [McC95c, Dar95t]. underground [Ano96-43]. Underneath [Has95]. understudy [Fra88]. unearthing [McC97c]. unearths [Fre97b]. Unfortunately [Ril96e]. Unicenter [Ano96-37, McC96h]. Unicycles [Ano95-57]. Unified [Fol96h]. UniForum [Ano93-326, Ano93-326]. Unify [Bau95e, Sch95k]. Unifying [Ano92-156]. Union [Ano94-273, Bau96b]. Unisy [Ano93-327]. Unisys [Ano94f, Car89e, Moa88a, Ano96n]. united [Ham90b, App96g]. unites [Sim97g]. Uniting [Ano92-157]. UniTree [Ano94-274]. unity [Bar90a, Tat89c]. Universal [Hal96, Fis96a]. University [Ano95-78]. Unix [Ano93-275, Car88b, Fra89d, Moa89k, Ale96b, Ale96c, Ano92-158, Ano93-47, Ano93-29, Ano93-45, Ano93-66, Ano93-89, Ano93-117, Ano93-196, Ano93-258, Ano93-270, Ano93-263, Ano93-314, Ano93-292, Ano93-326, Ano93-340, Ano94-107, Ano94-241, Ano94-285, Ano96-51, Ano97-48, Bal96b, Dav96b, Fra90e, Fra90i, Gil91, Gil96a, Gil96b, Hur94d, Jud90b, Lei90a, Lin95d, Lon90, Mel95, SF95b, Sem96f, Sha94a, Sim95c, Sim96f, Sim96i, Sim96t, Sim97h, Siv90e, Spe96, Tay96, Tuc98b, Var96a, Var96b, Wil94, Ano92-56, Ano92e, Ano92f, Ano93-86, Ano93-146, Ano93-328, Ano93-329, Ano93-331, Ano93-330, Ano93-332, Ano94-275, Ano95-129, McC95k, Ric94g]. UNIX-based [Gil96a]. UNIX-to-IBM [Ano93-47]. UNIX/NT [Ale96c]. UnixWare [Ano92-158]. Unleashed [Ano92-89]. unless [Ano95-134]. Unlocks [Ano92-97]. Unmask [Tuc97m]. unmasked [Ano97a]. unmovable [App96d]. Unplug [Ano93-346]. Unraveling [Hay96e]. Unruh [Ano93-327]. unscathed [Ano95-79]. Unsnarl [Dar96j]. Untangle [Sim97m]. until [Fly96b, Sim96u]. Unveils [Ano96p, Ano95-47]. upcoming [Sim96g]. Update [Hur96]. updates [Ano97t]. Upgrade [Ano92-46, Ric94e, McC90e, Ril96l]. Upgrades [Ril96j]. upmanship [Ano96h]. upper [Ano95-132, Kee97a]. upper-level [Kee97a]. ups [Ric92b, Sed90a, Ano97q]. upscale [Gil96b, Moa89k]. Upsizing [Var95d]. UPSs [Pic94d, Ano94-276]. upsurge [McC96a]. urgency [Tat89c]. US$1 [Ano94-71]. US$100 [CS95]. US$256B [Kel90c]. US$290 [Mar92]. US$50 [Dav89a]. USAA [Lei90e]. Usability [Ano93-262]. usage [HC90]. USAir [The95q, Mar95d]. USD [Jon97a, Kno97b]. Use [Ano92-59, Ano96u, App96i, Fis96b, Fly96b, McC96p, McC97d, TMS96, Ano94-278, Bar90c, Hay96e, McC96k, Ril96e, SF96e, Sem96e, SF96r, Str96b, Var96b, Wat96a, Ano93-242]. used [Ano96t, Ano97z, Dar96a, Gil96a, dJ96b, Ril97e, Sem96b]. useful [McC96m, Ros96, Spe96]. User [Ano92-44, Ano92-63, Ano92-109, Ano92-127, Ano92-159, Bla95b, Ker89l, Kol94, Ano88a, Ano96f, Car90a, Fol96b, Kay96e, Ker90c, Key89, Moa88c, Rot88a, Rot89, Tat89b, Ano92c, Dug96a, Gol95]. Users [Ano92-86, Ano92-160, Ano93z, Ano93-261, Ano94-271, Dug96a, Dug96b, Dug96c, Dug96d, Dug97a, Dug97b, Dug97c, Dug97d, Dug97e, McC90e, Mus88, Sch88b, Sne92a, Ano93-350, Ano95-96, Ano96s, App91, App96g, App96h, Asb97a, Dar96f, Dar96q, Eth90b, Fol96d, Fol96e, FGR88, Hal96, Hur96, Joy89a, Kno96a, McC96e, McC96j, McC96m, McM89a, Moa88c, Moa89a, Moa91b, Pan96a, Rad96c, Ril95i, Ril96h, Sim96g, Sim97n, Sta89d, The96d, Tuc97c, Ull94, Ano93-333, Sta95b, Sta95a, The94g, Ril95j]. Uses [Ano96r, Sco96, Ano96d, Ano97j, Ano97k, Ano97n, Dav92c, Ril96l]. Using [Ker90b, Mee96, Ale96b, AR96, Der92]. usual [Ano95w]. usually [The96b]. Utilities [Ano92-101, Ano92-170, Con92]. utility [CM89, McM90i].

V.42 [Emm89a]. V2K [de 98b]. valuable [McC97c, Rho95b, Str97d]. Value [Ano94-111, Ale92, Ano95-74, Asb97a, Gut96, Hay97k, Hen90, McC94b, Sno98, Wat98]. values [Ano95-132]. vast [Ano96b]. VAX [Lin92c, Sch90b]. VB4 [SF95k]. VDT [Sch88a]. VDTs [Meh88]. veer [Ano96-54]. Vega [Hen95]. vending [Ano95-136]. vendor [Bla96, Dav89d, Ger95, Hug90a, Inm97b]. Vendors [Ano92-120, Ano93-65, Gil96b, O’C91, Ano93-350, Ano94-268, Ano96a, Ano96-51, Ano97-27, Dar96c, Dav90f, Fly96b, Fol96d, Fol96e, Kay96c, Lin96c, Ril96f, SF96r, Sim97o, The96b, Ano93-334, Str94o, Ano95-32]. Version [Ano92-150, McC96h, Sim96j]. versus [Dav96b, Fra89d, Kno97j]. vertical [Fol96e, McC94e]. very [App96h, Lin96b, Sim96b]. via [Ano96f, Fis96b]. victims [Ano95s]. victory [Car89d]. video [Hal96, Maj90, McC97a]. Videoconferencing [Ano94-67, McC95d, Str94e, Ano94-282, Fra93g, Hin92, Ano92-162, Ano93-322]. View [App96e, Sem94g, McC90a, McM90b]. viewers [Dar96n]. viewing [Fly96b]. Views [Ano97-43]. VINES [Dar95u]. Viral [Ano96p]. Virtual [Ano93-320, Ano94-252, Ano94-281, Dur94b, Ker90c, Lei90c, LP96, McC95a, Pen97, Sem94f, Ano95-85, Ano96f, Ber95c, Fis96b, Ano93-335]. Virtues [Ano95-148, Ano95-149, Scr95b, Ano94-222, Ano94-223, Ano94-224, Ano94-225, Ano94-226, Ano94-227, Ano94-228, Ano94-229, Ano94-230, Ano94-231, Ano94-232, Ano94-233, Ano94-234, Ano94-235, Ano95-139, Ano95-140, Ano95-141, Ano95-142, Ano95-143, Ano95-144, Ano95-145, Ano95-146, Ano95-147, Scr95a]. Virus [Coh95, McC96i, Ale95c, Ano95w, McA89, Ril96a, Ano93-116]. Viruses [Ano96-30, Sch91]. Vision [Ano92-80, Ano95-58, Fra90b, Fra90d, Moa91a, SF95i, Ano93-90, Ano93-91, Ano93-92, Ano93-93, Ano93-94, Ano93-95, Ano93-96, Ano93-97, Ano93-98, Ano93-99, Ano93-100, Ano94-97, Ano94-98, Ano94-99, Ano94-100, Ano94-101, Ano94-102, Ano94-103, Ano94-104, Ano94-105, Ano96r]. visionary [Kni90]. visit [Ano95u]. visitors [Ano97t, SF95f]. Visual [Ano94-106, SF95c, Ano93-336, Ano93-337, Ano93-338, Vin95c]. VisualAge [Hay96c, Bau94d]. VisualWorks [Has95]. Vital [Jef97d, Car97]. Viz [Ano92v]. VM [Tat89b]. Voice [McC96d, Ano96s, Dav89e, Jen91]. volume [Ano95z, App96a, Dar95d, Dav92a, Fol89, Sim96u]. volumes [Pan96a]. Voodoo [Esk97l]. VPNs [Fis96b]. VRML [SF96i]. vs [Ano95-137, Sim95j, Sim96t, Sim96v]. VSE [Car89d]. wagon [Ano95-70]. Wait [The95c]. Waiting [Dav97j, Wil97f]. Walker [Ano92-98]. Walking [de 97h]. Wall [Ril96b, SF97c]. Wallet [Hur94b]. Wallington [Kno97b]. WAN [Ano96i, Cas96, Dar95h, Dar95p, McC95c, SF96m, SM94, MS94]. wanna [Mik97h]. WANs [Ano93-347, Ano94e, Ano96p, Dav89c, Hur94a, Ano94-277]. Want [Ano92-160, Ano93-72, Dar96s, Ril95j, Rot97, Sch94, Ano93-73, Ano95-75, Ano95-57, Ano95-114, App96c, Dar96f, Fis96b, Kno97f, McC96m, Ril95i, Ril96f, Ril97d, SF95b, Sne92a, Ste96b, Ano94-278, Mae95, Ril97f]. Wanted [Dar96t, Ano94-279]. wants [Ano97-35]. war [Fra89d, Hay96d, Kno97m]. Ware [Ano95-75, McC96k]. Warehouse [Ano96-27, App96i, CM95, Dar96m, Ano97j, App96e, App96h, Fre97a, Gre96, Kig96, Sem96e, SF96r, Sim96a, Str95i, Str96c, Ano94-30, App95b, App96e, Rad95]. warehouses [App96f, Asb97a, Inm97a]. Warehousing [Gre96, Hac95, Str96c, Dar96m, Kay96a, Kno96e, Mar95d, Var96g, Wat96b, Wat96c]. Warp [Kay95f, Sim96j, Var95b]. warrior [Ano97i, Pic94b, Ano94-184]. wars [Str97d]. wary [Fra89i, SF95j]. Was [Str94h, Ano94-215, Ano95-138, Ano97w, App96i, Dar96a, Hur96, Ril95f, The96a]. Washington [Sch90a]. waste [Hin93, Var96k]. wasting [Ril96a]. Watch [Ano94-215, Ano95-125, Ano95-127, Ano95-128, Ano95-134, Ano95-126, Ano95-131, Ano95-130, Ano95-132, Ano95-133, Ano95-135, Ano95-137, Ano95-136, Ano95-129, Ano96-48, Hen95, Ano93-110, Ano94-202, Ano94-203, Ano94-204, Ano94-205, Ano94-206, Ano94-207, Ano94-208, Ano94-209, Ano94-210, Ano94-211, Ano94-212, Ano94-213, Ano94-214, Ano95-115, Ano95-116, Ano95-117, Ano95-118, Ano95-119, Ano95-120, Ano95-121, Ano95-122, Ano95-123, Ano95-124]. waters [Fra89g, Moa89f]. watershed [McC96o]. wave [Ano95-127, Ker90a, Kno96b, Moa89i]. Wax [Pan90a]. Way [Ano92-109, Ano93-302, Ano93-79, Ano93-153, Ano93-191, Ano93-257, Ano93-335, Ano94-251, Ano96p, Dur94e, McC95k, Vin95b, Ano95-76, Ano96f, Bry89a, Cas90b, Dar96f, Dav90b, Dav90c, Esk97c, FC96, Fra91a, Hay97o, Jef97b, Kig96, McC96k, McC92, McC93b, Rad96b, Rea95, SF95f, Sco96, Sem96e, Sha93d, Vog96c, Moa95c, Ano95-96]. way-developers [Ano96f]. Ways [Ano93-281, Ano93-347, Ano95y]. weak [Moa88b, Tab92b]. Weakness [Var95b]. Weapon [Sem96e, Ger96a]. weapons [Str97d]. wearing [Ril95g]. Weaving [Lib88]. Web [Hay97h, Sem95h, Ano95-29, Ano95-135, Ano96b, Ano96h, Ano96k, Ano96n, Ano96r, Ano96f, Ano96-50, Ano97n, Ano97t, Ano97u, Ano97v, Ano97y, Ano97-35, Ano97-32, Ano97-36, Ano97-39, Ano97-44, Ano97-41, App96a, Cas95d, Dev97, Esk97b, Fly96b, Hab97, Hay96e, Hur96, Lin96c, McC96h, McC97d, McC97e, Mun97, Rad96a, SF97i, Sem95e, SF96q, Sim96p, Str97d, TMS96, Tuc97b, Tuc97i, Tuc97k, Tuc97n, VM96a, Var96e, Var96h, Var96i, Var96j, Var96d, Wat96c, Wil97c, Wil98, Cas95e]. Web-based [Ano97-41]. Web-enabled [Var96h]. Web-to-database [Tuc97i]. Webmeister [She95]. Wedding [Ano94-280]. Welcome [Esk97m, Ano94-281, Bau94d, Str94m]. Well [Kau96, Rad96a, SF96h, Str95j, Ano96-54, App96h, Inm97a, SF96k, Sim96b, Var96f]. Wells [Ano97-48]. went [App96g, Dar96a]. We’re [Ano93-327, Str95k, dJ96b, Sar95, Ano94-215, App96e, Sha96, The96a]. West [Mae95]. Whatever [Ril96k]. wheelers [Ano95-57]. Where [Ano92-164, Ano94c, Ano94-128, Ano94-284, Ano94-286, Ano96-34, Car88c, FR89, Kno97h, McC90f, Nee95, Sch89c, Sha92, Str96a, Ano94-215, Ano95-57, Ano96b, App96f, Str95h, Ano94-285, Bul95, Kay95f, Sim97n]. whether [Hur96, Rad96c]. Which [Ale96c, Ano94-73, Moa90h, Str97d, Jag96a, SF95g, The96b, BS94]. While [Fol96c, Gil96a, Pan96b, SF96k, Lin96c]. white [Siv90i]. white-collar [Siv90i]. Whither [SF97k]. whiz [Ano96f]. Who [App97b, Der92, Fra89j, Rub95a, Sch89d, Sem96f, Sha96, Sim96u, Sim97n, Sta88, VM96a, Ano94-278, Ano98c, Hen96, Kel90d, McC96j, McC96m, Ril96d, SF95d, Ste96b, Ano92-165, Mar95a]. Whole [OH95, Ano96f, App96f, Ril96f, SF96q]. Wholesale [Cas95d, Ano97-32, Cas95e]. Wide [Ano95-29, Ano96b, Hur96, Sem95h, Ano93b, Roc97a]. Wider [Ano92-50, Ano92-103]. wild [McC89b]. Will [Ano92-98, Ano92-121, Ano92-165, Ano94-283, Fra90k, Hay96d, Kre93c, Maj90, Sim97o, Str97a, Tuc96b, Var96k, Ano93-350, Ano95t, Ano95w, Ano95-57, Ano96s, App97b, Asb97b, Fly96a, Fry96, Gre96, Hay96c, Inm97a, Jag96a, dJ96a, Jon97a, Kay96a, Kno96a, Mee96, OHE94, Rad96c, SF96i, SF96k, Sco96, Sim96k, TB90, Var96g, Var96i, Ano93-342, Ano93-343, Jag96b]. win [Fre97e, Kno97i, Moa89l]. Win32s [Ano93-344]. WIN95 [McC95j, Ano95w, Ano95z, Ril95f, Sim96v, Str95n, McC95f, Sim95j]. Window [Dav92d]. Windows [Ano92-29, Ano92-98, Ano93-349, Ano94-238, Ano95-131, Ano95-129, Ril95l, SF95b, Var94a, Var95b, McM90b, Ano92-50, Ano92-167, Ano92-168, Ano92-169, Ano92-170, Ano93l, Ano93-345, Ano94-219, Ano94-286, Bal96b, Dav96e, Dav97k, Dri96, Hok96, McM91b, Moa90i, Rad96b, SF96k, Sem94h, Sim96b, Sim97n, The92a, TMS96, Var96b]. Windows-95 [SF96k]. Windows-AS [Ano93-345]. Windows-AS/400 [Ano93-345]. Windows-Tweaking [Ano92-170]. Windows95 [Ber95a, Kay95g, Ano95-81, Ano95-137, Bea96, Kay95k, Sch95b, Sim95a, Sim96n, Str95e, Vog95, Ber95b, Gar95c, Sch95c, Sim95k]. Windows’NT [Ano95-137]. winner [Fly96a, Sim96g, The96a, VM96a]. Winners [Ano94-66]. Winning [Ano93-297, Ano94-170, Ale95c, Ric94b, Sch89d]. Wins [VM96a, Fos90b]. Wireless [Ano94-220, Ano94a, Rub95c, Str94n, Ano93-346, Ano93-347]. Wires [Ano94-162]. Wisconsin [App96g]. wishes [Jon98]. within [Ale95a, Ano93-350, Foc95]. Without [Ano93r, Ano93-70, Ano93-255, Ano94-38, Ano94-69, Ano94-277, Lin94c, Ano95-30, Cox90, Dar96m, Hur94a, Kay96a, Lei90d, McC96l, SF96c, SF97e, Var96g]. Wiz [Ano97w]. wizardry [Str95g]. WizDOM [Ano92-91]. woes [Dar97b]. Wolcott [Ano92-43]. wondered [Ano97-37]. won’t [Ano95-136, Vog96a]. woos [Tat89b]. word [CF91, Var96c]. WordPerfect [Fol96b]. words [Kno97j]. Work [Ano93-74, Dri96, Ano94-278, Ano95-133, Ano97-39, Dar96f, Dar96k, Jon97c, Kir89, Moa93a, Ric93b, Sha96, Sim96l, Tuc96b, Wil90, Ano93-194, Ano93-196, Lee95]. Workbenches [Ano93-52, Ano93-258]. worker [MF96]. workers [Ano97-43]. Workflow [Ano93-192, McC93g, Sch95l, Dar96g, Ano93-234, Ano93-348, Ano94-287, The94e, The94i, The95w]. workgroup [Dar96r, Ano93-68]. Working [Ano93-155, Ano97-50, Owe89, Sem95e, Van90, Ano96-48, Hay96e, McC96k, McC96m, Mee96, Wat96b]. Works [Ano94-216, Moa93c, Pic94c, Vog96a, Ano93-113, Ano93-339, WP94]. Workstation [Ano93-118, Ano93-307, Sim96n, Sim96v, The95g, Cor95b, Hea90a, Sha93c, Ano93b, Sim95j]. Workstations [Ale95d, Ale95e, Ano89c, Moa89l, Fra90g, Hea90b, Hen96, Lei90a, Ano93h]. World [Ano92-133, Ano96b, Ano97e, Mar92, Sem95h, Ste90, Ano93-311, Ano95-129, Ano95-138, Ano97-27, MF96, McC96b, Moa89g, SF95f, SF95c, SF95j, Sim96n, Ano95-29, Ano96b, Hur96]. worlds [Sta89e]. worries [Dar96m]. Worth [Hur96, Ano88c, Ano97-40, Car89g, Fre97d]. would [Ano95-113, App96c, App96e, Sta96]. Wrap [Win95, OHE94]. wrapped [Ano95-69, Dav97a, OHE94]. Wrapper [AR96]. wrapping [AR96]. wring [Fol96d]. Write [Dar96q, Dug96b, McC96k, Sco96, Str94n]. writer [Ano95-71]. writers [Ale95c]. Wrong [Ano94-67, Ano94-284, Ano94-282, McC96q, McM90j]. WWW [Ano97e].

X [Ano92-171, Ano93-231, Dav97c, McM90e, Ano92-171, Ano93-349]. X-Plore [Ano92-171]. X-terminals [McM90e]. Xbase [Ano93-350, Ano93-350]. Xerox [Ano92-172, Kno97b]. XL [Ano92-173]. XL/ [Ano92-173]. XPM [Ano93-351, Ano93-351]. Xpress [Ano92-46]. XTerminals [Val96].

Y2K [Ano97r, Ano97q, FM97]. Year [Ale95d, Ale95e, Ano92a, Ano93-148, Ano94-66, Ano95-71, Ano95-79, Ano95-138, Cel96, Jag96a, dJ96b, Var96c, Ano93-311, Mae96, OHE94, Pat96, SF96a, Sem96c, Sim96d, Ano96-35, Ano97m, Ano97u, App97b, DB97, FM97, Hay97t, dJ96a, Jag96b, Jon97a, Lin95a, War97, de 97d, de 97e, de 97f, de 97a, de 97g, de 97b, de 97c, de 97h, de 97i, de 98a, de 98b, Ano93-311]. years [Ano95t, Ano95-128, Pan96c, Sta96, Str97b, Str95e, Str96a]. yen [Als88]. yep [Pan96c]. Yes [LP96]. Yesterday [Ric94d]. yield [Sch88b]. Yields [Ano93-69, Ker90c]. You’d [Mae96, Dar96f, McC96d]. You’ll [App94, Dar96s, Fry96, Gut96, MF96, Pan96c, SF96r]. young [Ano93-311]. You’re [Kay96b, App96c, Gut96, Kay96e, McC96d, Mik97b, SF95e, SF96b, Sim96c, Tay96, Val96]. yours [Dav96a]. Yourself [Ano94-163, Ano96u, Spe96, SF96h, Tay96, Wor90]. You’ve [Str96c, FM97]. Yup [LP96].

Zenith [Sne91c]. zip [Ano97e]. zone [Ets98, de 98a].

References

Abramson:1991:BYS

[Abr91] R. Abramson. Bulletproof your software contracts. Datamation, 37(24):61, December 1991. CO-DEN DTMNAT. ISSN 0011-6963.

Alexander:1992:VDG

[Ale92] J. J. Alexander, Jr. The value of desktop gear. Datamation, 38(11):117–118, 120, May 1992. CODEN DTMNAT. ISSN 0011-6963.

Alexander:1994:MGP

[Ale94] Michael Alexander. Mine for gold with parallel systems. Datamation, 40(22):65–??, November 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Alexander:1995:RST

[Ale95a] M. Alexander. The real security threat: the enemy within. Datamation, 41(13):30–31, 33, July 1995. CODEN DTMNAT. ISSN 0011-6963.

Alexander:1995:SPSc

[Ale95b] M. Alexander. Spell PC success with two P’s-pentium and power. Datamation, 41(10):62, June 1995. CODEN DTMNAT. ISSN 0011-6963.

Alexander:1995:THV

[Ale95c] M. Alexander. ThinkWrap — Hackers and virus writers are winning the battle. Datamation, 41(19):106–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Alexander:1995:ASYa

[Ale95d] Michael Alexander. Another stellar year for workstations. Datamation, 41(10):61–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Alexander:1995:ASYb

[Ale95e] Michael Alexander. Another stellar year for workstations. Datamation, 41(10):61–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Alexander:1995:SPSa

[Ale95f] Michael Alexander. Spell PC Success with Two P’s — Pentium and Power. Datamation, 41(10): 62–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Alexander:1995:SPSb

[Ale95g] Michael Alexander. Spell PC Success with Two P’s — Pentium And Power. Datamation, 41(10): 62–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Alexander:1996:FNB

[Ale96a] M. Alexander. Framework — no blue sky, no outlandish promises — just the hard facts that you need. Datamation, 42(2):5–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Alexander:1996:SSU

[Ale96b] Michael Alexander. Simulate security on UNIX networks — SA-TAN has popularized the notion of using attack simulators to automatically probe TCP/IP networks in search of. Datamation, 42(5):69–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Alexander:1996:TCU

[Ale96c] Michael Alexander. Tools to cross the UNIX/NT divide — the smart money says UNIX and NT are going to be the top contenders in the enterprise operating-system arena. which should you choose? why not both, at least for now? Datamation, 42(3):54–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Alster:1988:JMC

[Als88]  N. Alster.  The Japan 10

market close-up: yen drives growth. Datamation, 34(17): 48/1–10, September 1988. CO-DEN DTMNAT. ISSN 0011-6963.

Amaru:1995:DBA

[Ama95a] Chris Amaru. Digital Boosts AlphaServer Performance. Datamation, 41(9):69–??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Amaru:1995:FSD

[Ama95b] Chris Amaru. FSIOP Speeds AS/400 Data Access. Datamation, 41(13):62–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Amaru:1995:SSBa

[Ama95c] Chris Amaru. SQL Server Bundles Replication. Datamation, 41 (11):61–??, June 15, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Amaru:1995:SSBb

[Ama95d] Chris Amaru. SQL Server Bundles Replication. Datamation, 41 (11):61–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1988:DIU

[Ano88a] Anonymous. The DATAMA-TION international user group directory. Datamation, 34(25): 25–48, December 1988. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1988:SGT

[Ano88b] Anonymous. Shifts in global IS trade. Datamation, 34(19): 44/1–10, October 1988. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1988:WYW

[Ano88c] Anonymous. What are you worth in ’88? (salary survey). Datamation, 34(19):53–66, October 1988. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1989:TPD

[Ano89a] Anonymous. At the top of the IS peak (DP industry). Datamation, 35(12):35–148, June 1989. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1989:IMD

[Ano89b] Anonymous. Inside the markets DP industry. Datamation, 35 (12):149–162, June 1989. CO-DEN DTMNAT. ISSN 0011-

Anonymous:1989:WNP

[Ano89c] Anonymous. Workstations and networks are prime-time hits startup companies. Datamation, 35(9):24–28, 30, May 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1990:DSS

[Ano90a] Anonymous. Desktop systems in the 1990s. the strategic role of the client. Datamation, 36(18): 41–56, September 1990. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1990:GLT

[Ano90b] Anonymous. A GRiD lock on the top spot (product reviews). Datamation, 36(22):113, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1990:SM

[Ano90c] Anonymous. Suppliers of email. Datamation, 36(22):108– 111, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:PYC

[Ano92a] Anonymous. 1992 Products Of The Year Contest. Datamation, 38(21):33–??, October 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:CSMa

[Ano92b] Anonymous. A Client / Server 4GL With Mix And Match GUIs. Datamation, 38(20):49–??, October 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:DEU

[Ano92c] Anonymous. A DBMS For Every User. Datamation, 38(16):84–89, August 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:REA

[Ano92d] Anonymous. A RAID For Every App. Datamation, 38(18):111– ??, September 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:RU

[Ano92e] Anonymous. A RAID on UNIX.

Datamation, 38(19):106–??, September 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:SCU

[Ano92f] Anonymous. A Stronger CASE For UNIX. Datamation, 38(17): 71–??, August 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:AM

[Ano92g] Anonymous. An accounting masterpiece? Datamation, 38(21): 115–??, October 15, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:ASN

[Ano92h] Anonymous. Amalgamated Software of North America. Datamation, 38(??):S–25–S–??, November 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:AC

[Ano92i]  Anonymous.  Andrew Corp.

Datamation, 38(??):S–5–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:ADT

[Ano92j] Anonymous. Apple’s DAL Takes On New AS/ 400 Roles. Datamation, 38(21):72–??, October 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:NN

[Ano92k] Anonymous. Are notebooks noteworthy? Datamation, 38 (23):110–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:YRC

[Ano92l] Anonymous. Are You Ready For Cutting Edge COBOL? Datamation, 38(21):77–??, October 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:AP

[Ano92m] Anonymous. The Asian 25 profiles. Datamation, 38(18):89–??, September 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:AIP

[Ano92n] Anonymous. Asia’s IT Profits Tumble, Despite Revenue Rise. Datamation, 38(18):80–??, September 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:AO

[Ano92o] Anonymous. AS/ 400 Application Opportunities. Datamation, 38(??):S–18–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:PD

[Ano92p] Anonymous. AS/ 400 Product Directory. Datamation, 38(??):

S–37–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:APP

[Ano92q] Anonymous. Austin pushes 486 price point. Datamation, 38(??): 15–??, November 15, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:ADD

[Ano92r] Anonymous. Avoid distributed disasters. Datamation, 38(22): 99–??, November 01, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:BSI

[Ano92s] Anonymous. Bancom Systems, Inc. Datamation, 38(??):S–24–S– ??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:BDR

[Ano92t] Anonymous. Bandwidth Demands Revive I-Muxes. Datamation, 38(17):67–??, August 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:BGR

[Ano92u] Anonymous. Benchmarks get real. Datamation, 38(25):61–??, December 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:BVL

[Ano92v] Anonymous. Biz Viz: The latest analytical tool for business. Datamation, 38(17):55–??, August 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:BRS

[Ano92w] Anonymous. Blue Rainbow Software International Corp. Datamation, 38(??):S–8–S–??, November 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:BTC

[Ano92x] Anonymous. Borland’s Triple Crown: Paradox, DBase, Inter-Base. Datamation, 38(16):45–??, August 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:BCD

[Ano92y] Anonymous. Business computer design. Datamation, 38(??):S– 10–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:BDC

[Ano92z] Anonymous. Buttoned-up data collection. Datamation, 38(??): 31–??, November 15, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:CCG

[Ano92-27] Anonymous. Can Compression Get You Out Of a Squeeze? Datamation, 38(20):65–??, October 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:CIR

[Ano92-28] Anonymous. Can IBM Resurge In PCs? Datamation, 38(22):28– ??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:CWR

[Ano92-29] Anonymous. Can Windows Really Make The AS/ 400 Friendly?

Datamation, 38(23):85–??, November 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:CFSa

[Ano92-30] Anonymous. Channeling the flood of stored corporate data. Datamation, 38(19):93–??, September 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:CYA

[Ano92-31] Anonymous. Chart your app’s performance. Datamation, 38 (19):105–??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:CSMb

[Ano92-32] Anonymous. Client/server management arrives. Datamation, 38 (23):28–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:Ca

[Ano92-33] Anonymous. CODA. Datamation, 38(??):S–4–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:CPP

[Ano92-34] Anonymous. Color Printers Paint a New Path. Datamation, 38(22):65–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:CSG

[Ano92-35] Anonymous. Coming Soon:

GUIs With Smarts. Datamation, 38(24):71–??, December 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:CFSb

[Ano92-36] Anonymous. Compaq flexes its server muscle. Datamation, 38 (25):81–??, December 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:CNP

[Ano92-37] Anonymous. Compaq nets a printer. Datamation, 38(??):6– ??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:CMM

[Ano92-38] Anonymous. CONTROL More Manufacturing. Datamation, 38 (25):82–??, December 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:CPD

[Ano92-39] Anonymous. Controls For PC Development. Datamation, 38 (23):108–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:CSS

[Ano92-40] Anonymous. The corporate software scorecard. Datamation, 38 (??):27–??, July 15, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:CSO

[Ano92-41] Anonymous. Cultivating the smooth operator. Datamation, 38(19):98–??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DBB

[Ano92-42] Anonymous. D and B’s Big Play In Client/Server. Datamation, 38 (25):28–??, December 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DW

[Ano92-43] Anonymous. Daly and Wolcott. Datamation, 38(??):S–11– S–??, November 01, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DIU

[Ano92-44] Anonymous. Datamation’s 1993 International User Group Calendar. Datamation, 38(25):74–??, December 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DD

[Ano92-45] Anonymous. Decision Data.

Datamation, 38(??):S–18–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DUX

[Ano92-46] Anonymous. DEC’s Upgrade Xpress. Datamation, 38(??):13– ??, July 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:DLL

[Ano92-47] Anonymous. Dell’s lighter, leaner notebook. Datamation, 38(??):8– ??, July 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:DPG

[Ano92-48] Anonymous. Desktop publishing goes on line. Datamation, 38 (25):45–??, December 15, 1992. CODEN DTMNAT. ISSN 0011-

Anonymous:1992:DAT

[Ano92-49] Anonymous. Develop applications on time, every time. Datamation, 38(22):85–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DSW

[Ano92-50] Anonymous. Developers see wider Windows apps. Datamation, 38(??):24–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DKS

[Ano92-51] Anonymous. A different kind of server. Datamation, 38(16):82– ??, August 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DSP

[Ano92-52] Anonymous. DoD’s Strassmann: The Politics Of Downsizing. Datamation, 38(21):106– ??, October 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DYS

[Ano92-53] Anonymous. Does Your IS Shop Measure Up? Datamation, 38 (18):26–??, September 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DDS

[Ano92-54] Anonymous. Double data security. Datamation, 38(??):21–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DYDb

[Ano92-55] Anonymous. Double your drives with data compression. Datamation, 38(24):49–??, December 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DYDa

[Ano92-56] Anonymous. Downsize Your Database With UNIX. Datamation, 38(21):65–??, October 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DMS

[Ano92-57] Anonymous. Downsizing and midrange systems. Datamation, 38(??):S–4–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:DAM

[Ano92-58] Anonymous. Downsizing: The application migration. Datamation, 38(23):36–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:EUG

[Ano92-59] Anonymous. Ease of Use and The GUIs. Datamation, 38(??): S–25–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:ED

[Ano92-60] Anonymous. Easing AS/400 development. Datamation, 38(24): 97–??, December 01, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:FTS

[Ano92-61] Anonymous. Fast track for supercomputer data. Datamation, 38(24):87–??, December 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:FPC

[Ano92-62] Anonymous. Five-Platform COBOL. Datamation, 38(??): 17–??, September 15, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:FSD

[Ano92-63] Anonymous. Fujitsu’s Sekizawa: Dealing with changing user requirements. Datamation, 38(18): 87–??, September 01, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:GEF

[Ano92-64] Anonymous. Gateway to an enterprise future. Datamation, 38 (21):120–??, October 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:GDG

[Ano92-65] Anonymous. Geographic databases go relational. Datamation, 38(21):116–??, October 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:GGD

[Ano92-66] Anonymous. Get a Grip On DB2 Data Structure Changes. Datamation, 38(20):83–??, October 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:GTR

[Ano92-67] Anonymous. Go Totally RAD And Build Apps Faster. Datamation, 38(19):79–??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:GS

[Ano92-68] Anonymous. Group 1 software. Datamation, 38(??):S–46– S–??, November 01, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:GDI

[Ano92-69] Anonymous. Gupta’s double impact. Datamation, 38(16):28–??, August 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:HS

[Ano92-70] Anonymous. Help/systems.

Datamation, 38(??):S–35–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:HET

[Ano92-71] Anonymous. Here’s An Ethernet That’s Always a 10. Datamation, 38(20):73–??, October 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:HAL

[Ano92-72] Anonymous. How AMEX Leverages Technology Assets. Datamation, 38(24):95–??, December 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:HIR

[Ano92-73] Anonymous. How IBM Is Rebuilding SNA. Datamation, 38 (20):28–??, October 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:HMM

[Ano92-74] Anonymous. How to make millions with database marketing.

Datamation, 38(16):77–??, August 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:HCP

[Ano92-75] Anonymous. HP’s Comeback PCs. Datamation, 38(17):49–??, August 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:ICa

[Ano92-76] Anonymous. IBM Corp. Datamation, 38(??):S–28–S–??, November 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:ILF

[Ano92-77] Anonymous. IBM Legitimizes FDDI? Datamation, 38(17):87– ??, August 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:ISS

[Ano92-78] Anonymous. IBM Serves Up Speed. Datamation, 38(16):81– ??, August 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:IBC

[Ano92-79] Anonymous. IBM’s Big Changes Going Slowly. Datamation, 38 (25):39–??, December 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:IHV

[Ano92-80] Anonymous. IBM’s Hancock: a Vision Of Networks In The 21st Century. Datamation, 38(20):34– ??, October 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:ICb

[Ano92-81] Anonymous. IBS Conversions. Datamation, 38(??):S–38–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:IS

[Ano92-82] Anonymous. IM Shores Up The AS/ 400. Datamation, 38(20):91– ??, October 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:Cb

[Ano92-83] Anonymous. It’s in the cards.

Datamation, 38(??):23–??, September 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:JDE

[Ano92-84] Anonymous. J. D. Edwards.

Datamation, 38(??):S–21–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:JI

[Ano92-85] Anonymous. JBA International.

Datamation, 38(??):S–14–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:LGM

[Ano92-86] Anonymous. The Legent/Goal merger: What it means to users. Datamation, 38(22):91–??, November 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:LCM

[Ano92-87] Anonymous. Low-Cost Mainframe Storage For LANs. Datamation, 38(19):73–??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:MAM

[Ano92-88] Anonymous. Mainframe accounting moves down. Datamation, 38(22):112–??, November 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:MDP

[Ano92-89] Anonymous. Mainframe DBMS Power Unleashed! Datamation, 38(21):26–??, October 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:MBM

[Ano92-90] Anonymous. Making the best mainframe deal. Datamation, 38 (18):73–??, September 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:MW

[Ano92-91] Anonymous. Management Wiz-DOM. Datamation, 38(??):15– ??, July 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:MS

[Ano92-92] Anonymous. Managing AS/ 400 Systems. Datamation, 38(??): S–21–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:MMMb

[Ano92-93] Anonymous. Managing the move from midrange to client/server. Datamation, 38(21):87–??, October 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:MOS

[Ano92-94] Anonymous. Manufacturing open systems. Datamation, 38 (20):105–??, October 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:MYC

[Ano92-95] Anonymous. Mark Your Calendars With LAN Software. Datamation, 38(23):77–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:MMMc

[Ano92-96] Anonymous. Mastering E-Mail Management. Datamation, 38 (25):51–??, December 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:MUD

[Ano92-97] Anonymous. Microsoft unlocks database power. Datamation, 38 (23):107–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:MWH

[Ano92-98] Anonymous. Microsoft’s Walker On How Microsoft Will Roll Out Windows NT. Datamation, 38 (23):80–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:MMMa

[Ano92-99] Anonymous. Multimedia moves to the motherboard. Datamation, 38(20):57–??, October 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:MTD

[Ano92-100] Anonymous. Multimedia tools for development pros. Datamation, 38(17):90–??, August 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:NUM

[Ano92-101] Anonymous. New Utilities Make LAN Management Easier. Datamation, 38(23):69–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:NBB

[Ano92-102] Anonymous. Novell bids to buoy storage management. Datamation, 38(19):63–??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:NCW

[Ano92-103] Anonymous. Novell Casts a Wider Net. Datamation, 38(24): 28–??, December 01, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:NYP

[Ano92-104] Anonymous. Now Your PC Can “Read” Your Fax. Datamation, 38(22):79–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:OAG

[Ano92-105] Anonymous. Old Apps Get a New Image. Datamation, 38(24): 100–??, December 01, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:OEP

[Ano92-106] Anonymous. Oracle’s Ellison: The promise of parallelism. Datamation, 38(16):73–??, August 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:OPP

[Ano92-107] Anonymous. Oracle’s Parallel Punch For OLTP. Datamation, 38(16):67–??, August 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:OLB

[Ano92-108] Anonymous. Outsourcing: Look beyond the price tag. Datamation, 38(23):93–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:PWE

[Ano92-109] Anonymous. Paving The Way For End-User SQL Tools. Datamation, 38(23):53–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:PSGa

[Ano92-110] Anonymous. PC Security Grows Up. Datamation, 38(22):61–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:PS

[Ano92-111] Anonymous. Perle Systems.

Datamation, 38(??):S–16–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:PA

[Ano92-112] Anonymous. The Platt agenda.

Datamation, 38(23):100–??, November 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:PPG

[Ano92-113] Anonymous. PostScript Printing Gets a Boost. Datamation, 38 (19):89–??, September 15, 1992.

CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:PC

[Ano92-114] Anonymous. Premenos Corp. Datamation, 38(??):S–40–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:PAA

[Ano92-115] Anonymous. Print (almost) anything anywhere. Datamation, 38 (19):83–??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:PSGb

[Ano92-116] Anonymous. Publishing software gets serious. Datamation, 38(24): 85–??, December 01, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:PGF

[Ano92-117] Anonymous. Putting a Good Face on DBMS. Datamation, 38 (??):17–??, July 15, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:QJ

[Ano92-118] Anonymous. Quality is job #1. Datamation, 38(20):98–??, October 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:QC

[Ano92-119] Anonymous. Quixx Corp. Datamation, 38(??):S–48–S–??, November 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:RVF

[Ano92-120] Anonymous. RDBMS Vendors Face An Object Future. Datamation, 38(22):40–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:RTW

[Ano92-121] Anonymous. Real-time will get real standards. Datamation, 38 (21):97–??, October 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:RBS

[Ano92-122] Anonymous. Reengineer to boost software productivity. Datamation, 38(25):57–??, December 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:RL

[Ano92-123] Anonymous. Roaming The LANscape. Datamation, 38(24):98– ??, December 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SSf

[Ano92-124] Anonymous. Salaries scale up 5.5% in 1992. Datamation, 38 (17):38–??, August 15, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SSD

[Ano92-125] Anonymous. Save storage dollars with helical scan. Datamation, 38(17):83–??, August 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SSC

[Ano92-126] Anonymous. A shared source for color output. Datamation, 38 (??):19–??, July 15, 1992. CO-DEN DTMNAT. ISSN 0011-

Anonymous:1992:SUS

[Ano92-127] Anonymous. Sidestepping user support traps. Datamation, 38 (22):102–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SSe

[Ano92-128] Anonymous. Simultaneous sessions. Datamation, 38(??):13–??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SLD

[Ano92-129] Anonymous. A smart look for dumb terminals. Datamation, 38 (18):45–??, September 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SJ

[Ano92-130] Anonymous. Smoothing the jaggies. Datamation, 38(18):112–??, September 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SSb

[Ano92-131] Anonymous. SoftLanding Systems. Datamation, 38(??):S–37– S–??, November 01, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:S

[Ano92-132] Anonymous. Software 2000. Datamation, 38(??):S–45–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SDD

[Ano92-133] Anonymous. Software development in a downsized world. Datamation, 38(??):S–12–S–??, November 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:SKC

[Ano92-134] Anonymous. Software to keep the customer satisfied. Datamation, 38(22):109–??, November 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:SBM

[Ano92-135] Anonymous. Software’s big 50:

The magic is stronger. Datamation, 38(19):40–??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SBS

[Ano92-136] Anonymous. Software’s big 50: The Sorcerer’s Apprentices. Datamation, 38(19):50–??, September 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:SSM

[Ano92-137] Anonymous. Stop squawks with modem management. Datamation, 38(22):71–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SBN

[Ano92-138] Anonymous. The strategy behind the new Control Data. Datamation, 38(17):79–??, August 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:SSR

[Ano92-139] Anonymous. Stratus Spreads RISC Bet. Datamation, 38(19): 28–??, September 15, 1992. CO-DEN DTMNAT. ISSN 0011-

Anonymous:1992:STG

[Ano92-140] Anonymous. Stress Tests For GUI Programs. Datamation, 38 (18):37–??, September 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SKD

[Ano92-141] Anonymous. Supercomputers Knock At IS Doors. Datamation, 38(24):79–??, December 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:STN

[Ano92-142] Anonymous. Sybase Takes New Route with NLM. Datamation, 38(16):51–??, August 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SMR

[Ano92-143] Anonymous. System management by remote control. Datamation, 38(25):84–??, December 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SSA

[Ano92-144] Anonymous. System Software Associates. Datamation, 38(??): S–12–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:TG

[Ano92-145] Anonymous. Tangent Group. Datamation, 38(??):S–26–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:TDH

[Ano92-146] Anonymous. Terabytes of data — how to get at them? Datamation, 38(16):38–??, August 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:AK

[Ano92-147] Anonymous. The AcerPAC-kage.

Datamation, 38(17):88–??, August 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:OE

[Ano92-148] Anonymous. The AS/ 400 In An Open Environment. Datamation, 38(??):S–26–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:BD

[Ano92-149] Anonymous. The Best of DATA-MATION. Datamation, 38(17): 113–??, August 15, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:NIV

[Ano92-150] Anonymous. The New Improved Version of DB2. Datamation, 38(16):57–??, August 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:AHL

[Ano92-151] Anonymous. The OS/ 2 Apps Are Here At Last. Datamation, 38(24):43–??, December 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:PB

[Ano92-152] Anonymous. The PC Big 50.

Datamation, 38(24):36–??, December 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:TJF

[Ano92-153] Anonymous. Time To Join The FTAM Team? Datamation, 38 (18):53–??, September 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:TD

[Ano92-154] Anonymous. The traveling desktop. Datamation, 38(??):11–??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:TMP

[Ano92-155] Anonymous. Twin mainframes power Lufthansa’s reservations. Datamation, 38(20):95–??, October 01, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:URO

[Ano92-156] Anonymous. Unifying the relational and object models. Datamation, 38(22):56–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:UID

[Ano92-157] Anonymous. Uniting IBM/DEC Net Management. Datamation, 38(17):28–??, August 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:UUM

[Ano92-158] Anonymous. UnixWare: UNIX For The Masses. Datamation, 38 (??):10–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:UGG

[Ano92-159] Anonymous. User groups grow independent. Datamation, 38 (25):72–??, December 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:UWB

[Ano92-160] Anonymous. Users want the bare fax. Datamation, 38(18):116–??, September 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:WDF

[Ano92-161] Anonymous. What does frame relay really cost? Datamation, 38 (24):55–??, December 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:WMS

[Ano92-162] Anonymous. What MIS Should Know About Videoconferencing. Datamation, 38(17):60–??, August 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:WPM

[Ano92-163] Anonymous. What PC Managers Say About Client/Server. Datamation, 38(19):54–??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:WFE

[Ano92-164] Anonymous. Where to find expert help for downsizing. Datamation, 38(18):61–??, September 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:WWS

[Ano92-165] Anonymous. Who Will Support APPN? Datamation, 38(20):46– ??, October 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:WTI

[Ano92-166] Anonymous. Why TCP / IP Still Has The Edge. Datamation, 38(20):38–??, October 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:WDO

[Ano92-167] Anonymous. Windows Drives OOP On The Desktop. Datamation, 38(22):50–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:WGB

[Ano92-168] Anonymous. Windows Gets Down To BASICs. Datamation, 38(23):59–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:WKC

[Ano92-169] Anonymous. Windows keeps coming on strong. Datamation, 38(??):26–??, July 15, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:WTU

[Ano92-170] Anonymous. Windows-tweaking utilities. Datamation, 38(??):22– ??, July 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:XTX

[Ano92-171] Anonymous. X Terminals X-Plore New Territory. Datamation, 38(19):108–??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:XCB

[Ano92-172] Anonymous. Xerox’s CIO: Beyond Command And Control.

Datamation, 38(25):67–??, December 15, 1992. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1992:XD

[Ano92-173] Anonymous. XL/ Datacomp.

Datamation, 38(??):S–42–S–??, November 01, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:RL

[Ano93a] Anonymous. 4GLs Reclaim The Limelight. Datamation, 39(6): 88–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:DSE

[Ano93b] Anonymous. 4th Dimension Software: Enterprise-wide Automation from a Workstation. Datamation, 39(??):S–28–S–??, September 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:NPSa

[Ano93c] Anonymous. 85 New Products:

Show Report. Datamation, 39 (17):85–??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CPP

[Ano93d] Anonymous. A Capacity Planning Primer. Datamation, 39(6): 77–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:DGB

[Ano93e] Anonymous. A Database GUI Builder. Datamation, 39(21):98– ??, November 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PCC

[Ano93f] Anonymous. A Plan That Cured Chaos. Datamation, 39(1):77– ??, January 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SSSd

[Ano93g] Anonymous. A Souped Up Storage Server. Datamation, 39(21): 97–??, November 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TPP

[Ano93h] Anonymous. A Two-Pentium PC Matches Workstations. Datamation, 39(17):73–??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NDL

[Ano93i] Anonymous. And Now, A DB2 For The LAN. Datamation, 39 (7):23–??, April 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:GCS

[Ano93j] Anonymous. AS/400 gets client/ server facelift. Datamation, 39 (8):67–??, April 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:AUN

[Ano93k] Anonymous. ATM: Ultimate Network, Or Ultimate Hype? Datamation, 39(7):30–??, April 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:AW

[Ano93l] Anonymous. Audio for Windows. Datamation, 39(1):82–??, January 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:AYA

[Ano93m] Anonymous. Automate your asset management! Datamation, 39(11):47–??, June 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:ASM

[Ano93n] Anonymous. AViiON Serves More For Less. Datamation, 39 (15):69–??, August 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:APD

[Ano93o] Anonymous. Avoid the pitfalls of distributed development. Datamation, 39(9):69–??, May 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:APR

[Ano93p] Anonymous. Avoiding The Pitfalls Of Replacing SNA. Datamation, 39(16):38–??, August 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:BBB

[Ano93q] Anonymous. Battling BLOBs To Build Image Networks. Datamation, 39(11):65–??, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:BLB

[Ano93r] Anonymous. Big-LAN Backup Without Backaches. Datamation, 39(8):48–??, April 15, 1993. CO-DEN DTMNAT. ISSN 0011-

Anonymous:1993:BSC

[Ano93s] Anonymous. Big screens come to color notebooks. Datamation, 39 (22):89–??, November 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:BCT

[Ano93t] Anonymous. Bonding with customers through better service. Datamation, 39(21):69–??, November 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:BIM

[Ano93u] Anonymous. Bravo for inexpensive multimedia. Datamation, 39 (16):62–??, August 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:BCA

[Ano93v] Anonymous. Build custom apps at packaged prices. Datamation, 39(11):71–??, June 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:BGN

[Ano93w] Anonymous. Building a Global Network On a Shoestring. Datamation, 39(10):59–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:BPB

[Ano93x] Anonymous. Business profits from bulletin boards. Datamation, 39(9):45–??, May 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:BGO

[Ano93y] Anonymous. Buyer’s Guide:

One-Size-Fits-All GUI Tools? Datamation, 39(9):92–??, May 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:BGU

[Ano93z] Anonymous. Buyer’s Guide:

Users Rethink Life Cycle CASE. Datamation, 39(10):102–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CCH

[Ano93-27] Anonymous. CA Closes Hole in Client/Server Management. Datamation, 39(23):69–??, December 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:CCMa

[Ano93-28] Anonymous. Can CA Manage Open Systems? Datamation, 39(1):18–??, January 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CKR

[Ano93-29] Anonymous. Can Kerberos Really Make UNIX Secure? Datamation, 39(1):59–??, January 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:COa

[Ano93-30] Anonymous. Career Opportunities. Datamation, 39(6):106–??, March 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:COb

[Ano93-31] Anonymous. Career Opportunities. Datamation, 39(7):80–??, April 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:COc

[Ano93-32] Anonymous. Career Opportunities. Datamation, 39(8):99–??, April 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:COd

[Ano93-33] Anonymous. Career Opportunities. Datamation, 39(10):116–??, May 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:COe

[Ano93-34] Anonymous. Career Opportunities. Datamation, 39(14):90–??, July 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:COf

[Ano93-35] Anonymous. Career Opportunities. Datamation, 39(15):79–??, August 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:COg

[Ano93-36] Anonymous. Career Opportunities. Datamation, 39(16):82–??, August 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:COh

[Ano93-37] Anonymous. Career Opportunities. Datamation, 39(17):91–??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:COi

[Ano93-38] Anonymous. Career Opportunities. Datamation, 39(18):108–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:COj

[Ano93-39] Anonymous. Career Opportunities. Datamation, 39(19):104– ??, October 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:COk

[Ano93-40] Anonymous. Career Opportunities. Datamation, 39(20):105– ??, October 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:COo

[Ano93-41] Anonymous. Career Opportunities. Datamation, 39(21):118– ??, November 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:COp

[Ano93-42] Anonymous. Career Opportunities. Datamation, 39(22):112– ??, November 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:COq

[Ano93-43] Anonymous. Career Opportunities. Datamation, 39(23):118– ??, December 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CSN

[Ano93-44] Anonymous. CA’s Search For a New Pricing Paradigm. Datamation, 39(6):65–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CRE

[Ano93-45] Anonymous. Chase Research:

Excellence in UNIX I/O products. Datamation, 39(??):S–6–S– ??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CPB

[Ano93-46] Anonymous. CIOs Put the T Back in IT. Datamation, 39(23): 28–??, December 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CCU

[Ano93-47] Anonymous. CLEO Communications: UNIX-to-IBM Connectivity. Datamation, 39(??):S–22–S– ??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CSCa

[Ano93-48] Anonymous. Client/Server Comes To The Enterprise. Datamation, 39(2):101–??, January 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CSCc

[Ano93-49] Anonymous. Client/Server For CICS Finally Arrives. Datamation, 39(18):79–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CSI

[Ano93-50] Anonymous. Client/Server Integrators — Is Bigger Better? Datamation, 39(14):24–??, July 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CST

[Ano93-51] Anonymous. Client/Server Tools Revive CASE. Datamation, 39 (7):38–??, April 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CWG

[Ano93-52] Anonymous. Cobol Workbenches Get Down To Business. Datamation, 39(15):53–??, August 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CTE

[Ano93-53] Anonymous. Compaq Takes On The Enterprise. Datamation, 39(19):18–??, October 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CTD

[Ano93-54] Anonymous. Control Thousands of Desktop Apps! Datamation, 39(23):18–??, December 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CSSa

[Ano93-55] Anonymous. Corporate Software Scorecard. Datamation, 39(6): 104–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CSSb

[Ano93-56] Anonymous. Corporate Software Scorecard. Datamation, 39(10): 114–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CSSc

[Ano93-57] Anonymous. Corporate Software Scorecard. Datamation, 39(16): 78–??, August 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CSSd

[Ano93-58] Anonymous. Corporate Software Scorecard. Datamation, 39(18):

107–??, September 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CSSe

[Ano93-59] Anonymous. Corporate Software Scorecard. Datamation, 39(20): 100–??, October 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CSSf

[Ano93-60] Anonymous. Corporate Software Scorecard. Datamation, 39(22): 108–??, November 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CSSi

[Ano93-61] Anonymous. Corporate Software Scorecard. Datamation, 39(24): 77–??, December 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CSSh

[Ano93-62] Anonymous. Corporate Software Scorecard: Lotus-1-2-3 is top-selling spreadsheet package. Datamation, 39(23):114–??, December 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:CCMb

[Ano93-63] Anonymous. Credit card modems now do fax, too. Datamation, 39(16):65–??, August 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CDN

[Ano93-64] Anonymous. Customers drive new manufacturing software.

Datamation, 39(22):71–??, November 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:DVC

[Ano93-65] Anonymous. DBMS Vendors Chase Sybase For Client/Server. Datamation, 39(13):27–??, July 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:DUA

[Ano93-66] Anonymous. Developing UNIX Applications. Datamation, 39 (??):S–23–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:DSC

[Ano93-67] Anonymous. Diagnostic Software Comes To The Rescue. Datamation, 39(11):55–57, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:DMN

[Ano93-68] Anonymous. Digital Moves NT Beyond the Workgroup. Datamation, 39(24):53–??, December 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:DYI

[Ano93-69] Anonymous. Digital Yields To IBM In Net Management. Datamation, 39(20):59–??, October 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:DAB

[Ano93-70] Anonymous. Distributed Applications Without Bottlenecks. Datamation, 39(10):76–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:DDM

[Ano93-71] Anonymous. Distributed DBMSs Move Into The Trenches. Datamation, 39(8):59–??, April 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:DWPa

[Ano93-72] Anonymous. Does IS Want The PowerPC? Datamation, 39(21): 28–??, November 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:DWPb

[Ano93-73] Anonymous. Does IS want the PowerPC? Datamation, 39(21): 28–30, 35–36, November 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:DRR

[Ano93-74] Anonymous. Does Reengineering Really Work? Datamation, 39(15):22–??, August 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:ELF

[Ano93-75] Anonymous. Easing the LAN Fax Burden. Datamation, 39(14): 73–??, July 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:EIS

[Ano93-76] Anonymous. ECCS, Inc.: System Integration for Mass Storage. Datamation, 39(??):S–8–S– ??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:ETc

[Ano93-77] Anonymous. Economic Trends. Datamation, 39(6):82–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:ETd

[Ano93-78] Anonymous. Economic Trends.

Datamation, 39(10):89–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:ESS

[Ano93-79] Anonymous. EMASS Storage System Solutions: Leading the Way in Mass Data Storage. Datamation, 39(??):S–27–S– ??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:EPI

[Ano93-80] Anonymous. Emulex: Performance, Integrity and Low Cost. Datamation, 39(??):S–3– S–??, September 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:ETa

[Ano93-81] Anonymous. Enfin, Take 2.0.

Datamation, 39(2):102–??, January 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:ETP

[Ano93-82] Anonymous. Enterprise Technology at PC Prices. Datamation, 39(24):56–??, December 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:FFP

[Ano93-83] Anonymous. Fast 486s: Features Plus Speed. Datamation, 39(8): 82–??, April 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:FR

[Ano93-84] Anonymous. A faster, roomier AS/400. Datamation, 39(9):89– ??, May 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:FFS

[Ano93-85] Anonymous. FDDI: Not Fastest But Still Fit. Datamation, 39(7): 31–??, April 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:FLF

[Ano93-86] Anonymous. Finding Less Fault With UNIX. Datamation, 39 (20):77–??, October 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:FSF

[Ano93-87] Anonymous. FourGen Software: FourGen Customers Focus on Performance. pp S-4. Datamation, 39(??):??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:F

[Ano93-88] Anonymous. Framework. Datamation, 39(23):6–??, December 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:FSU

[Ano93-89] Anonymous. Full-Strength UNIX Finally Comes To PCs. Datamation, 39(14):47–??, July 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:FVa

[Ano93-90] Anonymous. Future Vision. Datamation, 39(6):112–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:FVb

[Ano93-91] Anonymous. Future Vision.

Datamation, 39(8):104–??, April 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:FVc

[Ano93-92] Anonymous. Future Vision.

Datamation, 39(10):124–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:FVd

[Ano93-93] Anonymous. Future Vision.

Datamation, 39(14):96–??, July 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:FVe

[Ano93-94] Anonymous. Future Vision.

Datamation, 39(16):88–??, August 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:FVf

[Ano93-95] Anonymous. Future Vision.

Datamation, 39(18):112–??, September 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:FVg

[Ano93-96] Anonymous. Future Vision.

Datamation, 39(19):108–??, October 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:FVh

[Ano93-97] Anonymous. Future Vision.

Datamation, 39(20):108–??, October 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:FVi

[Ano93-98] Anonymous. Future Vision. Datamation, 39(21):124–??, November 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:FVl

[Ano93-99] Anonymous. Future Vision. Datamation, 39(22):116–??, November 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:FVm

[Ano93-100] Anonymous. Future Vision. Datamation, 39(24):84–??, December 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:GCC

[Ano93-101] Anonymous. Get Control Of Cross-Platform Development. Datamation, 39(16):45–??, August 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:GSM

[Ano93-102] Anonymous. Get Smart Management For Your Smart Hubs. Datamation, 39(2):53–??, January 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:GLN

[Ano93-103] Anonymous. Global Leaders In Nine Product Categories. Datamation, 39(12):22–??, June 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:GRH

[Ano93-104] Anonymous. Growth Rate Highest Since 1988. Datamation, 39 (12):12–??, June 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HDS

[Ano93-105] Anonymous. Hard Demand For Soft Skills. Datamation, 39(2): 28–??, January 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HDI

[Ano93-106] Anonymous. HDS Delivers Improved Drives. Datamation, 39 (16):61–??, August 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HHD

[Ano93-107] Anonymous. Help For The Help Desk. Datamation, 39(7):59–??, April 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:HSR

[Ano93-108] Anonymous. Hierarchical Storage Reaches the LAN. Datamation, 39(24):73–??, December 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HSN

[Ano93-109] Anonymous. Honey, I Shrunk NetView! Datamation, 39(7): 16–??, April 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HCW

[Ano93-110] Anonymous. Hot Companies To Watch. Datamation, 39(17):20– 22, 26, 28, September 01, 1993. CODEN DTMNAT. ISSN 0011-

Anonymous:1993:HGC

[Ano93-111] Anonymous. Hot Growth Companies. Datamation, 39(12):15– ??, June 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HIQ

[Ano93-112] Anonymous. How ISO 9000 Quality Programs Affect IS. Datamation, 39(15):65–??, August 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:HCT

[Ano93-113] Anonymous. How To Certify That Training Works. Datamation, 39(17):67–??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HCY

[Ano93-114] Anonymous. How To Clean Up Your Dirty Data. Datamation, 39(16):51–??, August 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HGS

[Ano93-115] Anonymous. How To Get Started On The OOP Trail. Datamation, 39(21):40–??, November 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HKM

[Ano93-116] Anonymous. How To Kill a Mutant Virus. Datamation, 39(9): 53–??, May 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HUC

[Ano93-117] Anonymous. How UNIX Can Manage Systems and Networks.

Datamation, 39(??):S–16–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HLW

[Ano93-118] Anonymous. HP’s Latest Workstation Leap. Datamation, 39(1): 81–??, January 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:IS

[Ano93-119] Anonymous. IBM speeds up OS/2. Datamation, 39(11):91– ??, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:IHA

[Ano93-120] Anonymous. IBM/HP AnyLAN Is No SuperLAN. Datamation, 39(23):55–??, December 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:ISS

[Ano93-121] Anonymous. Imaging: Start Small, Think Big. Datamation, 39(13):47–??, July 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LC

[Ano93-122] Anonymous. In The Loss Column. Datamation, 39(12):16–??, June 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:IGE

[Ano93-123] Anonymous. Indigo Goes To Extremes. Datamation, 39(6):85– ??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:IFO

[Ano93-124] Anonymous. Information Foundation: The Open Systems Migration Experts. Datamation, 39(??):S–30–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Ia

[Ano93-125] Anonymous. Inside. Datamation, 39(6):11–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Ib

[Ano93-126] Anonymous. Inside. Datamation, 39(7):9–??, April 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Ic

[Ano93-127] Anonymous. Inside. Datamation, 39(8):9–??, April 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Id

[Ano93-128] Anonymous. Inside. Datamation, 39(10):11–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Ie

[Ano93-129] Anonymous. Inside. Datamation, 39(11):11–??, June 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:If

[Ano93-130] Anonymous. Inside. Datamation, 39(13):9–??, July 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Ig

[Ano93-131] Anonymous. Inside. Datamation, 39(14):9–??, July 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Ih

[Ano93-132] Anonymous. Inside. Datamation, 39(15):8–??, August 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Ii

[Ano93-133] Anonymous. Inside. Datamation, 39(16):8–??, August 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Ij

[Ano93-134] Anonymous. Inside. Datamation, 39(17):6–??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Ik

[Ano93-135] Anonymous. Inside. Datamation, 39(18):8–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Il

[Ano93-136] Anonymous. Inside. Datamation, 39(19):8–??, October 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Im

[Ano93-137] Anonymous. Inside. Datamation, 39(20):8–??, October 15, 1993. CODEN DTMNAT. ISSN 0011-

Anonymous:1993:In

[Ano93-138] Anonymous. Inside. Datamation, 39(21):10–??, November 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Ir

[Ano93-139] Anonymous. Inside. Datamation, 39(22):10–??, November 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:It

[Ano93-140] Anonymous. Inside. Datamation, 39(23):10–??, December 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Iu

[Ano93-141] Anonymous. Inside. Datamation, 39(24):8–??, December 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:IRB

[Ano93-142] Anonymous. Insoft: Revolutionizing Business with Desktop Conferencing. Datamation, 39(??): S–13–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:INP

[Ano93-143] Anonymous. International New Products. Datamation, 39(21): 117–??, November 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:ITI

[Ano93-144] Anonymous. Is It Time To Insource Your Financial Apps?

Datamation, 39(18):75–??, September 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:SC

[Ano93-145] Anonymous. IS Satisfies The Customer. Datamation, 39(19): 79–??, October 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SSU

[Ano93-146] Anonymous. IS Seeks Security In UNIX. Datamation, 39(15): 59–??, August 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SSA

[Ano93-147] Anonymous. Is Smart Software The Answer for Help Desks? Datamation, 39(23):42–??, December 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:IRP

[Ano93-148] Anonymous. IT Revenues By Product: a Five-Year Perspective. Datamation, 39(12):22–??, June 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:PFPa

[Ano93-149] Anonymous. It’s a Phone! It’s a Fax! It’s a Pen Computer! Datamation, 39(8):79–??, April 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:JNI

[Ano93-150] Anonymous. Jobs’ NeXTSTEP In Onto Intel PCs. Datamation, 39(10):97–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-

Anonymous:1993:KRC

[Ano93-151] Anonymous. KnowledgeWare Revs Up Its OS/2 CASE Tools. Datamation, 39(2):73–??, January 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:LSQ

[Ano93-152] Anonymous. LAGO Systems: Quality Storage Solutions. Datamation, 39(??):S–16–S–??, September 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:LBP

[Ano93-153] Anonymous. LAN Boom Paves The Way For Client/Server. Datamation, 39(22):40–??, November 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:LSG

[Ano93-154] Anonymous. LAN Storage Gets Serious. Datamation, 39(20):22– ??, October 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LTS

[Ano93-155] Anonymous. LAN Tools Start Working Together. Datamation, 39(21):61–??, November 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LSP

[Ano93-156] Anonymous. Landmark Systems: Performance Management Products. Datamation, 39(??):S–18– S–??, September 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:La

[Ano93-157] Anonymous. Letters. Datamation, 39(6):14–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Lb

[Ano93-158] Anonymous. Letters. Datamation, 39(7):10–??, April 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Lf

[Ano93-159] Anonymous. Letters. Datamation, 39(8):10–??, April 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Lg

[Ano93-160] Anonymous. Letters. Datamation, 39(10):14–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Le

[Ano93-161] Anonymous. Letters. Datamation, 39(11):14–??, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Lh

[Ano93-162] Anonymous. Letters. Datamation, 39(14):12–??, July 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Li

[Ano93-163] Anonymous. Letters. Datamation, 39(15):12–??, August 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Lj

[Ano93-164] Anonymous. Letters. Datamation, 39(16):12–??, August 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Lk

[Ano93-165] Anonymous. Letters. Datamation, 39(17):9–??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Ll

[Ano93-166] Anonymous. Letters. Datamation, 39(18):13–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Lm

[Ano93-167] Anonymous. Letters. Datamation, 39(19):12–??, October 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Ln

[Ano93-168] Anonymous. Letters. Datamation, 39(20):12–??, October 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:Lo

[Ano93-169] Anonymous. Letters. Datamation, 39(21):14–??, November 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAM

[Ano93-170] Anonymous. Life After The Merger At BankAmerica. Datamation, 39(11):85–??, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LCD

[Ano93-171] Anonymous. Link Corporate Data And Documents. Datamation, 39(7):65–??, April 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LYL

[Ano93-172] Anonymous. Link your LAN to your AS/400. Datamation, 39(19):85–??, October 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAa

[Ano93-173] Anonymous. Look Ahead. Datamation, 39(6):19–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAb

[Ano93-174] Anonymous. Look Ahead. Datamation, 39(7):13–??, April 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAc

[Ano93-175] Anonymous. Look Ahead. Datamation, 39(8):15–??, April 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAd

[Ano93-176] Anonymous. Look Ahead. Datamation, 39(10):19–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAe

[Ano93-177] Anonymous. Look Ahead. Datamation, 39(11):19–??, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAf

[Ano93-178] Anonymous. Look Ahead. Datamation, 39(13):13–??, July 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAg

[Ano93-179] Anonymous. Look Ahead. Datamation, 39(14):15–??, July 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAh

[Ano93-180] Anonymous. Look Ahead. Datamation, 39(15):17–??, August 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAi

[Ano93-181] Anonymous. Look Ahead. Datamation, 39(16):17–??, August 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAj

[Ano93-182] Anonymous. Look Ahead. Datamation, 39(17):13–??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAk

[Ano93-183] Anonymous. Look Ahead. Datamation, 39(19):15–??, October 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAl

[Ano93-184] Anonymous. Look Ahead. Datamation, 39(20):17–??, October 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAq

[Ano93-185] Anonymous. Look Ahead. Datamation, 39(21):19–??, November 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAs

[Ano93-186] Anonymous. Look Ahead. Datamation, 39(22):15–??, November 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAt

[Ano93-187] Anonymous. Look Ahead. Datamation, 39(23):13–??, December 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAu

[Ano93-188] Anonymous. Look Ahead. Datamation, 39(24):13–??, December 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LAR

[Ano93-189] Anonymous. Look Ahead:

Reengineering at Levi Strauss has created a tight fit with IT. Datamation, 39(18):19–??, September 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:MPL

[Ano93-190] Anonymous. Mainframe Power For 80% Less. Datamation, 39 (1):67–??, January 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MWB

[Ano93-191] Anonymous. Make Way For The 32-Bit Heavy Hitters. Datamation, 39(8):34–??, April 15, 1993.

CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MWC

[Ano93-192] Anonymous. Make Workflow Changes On The Fly. Datamation, 39(19):81–??, October 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MCM

[Ano93-193] Anonymous. Making C++ More Approachable. Datamation, 39 (2):37–??, January 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MMW

[Ano93-194] Anonymous. Making Multimedia Work. Datamation, 39(17):30– ??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MSL

[Ano93-195] Anonymous. Making The SQL Link To Data. Datamation, 39 (15):71–??, August 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MUW

[Ano93-196] Anonymous. Making UNIX Work. Datamation, 39(??):S–20– S–??, September 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MDM

[Ano93-197] Anonymous. Managing The Document Management Explosion. Datamation, 39(13):41–??, July 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MRa

[Ano93-198] Anonymous. Media Report.

Datamation, 39(6):22–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MRb

[Ano93-199] Anonymous. Media Report.

Datamation, 39(10):24–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MRd

[Ano93-200] Anonymous. Media Report.

Datamation, 39(11):24–??, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MMS

[Ano93-201] Anonymous. Memory Management Smartens Up. Datamation, 39(18):48–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MMM

[Ano93-202] Anonymous. MIS Meets Multimedia On The Network. Datamation, 39(14):32–??, July 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MST

[Ano93-203] Anonymous. Mixing SNA and TCP/IP? Not So Fast! Datamation, 39(24):36–??, December 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MMG

[Ano93-204] Anonymous. Modern Macros Gain Programming Power. Datamation, 39(15):40–??, August 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MGM

[Ano93-205] Anonymous. MS-DOS 6.0: Good But Not a Must. Datamation, 39 (8):42–??, April 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MFB

[Ano93-206] Anonymous. Multimedia Full Of Bravado. Datamation, 39(6):86– ??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MTE

[Ano93-207] Anonymous. Multimedia Tools Are Enterprise Ready. Datamation, 39(20):92–??, October 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MSP

[Ano93-208] Anonymous. Multiprocessor Software Plays Catch-Up. Datamation, 39(11):77–??, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NMF

[Ano93-209] Anonymous. Net Management Forges Stronger Enterprise Links. Datamation, 39(22):34– ??, November 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NLP

[Ano93-210] Anonymous. Network Laser Printers Tackle High-End Jobs. Datamation, 39(21):103–??, November 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:NSY

[Ano93-211] Anonymous. Network Security:

Is Your LAN A Sieve? Datamation, 39(17):79–??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NGB

[Ano93-212] Anonymous. New GUI Builders Make CICS Programming Fun! Datamation, 39(2):67–??, January 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:NLOa

[Ano93-213] Anonymous. New Life For Old General Ledgers. Datamation, 39 (17):55–??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NLOb

[Ano93-214] Anonymous. New life for old PS/2s. Datamation, 39(18):52– ??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NPa

[Ano93-215] Anonymous. New Products.

Datamation, 39(6):99–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NPb

[Ano93-216] Anonymous. New Products.

Datamation, 39(7):73–??, April 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NPc

[Ano93-217] Anonymous. New Products.

Datamation, 39(8):95–??, April 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NPd

[Ano93-218] Anonymous. New Products. Datamation, 39(10):107–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NPe

[Ano93-219] Anonymous. New Products. Datamation, 39(11):99–??, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NPf

[Ano93-220] Anonymous. New Products. Datamation, 39(13):55–??, July 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NPg

[Ano93-221] Anonymous. New Products. Datamation, 39(14):81–??, July 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NPh

[Ano93-222] Anonymous. New Products. Datamation, 39(15):74–??, August 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:NPi

[Ano93-223] Anonymous. New Products. Datamation, 39(16):69–??, August 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:NPj

[Ano93-224] Anonymous. New Products. Datamation, 39(18):93–??, September 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:NPn

[Ano93-225] Anonymous. New Products.

Datamation, 39(21):109–??, November 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:NPo

[Ano93-226] Anonymous. New Products.

Datamation, 39(22):95–??, November 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:NPp

[Ano93-227] Anonymous. New Products.

Datamation, 39(23):97–??, December 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:NPSb

[Ano93-228] Anonymous. New Products:

Show Report. Datamation, 39 (19):93–??, October 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NRN

[Ano93-229] Anonymous. New Rules, New Ratings As IS Reengineers. Datamation, 39(21):85–??, November 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NTE

[Ano93-230] Anonymous. New tools ease the move to AS/400s. Datamation, 39(6):59–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NSL

[Ano93-231] Anonymous. Non Standard Logics: Development Tools for X and Motif. Datamation, 39(??):

S–33–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NAC

[Ano93-232] Anonymous. North American CIOs Feel The Pressure. Datamation, 39(2):86–??, January 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:JSA

[Ano93-233] Anonymous. Not Just a Spreadsheet Anymore. Datamation, 39 (22):83–??, November 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NAW

[Ano93-234] Anonymous. Notes Adds Workflow. Datamation, 39(11):92–??, June 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:NBI

[Ano93-235] Anonymous. Novell Beefs Up Its IBM Connection. Datamation, 39(9):59–??, May 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NHS

[Ano93-236] Anonymous. NT: Hardware To Spare At High End. Datamation, 39(10):38–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NIR

[Ano93-237] Anonymous. NT: Is It Ready For Critical Apps? Datamation, 39(10):28–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-

Anonymous:1993:NPTa

[Ano93-238] Anonymous. NT: Power Tools Are Still “Not There.”. Datamation, 39(10):34–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:ODF

[Ano93-239] Anonymous. Object Databases Find Their Niche. Datamation, 39(18):56–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:OT

[Ano93-240] Anonymous. Object tools for OS/2. Datamation, 39(13):49– ??, July 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:SP

[Ano93-241] Anonymous. Of Standards and Protocols. Datamation, 39(??): S–4–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:OSG

[Ano93-242] Anonymous. OLE 2.0: Stronger Glue, Stickier To Use. Datamation, 39(20):65–??, October 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:OMP

[Ano93-243] Anonymous. Open Management Platforms Tackle NetWare. Datamation, 39(20):56–??, October 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:OSE

[Ano93-244] Anonymous. Operating System Evolution: From One To Several. Datamation, 39(??):S–2–S– ??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:OSN

[Ano93-245] Anonymous. Optical Shines For Near Line Storage. Datamation, 39(22):79–??, November 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:OSS

[Ano93-246] Anonymous. Our Secret Shopper Buys Mail Order PCs. Datamation, 39(10):41–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PFPb

[Ano93-247] Anonymous. Paper-Free Publishing Gets Real. Datamation, 39 (17):39–??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PMSb

[Ano93-248] Anonymous. PC Managers Speak Out About NT. Datamation, 39(19):68–??, October 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PBO

[Ano93-249] Anonymous. PCs Break Into The OLTP Ranks. Datamation, 39 (16):32–??, August 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PCF

[Ano93-250] Anonymous. Pen Computing: First, Pick An OS. Datamation, 39(19):36–??, October 01, 1993. CODEN DTMNAT. ISSN 0011-

Anonymous:1993:PHG

[Ano93-251] Anonymous. Pen Hardware: Get Practical. Datamation, 39(24): 30–??, December 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PCL

[Ano93-252] Anonymous. POSIX Cracks The Lock On MVS. Datamation, 39 (7):47–??, April 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PEA

[Ano93-253] Anonymous. PowerBuilding Enterprise Applications. Datamation, 39(14):18–??, July 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PAP

[Ano93-254] Anonymous. Pre-ATM Plus Savings. Datamation, 39(17):74–??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PPA

[Ano93-255] Anonymous. Prefab Programming: Apps Without Coding. Datamation, 39(10):46–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PCE

[Ano93-256] Anonymous. Premenos Corp.:

EDI/ e: a Competitive Maneuver. Datamation, 39(??):S–44–S– ??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PWP

[Ano93-257] Anonymous. Preparing The Way For Pentium. Datamation, 39(6):

36–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PSU

[Ano93-258] Anonymous. Prices Slashed For UNIX Workbenches. Datamation, 39(21):51–??, November 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PM

[Ano93-259] Anonymous. Project Management. Datamation, 39(1):41– ??, January 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PMSa

[Ano93-260] Anonymous. Project Management Software That’s IS Friendly. Datamation, 39(19): 55–??, October 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PMU

[Ano93-261] Anonymous. Proprietary Mini Users Plan Escape Routes. Datamation, 39(11):81–??, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PUT

[Ano93-262] Anonymous. Put Usability To The Test. Datamation, 39(14): 61–??, July 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:RRU

[Ano93-263] Anonymous. RAID Raises UNIX Reliability. Datamation, 39(17): 61–??, September 01, 1993. CO-DEN DTMNAT. ISSN 0011-

Anonymous:1993:RPF

[Ano93-264] Anonymous. RDBMS Prices Are In For a Fall. Datamation, 39 (15):30–??, August 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:RM

[Ano93-265] Anonymous. Recasting The Mainframe. Datamation, 39(20): 87–??, October 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:RDD

[Ano93-266] Anonymous. Recover From Distributed Disasters. Datamation, 39(23):73–??, December 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:RTC

[Ano93-267] Anonymous. Reengineering Tops CIO Menu. Datamation, 39(8): 73–??, April 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:RCM

[Ano93-268] Anonymous. Relational Considerations: Meeting Demand For Data. . Datamation, 39(??):S– 20–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:RLG

[Ano93-269] Anonymous. Remote LANs Get Outsourcing Option. Datamation, 39(14):36–??, July 15, 1993. CODEN DTMNAT. ISSN 0011-

Anonymous:1993:RBT

[Ano93-270] Anonymous. RISCy Business: Today’s UNIX Players. Datamation, 39(??):S–8–S–??, September 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:RTB

[Ano93-271] Anonymous. Routers Take On The Branch Office. Datamation, 39(18):83–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SCSa

[Ano93-272] Anonymous. SAP’s Client/ Server Battle Plan. Datamation, 39(6):26–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SME

[Ano93-273] Anonymous. Scaling E-Mail For The Enterprise. Datamation, 39 (8):20–??, April 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SMC

[Ano93-274] Anonymous. Secure E-Mail Cheaply With Software Encryption. Datamation, 39(23):48–??, December 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SUF

[Ano93-275] Anonymous. Securing The Unix Future. Datamation, 39(??):S– 24–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SS

[Ano93-276] Anonymous. Service For The Server. Datamation, 39(11):93– ??, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SSSa

[Ano93-277] Anonymous. Shopping for a Superior Superserver. Datamation, 39(1):49–??, January 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SMB

[Ano93-278] Anonymous. Shotgun Marriage For Battling Networks. Datamation, 39(6):52–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SRL

[Ano93-279] Anonymous. Should IS Ride The Local Bus? Datamation, 39(6): 47–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SSC

[Ano93-280] Anonymous. Six Sigma CASE:

Real Software Engineering Answers. . Datamation, 39(??): S–34–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SWI

[Ano93-281] Anonymous. Six Ways To Improve Mainframe Productivity. Datamation, 39(10):72–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SHH

[Ano93-282] Anonymous. Smart Hubs: Head of the Class? Datamation, 39 (1):86–??, January 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SFF

[Ano93-283] Anonymous. Smoothing The Flow Of Forms. Datamation, 39 (14):68–??, July 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SMD

[Ano93-284] Anonymous. Software Moguls: Data Protection — Across the Network. . Datamation, 39(??): S–11–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SSSb

[Ano93-285] Anonymous. Software, Services Soar. Datamation, 39(12):14–??, June 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:SB

[Ano93-286] Anonymous. Software’s Big 100. Datamation, 39(18):38–??, September 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:SRI

[Ano93-287] Anonymous. Special Report: 1993 Industry Outlook; IS Meets Business Challenges Head On. Datamation, 39(1):26–??, January 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:STP

[Ano93-288] Anonymous. SQL Testing In Pictures. Datamation, 39(18):80–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SSSc

[Ano93-289] Anonymous. Stretch Server Storage. Datamation, 39(20):88– ??, October 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SSE

[Ano93-290] Anonymous. Sybase Steps Up To The Enterprise. Datamation, 39 (13):18–??, July 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SYM

[Ano93-291] Anonymous. Synchronize Your E-Mail Directories. Datamation, 39(22):61–??, November 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SHP

[Ano93-292] Anonymous. Syncsort: High Performance UNIX Sorting. Datamation, 39(??):S–36–S–??, September 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:SHA

[Ano93-293] Anonymous. SynOptics Hub Has ATM Horsepower. Datamation, 39(23):81–??, December 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TCS

[Ano93-294] Anonymous. Tackle Client/ Server Apps With Born-Again 4GLs. Datamation, 39(23):61– ??, December 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TCA

[Ano93-295] Anonymous. Take Control Of Application Performance. Datamation, 39(10):67–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TCY

[Ano93-296] Anonymous. Take control of your AS/400 Systems! Datamation, 39(9):63–??, May 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TIW

[Ano93-297] Anonymous. TCP/IP Winning The Backbone Battle. Datamation, 39(19):26–??, October 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TTa

[Ano93-298] Anonymous. Technology Trends. Datamation, 39(6):81–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TTc

[Ano93-299] Anonymous. Technology Trends. Datamation, 39(10):88–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:BSP

[Ano93-300] Anonymous. The Battle Shifts To PC Services. Datamation, 39(2):47–??, January 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:BCF

[Ano93-301] Anonymous. The Best of COMDEX Fall ’92. Datamation, 39(2):109–??, January 15, 1993. CODEN DTMNAT. ISSN 0011-

Anonymous:1993:DWI

[Ano93-302] Anonymous. The Database Way To Integrate E-Mail. Datamation, 39(2):61–??, January 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:D

[Ano93-303] Anonymous. The Datamation 100. Datamation, 39(12):13–??, June 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:DIB

[Ano93-304] Anonymous. The Datamation 100 Information Bank. Datamation, 39(12):136–??, June 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:DNA

[Ano93-305] Anonymous. The Datamation North American 100. Datamation, 39(12):18–??, June 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HBL

[Ano93-306] Anonymous. The Heavy Burden Of Lan Costs. Datamation, 39 (11):60–??, June 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:LEW

[Ano93-307] Anonymous. The Low-End Workstation Race. Datamation, 39(2):104–??, January 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NST

[Ano93-308] Anonymous. The Network Server Takes Off. Datamation, 39(11):

40–??, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NII

[Ano93-309] Anonymous. The New IT Industry Takes Shape. Datamation, 39 (12):12–??, June 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:NPTb

[Ano93-310] Anonymous. The New Palmtops: Tools Or Toys? Datamation, 39(13):35–??, July 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:POR

[Ano93-311] Anonymous. The PC50: Our ranking of the leading PC companies reveals that the young turks of the PC world rose last year on the tide of plummeting prices. Ever-more-attractive prices continue to drive up sales of PCs this year. Datamation, 39(23):36–??, December 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PRK

[Ano93-312] Anonymous. The People Rollout: Key To Change. Datamation, 39(7):55–??, April 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SCSb

[Ano93-313] Anonymous. The Search For Client/Server Security. Datamation, 39(9):39–??, May 01, 1993. CODEN DTMNAT. ISSN 0011-

Anonymous:1993:SGA

[Ano93-314] Anonymous. The Software Group: Adding Reliability to UNIX Internetworking. Datamation, 39(??):S–14–S–??, September 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:SGN

[Ano93-315] Anonymous. The Struggle For Global Networks. Datamation, 39(18):26–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HEL

[Ano93-316] Anonymous. These Hubs Ease LAN Changes. Datamation, 39 (15):49–??, August 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:DCS

[Ano93-317] Anonymous. This Data Center Software Does Client/Server. Datamation, 39(24):45–??, December 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:TBH

[Ano93-318] Anonymous. Tight Budgets: It’s How You Cut ’Em. Datamation, 39(9):30–??, May 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TFE

[Ano93-319] Anonymous. Tools For Fast, Easy Access To Legacy Data. Datamation, 39(11):75–??, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TKP

[Ano93-320] Anonymous. Tools Keep Pace With Virtual Data Centers. Datamation, 39(2):81–??, January 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:TSS

[Ano93-321] Anonymous. Top Salaries Soar, Ranks Remain Flat. Datamation, 39(16):28–??, August 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TCV

[Ano93-322] Anonymous. Tune In To Cheap Videoconferencing. Datamation, 39(19):48–??, October 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TPO

[Ano93-323] Anonymous. Turmoil in Pen Operating Systems. Datamation, 39 (24):20–??, December 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TYM

[Ano93-324] Anonymous. Turn Your E-Mail Into Groupware. Datamation, 39 (8):28–??, April 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:TCO

[Ano93-325] Anonymous. Two Computers In One. Datamation, 39(14):67–??, July 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:UUE

[Ano93-326] Anonymous. UniForum: UniForum ’94: Exploring UNIX and Open Systems. pp S-40. Datamation, 39(??):??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:UUW

[Ano93-327] Anonymous. Unisys’ Unruh:

We’re Never Going To Be Done. Datamation, 39(20):83–??, October 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:UCO

[Ano93-328] Anonymous. UNIX Connectivity: Opening Client/Server Computing Opportunities. Datamation, 39(??):S–12–S–??, September 15, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:UPD

[Ano93-329] Anonymous. UNIX Product Directory. Datamation, 39(??):S– 28–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:USP

[Ano93-330] Anonymous. UNIX Server Power Grows In Clusters. Datamation, 39(16):57–??, August 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:USM

[Ano93-331] Anonymous. UNIX Storage Management: It’s Here Now! Datamation, 39(21):77–??, November 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:USS

[Ano93-332] Anonymous. UNIX Superserver Shoot-Out! Datamation, 39(22):

24–??, November 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:UPS

[Ano93-333] Anonymous. Users Praise SMP Price/Performance. Datamation, 39(20):73–??, October 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:VSS

[Ano93-334] Anonymous. Vendors Scramble To Support Open OLTP. Datamation, 39(18):67–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:VLP

[Ano93-335] Anonymous. Virtual LANs Pave The Way To ATM. Datamation, 39(16):20–??, August 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:VBG

[Ano93-336] Anonymous. Visual Basic Gets a Database Engine. Datamation, 39(15):70–??, August 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:VCM

[Ano93-337] Anonymous. Visual C++: a More Approachable C. Datamation, 39(15):36–??, August 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:VTG

[Ano93-338] Anonymous. Visual Tools Give COBOL A New Look. Datamation, 39(22):53–??, November 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:WOP

[Ano93-339] Anonymous. When Offshore Programming Works. Datamation, 39(14):61–??, July 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:WUM

[Ano93-340] Anonymous. When UNIX Makes a Difference: IS Success Stories. Datamation, 39(??):S–3–S– ??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:WAF

[Ano93-341] Anonymous. Why ATM Fascinates Network Chiefs. Datamation, 39(10):82–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:WCM

[Ano93-342] Anonymous. Will Cheaper Mainframes Slow Downsizing? Datamation, 39(19):63–??, October 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:WSR

[Ano93-343] Anonymous. Will Sun Also Rise In The Data Center? Datamation, 39(11):28–??, June 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:WNO

[Ano93-344] Anonymous. Win32s: Now One Code Fits All. Datamation, 39 (19):40–??, October 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:WR

[Ano93-345] Anonymous. A Windows-AS/400 router. Datamation, 39(9):90–??, May 01, 1993. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1993:WNY

[Ano93-346] Anonymous. Wireless Nets Not Yet Unplug and Play. Datamation, 39(17):47–??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:WWW

[Ano93-347] Anonymous. Wireless WANs Have a Ways To Go. Datamation, 39(10):52–??, May 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:WTE

[Ano93-348] Anonymous. Workflow Takes On the Enterprise. Datamation, 39 (23):88–??, December 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:XWT

[Ano93-349] Anonymous. X Windows Takes Hold In Commercial IS. Datamation, 39(20):50–??, October 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:XSS

[Ano93-350] Anonymous. Xbase: Standard Or Stopgap? An ANSIsupported group hopes to anoint a standard Xbase language within 18 to 36 months. That should be relatively easy to achieve, but getting vendors and users to bless a number of objectoriented extensions will be more of a challenge. Datamation, 39 (6):41–??, March 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:XPX

[Ano93-351] Anonymous. XPM Partners:

XPM: How to Do More with Less. Datamation, 39(??):S–24– S–??, September 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:YNL

[Ano93-352] Anonymous. Your Next LAN Storage Could Be a Mainframe. Datamation, 39(6):71–??, March 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:TWC

[Ano94a] Anonymous. A Tough Wireless Communicator. Datamation, 40 (9):67–??, May 01, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:AYD

[Ano94b] Anonymous. Access Your Data Through Pictures. Datamation, 40(2):61–??, January 21, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:ASW

[Ano94c] Anonymous. Alpha Sizzles — But Where’s The Steak? Datamation, 40(9):63–??, May 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PTC

[Ano94d] Anonymous. AS/400 progresses toward client/server OLTP.

Datamation, 40(5):57–??, March 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:AWR

[Ano94e] Anonymous. ATM WANs: Rocket Science Or Lost In Space? Datamation, 40(2):20–??, January 21, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:BTR

[Ano94f] Anonymous. Bass taverns reengineers dog and duck with unisys. Datamation, 40(23):39–40, December 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:Ba

[Ano94g] Anonymous. Benchmarks. Datamation, 40(5):125–??, March 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bb

[Ano94h] Anonymous. Benchmarks. Datamation, 40(6):81–??, March 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bc

[Ano94i] Anonymous. Benchmarks. Datamation, 40(7):79–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bd

[Ano94j] Anonymous. Benchmarks. Datamation, 40(9):82–??, May 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bf

[Ano94k] Anonymous. Benchmarks. Datamation, 40(10):94–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bh

[Ano94l] Anonymous. Benchmarks. Datamation, 40(11):88–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bk

[Ano94m] Anonymous. Benchmarks. Datamation, 40(14):84–??, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bm

[Ano94n] Anonymous. Benchmarks. Datamation, 40(15):70–??, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bp

[Ano94o] Anonymous. Benchmarks. Datamation, 40(17):86–??, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Br

[Ano94p] Anonymous. Benchmarks. Datamation, 40(18):92–??, September 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bt

[Ano94q] Anonymous. Benchmarks. Datamation, 40(19):77–??, October 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bv

[Ano94r] Anonymous. Benchmarks. Datamation, 40(20):82–??, October 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bx

[Ano94s] Anonymous. Benchmarks. Datamation, 40(21):97–??, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bz

[Ano94t] Anonymous. Benchmarks. Datamation, 40(22):97–??, November 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bbb

[Ano94u] Anonymous. Benchmarks. Datamation, 40(23):82–??, December 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bbd

[Ano94v] Anonymous. Benchmarks. Datamation, 40(24):77–??, December 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:BLM

[Ano94w] Anonymous. Build Links To Multiple Databases. Datamation, 40(2):65–??, January 21, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:BYL

[Ano94x] Anonymous. Button Up Your LAN Security. Datamation, 40 (3):59–??, February 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CSF

[Ano94y] Anonymous. Can Suites Fill Groupware Needs? Datamation, 40(1):27–??, January 07, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CRP

[Ano94z] Anonymous. CD-ROM Puts a Spin On Software Distribution. Datamation, 40(4):61–??, February 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:CSG

[Ano94-27] Anonymous. CeBIT ’94 Show Guide. Datamation, 40(5):65–??, March 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:CSE

[Ano94-28] Anonymous. Clean slate equals reduced lead times (outsourcing). Datamation, 40(23):41, December 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CSC

[Ano94-29] Anonymous. Client/Server Costs: Don’t Get Taken For a Ride. Datamation, 40(4):34–??, February 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CSW

[Ano94-30] Anonymous. Client/Server For The Warehouse. Datamation, 40 (6):66–??, March 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:C

[Ano94-31] Anonymous. Commentary. Datamation, 40(3):14–??, February 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CTM

[Ano94-32] Anonymous. Computer: Take a Memo. Datamation, 40(1):51– ??, January 07, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CSSa

[Ano94-33] Anonymous. Corporate Software Scorecard. Datamation, 40(2): 74–??, January 21, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CSSb

[Ano94-34] Anonymous. Corporate Software Scorecard. Datamation, 40(3): 75–??, February 01, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CSSc

[Ano94-35] Anonymous. Corporate Software Scorecard. Datamation, 40(4): 92–??, February 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CSSd

[Ano94-36] Anonymous. Corporate Software Scorecard. Datamation, 40(6): 82–??, March 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CSSe

[Ano94-37] Anonymous. Corporate Software Scorecard. Datamation, 40(8): 92–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CAH

[Ano94-38] Anonymous. C/S Apps Without the Heavy Lifting. Datamation, 40(3):22–??, February 01, 1994. CODEN DTMNAT. ISSN 0011-

Anonymous:1994:CEa

[Ano94-39] Anonymous. Cutting Edge.

Datamation, 40(10):23–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CEb

[Ano94-40] Anonymous. Cutting Edge.

Datamation, 40(11):23–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CEc

[Ano94-41] Anonymous. Cutting Edge.

Datamation, 40(13):11–??, July 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CEd

[Ano94-42] Anonymous. Cutting Edge.

Datamation, 40(14):21–??, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CEe

[Ano94-43] Anonymous. Cutting Edge.

Datamation, 40(15):17–??, August 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:CEf

[Ano94-44] Anonymous. Cutting Edge.

Datamation, 40(16):19–??, August 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:CEg

[Ano94-45] Anonymous. Cutting Edge.

Datamation, 40(17):17–??, September 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:CEh

[Ano94-46] Anonymous. Cutting Edge. Datamation, 40(18):21–??, September 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:CEi

[Ano94-47] Anonymous. Cutting Edge. Datamation, 40(19):19–??, October 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:CEj

[Ano94-48] Anonymous. Cutting Edge. Datamation, 40(20):21–??, October 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:CEk

[Ano94-49] Anonymous. Cutting Edge. Datamation, 40(21):19–??, November 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:CEl

[Ano94-50] Anonymous. Cutting Edge. Datamation, 40(22):17–??, November 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:CEm

[Ano94-51] Anonymous. Cutting Edge. Datamation, 40(23):17–??, December 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:CEn

[Ano94-52] Anonymous. Cutting Edge. Datamation, 40(24):17–??, December 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:DPa

[Ano94-53] Anonymous. Datamation Picks.

Datamation, 40(10):89–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:DPb

[Ano94-54] Anonymous. Datamation Picks.

Datamation, 40(11):82–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:DPc

[Ano94-55] Anonymous. Datamation Picks.

Datamation, 40(13):72–??, July 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:DPd

[Ano94-56] Anonymous. Datamation Picks.

Datamation, 40(14):79–??, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:DPe

[Ano94-57] Anonymous. DATAMATION Picks. Datamation, 40(15):67–??, August 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:DPf

[Ano94-58] Anonymous. DATAMATION Picks. Datamation, 40(16):72–??, August 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:DPg

[Ano94-59] Anonymous. DATAMATION Picks. Datamation, 40(17):82– ??, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:DPh

[Ano94-60] Anonymous. Datamation Picks. Datamation, 40(18):89–??, September 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:DPi

[Ano94-61] Anonymous. DATAMATION Picks. Datamation, 40(19):73– ??, October 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:DPj

[Ano94-62] Anonymous. Datamation picks. Datamation, 40(20):79–??, October 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:DPk

[Ano94-63] Anonymous. Datamation Picks. Datamation, 40(21):89–??, November 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:DPl

[Ano94-64] Anonymous. Datamation Picks. Datamation, 40(22):91–??, November 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:DPm

[Ano94-65] Anonymous. Datamation Picks. Datamation, 40(23):77–??, December 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:DPY

[Ano94-66] Anonymous. Datamation’s 1993 Products Of The Year Winners. Datamation, 40(4):30–??, February 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:DVW

[Ano94-67] Anonymous. Desktop Videoconferencing: What’s Wrong With This Picture? Datamation, 40 (9):28–??, May 01, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:DSA

[Ano94-68] Anonymous. Developing Serious Apps With Notes. Datamation, 40(8):28–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:DDC

[Ano94-69] Anonymous. Distribute Data Without Choking The Net. Datamation, 40(1):35–??, January 07, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:DDA

[Ano94-70] Anonymous. Does Downsizing Add Up For Financials? Datamation, 40(6):51–??, March 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:EGT

[Ano94-71] Anonymous. Ethernet Gets 10 Times Faster For Under US$1,000. Datamation, 40(8): 64–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:FRE

[Ano94-72] Anonymous. Frame Relay Eases SNA Integration. Datamation, 40(4):52–??, February 15, 1994. CODEN DTMNAT. ISSN 0011-

Anonymous:1994:FRL

[Ano94-73] Anonymous. Frame Relay Or Leased Lines — Which Is Cheaper? Datamation, 40(7): 39–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fa

[Ano94-74] Anonymous. Framework. Datamation, 40(1):7–??, January 07, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fb

[Ano94-75] Anonymous. Framework. Datamation, 40(2):8–??, January 21, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fc

[Ano94-76] Anonymous. Framework. Datamation, 40(3):6–??, February 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fd

[Ano94-77] Anonymous. Framework. Datamation, 40(4):8–??, February 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fe

[Ano94-78] Anonymous. Framework. Datamation, 40(5):9–??, March 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Ff

[Ano94-79] Anonymous. Framework. Datamation, 40(6):9–??, March 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fg

[Ano94-80] Anonymous. Framework. Datamation, 40(7):6–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fh

[Ano94-81] Anonymous. Framework. Datamation, 40(8):11–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fi

[Ano94-82] Anonymous. Framework. Datamation, 40(9):13–??, May 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fj

[Ano94-83] Anonymous. Framework. Datamation, 40(10):16–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fl

[Ano94-84] Anonymous. Framework. Datamation, 40(11):15–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fm

[Ano94-85] Anonymous. Framework. Datamation, 40(13):7–??, July 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fn

[Ano94-86] Anonymous. Framework. Datamation, 40(14):11–??, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fo

[Ano94-87] Anonymous. Framework. Datamation, 40(15):11–??, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fp

[Ano94-88] Anonymous. Framework. Datamation, 40(16):11–??, August 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fx

[Ano94-89] Anonymous. Framework. Datamation, 40(17):11–??, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fy

[Ano94-90] Anonymous. Framework. Datamation, 40(18):11–??, September 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fz

[Ano94-91] Anonymous. Framework. Datamation, 40(19):11–??, October 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fba

[Ano94-92] Anonymous. Framework. Datamation, 40(20):11–??, October 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fbb

[Ano94-93] Anonymous. Framework. Datamation, 40(21):11–??, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fbc

[Ano94-94] Anonymous. Framework. Datamation, 40(22):13–??, November 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fbd

[Ano94-95] Anonymous. Framework. Datamation, 40(23):11–??, December 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Fbe

[Ano94-96] Anonymous. Framework. Datamation, 40(24):11–??, December 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:FVa

[Ano94-97] Anonymous. Future Vision.

Datamation, 40(1):76–??, January 07, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:FVb

[Ano94-98] Anonymous. Future Vision.

Datamation, 40(2):80–??, January 21, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:FVc

[Ano94-99] Anonymous. Future Vision.

Datamation, 40(3):80–??, February 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:FVd

[Ano94-100] Anonymous. Future Vision.

Datamation, 40(4):96–??, February 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:FVe

[Ano94-101] Anonymous. Future Vision. Datamation, 40(5):132–??, March 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:FVf

[Ano94-102] Anonymous. Future Vision. Datamation, 40(6):88–??, March 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:FVg

[Ano94-103] Anonymous. Future Vision. Datamation, 40(7):84–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:FVh

[Ano94-104] Anonymous. Future Vision. Datamation, 40(8):96–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:FVi

[Ano94-105] Anonymous. Future Vision. Datamation, 40(9):88–??, May 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:GIT

[Ano94-106] Anonymous. Get It Together With Visual Basic. Datamation, 40(2):45–??, January 21, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:GYC

[Ano94-107] Anonymous. Get Your CICS On Anyone’s UNIX? Datamation, 40(4):45–??, February 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:GLD

[Ano94-108] Anonymous. Global LAN Directories: The Pain, The Payoff. Datamation, 40(7):30–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:HHQ

[Ano94-109] Anonymous. Here’s How To Quantify IT Investment Benefits. Datamation, 40(1):45–??, January 07, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:HMH

[Ano94-110] Anonymous. How Much Help Is Enough? Datamation, 40(7): 49–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:HGM

[Ano94-111] Anonymous. How to Get More Value From EDI. Datamation, 40(9):56–??, May 01, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:HRC

[Ano94-112] Anonymous. Hub/Router Combos: Don’t Overwork Them. Datamation, 40(5):42–??, March 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Be

[Ano94-113] Anonymous. In Box. Datamation, 40(10):19–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bg

[Ano94-114] Anonymous. In Box. Datamation, 40(11):17–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bi

[Ano94-115] Anonymous. In Box. Datamation, 40(13):9–??, July 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bj

[Ano94-116] Anonymous. In Box. Datamation, 40(14):15–??, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bl

[Ano94-117] Anonymous. In Box. Datamation, 40(15):15–??, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bn

[Ano94-118] Anonymous. In Box. Datamation, 40(16):13–??, August 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bo

[Ano94-119] Anonymous. In Box. Datamation, 40(17):15–??, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bq

[Ano94-120] Anonymous. In Box. Datamation, 40(18):15–??, September 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bs

[Ano94-121] Anonymous. In Box. Datamation, 40(19):13–??, October 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bu

[Ano94-122] Anonymous. In Box. Datamation, 40(20):15–??, October 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bw

[Ano94-123] Anonymous. In Box. Datamation, 40(21):13–??, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:By

[Ano94-124] Anonymous. In Box. Datamation, 40(22):15–??, November 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bba

[Ano94-125] Anonymous. In Box. Datamation, 40(23):12–??, December 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Bbc

[Ano94-126] Anonymous. In Box. Datamation, 40(24):13–??, December 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:RCC

[Ano94-127] Anonymous. IS Rises To The Competitiveness Challenge. Datamation, 40(1):16–??, January 07, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:PDY

[Ano94-128] Anonymous. It’s 5:00 PM: Do You Know Where All Your PCs Are? Datamation, 40(6):56–??, March 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:LHF

[Ano94-129] Anonymous. LAN Hub’s Future Is In Switching. Datamation, 40 (1):55–??, January 07, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lj

[Ano94-130] Anonymous. LANscape. Datamation, 40(10):40–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Ll

[Ano94-131] Anonymous. LANscape. Datamation, 40(11):36–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lm

[Ano94-132] Anonymous. LANScape. Datamation, 40(13):22–??, July 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Ln

[Ano94-133] Anonymous. LANScape. Datamation, 40(14):32–??, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lo

[Ano94-134] Anonymous. LANscape. Datamation, 40(15):28–??, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lp

[Ano94-135] Anonymous. LANscape. Datamation, 40(16):34–??, August 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lx

[Ano94-136] Anonymous. LanScape. Datamation, 40(17):36–??, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Ly

[Ano94-137] Anonymous. LANScape. Datamation, 40(18):40–??, September 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lz

[Ano94-138] Anonymous. LANScape. Datamation, 40(19):34–??, October 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lba

[Ano94-139] Anonymous. LANScape. Datamation, 40(20):38–??, October 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lbb

[Ano94-140] Anonymous. LANScape. Datamation, 40(21):36–??, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lbc

[Ano94-141] Anonymous. LANScape. Datamation, 40(22):34–??, November 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lbd

[Ano94-142] Anonymous. LANScape. Datamation, 40(23):35–??, December 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lbe

[Ano94-143] Anonymous. LANScape. Datamation, 40(24):34–??, December 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:La

[Ano94-144] Anonymous. Letters. Datamation, 40(1):10–??, January 07, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lb

[Ano94-145] Anonymous. Letters. Datamation, 40(2):12–??, January 21, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lc

[Ano94-146] Anonymous. Letters. Datamation, 40(3):10–??, February 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Ld

[Ano94-147] Anonymous. Letters. Datamation, 40(4):12–??, February 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Le

[Ano94-148] Anonymous. Letters. Datamation, 40(5):10–??, March 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lf

[Ano94-149] Anonymous. Letters. Datamation, 40(6):12–??, March 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lg

[Ano94-150] Anonymous. Letters. Datamation, 40(7):11–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Lh

[Ano94-151] Anonymous. Letters. Datamation, 40(8):14–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Li

[Ano94-152] Anonymous. Letters. Datamation, 40(9):14–??, May 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:LAb

[Ano94-153] Anonymous. Look Ahead. Datamation, 40(1):13–??, January 07, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:LAc

[Ano94-154] Anonymous. Look Ahead. Datamation, 40(2):17–??, January 21, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:LAd

[Ano94-155] Anonymous. Look Ahead. Datamation, 40(3):19–??, February 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:LAe

[Ano94-156] Anonymous. Look Ahead. Datamation, 40(4):15–??, February 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:LAf

[Ano94-157] Anonymous. Look Ahead. Datamation, 40(5):13–??, March 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:LAg

[Ano94-158] Anonymous. Look Ahead. Datamation, 40(6):15–??, March 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:LAh

[Ano94-159] Anonymous. Look Ahead. Datamation, 40(7):15–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:LAi

[Ano94-160] Anonymous. Look Ahead. Datamation, 40(8):19–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:LAj

[Ano94-161] Anonymous. Look Ahead. Datamation, 40(9):19–??, May 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:LMN

[Ano94-162] Anonymous. Look Ma, No Wires. Datamation, 40(8):72–??, April 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:LCC

[Ano94-163] Anonymous. Low-Cost CD-ROMs — Make Them Yourself! Datamation, 40(7):45–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:MRa

[Ano94-164] Anonymous. Media Report.

Datamation, 40(5):21–??, March 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:MRb

[Ano94-165] Anonymous. Media Report.

Datamation, 40(6):21–??, March 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:MRc

[Ano94-166] Anonymous. Media Report.

Datamation, 40(7):21–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:MRd

[Ano94-167] Anonymous. Media Report.

Datamation, 40(8):23–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:MRe

[Ano94-168] Anonymous. Media Report.

Datamation, 40(9):23–??, May 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:MTN

[Ano94-169] Anonymous. Microsoft Tunes NT For Daytona. Datamation, 40(9):39–??, May 01, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:MAW

[Ano94-170] Anonymous. Multidimensional Analysis: Winning The Competitive Game. Datamation, 40(4): 21–??, February 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:MM

[Ano94-171] Anonymous. Multimedia For The Masses. Datamation, 40(4): 73–??, February 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:NSS

[Ano94-172] Anonymous. NetWare Storage: Starting To Be Manageable. Datamation, 40(8):54–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:NPa

[Ano94-173] Anonymous. New Products. Datamation, 40(1):61–??, January 07, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:NPb

[Ano94-174] Anonymous. New Products. Datamation, 40(2):69–??, January 21, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:NPc

[Ano94-175] Anonymous. New Products. Datamation, 40(3):67–??, February 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:NPd

[Ano94-176] Anonymous. New Products. Datamation, 40(4):85–??, February 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:NPe

[Ano94-177] Anonymous. New Products. Datamation, 40(5):115–??, March 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:NPf

[Ano94-178] Anonymous. New Products.

Datamation, 40(6):73–??, March 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:NPg

[Ano94-179] Anonymous. New Products.

Datamation, 40(7):68–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:NPh

[Ano94-180] Anonymous. New Products.

Datamation, 40(8):81–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:NPS

[Ano94-181] Anonymous. New Products:

SunSelect’s SolarNet network operating system alternative. Datamation, 40(9):76–??, May 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:NAS

[Ano94-182] Anonymous. NT Advanced Server: Not A NetWare Killer. Datamation, 40(1):30–??, January 07, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:ODA

[Ano94-183] Anonymous. Open Documents Across The Enterprise. Datamation, 40(6):70–??, March 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:OUR

[Ano94-184] Anonymous. Outfitting The Ultimate Road Warrior. Datamation, 40(5):37–??, March 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:OST

[Ano94-185] Anonymous. Outsourcing As a Survival Tactic. Datamation, 40 (8):48–??, April 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PSa

[Ano94-186] Anonymous. Paradigm Shift. Datamation, 40(10):37–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PSc

[Ano94-187] Anonymous. Paradigm Shift. Datamation, 40(11):32–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PSd

[Ano94-188] Anonymous. Paradigm Shift. Datamation, 40(13):19–??, July 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PSe

[Ano94-189] Anonymous. Paradigm Shift. Datamation, 40(14):29–??, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PSf

[Ano94-190] Anonymous. Paradigm Shift. Datamation, 40(15):23–??, August 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:PSg

[Ano94-191] Anonymous. Paradigm Shift. Datamation, 40(16):29–??, August 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:PSh

[Ano94-192] Anonymous. Paradigm Shift.

Datamation, 40(17):31–??, September 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:PSi

[Ano94-193] Anonymous. Paradigm Shift.

Datamation, 40(18):31–??, September 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:PSj

[Ano94-194] Anonymous. Paradigm Shift.

Datamation, 40(19):29–??, October 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:PSk

[Ano94-195] Anonymous. Paradigm Shift.

Datamation, 40(20):31–??, October 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:PSl

[Ano94-196] Anonymous. Paradigm Shift.

Datamation, 40(21):31–??, November 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:PSm

[Ano94-197] Anonymous. Paradigm Shift.

Datamation, 40(22):27–??, November 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:PSn

[Ano94-198] Anonymous. Paradigm Shift.

Datamation, 40(23):29–??, December 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:PSo

[Ano94-199] Anonymous. Paradigm Shift. Datamation, 40(24):29–??, December 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:PSB

[Ano94-200] Anonymous. PCI: Server Bus For The Fast LANe. Datamation, 40 (6):61–??, March 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PRT

[Ano94-201] Anonymous. Pick The RAID That’s Right For You. Datamation, 40(4):79–??, February 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWa

[Ano94-202] Anonymous. Press Watch. Datamation, 40(10):95–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWb

[Ano94-203] Anonymous. Press Watch. Datamation, 40(11):89–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWc

[Ano94-204] Anonymous. Press Watch. Datamation, 40(13):75–??, July 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWd

[Ano94-205] Anonymous. Press Watch. Datamation, 40(14):85–??, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWe

[Ano94-206] Anonymous. Press Watch. Datamation, 40(15):71–??, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWf

[Ano94-207] Anonymous. Press Watch. Datamation, 40(16):77–??, August 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWg

[Ano94-208] Anonymous. Press Watch. Datamation, 40(17):87–??, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWh

[Ano94-209] Anonymous. Press Watch. Datamation, 40(18):95–??, September 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWi

[Ano94-210] Anonymous. Press Watch. Datamation, 40(19):79–??, October 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWj

[Ano94-211] Anonymous. Press Watch. Datamation, 40(20):83–??, October 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWk

[Ano94-212] Anonymous. Press Watch. Datamation, 40(21):99–??, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWl

[Ano94-213] Anonymous. Press Watch. Datamation, 40(22):99–??, November 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWm

[Ano94-214] Anonymous. Press Watch. Datamation, 40(24):79–??, December 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWR

[Ano94-215] Anonymous. Press Watch, Remember where you were when JFK was shot? You may be too old for IS? Datamation, 40 (23):85–??, December 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PWO

[Ano94-216] Anonymous. Project 4.0 Works In the Office. Datamation, 40(6): 64–??, March 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PMG

[Ano94-217] Anonymous. Project Management Goes Client/Server. Datamation, 40(5):109–??, March 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:QC

[Ano94-218] Anonymous. Quest contest. Datamation, 40(16):16–??, August 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:RTD

[Ano94-219] Anonymous. Real-Time Data Under Windows. Datamation, 40(4):74–??, February 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:RCW

[Ano94-220] Anonymous. Reality Check:

Wireless, Pay More. Datamation, 40(8):60–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SS

[Ano94-221] Anonymous. Sequent Scales Up.

Datamation, 40(8):71–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SVa

[Ano94-222] Anonymous. Services and Virtues. Datamation, 40(10): 38–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SVb

[Ano94-223] Anonymous. SerVices and Virtues. Datamation, 40(11): 34–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SVc

[Ano94-224] Anonymous. SerVices and Virtues. Datamation, 40(13):20– ??, July 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:SVd

[Ano94-225] Anonymous. SerVices and Virtues. Datamation, 40(14):30– ??, July 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:SVe

[Ano94-226] Anonymous. SerVices and Virtues. Datamation, 40(15):26– ??, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SVf

[Ano94-227] Anonymous. SerVices and Virtues. Datamation, 40(16):32– ??, August 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SVg

[Ano94-228] Anonymous. SerVices and Virtues. Datamation, 40(17):34– ??, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SVh

[Ano94-229] Anonymous. SerVices and Virtues. Datamation, 40(18):36– ??, September 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SVi

[Ano94-230] Anonymous. SerVices and Virtues. Datamation, 40(19):32– ??, October 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SVj

[Ano94-231] Anonymous. SerVices and Virtues. Datamation, 40(20):36– ??, October 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SVk

[Ano94-232] Anonymous. SerVices and Virtues. Datamation, 40(21):34– ??, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SVl

[Ano94-233] Anonymous. SerVices and Virtues. Datamation, 40(22):32– ??, November 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SVm

[Ano94-234] Anonymous. SerVices and Virtues. Datamation, 40(23):33– ??, December 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SVn

[Ano94-235] Anonymous. SerVices and Virtues. Datamation, 40(24):32– ??, December 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SRa

[Ano94-236] Anonymous. Show Report. Datamation, 40(8):79–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SRc

[Ano94-237] Anonymous. Show Report. Datamation, 40(21):87–??, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SSW

[Ano94-238] Anonymous. SMP Server For Windows NT. Datamation, 40 (5):105–??, March 01, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SSS

[Ano94-239] Anonymous. SMP Soars With Some More Pentiums. Datamation, 40(1):41–??, January 07, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SC

[Ano94-240] Anonymous. Snap-On Code.

Datamation, 40(3):63–??, February 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:SUM

[Ano94-241] Anonymous. Standard UNIX Management: What’s The Holdup? Datamation, 40(4):67– ??, February 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SBT

[Ano94-242] Anonymous. Successful Benchmarking With The Team Approach. Datamation, 40(3):53– ??, February 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:ST

[Ano94-243] Anonymous. Surprise! OS/2 is taking off. Datamation, 40(8): 43–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:SMA

[Ano94-244] Anonymous. System Management Across The Enterprise. Datamation, 40(8):74–??, April 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:TCR

[Ano94-245] Anonymous. Take Control Of Remote Access. Datamation, 40 (7):62–??, April 01, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:TCS

[Ano94-246] Anonymous. Take Control Of Support. Datamation, 40(2):62– ??, January 21, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:TTa

[Ano94-247] Anonymous. Technology Trends. Datamation, 40(5):126–??, March 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:TTb

[Ano94-248] Anonymous. Technology Trends.

Datamation, 40(7):80–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:TTc

[Ano94-249] Anonymous. Technology Trends.

Datamation, 40(9):84–??, May 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:BCS

[Ano94-250] Anonymous. The Best In Client/ Server Computing. Datamation, 40(5):26–??, March 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:CEW

[Ano94-251] Anonymous. The Cheap, Easy Way To Harness Imaging’s Power. Datamation, 40(2):55– ??, January 21, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:VRS

[Ano94-252] Anonymous. The Virtual Reality Of Switched Networks. Datamation, 40(9):34–??, May 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Ta

[Ano94-253] Anonymous. ThinkWrap. Datamation, 40(10):100–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Tb

[Ano94-254] Anonymous. ThinkWrap. Datamation, 40(11):92–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Tc

[Ano94-255] Anonymous. ThinkWrap. Datamation, 40(13):80–??, July 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Td

[Ano94-256] Anonymous. ThinkWrap. Datamation, 40(14):88–??, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Te

[Ano94-257] Anonymous. ThinkWrap. Datamation, 40(15):76–??, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Tf

[Ano94-258] Anonymous. ThinkWrap. Datamation, 40(16):80–??, August 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Tg

[Ano94-259] Anonymous. ThinkWrap. Datamation, 40(17):92–??, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Th

[Ano94-260] Anonymous. ThinkWrap. Datamation, 40(18):100–??, September 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:Ti

[Ano94-261] Anonymous. ThinkWrap. Datamation, 40(19):84–??, October 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Tj

[Ano94-262] Anonymous. ThinkWrap. Datamation, 40(20):90–??, October 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Tk

[Ano94-263] Anonymous. ThinkWrap. Datamation, 40(21):104–??, November 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:Tl

[Ano94-264] Anonymous. ThinkWrap. Datamation, 40(22):104–??, November 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:Tm

[Ano94-265] Anonymous. ThinkWrap. Datamation, 40(23):92–??, December 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:Tn

[Ano94-266] Anonymous. ThinkWrap. Datamation, 40(24):84–??, December 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:TNC

[Ano94-267] Anonymous. Third NetView Not The Charm. Datamation, 40(7): 55–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:TVH

[Ano94-268] Anonymous. Three vendors help ryder reengineer refueling. Datamation, 40(23):43–44, December 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:PCR

[Ano94-269] Anonymous. TO: mis@corporate. com RE: Internet E-Mail. Datamation, 40(3):39–??, February 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:TRA

[Ano94-270] Anonymous. Tools For Remote Access. Datamation, 40(7):66– ??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:TPU

[Ano94-271] Anonymous. Turn Power Users Into Power Savers. Datamation, 40(9):53–??, May 01, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:UFW

[Ano94-272] Anonymous. Uncertain Future: What’s The Next Step After Client/Server? Datamation, 40 (6):26–??, March 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:UPG

[Ano94-273] Anonymous. Union Pacific gets maintenance on track. Datamation, 40(23):42–43, December 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:UB

[Ano94-274] Anonymous. UniTree Branches Out. Datamation, 40(9):66–??, May 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:UTR

[Ano94-275] Anonymous. UNIX Takes To The Road. Datamation, 40(7):59–??, April 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1994:UGS

[Ano94-276] Anonymous. UPSs Get Smarter.

Datamation, 40(9):69–??, May 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:WHA

[Ano94-277] Anonymous. WANs: Hot To Add Branches Without Going Bankrupt. Datamation, 40(6): 42–??, March 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:WDM

[Ano94-278] Anonymous. Want To Do Multimedia Apps? Here Are The Tools: Multimedia development is no longer an obscure craft practiced by graphics arts specialists. Tools abound for typical business use. One promising application is performance support — just-in-time training for employees who have to rapidly adopt new work methods or learn about new products. Datamation, 40(9):45–??, May 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:WMC

[Ano94-279] Anonymous. Wanted: The Multithreaded CIO. Datamation, 40 (8):34–??, April 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:WBS

[Ano94-280] Anonymous. Wedding Bells Sound For Objects and Relational Data. Datamation, 40(5):

49–??, March 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:WVO

[Ano94-281] Anonymous. Welcome To The Virtual IS Organization. Datamation, 40(3):32–??, February 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:WWPc

[Ano94-282] Anonymous. What’s wrong with this picture? (videoconferencing). Datamation, 40(9):28–32, May 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:WWPa

[Ano94-283] Anonymous. When will PCM-CIA be plug and play? Datamation, 40(2):39–??, January 21, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:WDC

[Ano94-284] Anonymous. Where do C/S apps go wrong? Datamation, 40(2): 30–??, January 21, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:WUH

[Ano94-285] Anonymous. Where’s UNIX Headed? Datamation, 40(7): 24–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:WNW

[Ano94-286] Anonymous. Windows NT: Where Are The Tools? Datamation, 40(3):45–??, February 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:WTN

[Ano94-287] Anonymous. Workflow Through Notes. Datamation, 40(3):60– ??, February 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:AIS

[Ano95a] Anonymous. Advertisers Index/ Sales masthead. Datamation, 41 (14):83–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:Bb

[Ano95b] Anonymous. Benchmarks. Datamation, 41(1):82–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Bc

[Ano95c] Anonymous. Benchmarks. Datamation, 41(2):78–??, February 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Bg

[Ano95d] Anonymous. Benchmarks. Datamation, 41(3):82–??, February 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Bi

[Ano95e] Anonymous. Benchmarks. Datamation, 41(4):70–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Bk

[Ano95f] Anonymous. Benchmarks. Datamation, 41(5):77–??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Bm

[Ano95g] Anonymous. Benchmarks. Datamation, 41(6):70–??, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEa

[Ano95h] Anonymous. Cutting Edge. Datamation, 41(1):17–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEba

[Ano95i]  Anonymous.  Cutting Edge.

Datamation, 41(2):17–??, February 01, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:CEca

[Ano95j]  Anonymous.  Cutting Edge.

Datamation, 41(3):17–??, February 15, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:CEfa

[Ano95k] Anonymous. Cutting Edge. Datamation, 41(4):19–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEga

[Ano95l]  Anonymous.  Cutting Edge.

Datamation, 41(5):17–??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEh

[Ano95m] Anonymous. Cutting Edge. Datamation, 41(6):15–??, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEia

[Ano95n] Anonymous. Cutting Edge.

Datamation, 41(8):15–??, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEj

[Ano95o] Anonymous. Cutting Edge.

Datamation, 41(9):15–??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEk

[Ano95p] Anonymous. Cutting Edge.

Datamation, 41(10):17–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEla

[Ano95q] Anonymous. Cutting Edge.

Datamation, 41(11):15–??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEB

[Ano95r] Anonymous. Cutting Edge — Believe it or not, Microsoft has put its name on a product created by Computer Associates and has started selling it. Datamation, 41(11):15–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEC

[Ano95s] Anonymous. Cutting Edge — Cyber crooks are as busy as ever, and their victims are talking about it. Datamation, 41(23): 17–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEF

[Ano95t] Anonymous. Cutting Edge — Five years from now, 50% of all methodology gums will be publishing books on objects. Datamation, 41(15):17–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEG

[Ano95u] Anonymous. Cutting Edge — Going off to visit the engineer down the hall? Here’s a handy EngineerSpeak translation guide. Datamation, 41(17):15–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEIb

[Ano95v] Anonymous. Cutting edge — IBM’s new I/T Security Consulting Practice features emergency response to hacker intension. Datamation, 41(14):15–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEIc

[Ano95w] Anonymous. Cutting Edge — It will be business as usual when Win95 hits the streets, at least as far as virus infections go. Datamation, 41(16):15–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEL

[Ano95x] Anonymous. Cutting Edge — Looking for a good place to hold ATM trials? Try Finland. Really. Datamation, 41(21):15–??, 1995. CODEN DTMNAT. ISSN 0011-

Anonymous:1995:CEM

[Ano95y] Anonymous. Cutting Edge — Maybe there aren’t 50, but we can give you five ways to leave your mainframe. Datamation, 41 (10):17–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:CEN

[Ano95z] Anonymous. Cutting Edge — Need a few tips on how to negotiate a large-volume purchase of Microsoft’s Win95? Datamation, 41(19):21–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEO

[Ano95-27] Anonymous. Cutting Edge — Our reporter is still scratching his head over the stock market frenzy on Netscape’s IPO. Datamation, 41(18):17–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEP

[Ano95-28] Anonymous. Cutting Edge — Pricing mistakes by leading outsourcing companies have resulted in good deals for the IS shops. Datamation, 41(20):15–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CES

[Ano95-29] Anonymous. Cutting Edge — Setting up and maintaining World Wide Web sites for corporate customers is a booming business. Datamation, 41(13):15–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:CEY

[Ano95-30] Anonymous. Cutting Edge — Your enterprise ain’t up to speed without a centralized directory of all your nodes and their privileges, connections, and even content. Datamation, 41(12):15–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:DCIb

[Ano95-31] Anonymous. Datamation 100 Company Index. Datamation, 41 (10):5–??, June 01, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:DCIa

[Ano95-32] Anonymous. Datamation 100 Company Index — Alphabetical listing of profiles for top 100 North Amencan IT vendors. Datamation, 41(10):5–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:DPIb

[Ano95-33] Anonymous. Datamation 100 Product Index. Datamation, 41 (10):107–??, June 01, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:DPIa

[Ano95-34] Anonymous. Datamation 100 Product Index — Comprehensive listing of all products mentioned in this issue. Datamation, 41(10): 107–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:DGa

[Ano95-35] Anonymous. DATAMATION Global 100. Datamation, 41(10):

45–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:DGb

[Ano95-36] Anonymous. DATAMATION Global 100. Datamation, 41(10): 45–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:DNAa

[Ano95-37] Anonymous. DATAMATION North American 100. Datamation, 41(10):44–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:DNAb

[Ano95-38] Anonymous. DATAMATION North American 100. Datamation, 41(10):44–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:DPa

[Ano95-39] Anonymous. Datamation Picks.

Datamation, 41(1):79–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:DPb

[Ano95-40] Anonymous. Datamation Picks.

Datamation, 41(2):75–??, February 01, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:DPc

[Ano95-41] Anonymous. Datamation Picks.

Datamation, 41(3):75–??, February 15, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:DPd

[Ano95-42] Anonymous. Datamation Picks.

Datamation, 41(4):67–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:DPe

[Ano95-43] Anonymous. DATAMATION Picks. Datamation, 41(5):73–??, March 15, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:DPf

[Ano95-44] Anonymous. Datamation Picks. Datamation, 41(6):67–??, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:DPg

[Ano95-45] Anonymous. Datamation Picks. Datamation, 41(8):75–??, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:DPh

[Ano95-46] Anonymous. Datamation Picks. Datamation, 41(9):85–??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:DPT

[Ano95-47] Anonymous. DATAMATION Picks — Toshiba introduces a 4.8-pound multimedia notebook based on a 2.9V 90MHz Pentium chip; Micro Focus unveils Bridge-Ware. Datamation, 41(12):74–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Fa

[Ano95-48] Anonymous. Framework. Datamation, 41(1):11–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Fb

[Ano95-49] Anonymous. Framework. Datamation, 41(2):11–??, February 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Fd

[Ano95-50] Anonymous. Framework. Datamation, 41(3):11–??, February 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Fe

[Ano95-51] Anonymous. Framework. Datamation, 41(4):11–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Ff

[Ano95-52] Anonymous. Framework. Datamation, 41(5):11–??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Fg

[Ano95-53] Anonymous. Framework. Datamation, 41(6):8–??, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Fh

[Ano95-54] Anonymous. Framework. Datamation, 41(8):9–??, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Fj

[Ano95-55] Anonymous. Framework. Datamation, 41(10):7–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:FBL

[Ano95-56] Anonymous. FRAMEWORK — Being late adopting ATM may be as big a mistake as being late adopting PCs. Datamation, 41 (11):7–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:FUC

[Ano95-57] Anonymous. Framework — Unicycles are cool. but two-wheelers will take you where you want to go. Datamation, 41(14):7–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:FWH

[Ano95-58] Anonymous. Framework — What happens when Vision meets Reality head on? Datamation, 41(12):7–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:FWL

[Ano95-59] Anonymous. Framework — What’s the likelihood of Microsoft pulling off its enterprisecomputing scheme? Datamation, 41(10):7–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Fi

[Ano95-60] Anonymous. Framwork. Datamation, 41(9):7–??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Ba

[Ano95-61] Anonymous. In Box. Datamation, 41(1):15–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Bd

[Ano95-62] Anonymous. In Box. Datamation, 41(3):13–??, February 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Bh

[Ano95-63] Anonymous. In Box. Datamation, 41(4):15–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Bj

[Ano95-64] Anonymous. In Box. Datamation, 41(5):13–??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Bl

[Ano95-65] Anonymous. In Box. Datamation, 41(6):11–??, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Bn

[Ano95-66] Anonymous. In Box. Datamation, 41(8):10–??, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Bo

[Ano95-67] Anonymous. In Box. Datamation, 41(10):9–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Bp

[Ano95-68] Anonymous. In Box. Datamation, 41(11):10–??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:BSW

[Ano95-69] Anonymous. In Box — a shrinkwrapped alternative to custom salesforce automation systems. Datamation, 41(11):10–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:BEL

[Ano95-70] Anonymous. In Box — Eli Lilly has not just missed the boat, it’s hopped on a covered wagon heading in the other direction. Datamation, 41(20):13–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:BEY

[Ano95-71] Anonymous. In Box — Embattled Year 2000 letter writer clarifies his position. Datamation, 41 (19):12–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:BNR

[Ano95-72] Anonymous. In Box — Is Notes ROI Getting Better? Datamation, 41(10):9–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:BMS

[Ano95-73] Anonymous. In Box — More security is the order of the day. Datamation, 41(15):11–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:BPD

[Ano95-74] Anonymous. In Box — Protecting data is the real value of hierarchical storage management. Datamation, 41(17):9–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:BRW

[Ano95-75] Anonymous. In Box — Readers want the real inside dope about Bridge Ware. Datamation, 41 (12):9–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:BSE

[Ano95-76] Anonymous. In Box — Should everyone run their business the Andersen way? Datamation, 41 (14):11–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:BTM

[Ano95-77] Anonymous. In Box — There are multitudes of commercial class libraries that provide solid application development frameworks. Datamation, 41(18):11–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:BTT

[Ano95-78] Anonymous. In Box — Think there’s no such thing as IT-savvy MBAs? Check out the University of California at Irvine. Datamation, 41(23):14–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:BTY

[Ano95-79] Anonymous. In Box — Think you can breeze into the Year 2000 unscathed? Read on. Datamation, 41(16):9–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:BWP

[Ano95-80] Anonymous. In Box — What’s the point about the turn-of-the-century data calculation problem? Datamation, 41(13):9–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:BWD

[Ano95-81] Anonymous. In Box — Windows95 defenders fight back. Datamation, 41(21):11–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:BII

[Ano95-82] Anonymous. In Box: It’s impossible to idiot-proof a system because idiots are so darned ingenious! Datamation, 41(2):13– ??, February 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Ia

[Ano95-83] Anonymous. International. Datamation, 41(1):64–1–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Ib

[Ano95-84] Anonymous. International. Datamation, 41(11):60–1–??, June 15, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:ICV

[Ano95-85] Anonymous. International — Can virtual reality help sell kitchen appliances? Datamation, 41(13):60–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:IFF

[Ano95-86] Anonymous. INTERNATIONAL — This follow-on feature to our DATAMATION 100, focuses on the European IT market and includes profiles of the top 10 IT suppliers. (a special supplement included in International Edition issues, only.). Datamation, 41 (11):60–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:La

[Ano95-87] Anonymous. LANScape. Datamation, 41(1):34–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Lb

[Ano95-88] Anonymous. LANScape. Datamation, 41(2):34–??, February 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Ld

[Ano95-89] Anonymous. LANScape. Datamation, 41(3):35–??, February 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Le

[Ano95-90] Anonymous. LANScape. Datamation, 41(4):34–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Lf

[Ano95-91] Anonymous. LANScape. Datamation, 41(5):34–??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Lg

[Ano95-92] Anonymous. LANScape. Datamation, 41(6):32–??, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Lh

[Ano95-93] Anonymous. LANScape. Datamation, 41(8):34–??, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Li

[Ano95-94] Anonymous. LANScape. Datamation, 41(9):32–??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Lj

[Ano95-95] Anonymous. LANScape. Datamation, 41(10):32–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:LYE

[Ano95-96] Anonymous. LANScape — Are your end-users making up stupid passwords? Help is on the way. Datamation, 41(12):32–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:LSE

[Ano95-97] Anonymous. LANScape — Specialization is the enemy of progress. Datamation, 41(10): 32–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:LSY

[Ano95-98] Anonymous. LanScape — Sure you can lie with statistics. But you can lie with surveys even better. Datamation, 41(14):28–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:LOa

[Ano95-99] Anonymous. Lost in the Ozone. Datamation, 41(5):24–??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:LOc

[Ano95-100] Anonymous. Lost in the Ozone.

Datamation, 41(8):22–??, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:LOd

[Ano95-101] Anonymous. Lost in the Ozone.

Datamation, 41(10):24–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:LOJ

[Ano95-102] Anonymous. Lost in the Ozone — Jonesy builds firewall to tee off his boss. Datamation, 41(10): 24–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PSa

[Ano95-103] Anonymous. Paradigm Shift.

Datamation, 41(1):29–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PSb

[Ano95-104] Anonymous. Paradigm Shift.

Datamation, 41(2):29–??, February 01, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:PSd

[Ano95-105] Anonymous. Paradigm Shift.

Datamation, 41(3):29–??, February 15, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:PSea

[Ano95-106] Anonymous. Paradigm Shift.

Datamation, 41(4):31–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PSf

[Ano95-107] Anonymous. Paradigm Shift. Datamation, 41(5):29–??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PSg

[Ano95-108] Anonymous. Paradigm Shift. Datamation, 41(6):27–??, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PSh

[Ano95-109] Anonymous. Paradigm Shift. Datamation, 41(8):27–??, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PSi

[Ano95-110] Anonymous. Paradigm Shift. Datamation, 41(9):27–??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PSj

[Ano95-111] Anonymous. Paradigm Shift. Datamation, 41(10):29–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PSE

[Ano95-112] Anonymous. Paradigm shift — Encapsulation, inheritance, and polymorphism are the hobgoblins of small OO purists minds. Datamation, 41(14):25–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PSO

[Ano95-113] Anonymous. Paradigm Shift — Our seers predicted back in January that Lotus would be acquired. Datamation, 41(12):25– ??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PSY

[Ano95-114] Anonymous. Paradigm Shift — You want proof that OLE’s for real? We got it. Datamation, 41 (10):29–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:PWa

[Ano95-115] Anonymous. Press Watch. Datamation, 41(1):83–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWb

[Ano95-116] Anonymous. Press Watch. Datamation, 41(2):81–??, February 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWca

[Ano95-117] Anonymous. Press Watch. Datamation, 41(3):83–??, February 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWf

[Ano95-118] Anonymous. Press Watch. Datamation, 41(4):71–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWga

[Ano95-119] Anonymous. Press Watch. Datamation, 41(5):79–??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWha

[Ano95-120] Anonymous. Press Watch. Datamation, 41(6):71–??, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWia

[Ano95-121] Anonymous. Press Watch. Datamation, 41(8):79–??, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWj

[Ano95-122] Anonymous. Press Watch. Datamation, 41(9):87–??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWk

[Ano95-123] Anonymous. Press Watch. Datamation, 41(10):111–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWl

[Ano95-124] Anonymous. Press Watch. Datamation, 41(11):77–??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWC

[Ano95-125] Anonymous. Press Watch — As a computing element, DNA is slower than silicon, but much more massively parallel. Datamation, 41(10):111–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWIc

[Ano95-126] Anonymous. Press Watch — At MIT, intelligent software agents are becoming even smarter. Datamation, 41(20):89–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWG

[Ano95-127] Anonymous. Press Watch — Get your boards and get ready to surf the fifth wave of computers: public network computing. Datamation, 41(19):99–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWH

[Ano95-128] Anonymous. Press Watch — Holography has been around for years, but now it is being considered for an extremely swift and massive form of storage. Datamation, 41(23):81–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWWd

[Ano95-129] Anonymous. Press Watch — In a world of Windows clients, NetWare departmental servers, and the looming presence of NT, what’s the role of UNIX. Datamation, 41(21):94–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWPa

[Ano95-130] Anonymous. Press Watch — Is the Pentagon holdingback on the release of an experimental missile that has a cockroach at the controls? Datamation, 41(11): 77–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWN

[Ano95-131] Anonymous. Press Watch — Novell has to expand NetWare’s support for apps or succumb to Windows NTs attacks. Datamation, 41(12):77–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWPb

[Ano95-132] Anonymous. Press Watch — Putting monetary values on person-hours in the IT department helps when talking bottomline with upper management. Datamation, 41(18):93–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWS

[Ano95-133] Anonymous. Press Watch — Some people really do know how to get a C/S development team to work right. Datamation, 41 (13):77–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:PWIb

[Ano95-134] Anonymous. Press Watch — The Internet may run out of room soon, unless the next-generation Internet Protocol gets off the ground. Datamation, 41(15):75– ??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWWa

[Ano95-135] Anonymous. Press Watch — The Web is at least one man’s answer to alienation in this society. Datamation, 41(14):77–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWWc

[Ano95-136] Anonymous. Press Watch — Watch out all you impulse buyers. LaserVend is strategically placing software-vending machines in your path. And soon you won’t even need exact change. Datamation, 41(17):97– ??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PWWb

[Ano95-137] Anonymous. Press Watch — Windows95 or Windows’NT? It’s Microsoft vs Microsoft. Datamation, 41(16):79–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PYW

[Ano95-138] Anonymous. Products of the Year — What was the best IS product of the year? Tell us. Tell the world. Cast your ballot in DATAMATION’s 1995 Products of the Year contest. Datamation, 41(23):16–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:SVa

[Ano95-139] Anonymous. SerVices and Virtues. Datamation, 41(1): 32–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:SVb

[Ano95-140] Anonymous. SerVices and Virtues. Datamation, 41(2):32– ??, February 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:SVd

[Ano95-141] Anonymous. SerVices and Virtues. Datamation, 41(3):33– ??, February 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:SVe

[Ano95-142] Anonymous. SerVices and Virtues. Datamation, 41(4):32– ??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:SVf

[Ano95-143] Anonymous. SerVices and Virtues. Datamation, 41(5):32– ??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:SVg

[Ano95-144] Anonymous. SerVices and Virtues. Datamation, 41(6):30– ??, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:SVh

[Ano95-145] Anonymous. SerVices and Virtues. Datamation, 41(8):32– ??, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:SVi

[Ano95-146] Anonymous. SerVices and Virtues. Datamation, 41(9):30– ??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:SVj

[Ano95-147] Anonymous. Services and Virtues. Datamation, 41(10): 37–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:SVS

[Ano95-148] Anonymous. SerVices and Virtues — Sure, everyone loves to hates Andersen Consulting It’s still the best. Datamation, 41 (10):37–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:SVW

[Ano95-149] Anonymous. SerVices and Virtues — When no news is bad news, bad news is good news. Make sense? No, but this is ISSC. Datamation, 41(12):28–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Ta

[Ano95-150] Anonymous. ThinkWrap. Datamation, 41(1):88–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Tb

[Ano95-151] Anonymous. ThinkWrap. Datamation, 41(2):88–??, February 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Td

[Ano95-152] Anonymous. ThinkWrap. Datamation, 41(3):88–??, February 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Te

[Ano95-153] Anonymous. ThinkWrap. Datamation, 41(4):76–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Tf

[Ano95-154] Anonymous. ThinkWrap. Datamation, 41(5):84–??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Tg

[Ano95-155] Anonymous. ThinkWrap. Datamation, 41(6):76–??, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Th

[Ano95-156] Anonymous. ThinkWrap. Datamation, 41(8):84–??, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Ti

[Ano95-157] Anonymous. ThinkWrap. Datamation, 41(9):92–??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:Tj

[Ano95-158] Anonymous. ThinkWrap. Datamation, 41(10):116–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:TDO

[Ano95-159] Anonymous. ThinkWrap — Dealing with an outsourcer? Here’s what you need to put in the contract. Datamation, 41 (14):84–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1995:TIS

[Ano95-160] Anonymous. ThinkWrap — IBM sets its sights on client/server database market. Datamation, 41(10):116–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:TRS

[Ano95-161] Anonymous. ThinkWrap — In reengineering, speed is everything. Datamation, 41(12):84–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:TWC

[Ano95-162] Anonymous. ThinkWrap — Why chose between quick-and-dirty RAD and slower-than-death CASE? Take the best of both. Datamation, 41(16):84–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:ADE

[Ano96a] Anonymous. Application development environments for enterprise projects — Datamation’s feature summary of enterprise app dev environments from leading vendors. Datamation, 42(13): 106–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:AWT

[Ano96b] Anonymous. AS/400: Around the world of the AS/400 — there’s a vast amount of AS/ 400 information literally at your fingertips if you know where to look. recently, we hopped on the World Wide Web to take a look at the growing number of Web sites offering AS/400 administrators information on overseeing their systems. Datamation, 42 (2):73–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:BA

[Ano96c] Anonymous. Business advantage.

Datamation, 42(18):10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:BAA

[Ano96d] Anonymous. Business advantage — Anna Jaques Hospital automates its software testing. Indiana U. uses ATM link. Datamation, 42(15):16–??, ???? 1996.

CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CSI

[Ano96e] Anonymous. Case study: The Internet shopping network. Datamation, 42(7):40–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CSS

[Ano96f] Anonymous. Client/server: Surf C/S systems via the Intranet — thanks to a whole series of new tools-and with more on the waydevelopers are now building Web pages that include gee-whiz enhancements such as 3-D virtual reality, animation, and real-time interactivity based on the user’s profile. Datamation, 42(2):67–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CEBa

[Ano96g] Anonymous. Cutting edge — business advantage: Fruit of the Loom helps distributors sell on line. NatWest Bank slashing app dev costs. Datamation, 42(13): 10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CEF

[Ano96h] Anonymous. Cutting edge — fast Ethernet, OLAP on the Web, datawarehousing oneupmanship, and more. Datamation, 42(7):13–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CEG

[Ano96i] Anonymous. Cutting edge — get ready for the latest step in WAN bandwidth, the cable TV modem. Datamation, 42(5):14– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1996:CEIa

[Ano96j] Anonymous. Cutting edge — IBI steps into the future with a new breed of middleware. Datamation, 42(3):14–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CEIb

[Ano96k] Anonymous. Cutting edge — Internet stations may make sense for kids, Web junkies, and households, but they don’t make sense in the corporate arena. Datamation, 42(4):13–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CENc

[Ano96l] Anonymous. Cutting edge — news analysis — why is Cisco gobbling up all those networking firms? the honor system for Live Object (OpenDoc) parts. SAP’s R/3 for dummies? Datamation, 42(15):16–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CENb

[Ano96m] Anonymous. Cutting edge — news analysis: Microsoft and Siemens Nixdorf to build object framework. reader survey: Net PCs. Datamation, 42(13):10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CENa

[Ano96n] Anonymous. Cutting edge — NT to run as AS/400 guest?. the Cisco/StrataCom deal. new OLAP benchmark. Web content under Microscope. Merlin takes shape. put some luster on your clusters. A clear path from Unisys. Intermezzo searches several servers. cost-management hel. Datamation, 42(10):11–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CEO

[Ano96o] Anonymous. Cutting edge — object frameworks, more server suites, say goodbye to 1.44MB floppies, and more. Datamation, 42(8):6–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CEPa

[Ano96p] Anonymous. Cutting edge — plug-and-play WANs. another step toward electronic commerce. firewalls to offer viral protection. inso unveils on-line query refiner. Amdahl first with eight-way NT clusters. need 1TB on a 5.25-Inch disk? Datamation, 42(12):11–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CEPb

[Ano96q] Anonymous. Cutting edge — products: Compatibility server enables docs to be shipped. Datamation, 42(13):10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-

Anonymous:1996:CES

[Ano96r] Anonymous. Cutting edge — Sequel Net controls Web cruisers. Sun guns for Hewlett–Packard, IBM. Sagister helps Bull open up. hope for American programmers. dig ever deeper with datamining. new meaning to “star” topology. EDA/SQL loses a little and gains a lot. Vision uses rules to build C/S apps. pilot. Datamation, 42(9):9–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CEL

[Ano96s] Anonymous. Cutting edge — the latest fix-it feature for telecommuters is a service that will let home users transfer their voice mail messages to coworkers at the office. Datamation, 42(1):15– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1996:CET

[Ano96t] Anonymous. Cutting edge — the technologies used in IBM’s Deep Blue may be coming to a computer near you sooner than you think. Datamation, 42(6):13–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CEU

[Ano96u] Anonymous. Cutting edge — use DMI to reduce desktop support costs summer school for IS professionals. Java, Java everywhere. first Plan 9, now inferno. go beyond RAID-5. downsizing of IS slows in ’96. how to link relational and object databases. brace yourself for a price hike for ISDN service. Datamation, 42 (11):10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CEBb

[Ano96v] Anonymous. Cutting edge: Business advantage: Pioneer implements SAP R/3 with HP’s help. Database marketing at NYNEX. Datamation, 42(16):10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CENd

[Ano96w] Anonymous. Cutting edge: News analysis: Here comes your datamart in a box. Datamation exclusive! Diary of an SAP R/3 project. Intranets meet doc management. Datamation, 42(16): 10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CEEb

[Ano96x] Anonymous. Cutting edge: Over the edge: An applet for picking garbage. The Salem bench trials. Datamation, 42(16):10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CEEa

[Ano96y] Anonymous. Cutting edge: Over the edge: Scott Adams on how to deal with stupid managers. Datamation, 42(13):10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:CEPc

[Ano96z] Anonymous. Cutting edge: Products: Microsoft’s new app dev environment: Office 97. Oracle ships cluster software. Datamation, 42(16):10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:DWT

[Ano96-27] Anonymous. Data warehouse tools. Datamation, 42(10):44–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:DSB

[Ano96-28] Anonymous. DATAMATION 100 section breakout — includes each company’s total IT revenue and percentage of revenue by section-from large-scale systems to peripherals. Datamation, 42 (12):70–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:DDT

[Ano96-29] Anonymous. Dataweb development tools. Datamation, 42(7): 46–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:HMA

[Ano96-30] Anonymous. How Microsoft’s ActiveX prevents Internet viruses. Datamation, 42(9):30–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:BOA

[Ano96-31] Anonymous. In box — an OLE automation fan speaks. Datamation, 42(3):9–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:BOO

[Ano96-32] Anonymous. In box — opinions offered on Internet terminals, buzzwords, outsourcing IT services, and more. Datamation, 42(6):7–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:BSA

[Ano96-33] Anonymous. In box — save the accolades of NUMA for Dirk Pitt fans. Datamation, 42(1):9–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:BWS

[Ano96-34] Anonymous. In box — where to spend time looking for hard facts. Datamation, 42(4):7–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:FYP

[Ano96-35] Anonymous. In fox — the Year 2000 problem is only the tip of the iceberg. Datamation, 42(5): 7–??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1996:IDL

[Ano96-36] Anonymous. The IT dozen: Leading tactical consulting companies. Datamation, 42(5):90–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:MFP

[Ano96-37] Anonymous. Manage factory production with CA-Unicenter. Datamation, 42(8):34–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:NA

[Ano96-38] Anonymous. News analysis. Datamation, 42(18):10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:OD

[Ano96-39] Anonymous. OLAP for datawebs.

Datamation, 42(7):43–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:EDM

[Ano96-40] Anonymous. Over the edge — documentation mills on the rise. the MISfortune 500. Datamation, 42(12):22–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:EPO

[Ano96-41] Anonymous. Over the edge — PREMIERE! only kidding — DATAMATION’s new humor section. Datamation, 42(7):24– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1996:ERL

[Ano96-42] Anonymous. Over the edge — read the latest on Internet cheapboxes and data moths (just kidding) in DATAMATION’s new humor section. Datamation, 42 (8):19–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:ESU

[Ano96-43] Anonymous. Over the edge — salary of underground IS. Datamation, 42(15):16–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:ETC

[Ano96-44] Anonymous. Over the edge — telecom complexity competition. high-tech ad debugger. Datamation, 42(9):20–??, ???? 1996.

CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:EFT

[Ano96-45] Anonymous. Over the edge — the first truly bug-free software. top 10 signs of IS project failure. Datamation, 42(11):18–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:EMB

[Ano96-46] Anonymous. Over the edge — the museum of bugs. marketing ploy of the month. Datamation, 42(10):23–??, ???? 1996. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:EWI

[Ano96-47] Anonymous. Over the edge: What if they held an Expo and nobody came? Datamation, 42 (18):10–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:PWN

[Ano96-48] Anonymous. Press watch — the National Science Foundation and MCI are working on an ultra high-speed backbone network service for the Internet. Datamation, 42(1):73–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:PHS

[Ano96-49] Anonymous. Products — hot switches heat up the networking industry Compuware’s Eco-CLlENT. Datamation, 42(15): 16–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:PSA

[Ano96-50] Anonymous. Products: SQL Anywhere lets you leave your intranet (but not your data). JWorks builds Web apps for Informix databases. Datamation, 42(18):10–??, ???? 1996. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:SNM

[Ano96-51] Anonymous. Sun’s new mainframekilling 64Bit UltraSPARC servers — Sun’s new multi-CPU Ultra-SPARC Enterprise line of servers shows what UNIX hardware vendors have in store for the future: bigger, better, faster, more. Datamation, 42(11):36–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:TCS

[Ano96-52] Anonymous. ThinkWrap — client/server technology is improving at a breakneck pace. start now, or be left out. Datamation, 42(2):78–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:TCB

[Ano96-53] Anonymous. Tools to control business processes. Datamation, 42(8):40–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:TNP

[Ano96-54] Anonymous. Top new products for 1996 — the best new products for IS managers are the type that not only help you do your job but help you do it well: products you can bet your business on or products your can count on when your needs veer away from the mainstream. Datamation, 42 (4):30–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:AU

[Ano97a] Anonymous. ActiveX unmasked. Datamation, 43(1):134–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:AIS

[Ano97b] Anonymous. Advertiser’s index, sales staff. Datamation, 43(7): 121–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:BA

[Ano97c] Anonymous. Business advantage. Datamation, 43(1):12–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:BAF

[Ano97d] Anonymous. Business advantage: F&D Reports offers on-line credit information. Datamation, 43(8): 10–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:BAH

[Ano97e] Anonymous. Business advantage: Hilton guests zip from airport to their rooms. The World Cup meets the WWW. Datamation, 43(9):12–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CO

[Ano97f] Anonymous. Career opportunities. Datamation, 43(7):115–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CRP

[Ano97g] Anonymous. Choose the right platform and NOS for your intranet. Datamation, 43(1):118– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1997:CS

[Ano97h] Anonymous. Cover story. Datamation, 43(8):40–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CEBc

[Ano97i] Anonymous. Cutting edge: Business advantage — E-payments for road-warrior expenses. Sabre gets serious about scheduling. Datamation, 43(6):14–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CEBa

[Ano97j] Anonymous. Cutting edge: Business advantage: Bank of Ireland uses data warehouse to exploit regulatory changes. Datamation, 43(3):14–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CEBd

[Ano97k] Anonymous. Cutting edge: Business advantage: Mobil uses IT to speed past competition. This is not your father’s insurance agent. Datamation, 43(10):10–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CEBb

[Ano97l] Anonymous. Cutting edge: Business advantage: Siemens Nixdorf aims to bring the European market up to U.S. speed. Datamation, 43(5):14–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CENa

[Ano97m] Anonymous. Cutting edge: News analysis — JavaMan. ActiveXecutables. Are they friend or foe? AS/400 shops tackle Year 2000 problem. Datamation, 43(1):12– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1997:CENb

[Ano97n] Anonymous. Cutting edge: News analysis: At SIM, an educated IS manager is best. SAP opens R/3 to the net. Business advantage: Online Interactive uses Litlenet to deliver Web commerce. Lotus launches Approach 97. QMS delivers network printer. Datamation, 43(2):14–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CENd

[Ano97o] Anonymous. Cutting edge: News analysis. Business advantage. Products. Datamation, 43 (4):14–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CENc

[Ano97p] Anonymous. Cutting edge: News analysis: How much can NCs really save? Datamation, 43(3):14– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1997:CENf

[Ano97q] Anonymous. Cutting edge: News analysis: IBM’s e-commerce server delivers for UPS. Client/ server susceptible to Y2K. Datamation, 43(6):14–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CENe

[Ano97r] Anonymous. Cutting edge: News analysis: The Big Six face the Y2K problem. BEA Systems strengthens its middleware hold. Datamation, 43(5):14–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CEPa

[Ano97s] Anonymous. Cutting edge:

Products: At CeBIT ’97 the Internet preps for the enterprise. Datamation, 43(3):14–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CEPb

[Ano97t] Anonymous. Cutting edge:

Products: Engage releases a suite to collect info on Web site visitors. ESS software updates self-service HR app. InfoVista tracks conformance. Datamation, 43(5):14–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CEPc

[Ano97u] Anonymous. Cutting edge:

Products: Scan documents directly to the Web cross-platform Year 2000 conversion. Datamation, 43(6):14–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CEPd

[Ano97v] Anonymous. Cutting edge:

Products: Tandem’s NT clustered server. IBM brings its servers to the Web. Analyze network resources automatically. Datamation, 43(10):10–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CETc

[Ano97w] Anonymous. Cutting edge: Talk of the town: a minister for the millennium. The Wiz that was. Datamation, 43(10):10–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CETd

[Ano97x] Anonymous. Cutting edge: Talk of the town: a pride of puzzles. Business Land O’Lakes spreads OLAP around. Amoco improves accounts payable system. Datamation, 43(11):14–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CETb

[Ano97y] Anonymous. Cutting edge: Talk of the town: Database tool makes sense of “the tribes of TV”. The sound of money on the Web. Datamation, 43(9):12–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:CETa

[Ano97z] Anonymous. Cutting edge: Talk of the town: Intellectual real estate: New, used, and listed. New group to standardize system availability terms. Datamation, 43(8):10–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:DAL

[Ano97-27] Anonymous. The Datamation 100 is an annual listing of the top IT vendors in the world, ranked by revenue. This section lists the vendors alphabetically, with information on the breakdown of their IT revenues. Datamation, 43(7):74–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:DAN

[Ano97-28] Anonymous. Dealing with the aftershock of a new SAP implementation. Datamation, 43(4): 105–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:EM

[Ano97-29] Anonymous. Editorial masthead. Datamation, 43(7):5–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:HYT

[Ano97-30] Anonymous. Hang on to your travelling customers. Datamation, 43(5):111–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:C

[Ano97-31] Anonymous. IS by committee.

Datamation, 43(6):48–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:LNW

[Ano97-32] Anonymous. Lee’s new Web app, anchored on a mainframe, provides wholesale distributors with on-line product information. Datamation, 43(4):52–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:MC

[Ano97-33] Anonymous. Mission control. Datamation, 43(7):36–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:EDH

[Ano97-34] Anonymous. Over the edge: Datamation’s humor section has moved to page 148. Datamation, 43(1):12–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:EIW

[Ano97-35] Anonymous. Over the edge: Interactive Web site wants to pump you up. Datamation, 43(1):148– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1997:PBB

[Ano97-36] Anonymous. Plugin: Bigger is better on the Web. Datamation, 43(11):111–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:PYE

[Ano97-37] Anonymous. Plugin: Have you ever wondered. Datamation, 43 (10):126–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:PHR

[Ano97-38] Anonymous. Plugin: Help for the R13 blues. Datamation, 43 (9):128–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:PMW

[Ano97-39] Anonymous. Plugin: Make the Web work for you. Datamation, 43(5):128–??, ???? 1997. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:PMY

[Ano97-40] Anonymous. Plugin: Measure your worth. Datamation, 43(6): 122–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:PWB

[Ano97-41] Anonymous. Plugin: Web-based training. Datamation, 43(8):84– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Anonymous:1997:PBN

[Ano97-42] Anonymous. Products: Build networks quickly with plug and stor 100EX. Datamation, 43(8): 10–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:PDV

[Ano97-43] Anonymous. Products: Data Views’ dynamic data. Mobile workers stay in touch. Datamation, 43(9):12–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:PWA

[Ano97-44] Anonymous. Products: Web app dev made easier with Microsoft’s InterDev Pyramid pushes SMP limits. Datamation, 43(1):12–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:RDDa

[Ano97-45] Anonymous. Raising dough from data: The decline and fall of the average buyer. Datamation, 43 (5):52–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:RDDb

[Ano97-46] Anonymous. Raising dough from data: Why DO they do it? Datamation, 43(5):54–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:SYD

[Ano97-47] Anonymous. Stalking your data. Datamation, 43(9):60–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:WFL

[Ano97-48] Anonymous. Wells Fargo’s online customer banking system is based on S/390 mainframes connected to UNIX servers. Datamation, 43(4):48–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:WSM

[Ano97-49] Anonymous. When speed means profit, consider going paperless. Datamation, 43(6):101–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1997:WKM

[Ano97-50] Anonymous. Working out the kinks in a megadeal. Datamation, 43(6):45–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1998:BAA

[Ano98a] Anonymous. Business advantage: Applix’s Enterprise Anyware eases customers’ angst. Datamation, 44(2):12–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1998:CETb

[Ano98b] Anonymous. Cutting edge: Talk of the town: Food in the net. Datamation, 44(2):12–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1998:CETa

[Ano98c] Anonymous. Cutting edge: Talk of the town: The man who gets gotchas Fleer/Skybox takes trading-card production to the Net. Maestro helps. Datamation, 44(1):10–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1998:FF

[Ano98d] Anonymous. A final farewell.

Datamation, 44(2):104–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1998:PDF

[Ano98e] Anonymous. Plugin: Datamation on the e-frontier. Datamation, 44(2):93–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1998:PFR

[Ano98f] Anonymous. Plugin: Find the right candidate on line. Datamation, 44(1):125–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1998:PSA

[Ano98g] Anonymous. Products: See applications better with LabVIEW 5.0. Ordering office supplies just get easier. Datamation, 44(2):12– ??, ???? 1998. CODEN DTM-NAT. ISSN 0011-6963.

Appleton:1991:DSU

[App91] E. L. Appleton. DEC, sun users eye minisupers. Datamation, 36 (24):64, December 1991. CO-DEN DTMNAT. ISSN 0011-6963.

Appleton:1992:IBC

[App92a] E. L. Appleton. IBM’s big changes going slowly. Datamation, 38(25):39–43, December 1992. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1992:SDR

[App92b] E. L. Appleton. Smart document retrieval. Datamation, 38 (2):20–23, January 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Appleton:1993:LCD

[App93a] E. L. Appleton. Link corporate data and documents. Datamation, 39(7):65, 67–70, April 1993. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1993:NSY

[App93b] E. L. Appleton. Network security: is your LAN a sieve? Datamation, 39(17):79–81, September 1993. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1994:SHW

[App94] Elaine L. Appleton. Staffing Up? Here’s What You’ll Pay. Datamation, 40(20):53–??, October 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Appleton:1995:SSMc

[App95a] E. L. Appleton. Sales surge as mainframes find a role in client/ server. Datamation, 41(10):48, June 1995. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1995:RSY

[App95b] Elaine Appleton. The Right Server for Your Data Warehouse. Datamation, 41(5):56–??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1995:RI

[App95c] Elaine L. Appleton. Recruiting on the Internet. Datamation, 41 (14):39–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Appleton:1995:SSMa

[App95d] Elaine L. Appleton. Sales Surge as Mainframes Find a Role in Client/Server. Datamation, 41 (10):48–??, June 01, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Appleton:1995:SSMb

[App95e] Elaine L. Appleton. Sales Surge as Mainframes Find a Role in Client/Server. Datamation, 41 (10):48–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Appleton:1995:DET

[App95f] Elaine L. Appleton. The Document Explosion: Take Control! Datamation, 41(20):64–65, 67– 68, November 1995. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1996:BTB

[App96a] Elaine L. Appleton. Big time, big-league Web outsourcing — egads! your company is graduating from a simple home page to a big-deal, big-volume, big-bucks-generating Web site. what to do? outsource, of course. Datamation, 42(15):84–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1996:CCB

[App96b] Elaine L. Appleton. Call in the cavalry before 2000. Datamation, 42(1):42–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1996:CMS

[App96c] Elaine L. Appleton. Can Microsoft support the enterprise? Microsoft would certainly want you to think so. but if your IS shop manages thousands of desktops, dozens of file servers, and a fistful of operating systems, you’re probably thinking twice about going mission critical with Microsoft. Datamation, 42(6): 85–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1996:COU

[App96d] Elaine L. Appleton. Crisis outsourcing — under the gun? sometimes you just gotta do crisis staffing. here’s how one Texas agency got the help it needed when it faced an unmovable deadline. Datamation, 42 (11):54–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1996:DWO

[App96e] Elaine L. Appleton. Data Warehouse with an OLAP view — the folks at PacifiCare were committed to building a data warehouse. A good OLAP tool would let them peek inside and manipulate the data. Datamation, 42 (8):94–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1996:HES

[App96f] Elaine L. Appleton. High-end storage takes off — direct access storage devices is where the data lives, and with the growth of data warehouses, everyone’s data needs a whole lotta elbow room. Datamation, 42(12):48–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1996:MCO

[App96g] Elaine L. Appleton. Managed care outsourcing offers flexibility — when it went looking for a desktop outsourcer, United Wisconsin Services needed more than a single service plan for its users. it found three. Datamation, 42 (7):84–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1996:SCS

[App96h] Elaine L. Appleton. On-site:

Chase scales reporting to the high end — most query- andreporting tools don’t scale up very well. so Chase Manhattan turned to Actuate when 3,000 of it users needed reports from a data warehouse. Datamation, 42 (9):84–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1996:UYD

[App96i] Elaine L. Appleton. Use your data warehouse to compete! the IT managers at ReliaStar Financial knew what technologies they needed to stay competitive, but finding the data warehouse system to make it happen was another story. Datamation, 42(10): 34–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1997:HSE

[App97a] Elaine L. Appleton. How to survive ERP. Datamation, 43(3): 50–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Appleton:1997:WWP

[App97b] Elaine L. Appleton. Who will pay for Year 2000 fixes? It’s an issue so cloaked in secrecy that almost no one but attorneys — who stand to make money on the problem — will talk about it on the record. Datamation, 43 (1):63–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Aronica:1996:WYL

[AR96] Ronald C. Aronica and Donald E. Rimel, Jr. Wrapper your legacy systems — you may be tempted to just bulldoze your legacy apps. don’t. you can integrate them into your next-generation systems architecture using object-wrapping techniques. Datamation, 42(12): 83–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Armstrong:1990:PFE

[Arm90a] D. Armstrong. The people factor in EIS success. Datamation, 36(7):73–75, 78–79, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Armstrong:1990:HRL

[Arm90b] D. A. Armstrong. How Rockwell launched its EIS. Datamation, 36 (5):69–72, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Arnum:1990:PPD

[Arn90] E. Arnum. PC power drives demand for token nets. Datamation, 36(10):84–89, May 1990. CODEN DTMNAT. ISSN 0011-6963.

Asbrand:1994:SSS

[Asb94] Debbie Asbrand. Step Up to Scalable Servers. Datamation, 40 (24):63–??, December 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Asbrand:1997:DRM

[Asb97a] Deborah Asbrand. Is datamining ready for the masses? Forget the hokey headlines about “nuggets.” Opening data warehouses to new users has real business value, but the . . .. Datamation, 43(11):66–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Asbrand:1997:YAC

[Asb97b] Deborah Asbrand. Is your automated customer service killing you? Too much automation in your call center will send customers fleeing to the competition. Here’s how to. Datamation, 43(5):62–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Asbrand:1997:OYM

[Asb97c] Deborah Asbrand. Outsource your maintenance migraines. Datamation, 43(6):50–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Asbrand:1997:OBS

[Asb97d] Deborah Asbrand. Outsourcing becomes strategic. Datamation, 43(7):73–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Asbrand:1997:RT

[Asb97e] Deborah Asbrand. A rising tide. Datamation, 43(8):44–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Asbrand:1997:SEC

[Asb97f] Deborah Asbrand. Squeeze out excess costs with supply-chain solutions. Datamation, 43(3):62– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Asbrand:1997:WSA

[Asb97g] Deborah Asbrand. Why sales automation is now succeeding. Datamation, 43(1):80–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Asbrand:1998:NSF

[Asb98a] Deborah Asbrand. The NT service factory. Datamation, 44(1): 66–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Asbrand:1998:TMM

[Asb98b] Deborah Asbrand. Thoroughly modern manufacturing. Datamation, 44(2):64–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Balban:1990:OWN

[Bal90] M. Balban. OCR: what’s needed next. Datamation, 36(8):98–99, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Ball:1996:TNB

[Bal96a] Cheryl Ball. Time for new blueprints — multitier client/ server environments provide a more effective modularity for development. Datamation, 42(6): 97–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Ball:1996:UWN

[Bal96b] Cheryl Ball. UNIX and Windows NT square off — the fate of UNIX as the database client/ server operating environment of choice is becoming increasingly clouded. Datamation, 42(5):75– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Barnat:1990:EUC

[Bar90a] C. Barnat. European unity creates new superpowers (in the computer industry). Datamation, 36(12):118–123, 127–128, 130, 132, 134, 136, 138, 140, 143– 146, June 1990. CODEN DTM-NAT. ISSN 0011-6963.

Baronas:1990:GQS

[Bar90b] J. Baronas. A guide to quality scanning. Datamation, 36(8):96– 97, April 1990. CODEN DTM-NAT. ISSN 0011-6963.

Barron:1990:EUC

[Bar90c] D. Barron. Ease of use coming to PC graphics. Datamation, 36(10):37–38, 40, 42, May 1990. CODEN DTMNAT. ISSN 0011-6963.

Baum:1994:BCS

[Bau94a] David Baum. Backing into Client/Server. Datamation, 40 (20):71–??, October 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Baum:1994:GGI

[Bau94b] David Baum. Graphical 4GLs Get It in Gear. Datamation, 40 (10):75–??, May 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Baum:1994:OM

[Bau94c] David Baum. Objects for the Mainframe. Datamation, 40(15): 49–??, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Baum:1994:WV

[Bau94d] David Baum. Welcome to the VisualAge. Datamation, 40(17):55– ??, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Baum:1995:GINb

[Bau95a] D. Baum. Groupware: is it notes or nothing? Datamation, 41 (8):45–48, May 1995. CODEN DTMNAT. ISSN 0011-6963.

Baum:1995:GINa

[Bau95b] David Baum. Groupware: Is It Notes or Nothing? Datamation, 41(8):45–??, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Baum:1995:NTB

[Bau95c] David Baum. New Technology Burnout? Here’s How To Cope. Datamation, 41(21):45–47, November 1995. CODEN DTM-NAT. ISSN 0011-6963.

Baum:1995:RTC

[Bau95d] David Baum. The Right Tools for Coding Business Rules. Datamation, 41(4):36–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Baum:1995:UDD

[Bau95e] David Baum. Unify Dispersed Development Teams. Datamation, 41(15):37–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Baum:1996:T

[Bau96a] David Baum. Tool up for 2000.

Datamation, 42(1):49–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Baum:1996:UPS

[Bau96b] David Baum. Union Pacific stays on track for 2000. Datamation, 42(1):39–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Beaulieu:1996:PBD

[Bea96] Greg Beaulieu. Perspective — I became a dinosaur the day I got a notebook PC running Windows95. Datamation, 42(7):33– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Beckley:1994:TFR

[Bec94] Glen B. Beckley. TQM: Find the Red Flags Hiding in Existing Systems. Datamation, 40 (17):63–64, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Beech:1989:FS

[Bee89a] D. Beech. The future of SQL. Datamation, 35(4):45–48, February 1989. CODEN DTMNAT. ISSN 0011-6963.

Beech:1989:NLS

[Bee89b] D. Beech. New life for SQL. Datamation, 35(3):29–31, 33, 35– 36, February 1989. CODEN DTMNAT. ISSN 0011-6963.

Belitos:1990:CSS

[Bel90] B. Belitos. Client/server systems attract retail bankers. Datamation, 36(11):77–78, 80, June 1990. CODEN DTMNAT. ISSN 0011-6963.

Benson:1998:BPI

[Ben98] Robert J. Benson. Building profitable IT. Datamation, 44(1):74– ??, ???? 1998. CODEN DTM-NAT. ISSN 0011-6963.

Bernstein:1994:IAI

[Ber94]  David Bernstein.  Insulate

Against Internet Intruders. Datamation, 40(19):49–50, 52, October 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Bermant:1995:ANI

[Ber95a] C. Bermant. Applications ARE new and improved (windows95).

Datamation, 41(16):57–58, 60, September 1995. CODEN DTM-NAT. ISSN 0011-6963.

Bermant:1995:WAN

[Ber95b] Charles Bermant. Windows95 Applications Are New and Improved. Datamation, 41(16):57– ??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Bernstein:1995:SVO

[Ber95c] David S. Bernstein. Secure the virtual office. Datamation, 41(1): 49–51 (or 53–??), January 1995. CODEN DTMNAT. ISSN 0011-6963.

Best:1990:BSS

[Bes90] L. J. Best. Building software skyscrapers. Datamation, 36(6): 67–74, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Bernacki:1990:MPB

[BF90] W. Bernacki and M. Freely.

Modems: performance at a better price. Datamation, 36 (21):107–108, 110–111, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Blankenhorn:1988:POP

[Bla88]  D. Blankenhorn.  Pursuing

one peripheral (printer/copier/ scanner/fax). Datamation, 34 (20):71, 74, 76, October 1988. CODEN DTMNAT. ISSN 0011-6963.

Black:1990:EPP

[Bla90] N. H. Black. Escape from the paper pit (image processing). Datamation, 36(8):86–88, April 1990.

CODEN DTMNAT. ISSN 0011-6963.

Black:1995:GYD

[Bla95a] George Black. Get Your Data Back! Datamation, 41(23):54– ??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Black:1995:SEU

[Bla95b] George Black. Simplify End-User Computing: Outsource It. Datamation, 41(17):67–69, September 1995. CODEN DTMNAT. ISSN 0011-6963.

Black:1996:SMT

[Bla96] George Black. Save money through asset management — OK, so you saved 10% on your PC hardware by cutting a killer deal with your vendor. but that’s not going to mean anything if you don’t control PC management costs. here’s how. Datamation, 42(7):62–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Bloombecker:1989:M

[Blo89a] J. J. Bloombecker. Malpractice in IS? Datamation, 35(20):85–86, October 1989. CODEN DTM-NAT. ISSN 0011-6963.

Bloombecker:1989:SCC

[Blo89b]  J. J. Bloombecker.  Short-

circuiting computer crime. Datamation, 35(19):71–72, October 1989. CODEN DTMNAT. ISSN 0011-6963.

Bell:1990:SS

[BM90] C. G. Bell and G. Miranker. Shopping for supercomputers.

Datamation, 36(20):76–78, 80, October 1990. CODEN DTM-NAT. ISSN 0011-6963.

Bobrow:1989:ODO

[Bob89] D. G. Bobrow. The object of desire object-oriented programming. Datamation, 35(9):37–38, 41, May 1989. CODEN DTM-NAT. ISSN 0011-6963.

Bookman:1990:TKT

[Boo90] H. Bookman. Training for a key transition. Datamation, 36(6): 109–113, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Boult:1990:CP

[Bou90] R. Boult. CIM in the pipeline.

Datamation, 36(5):96–95, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Bowen:1994:MSD

[Bow94] Barry D. Bowen. MP Servers:

Don’t Buy by the Numbers. Datamation, 40(11):76–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Bradley:1989:EEA

[Bra89] J. Bradley. An educational experience at aetna. Datamation, 35 (7):46–47, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Brousell:1988:NED

[Bro88] D. R. Brousell. The new era of DBMS integration. Datamation, 34(16):57–58, August 1988. CODEN DTMNAT. ISSN 0011-

Brookman:1989:TBP

[Bro89] H. Brookman. Testing for the best programmer. Datamation, 35(7):83–86, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Brody:1990:NPH

[Bro90] A. Brody. The new power of the HP LaserJet III. Datamation, 36 (19):103–104, October 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Bryan:1988:HSN

[Bry88] M. Bryan. With high speed, new applications (modems). Datamation, 34(24):59–60, 66, 72, December 1988. CODEN DTM-NAT. ISSN 0011-6963.

Bryan:1989:HWC

[Bry89a] M. Bryan. Help is on the way (consultants and capacity planning). Datamation, 35(6):45–51, March 1989. CODEN DTMNAT. ISSN 0011-6963.

Bryan:1989:SPD

[Bry89b] M. Bryan. Shopping for a PC DBMS. Datamation, 35(6):25– 35, March 1989. CODEN DTM-NAT. ISSN 0011-6963.

Baum:1994:WRC

[BS94] David Baum and Kevin Strehlo. Which RDBMS Can Handle Your Toughest Jobs? Datamation, 40(13):59–??, July 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Bulkeley:1995:WRY

[Bul95] Debra Bulkeley. Where To Recruit Your Next Rising Star.

Datamation, 41(19):56–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Bunker:1990:IFB

[Bun90] T. Bunker. IBM fiber breaks data center barrier. Datamation, 36 (23):85–86, December 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Bunker:1991:WSF

[Bun91] T. Bunker. What’s slowing fast token ring. Datamation, 36(24): 51–52, December 1991. CODEN DTMNAT. ISSN 0011-6963.

Buse:1988:CSC

[Bus88] J. Buse. Chargeback systems come of age (finance management). Datamation, 34(21):47– 57, November 1988. CODEN DTMNAT. ISSN 0011-6963.

Caniano:1988:ATC

[Can88] S. Caniano. All TP1s are not created equal (transaction processing benchmark). Datamation, 34(16):51–53, August 1988. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1988:ATG

[Car88a] R. Emmett Carlyle. Advanced technology groups (IS management). Datamation, 34(21):18– 24, November 1988. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1988:RUS

[Car88b] R. Emmett Carlyle. RISC and unix: the signs of transitional times (IS industry). Datamation, 34(17):38–44, September 1988.

CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1988:WMF

[Car88c] R. Emmett Carlyle. Where methodology falls short (CASE). Datamation, 34(24):43–44, 46– 47, December 1988. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1989:LAS

[Car89a] R. Carlyle. Leaping ahead in software productivity (programming languages). Datamation, 35(23):22–24, 31, 32, December 1989. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1989:CAC

[Car89b] R. E. Carlyle. Computer associates: at a crossroads. Datamation, 35(8):61–63, 65, 67, 69, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1989:FCA

[Car89c] R. E. Carlyle. Fighting corporate amnesia (programming management). Datamation, 35(3):41–44, February 1989. CODEN DTM-NAT. ISSN 0011-6963.

Carlyle:1989:IVV

[Car89d] R. E. Carlyle. IBM’s VSE: a victory for the techies? Datamation, 35(14):25–27, July 1989. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1989:NUA

[Car89e] R. E. Carlyle. The new unisys approach. Datamation, 35(19): 53, 56, October 1989. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1989:S

[Car89f] R. E. Carlyle. The selling of IS. Datamation, 35(13):22–24, 26, July 1989. CODEN DTM-NAT. ISSN 0011-6963.

Carlyle:1989:WYW

[Car89g] R. E. Carlyle. What are you worth? (DATAMATION salary survey). Datamation, 35(19):22– 25, 28–30, October 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Carlyle:1990:CUG

[Car90a] R. Carlyle. CA user group sells its software (computer associates). Datamation, 36(18):57– 58, 60, September 1990. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1990:FSB

[Car90b] R. Carlyle. Financial systems that bust borders. Datamation, 36(22):77–78, 80, 82, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1990:YDR

[Car90c] R. Carlyle. Is your data ready for the repository? (IBM AD/ cycle CASE environment). Datamation, 36(1):43–44, 46, 48, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1990:RS

[Car90d] R. Carlyle. Recovery! (IS spending). Datamation, 36(7):34–39, 42, 44, 48, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1990:SAL

[Car90e] R. Carlyle. Software AG leaps into server systems. Datamation, 36(19):99, 101, October 1990. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1990:TO

[Car90f] R. Carlyle. The tomorrow organization. Datamation, 36(3):22– 25, 28–29, February 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Carlyle:1990:WRU

[Car90g] R. Carlyle. What’s right with the US market (for information systems). Datamation, 36(12):40– 46, 53–56, 58, 60, 64, 66–69, 72, 74, 76, 78–81, 84, 86, June 1990. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1990:WCM

[Car90h] R. Carlyle. Why consolidation makes sense. Datamation, 36(8): 24–27, 30–31, April 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Caron:1991:TPM

[Car91] J. Caron. Tracking project management. Datamation, 37(19): 95–97, October 1991. CODEN DTMNAT. ISSN 0011-6963.

Carson:1997:TVS

[Car97] Barbara H. Carson. Ten vital steps to save your C/S system from disaster. Datamation, 43 (2):123–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Caswell:1990:MMC

[Cas90a] S. A. Caswell. Mail messages that cut red tape. Datamation, 36(20):47–48, 50, October 1990. CODEN DTMNAT. ISSN 0011-6963.

Caswell:1990:NLS

[Cas90b] S. A. Caswell. A new LAN standard lights the way. Datamation, 36(9):75–80, May 1990. CODEN DTMNAT. ISSN 0011-6963.

Cassidy:1995:KBGc

[Cas95a] P. Cassidy. Keep the barbarians at the gateway (network security). Datamation, 41(11):51, June 1995. CODEN DTMNAT. ISSN 0011-6963.

Cassidy:1995:KBGa

[Cas95b] Peter Cassidy. Keep the Barbarians at the Gateway. Datamation, 41(11):51–??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Cassidy:1995:KBGb

[Cas95c] Peter Cassidy. Keep the Barbarians at the Gateway. Datamation, 41(11):51–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Cassidy:1995:WSWa

[Cas95d] Peter Cassidy. Wholesale success on the Web. Datamation, 41(11): 48–??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Cassidy:1995:WSWb

[Cas95e] Peter Cassidy. Wholesale Success on the web. Datamation, 41(11):

48–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Cassidy:1996:NRM

[Cas96] Peter Cassidy. The network revolution — the melting of boundaries between LAN, WAN, and Internet technologies, plus consolidation of key industry players, are combining to remake the networked landscape. Datamation, 42(12):62–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Carey:1989:PC

[CC89] J. M. Carey and J. D. Currey. The prototyping conundrum. Datamation, 35(11):29–33, June 1989. CODEN DTMNAT. ISSN 0011-6963.

Celko:1996:SFY

[Cel96] Joe Celko. Start fixing year 2000 problems now! Datamation, 42 (1):36–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Caron:1991:HLW

[CF91] J. Caron and M. Freely. The highs and lows of word processing. Datamation, 36(24):79–81, December 1991. CODEN DTM-NAT. ISSN 0011-6963.

Chaddha:1995:DDC

[Cha95] Parmeet Sinch Chaddha. Dancing with Data: Check Your Choreography. Datamation, 41 (14):59–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Childs:1991:SSS

[Chi91] R. G. Childs. Seven steps to successful E-mail. Datamation, 37(18):56, 58, September 1991. CODEN DTMNAT. ISSN 0011-6963.

Childs:1992:HLM

[Chi92] R. G. Childs. How LTV mailenabled its applications. Datamation, 38(14):33–35, July 1992. CODEN DTMNAT. ISSN 0011-6963.

Clark:1990:HRT

[Cla90] S. Clark. How to recruit trading partners. Datamation, 36(21): 75–76, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Carlyle:1989:RIU

[CM89] R. E. Carlyle and J. Moad. The rise of an information utility (IBM). Datamation, 34(26):26– 31, 34, January 1989. CODEN DTMNAT. ISSN 0011-6963.

Celko:1995:DWD

[CM95] Joe Celko and Jackie McDonald. Don’t Warehouse Dirty Data. Datamation, 41(19):42–44, 51, 53, October 1995. CODEN DTMNAT. ISSN 0011-6963.

Codd:1988:FFSa

[Cod88a] E. F. Codd. Fatal flaws in SQL.

1. Datamation, 34(16):45–48, August 1988. CODEN DTM-NAT. ISSN 0011-6963.

Codd:1988:FFSb

[Cod88b] E. F. Codd. Fatal flaws in SQL.

2. Datamation, 34(17):71–74, September 1988. CODEN DTM-NAT. ISSN 0011-6963.

Codd:1990:MADa

[Cod90a] E. F. Codd. Mastering the art of database fusion. Datamation, 36(20):61–62, 64, October 1990. CODEN DTMNAT. ISSN 0011-6963.

Codd:1990:MADb

[Cod90b] E. F. Codd. Mastering the art of database fusion. II. Datamation, 36(21):81–82, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Cohen:1995:WBD

[Coh95] Fred Cohen. Why 32-Bit Desktops Need Virus Protection. Datamation, 41(13):41–43, July 1995. CODEN DTMNAT. ISSN 0011-6963.

Connor:1992:NUM

[Con92] D. Connor. New utilities make LAN management easier. Datamation, 38(23):69–72, November 1992. CODEN DTMNAT. ISSN 0011-6963.

Connolly:1997:BTR

[Con97] Paul Connolly. Blood test: THE ROI of testing tools. Datamation, 43(5):86–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Cook:1991:RS

[Coo91] R. Cook. The rise of the superfloppy. Datamation, 37(19): 39–42, October 1991. CODEN DTMNAT. ISSN 0011-6963.

Cook:1992:CTC

[Coo92a] R. Cook. Cutting through the 486 confusion (microprocessors).

Datamation, 38(4):43–45, February 1992. CODEN DTMNAT. ISSN 0011-6963.

Cook:1992:WAP

[Coo92b] R. Cook. What’s ahead for PC/ fax. Datamation, 38(14):27–29, July 1992. CODEN DTMNAT. ISSN 0011-6963.

Corcoran:1995:PCN

[Cor95a] C. T. Corcoran. PowerPC: close, but no cigar. Datamation, 41(7): 36–38, 40, April 1995. CODEN DTMNAT. ISSN 0011-6963.

Corcoran:1995:CCC

[Cor95b] Cate T. Corcoran. Consolidate computers and conquer desktop costs (dual Pentium PC/ workstation systems). Datamation, 41(18):61, 63–64, October 1995. CODEN DTMNAT. ISSN 0011-6963.

Corcoran:1995:DPP

[Cor95c] Cate T. Corcoran. Does plug and play really mean buy all new desktops? Datamation, 41(4): 49–50, 52, March 01, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Corcoran:1996:SDM

[Cor96] Cate T. Corcoran. SMS: a desktop manager for the Enterprise? Microsoft’s Systems Management Server can’t manage all the systems across an enterprise, but it talks to all the products that can. Datamation, 42(6):71– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Cougher:1988:MP

[Cou88] J. D. Cougher. Motivating IS personnel. Datamation, 34(18): 59–64, September 1988. CODEN DTMNAT. ISSN 0011-6963.

Cox:1990:TRR

[Cox90] G. H. Cox. Technology’s rewards without the risks. Datamation, 36(3):69–70, 72, 75, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Corcoran:1995:ISU

[CS95] Cate T. Corcoran and Mark Schlack. Internet servers for under US$100? here’s one! (caldera network desktop). Datamation, 41(12):41, 43, July 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Cusack:1993:PRK

[Cus93] S. Cusack. The people rollout: key to change. Datamation, 39 (7):55–56, April 1993. CODEN DTMNAT. ISSN 0011-6963.

Daily:1995:WIT

[Dai95] John C. Daily. What It Takes To Be C10 -. Datamation, 41(20): 61–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Dallis:1998:LRH

[Dal98] Dwight B. Dallis. Letter from Redmond: Hydra: beast or hero? Datamation, 44(1):42–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Dambrot:1989:JPS

[Dam89] S. M. Dambrot. Japan prepares for software crisis. Datamation, 35(9):80/13–14, 16, May 1989. CODEN DTMNAT. ISSN 0011-6963.

Darst:1990:BPQ

[Dar90] D. Darst. Balancing productivity and quality (IS management). Datamation, 36(18):117– 119, September 1990. CODEN DTMNAT. ISSN 0011-6963.

Darling:1995:LRA

[Dar95a] C. B. Darling. Leaping into remote access (server acquisition). Datamation, 41(7):59–66, April 1995. CODEN DTMNAT. ISSN 0011-6963.

Darling:1995:RTY

[Dar95b] C. B. Darling. Reach out and touch your LAN (remote access servers). Datamation, 41(7):57– 59, April 1995. CODEN DTM-NAT. ISSN 0011-6963.

Darling:1995:RAA

[Dar95c] C. B. Darling. Reinventing an airline’s IS (ATA internal systems). Datamation, 41(14):67, 69–74, August 1995. CODEN DTMNAT. ISSN 0011-6963.

Darling:1995:BFP

[Dar95d] Charles B. Darling. Baxter finds partners for volume purchases (healthcare business PC buying). Datamation, 41(18):85, 87–88, 90–91, October 1995. CODEN DTMNAT. ISSN 0011-6963.

Darling:1995:BLL

[Dar95e] Charles B. Darling. Building links to legacy apps. Datamation, 41(1):70–??, January 1995.

CODEN DTMNAT. ISSN 0011-6963.

Darling:1995:CPS

[Dar95f] Charles B. Darling. Control the PC shopping spree (enterprise resource management). Datamation, 41(18):82–85, October 1995. CODEN DTMNAT. ISSN 0011-6963.

Darling:1995:DDS

[Dar95g] Charles B. Darling. Dig Deep To Strike Data Access Gold. Datamation, 41(20):76–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Darling:1995:GDF

[Dar95h] Charles B. Darling. Go with the (data) flow (bank regional WAN). Datamation, 41(12):67– 68, 70–71, 73, July 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Darling:1995:IYA

[Dar95i] Charles B. Darling. Immortalize Your Apps. Datamation, 41(14): 65–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Darling:1995:IPI

[Dar95j] Charles B. Darling. Integrate Printing into IT Strategy. Datamation, 41(9):72–??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Darling:1995:ITC

[Dar95k] Charles B. Darling. Is It Time To Change Your NOS? Datamation, 41(16):71–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Darling:1995:LUS

[Dar95l] Charles B. Darling. Look Under the SQL Hood -. Datamation, 41 (20):79–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Darling:1995:MRM

[Dar95m] Charles B. Darling. Mainframe RAID Must Match Your App Needs. Datamation, 41(5):65–??, March 15, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Darling:1995:OSD

[Dar95n] Charles B. Darling. One Size DASD Does Not Fit All. Datamation, 41(5):67–??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Darling:1995:RA

[Dar95o] Charles B. Darling. Reinventing an Airine’s IS. Datamation, 41 (14):67–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Darling:1995:RCS

[Dar95p] Charles B. Darling. Routers Can Save Your WAN Dollars. Datamation, 41(12):64–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Darling:1995:TRL

[Dar95q] Charles B. Darling. Tap into a Rich Legacy of Data. Datamation, 41(1):65–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Darling:1995:IOT

[Dar95r] Charles B. Darling. The Ins and Outs of Testing Switches.

Datamation, 41(3):69–??, February 15, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Darling:1995:TEO

[Dar95s] Charles B. Darling. Two Evaluations, One Printer. Datamation, 41(9):75–??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Darling:1995:UYL

[Dar95t] Charles B. Darling. Unclog Your Local Backbone with ATM. Datamation, 41(3):67–??, February 15, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Darling:1995:VMG

[Dar95u] Charles B. Darling. VINES makes the grade (in school district). Datamation, 41(16):74– 77, September 1995. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:NIB

[Dar96a] Charles Darling. Need an integrator? build a relationship! it used to be that when you went outside to get a system built, you looked for a firm that was good at build ng systems. today, an outsourcer must develop a system that is thoroughly. Datamation, 42(5):84–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:NPP

[Dar96b] Charles Darling. A new procurement paradigm builds revenue — partnering can be more than just marketing hype. at California’s Franchise Tax Board, partnership is the rocket that business, IS, and integrator ride to effective, business-drive systems. Datamation, 42(5):88–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:BAD

[Dar96c] Charles B. Darling. Build an App Dev Framework for DSS — it’s a three-ring circus out there as vendors make outrageous claims about bogus issues and expect customers to cheer about their Decision Support System Products. ignore them. Datamation, 42(8):82–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:DOP

[Dar96d] Charles B. Darling. Desktop outsourcing: Pry open the black box — if your methodologies for acquiring, installing, and supporting PCs are perfect, if your execution is consistently flawless, if your desk-top resources are in perfect control, then skip this article-there’s nothing in it for you. Datamation, 42(7):82–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:DYO

[Dar96e] Charles B. Darling. Does your OO app need an OO database? Datamation, 42(18):92–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:DTP

[Dar96f] Charles B. Darling. A DSS trilogy: Part I — think outside the OLAP box — if you think OLAP is a tool that let users play with data, think again it’s an application development environment, and if you want it to work, you’d better think of it that way. Datamation, 42(8):88–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:ECW

[Dar96g] Charles B. Darling. Embrace change with workflow tools. Datamation, 42(16):102–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:EAG

[Dar96h] Charles B. Darling. Enterprise apps: Get a tough tool — “Enterprise” is easy to say, hard to do. you need a muscular tool to develop mission-critical, n-tier applications that span the enterprise. here’s what to look for. Datamation, 42(13):102–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:EBS

[Dar96i] Charles B. Darling. Ethernet backbone switches: On the road to ATM. Datamation, 42(1):64– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Darling:1996:ESH

[Dar96j] Charles B. Darling. Ethernet switches help unsnarl Boston traffic. Datamation, 42(1):66–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:FYE

[Dar96k] Charles B. Darling. Find your enterprise Internet connection — everyone has an Internet product to sell. most are trash, but some of them even work. here’s how to sort through the dreck to find what you need to link your enterprise to the future. Datamation, 42(3):42–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:HYL

[Dar96l] Charles B. Darling. Hook your legacy LAN to ATM — speed your Ethernet traffic across your ATM backbone? LANF, makes that a practical reality — practically. Datamation, 42(15):98–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:HIY

[Dar96m] Charles B. Darling. How to integrate your data warehouse — data warehousing is tricky enough without worries over products that don’t cooperate. keep a close eye on integration to assemble a smooth, efficient toolkit. Datamation, 42(10):40– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Darling:1996:MYR

[Dar96n] Charles B. Darling. Manage your reporting environment — beef up your data access physique with intelligent viewers, data cupboards, seamless report development, and-most. Datamation, 42 (9):78–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:NTN

[Dar96o] Charles B. Darling. A new tool for a new IS mission — Piper Jaffray Cos. picked an enterprise application development environment as just one step in its strategic-retooling effort to better exploit technology. Datamation, 42(13):98–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:PPE

[Dar96p] Charles B. Darling. Partitions power enterprise client/server. Datamation, 42(18):124–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:TDW

[Dar96q] Charles B. Darling. THINKWRAP — don’t write off the Metadata Coalition… yet — pundits may dismiss the coalition, but users are interested and guardedly optimistic. Datamation, 42(9):96– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Darling:1996:TTC

[Dar96r] Charles B. Darling. The twists and turns of client/server migration — the move from mainframes to client/server is never an easy task, and sometimes a workgroup project can balloon into enterprisewide significance. just ask Banco Internacional. Datamation, 42(15):121– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Darling:1996:WFI

[Dar96s] Charles B. Darling. Want fast Internet access? go ISDN — when it’s time to accommodate digital dialup in your enterprise, you’ll find there’s more to evaluating ISDN devices than just delving into their features and technology. Datamation, 42(11):78–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1996:WOR

[Dar96t] Charles B. Darling. Wanted:

One reliable Internet connection — MathWorks relies on the Internet as a link to its customers. switching ISPs gave it the reliability and pricing it needed. Datamation, 42(3):48–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Darling:1997:DM

[Dar97a] Charles B. Darling. Datamining for the masses. Datamation, 43 (2):52–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Darling:1997:EIW

[Dar97b] Charles B. Darling. Ease implementation woes with packaged datamarts. Datamation, 43(3): 94–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Staff:1994:MST

[Dat94]  Datamation Staff.  More of

the Same (and then some) at COMDEX. Datamation, 40(24): 72–??, December 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Davis:1989:UGR

[Dav89a] D. Davis. U.S. giants run a US$50 billion IS tab (industryby-industry spending survey).

Datamation, 35(22):42–80, November 1989. CODEN DTM-NAT. ISSN 0011-6963.

Davis:1989:CEM

[Dav89b] L. Davis. Can education meet IS career demands? Datamation, 35 (6):65–72, March 1989. CODEN DTMNAT. ISSN 0011-6963.

Davis:1989:LW

[Dav89c] L. Davis. From LANs to WANs. Datamation, 35(9):73–74, May 1989. CODEN DTMNAT. ISSN 0011-6963.

Davis:1989:ISP

[Dav89d] L. Davis. Initiating systems projects. l. managing your vendor. Datamation, 35(20):81– 82, 84, October 1989. CODEN DTMNAT. ISSN 0011-6963.

Davis:1989:PMG

[Dav89e] L. Davis. Phone mail gets stamp of approval (voice messaging). Datamation, 35(7):69–78, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Davis:1989:PAO

[Dav89f] L. Davis. Point of access (online services). Datamation, 35 (14):35–37, July 1989. CODEN DTMNAT. ISSN 0011-6963.

Davis:1989:RGS

[Dav89g] L. Davis. Retailers go shopping for EDI. Datamation, 35(5):53– 54, 56, March 1989. CODEN DTMNAT. ISSN 0011-6963.

Davis:1989:CSA

[Dav89h] S. G. Davis. Can stand-alone EISs stand up? Datamation, 35 (13):41–42, 44, July 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Davis:1990:ADC

[Dav90a] D. B. Davis. IS automates for data center survival. Datamation, 36(22):95–96, 98, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Davis:1990:DMP

[Dav90b] L. Davis. The dawn of MAN: a promising way to connect LANs. Datamation, 36(11):85–86, 88, 90, June 1990. CODEN DTM-NAT. ISSN 0011-6963.

Davis:1990:LWD

[Dav90c] L. Davis. A logical way to distribute computing (communications packages). Datamation, 36 (6):95–102, March 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Davis:1990:LAG

[Dav90d] L. Davis. On-line applications grow up. Datamation, 36(1): 61–63, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Davis:1990:PLC

[Dav90e] L. Davis. Peering at the LU 6.2 choice. Datamation, 36(3):49– 50, 52, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Davis:1990:SBT

[Dav90f] L. Davis. Software’s big 50 (top independent software vendors). Datamation, 36(23):67–71, December 1990. CODEN DTM-NAT. ISSN 0011-6963.

Davis:1990:STN

[Dav90g] L. Davis. Staying on top of network control. Datamation, 36 (8):53–56, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Davis:1991:FTH

[Dav91] L. Davis. On the fast track to HR integration. Datamation, 37 (18):61–63, 65, September 1991. CODEN DTMNAT. ISSN 0011-6963.

Davis:1992:ACY

[Dav92a] D. Davis. Add color to your volume printing. Datamation, 38(4): 69–71, February 1992. CODEN DTMNAT. ISSN 0011-6963.

Davis:1992:HAE

[Dav92b] L. Davis. How to automate EPA reports. Datamation, 38(1): 57–58, January 1992. CODEN DTMNAT. ISSN 0011-6963.

Davis:1992:NUO

[Dav92c] L. Davis. New uses for optical storage. Datamation, 38(6):78– 81, March 1992. CODEN DTM-NAT. ISSN 0011-6963.

Davis:1992:WDM

[Dav92d] L. Davis. A window into document management. Datamation, 38(9):90–92, April 1992. CODEN DTMNAT. ISSN 0011-6963.

Davis:1993:BGN

[Dav93a] D. B. Davis. Building a global network on a shoestring. Datamation, 39(10):59–63, May 1993. CODEN DTMNAT. ISSN 0011-6963.

Davis:1993:TPO

[Dav93b] D. B. Davis. Turmoil in pen operating systems. Datamation, 39 (24):20–22, 26, December 1993. CODEN DTMNAT. ISSN 0011-6963.

Davis:1994:GMA

[Dav94] Arnold Davis. General Magic’s Agents: a More flexible EDI. Datamation, 40(16):51–??, August 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Davis:1996:LRMa

[Dav96a] Dwight B. Davis. Letter from Redmond — my directory’s better than yours. Datamation, 42 (15):146–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Davis:1996:LRNa

[Dav96b] Dwight B. Davis. Letter from Redmond — NT versus UNIX? no contest. Datamation, 42(14): 140–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Davis:1996:LRH

[Dav96c] Dwight B. Davis. Letter from Redmond: How enterprising is Microsoft? Datamation, 42(13): 128–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Davis:1996:LRMb

[Dav96d] Dwight B. Davis. Letter from Redmond: Microsoft’s object lock is rusting. Datamation, 42 (16):148–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Davis:1996:LRNb

[Dav96e] Dwight B. Davis. Letter from Redmond: NCs threaten Windows PCs. Datamation, 42(18): 134–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Davis:1997:LRSb

[Dav97a] Dwight B. Davis. Letter from Redmond: a shrink-wrapped enterprise? Datamation, 43(10): 40–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Davis:1997:LRB

[Dav97b] Dwight B. Davis. Letter from Redmond: Beyond the Internet. Datamation, 43(1):150–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Davis:1997:LRD

[Dav97c] Dwight B. Davis. Letter from Redmond: Developers first? or Active X? Datamation, 43(11): 44–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Davis:1997:LRJ

[Dav97d] Dwight B. Davis. Letter from Redmond: Java: To be or not to be. Datamation, 43(2):136–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Davis:1997:LRSa

[Dav97e] Dwight B. Davis. Letter from Redmond: Sold on (or by) support. Datamation, 43(5):130–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Davis:1997:LRE

[Dav97f] Dwight B. Davis. Letter from Redmond: The end of cheap software? Datamation, 43(9):130– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Davis:1997:LRM

[Dav97g] Dwight B. Davis. Letter from Redmond: The myth of heterogeneity. Datamation, 43(7):122– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Davis:1997:LRTa

[Dav97h] Dwight B. Davis. Letter from Redmond: Time to choose your objects. Datamation, 43(3):130– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Davis:1997:LRTb

[Dav97i] Dwight B. Davis. Letter from Redmond: Tooling up for the enterprise. Datamation, 43(4):130– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Davis:1997:LRWb

[Dav97j] Dwight B. Davis. Letter from Redmond: Waiting for Active Directory. Datamation, 43(8): 86–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Davis:1997:LRWa

[Dav97k] Dwight B. Davis. Letter from Redmond: Windows terminals gain steam. Datamation, 43(6): 124–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Delohery:1997:YBD

[DB97] Pat D. Delohery and Jack Bucsko. Year 2000: a blessing in disguise. Datamation, 43(10):29–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Doll:1989:PPO

[DD89] M. W. Doll and W. J. Doll. A productive PC operation. Datamation, 35(20):77–78, October 1989. CODEN DTMNAT. ISSN 0011-6963.

deJager:1997:YMH

[de 97a] Peter de Jager. Year 2000: a matter of honor. Datamation, 43 (3):30–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

deJager:1997:YSS

[de 97b] Peter de Jager. Year 2000: a snake in the silicon. Datamation, 43(6):31–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

deJager:1997:YTT

[de 97c] Peter de Jager. Year 2000: a taste of testing. Datamation, 43 (2):39–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

deJager:1997:YCB

[de 97d] Peter de Jager. Year 2000: Cellateral benefits. Datamation, 43 (9):27–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

deJager:1997:YDM

[de 97e] Peter de Jager. Year 2000: Dear Mr. Gates. Datamation, 43(8): 27–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

deJager:1997:YGA

[de 97f] Peter de Jager. Year 2000: Government in action. Datamation, 43(7):27–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

deJager:1997:YSB

[de 97g] Peter de Jager. Year 2000: The silver bullet solution. Datamation, 43(5):33–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

deJager:1997:YWT

[de 97h] Peter de Jager. Year 2000: Walking on thin ice. Datamation, 43 (4):31–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

deJager:1997:YYS

[de 97i] Peter de Jager. Year 2000: Year 2000’s silver lining. Datamation, 43(11):31–??, ???? 1997. CO-DEN DTMNAT. ISSN 0011-6963.

deJager:1998:YEB

[de 98a] Peter de Jager. Year 2000: Entering the Black Zone. Datamation, 44(1):39–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

deJager:1998:YRV

[de 98b] Peter de Jager. Year 2000: Revisiting V2K certification. Datamation, 44(2):25–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Delottinville:1994:OOM

[Del94] Paul Delottinville. Open OLTP:

To Monitor or Not To Monitor.

Datamation, 40(21):59–??, November 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Dern:1992:WUY

[Der92] D. P. Dern. Who’s using your LAN? (security software buyer’s guide). Datamation, 38(12):104– 106, June 1992. CODEN DTM-NAT. ISSN 0011-6963.

Devoney:1997:EWD

[Dev97]  Chris Devoney.  Enterprise

Web development environments. Datamation, 43(4):98–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Ditlea:1992:HAL

[Dit92] S. Ditlea. How AMEX leverages technology assets. Datamation, 38(24):95–96, December 1992. CODEN DTMNAT. ISSN 0011-6963.

Jager:1996:TGP

[dJ96a] P. de Jager. ThinkWrap — the general press needs to shout about the Year 2000 problem in banner headlines, so your idiot managers will take you seriously. Datamation, 42(1):76–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Jager:1996:TLY

[dJ96b] Peter de Jager. THINKWRAP — lock up your year 2000 staff the year 2000 deadline is different from the projects and deadlines we’re used to. Datamation, 42 (10):102–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Drips:1996:WDT

[Dri96] Michael Drips. Windows development tools that work — can a lowly Windows developer find affordable tools that meet all his stand-alone, client/server, and Internet development needs? no problem, he says — but with some caveats. Datamation, 42 (9):36–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Dugger:1996:MUC

[Dug96a] Randy Dugger. Managing users:

Close the IS-user gap. Datamation, 42(16):27–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Dugger:1996:MUD

[Dug96b] Randy Dugger. Managing users:

Don’t write off business managers. Datamation, 42(18):29–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Dugger:1996:MUH

[Dug96c] Randy Dugger. Managing users:

How’re your PR skills? Datamation, 42(15):30–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Dugger:1996:MUS

[Dug96d] Randy Dugger. Managing users:

Selling the decision maker. Datamation, 42(13):89–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Dugger:1997:MUF

[Dug97a] Randy Dugger. Managing users:

Face to face with users. Datamation, 43(2):29–??, ???? 1997.

CODEN DTMNAT. ISSN 0011-6963.

Dugger:1997:MUG

[Dug97b] Randy Dugger. Managing users: Guaranteed customer satisfaction. Datamation, 43(4):41–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Dugger:1997:MUH

[Dug97c] Randy Dugger. Managing users: How to outsource strategically. Datamation, 43(5):43–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Dugger:1997:MUL

[Dug97d] Randy Dugger. Managing users: Lower the risk of outsourcing. Datamation, 43(1):29–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Dugger:1997:MUS

[Dug97e] Randy Dugger. Managing users: Supporting your support people. Datamation, 43(3):39–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Durr:1994:FIS

[Dur94a] Michael Durr. Fast Imaging on Slow Networks. Datamation, 40 (24):61–??, December 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Durr:1994:MSV

[Dur94b] Michael Durr. Managing the Switch to Virtual Networks. Datamation, 40(11):65–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Durr:1994:RCM

[Dur94c] Michael Durr. Remote Connectivity Making You Crazy? Consolidate! Datamation, 40(24):51– ??, December 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Durr:1994:RNR

[Dur94d] Michael Durr. Remote Node or Remote Control? Datamation, 40(17):49–51, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Durr:1994:FCW

[Dur94e] Michael Durr. The Fastest, Cheapest Way To Link Remote Offices. Datamation, 40(18):44– 46, September 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Durr:1994:TRN

[Dur94f] Michael Durr. Tuning Remote Nodes in 10 Easy Steps. Datamation, 40(17):51–??, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Durr:1995:RSM

[Dur95] Michael Durr. Remote Site Management Tricks. Datamation, 41 (6):49–??, April 01, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Dyson:1990:SIG

[Dys90a] E. Dyson. Shearson’s image of groupware (electronic imaging). Datamation, 36(7):61, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Dyson:1990:WGG

[Dys90b] E. Dyson. Why groupware is gaining ground. Datamation, 36 (5):52–56, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Eldred:1995:SOB

[Eld95] Eric Eldred. Sun Opens Up the Bus Bottleneck. Datamation, 41 (21):83–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Ellinger:1988:HSM

[Ell88] R. S. Ellinger. A head start down the MAP/ TOP 3.0 road. Datamation, 34(20):63–64, 68, October 1988. CODEN DTMNAT. ISSN 0011-6963.

Emmett:1989:DPM

[Emm89a] A. Emmett. The dual-protocol modem duel (CCITT V.42). Datamation, 35(4):81–82, 86–88, February 1989. CODEN DTM-NAT. ISSN 0011-6963.

Emmett:1989:STM

[Emm89b] A. Emmett. Serving two masters (modems). Datamation, 35(14): 47–48, 52, July 1989. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1997:SDE

[ES97] Dennis Eskow and Michelle Schoenung. Snap-on datamarts end business chaos. Datamation, 43(6):72–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Esh:1989:BBN

[Esh89] T. Esh. Beyond the borders of the new Europe. Datamation, 35 (5):76/25–26, March 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Eskow:1997:YPS

[Esk97a] Dennis Eskow. Are you pay scales right? Datamation, 43 (7):102–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1997:HPW

[Esk97b] Dennis Eskow. Harvest profits from Web farms. Datamation, 43 (3):44–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1997:OAS

[Esk97c] Dennis Eskow. Our all-stars make money the new way. Datamation, 43(8):56–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1997:SPA

[Esk97d] Dennis Eskow. Starting point:

Another roadside attraction. Datamation, 43(3):5–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1997:SPC

[Esk97e] Dennis Eskow. Starting point:

Counting on technology. Datamation, 43(9):5–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1997:SPDb

[Esk97f] Dennis Eskow. Starting point:

Do the math. Datamation, 43 (11):5–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1997:SPH

[Esk97g] Dennis Eskow. Starting point: How am I doing? Datamation, 43(6):5–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1997:SPN

[Esk97h] Dennis Eskow. Starting point: Now Bill Gates changes all the business rules. Datamation, 43 (7):5–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1997:SPO

[Esk97i] Dennis Eskow. Starting point: Our new services are just for you. Datamation, 43(4):5–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1997:SPDa

[Esk97j] Dennis Eskow. Starting point: S-Day minus 6.7. Datamation, 43 (8):5–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1997:SPI

[Esk97k] Dennis Eskow. Starting point: The idiot’s guide to corporate IT leadership. Datamation, 43(5):5– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Eskow:1997:SPV

[Esk97l] Dennis Eskow. Starting point: Voodoo IT management and productivity gods. Datamation, 43 (2):5–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1997:SPW

[Esk97m] Dennis Eskow. Starting point: Welcome, Mr. Jones. Datamation, 43(10):5–??, ???? 1997.

CODEN DTMNAT. ISSN 0011-6963.

Eskow:1997:PMA

[Esk97n] Simon Eskow. Pump more out of applications. Datamation, 43 (5):78–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1998:SPP

[Esk98a] Dennis Eskow. Starting point:

Proof of prom. Datamation, 44 (1):5–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Eskow:1998:SPR

[Esk98b] Dennis Eskow. Starting point:

RIP, papermate. Datamation, 44 (2):5–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Etheridge:1988:EE

[Eth88]  J. Etheridge.  EDI in eu-

rope. Datamation, 34(19):44/15– 18, October 1988. CODEN DTMNAT. ISSN 0011-6963.

Etheridge:1989:IRR

[Eth89] J. Etheridge. Interest rate on the rise (european 1992 market, impact on banking). Datamation, 35(5):76/15, 26, March 1989. CODEN DTMNAT. ISSN 0011-6963.

Etheridge:1990:PDI

[Eth90a] J. Etheridge. Peugeot drives integration. Datamation, 36(1):88– 81, 88–84, January 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Etheridge:1990:PFM

[Eth90b] J. Etheridge. A proprietary faith for midrange users (operating systems). Datamation, 36 (9):102–104, May 1990. CODEN DTMNAT. ISSN 0011-6963.

Etsten:1998:TFS

[Ets98] Deborah Etsten. TCO finds the strike zone. Datamation, 44(1): 97–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Flynn:1996:HJM

[FC96] Jim Flynn and Bill Clarke. How Java makes network-centric computing real — Java is likely to have a profound impact on the way businesses manage application development by bringing about an entirely new model — the network-centric model — of C/S computing. Datamation, 42(5):42–43, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Fosdick:1988:DUS

[FGR88] H. Fosdick and L. Garcia-Rose. DB2 users stand up to be counted. Datamation, 34(20):45, 48, 54, 58, October 1988. CO-DEN DTMNAT. ISSN 0011-6963.

Frazier:1994:GRP

[FH94] Donald Frazier and Kris Herbst. Get Ready To Profit from the InfoBahn. Datamation, 40(10): 50–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Fisher:1989:DIT

[Fis89] M. J. Fisher. Digging out with image technology. Datamation, 35(8):18–21, 24, 26–27, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Fisher:1996:GRP

[Fis96a] Susan E. Fisher. Get ready for plug and play — PnP hardware and software are coming sooner than you think, and the Universal Serial Bus is going to help make it happen. Datamation, 42 (9):62–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Fisher:1996:VUT

[Fis96b] Susan E. Fisher. VPNs use tunneling to build business links — if you just want to give company insiders easy and relatively secure remote access to company data via the Internet, you can set up a virtual private network. Datamation, 42(11):66–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Flynn:1996:BII

[Fly96a] Jim Flynn. Battle for the IN-TERNET infrastructure — Microsoft and Netscape are fighting for the hearts and minds of Internet developers everywhere. the winner will determine how you develop apps for the Internet and your corporate intranet into the future. Datamation, 42 (9):28–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Flynn:1996:UWI

[Fly96b] Jim Flynn. Use the Web for imaging! Web add-ons from imaging vendors make basic document viewing possible, but sophisticated support is on hold until object standards are settled. Datamation, 42(11):62–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Farhoomand:1989:MCS

[FM89a] A. F. Farhoomand and M. Murphy. Managing computer security. Datamation, 34(26):67–68, January 1989. CODEN DTM-NAT. ISSN 0011-6963.

Francis:1989:NDP

[FM89b] B. Francis and J. McMullen. New demands on PC support. Datamation, 35(21):101–102, November 1989. CODEN DTMNAT. ISSN 0011-6963.

Freeman:1997:YGY

[FM97] Leland G. Freeman and C. Lawrence Meador. So you’ve got your Year 2000 project under control. But what about your suppliers, partners, and the other companies you depend on? If they don’t have their Y2K act together, they could take you down with them. Datamation, 43(1):40–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Focus:1995:RST

[Foc95] Michael Alexander Focus. The Real Security Threat: The Enemy Within. Datamation, 41 (13):30–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Foley:1988:PMI

[Fol88a] M. J. Foley. Post-MRP II: what comes next? Datamation, 34 (24):24, 32, 36, December 1988. CODEN DTMNAT. ISSN 0011-6963.

Foley:1988:TPP

[Fol88b] M. J. Foley. The traveling programmer’s popular show (third party programmers). Datamation, 34(16):63–65, August 1988. CODEN DTMNAT. ISSN 0011-6963.

Foley:1989:HVP

[Fol89] M. J. Foley. The high-volume printer option. Datamation, 35 (20):59–60, October 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Foley:1993:PFP

[Fol93] M. J. Foley. It’s a phone! it’s a fax! it’s a pen computer! Datamation, 39(8):79–80, April 1993. CODEN DTMNAT. ISSN 0011-6963.

Foley:1996:CSS

[Fol96a] Mary Jo Foley. Can server suites play in the enterprise? Microsoft’s back-office, IBM’s Project Eagle, Oracle’s InterOffice, and others are all coming on as bundled. Datamation, 42 (6):46–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Foley:1996:DSM

[Fol96b] Mary Jo Foley. Desktop suites moonlight as development tools — the latest releases of Office, SmartSuite, and WordPerfect suite are 32-bit, Internetenabled app builders and cool end-user office products, too. Datamation, 42(11):48–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Foley:1996:FWM

[Fol96c] Mary Jo Foley. Framework — while Microsoft still has a lot of growing up to do to fill the enterprise shoes of its competitors, it is on the right track. Datamation, 42(6):5–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Foley:1996:HND

[Fol96d] Mary Jo Foley. Hot new database technologies — the tide has shifted from getting data in to getting data out of a database. the Big Six database vendors are developing major new technologies to help users wring every bit of information out. Datamation, 42(15):44–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Foley:1996:OBA

[Fol96e] Mary Jo Foley. Objects for business apps are coming! slowly — as users and vendors continue their quests for more tailorable applications comprised of vertical business objects, early implementations are starting to debut. Datamation, 42(10):52–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Foley:1996:RYM

[Fol96f] Mary Jo Foley. Ready yet? Microsoft grows into the enterprise — the kids from Redmond have donned their tasseltopped loafers and think they’re ready to give you missioncritical, enterprise-quality systems and support. sure, NT and BackOffice are top notch. but what about things like reliability, availability, and serviceability? Datamation, 42(6):34–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Foley:1996:SCI

[Fol96g] Mary Jo Foley. Second chance for IBM? IBM’s delivering great hardware, operating systems, ools, and killer middleware. so why are all its long-term customers going. Datamation, 42 (13):34–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Foley:1996:SEU

[Fol96h] Mary Jo Foley. Stalking the elusive unified directory service — don’t expect IBM, Microsoft, or Novell to make your life easier by linking their e-mail, networking. Datamation, 42(9):48–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Foley:1996:SPS

[Fol96i] Mary Jo Foley. Starting point — should you give IBM a second chance? Datamation, 42(13):5– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Fosdick:1990:TSA

[Fos90a] H. Fosdick. Ten steps to AD/ cycle. Datamation, 36(23):59– 60, 64, December 1990. CODEN DTMNAT. ISSN 0011-6963.

Foss:1990:EWK

[Fos90b] W. B. Foss. Early wins are key to system success management. Datamation, 36(2):79–82, January 1990. CODEN DTM-NAT. ISSN 0011-6963.

Fleischer:1988:SYD

[FP88] M. Fleischer and S. Patel. Shaping up your data center. Datamation, 34(23):89–99, November 1988. CODEN DTMNAT. ISSN 0011-6963.

Floyd:1989:WBS

[FR89] B. Floyd and B. Ronen. Where best to system invest (office automation). Datamation, 35(22): 111, 114, November 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Francis:1988:PBU

[Fra88] B. Francis. PC back-up’s optical understudy. Datamation, 34(25): 57–60, December 1988. CODEN DTMNAT. ISSN 0011-6963.

Francis:1989:CPL

[Fra89a] B. Francis. Climbing the PC ladder career paths. Datamation, 35 (9):77–78, May 1989. CODEN DTMNAT. ISSN 0011-6963.

Francis:1989:CNS

[Fra89b] B. Francis. Competitiveness: a new standard. Datamation, 35 (14):53–55, July 1989. CODEN DTMNAT. ISSN 0011-6963.

Francis:1989:DD

[Fra89c] B. Francis. The desktop dimension. Datamation, 35(21):28–31, 34–35, 38, 40, November 1989. CODEN DTMNAT. ISSN 0011-6963.

Francis:1989:DTW

[Fra89d] B. Francis. Desktop tug of war:

OS/2 versus unix. Datamation, 35(4):18–21, 24–25, February 1989. CODEN DTMNAT. ISSN 0011-6963.

Francis:1989:FLN

[Fra89e] B. Francis. Frito lays a new IS bet (DSS). Datamation, 35(4): 75, 78, February 1989. CODEN DTMNAT. ISSN 0011-6963.

Francis:1989:PP

[Fra89f] B. Francis. The IS push for 386 power. Datamation, 35(6):39–44, March 1989. CODEN DTMNAT. ISSN 0011-6963.

Francis:1989:MWP

[Fra89g] B. Francis. Muddy waters (PC spreadsheets). Datamation, 35 (14):31–32, July 1989. CODEN DTMNAT. ISSN 0011-6963.

Francis:1989:JMA

[Fra89h] B. Francis. Not just the Mac-Maverick anymore? (Apple Macintosh). Datamation, 35(11):43– 47, June 1989. CODEN DTM-NAT. ISSN 0011-6963.

Francis:1989:WRN

[Fra89i] B. Francis. A wary reception to new PC bus? Datamation, 35 (21):57–58, 60, November 1989. CODEN DTMNAT. ISSN 0011-6963.

Francis:1989:WD

[Fra89j] B. Francis. Who’s on deck with OS/2? Datamation, 35(9):54–56, May 1989. CODEN DTMNAT. ISSN 0011-6963.

Francis:1990:CSM

[Fra90a] B. Francis. Client/server: the model for the 1990s. Datamation, 36(4):34–36, 38, 40, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Francis:1990:CVP

[Fra90b] B. Francis. A corporate vision for publishing. Datamation, 36(2): 65–67, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Francis:1990:KCS

[Fra90c] B. Francis. Kimberly-clark sees profits in IS. Datamation, 36 (10):53–54, May 1990. CODEN DTMNAT. ISSN 0011-6963.

Francis:1990:NVQ

[Fra90d] B. Francis. A new vision of quality control. Datamation, 36 (7):70–72, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Francis:1990:OFS

[Fra90e] B. Francis. Oil firms seek UNIX standards. Datamation, 36(18): 63–64, September 1990. CODEN DTMNAT. ISSN 0011-6963.

Francis:1990:MM

[Fra90f] B. Francis. OS/2’s multiuser migration. Datamation, 36(22):63– 64, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Francis:1990:SFD

[Fra90g] B. Francis. A secure future for diskless workstations. Datamation, 36(22):53–54, 56, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Francis:1990:SPR

[Fra90h] B. Francis. System PRO revs up networks. Datamation, 36(21): 30–32, 34, 38, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Francis:1990:UCF

[Fra90i] B. Francis. UNIX checks in at four seasons. Datamation, 36 (8):47–49, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Francis:1990:WTR

[Fra90j] B. Francis. What’s taking the risk out of RISC. Datamation, 36(2):61–62, 64, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Francis:1990:WMG

[Fra90k] B. Francis. Will multimedia get down to business? Datamation, 36(23):43–44, December 1990. CODEN DTMNAT. ISSN 0011-6963.

Francis:1991:MWM

[Fra91a] B. Francis. Make way for modular PCs. Datamation, 37(22): 43–44, November 1991. CODEN DTMNAT. ISSN 0011-6963.

Francis:1991:OCD

[Fra91b] B. Francis. OCR comes down to the desktop. Datamation, 37(18): 44, 46, September 1991. CODEN DTMNAT. ISSN 0011-6963.

Francis:1991:PBN

[Fra91c] B. Francis. Pen-based notebooks find their niche. Datamation, 37 (19):44, 46, October 1991. CO-DEN DTMNAT. ISSN 0011-6963.

Francis:1992:HSP

[Fra92a] B. Francis. How to secure a portable PC. Datamation, 38 (6):43–44, March 1992. CODEN DTMNAT. ISSN 0011-6963.

Francis:1992:PSG

[Fra92b] B. Francis. PC security grows up. Datamation, 38(22):61–62, 64, November 1992. CODEN DTMNAT. ISSN 0011-6963.

Francis:1993:AYA

[Fra93a] B. Francis. Automate your asset management! Datamation, 39 (11):47–48, 50, June 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Francis:1993:CRD

[Fra93b] B. Francis. CD-ROMs drive toward new standards. Datamation, 39(4):57–58, 60, February 1993. CODEN DTMNAT. ISSN 0011-6963.

Francis:1993:MMM

[Fra93c] B. Francis. MIS meets multimedia on the network. Datamation, 39(14):32, 34, July 1993. CODEN DTMNAT. ISSN 0011-6963.

Francis:1993:OSS

[Fra93d] B. Francis. Our secret shopper buys mail order PCs. Datamation, 39(10):41–43, May 1993. CODEN DTMNAT. ISSN 0011-6963.

Francis:1993:RDD

[Fra93e] B. Francis. Recover from distributed disasters. Datamation, 39(23):73–74, December 1993. CODEN DTMNAT. ISSN 0011-6963.

Francis:1993:SCS

[Fra93f] B. Francis. The search for client/ server security. Datamation, 39 (9):39–40, 43, May 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Francis:1993:TCV

[Fra93g] B. Francis. Tune in to cheaper videoconferencing. Datamation, 39(19):48–50, 51, October 1993. CODEN DTMNAT. ISSN 0011-6963.

Frank:1997:NRP

[Fra97] Diane Frank. The new ROI in point of sale. Datamation, 43 (11):73–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Freeman:1996:HMC

[Fre96a] Eva Freeman. How to move ecash around the Internet. Datamation, 42(16):58–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Freeman:1996:MLE

[Fre96b] Eva Freeman. Middleware: Link everything to anything. Datamation, 42(16):119–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Freeman:1997:BTD

[Fre97a] Eva Freeman. Birth of a terabyte data warehouse. Datamation, 43 (4):80–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Freeman:1997:DUD

[Fre97b] Eva Freeman. Datamining unearths dollars from data. Datamation, 43(7):84–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Freeman:1997:EDR

[Fre97c] Eva Freeman. Eye on desktop reporting tools. Datamation, 43 (6):94–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Freeman:1997:OAD

[Fre97d] Eva Freeman. Is OO app dev worth the cost? True objectoriented application development has been difficult to master. Do the results justify the steep learning curve, or is component-based development the better solution for your IS shop? Datamation, 43(3):82–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Freeman:1997:JHW

[Fre97e] Eva Freeman. Jackpot: How to win the migration game. Datamation, 43(3):76–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Freeman:1997:MAW

[Fre97f] Eva Freeman. Manufacturing applications: What’s the right recipe? Datamation, 43(8):60– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Freeman:1997:SCM

[Fre97g] Eva Freeman. Supply chain: Modeling makes the difference.

Datamation, 43(10):64–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Freeman:1997:TYB

[Fre97h] Eva Freeman. Turn your budgeting operation into a profit center. Datamation, 43(1):90–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Freeman:1997:PPP

[Fre97i] Leland G. Freeman. Pick a portfolio and a partner. Datamation, 43(1):46–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Freeman:1998:GYO

[Fre98a] Eva Freeman. Get out of your office! Datamation, 44(2):72–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Freeman:1998:NTH

[Fre98b] Eva Freeman. New tools help server farms blossom. Datamation, 44(2):86–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Fryer:1996:IMA

[Fry96] Bronwyn Fryer. Internet mania — ask just about any software company executive what the biggest IT trend of 1997 will be, and you’ll get a single answer: corporate intranets. Datamation, 42(12):54–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Francis:1989:IMT

[FS89] B. Francis and W. Schatz. IBM mends third-party fences. Datamation, 34(26):41–42, 44, January 1989. CODEN DTMNAT. ISSN 0011-6963.

Ferreira:1988:MTJ

[FT88] J. Ferreira and M. E. Treacy. It’s more than just laptops (sales automation management). Datamation, 34(21):127–131, November 1988. CODEN DTMNAT. ISSN 0011-6963.

Gadacz:1989:SGP

[Gad89] O. Gadacz. Steel giant pioneers Korean IS. Datamation, 35 (11):64g–h, June 1989. CODEN DTMNAT. ISSN 0011-6963.

Garry:1995:CSHb

[Gar95a] G. Garry. Client/server hits the road (middleware for remote client/server apps). Datamation, 41(14):43–45, August 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Garry:1995:CSHa

[Gar95b] Greg Garry. Client/Server Hits the Road. Datamation, 41(14): 43–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Garry:1995:WRJ

[Gar95c] Greg Garry. Windows95 Refuseniks Just Say No! Datamation, 41(16):69–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Garner:1998:SOY

[Gar98] Rochelle Garner. Strategic outsourcing: It’s your move. Datamation, 44(2):32–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Gershkoff:1990:MBG

[Ger90] I. Gershkoff. The make or buy game (software applications). Datamation, 36(4):73–77, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Gerber:1995:DMY

[Ger95] Cheryl Gerber. Demand more from your outsourcer (outsourcing vendor evaluation). Datamation, 41(5):62–??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Gerber:1996:EYD

[Ger96a] Cheryl Gerber. Excavate your data — datamining could be your no. 1 strategic weapon — and source of profits — in dissecting archival information. but it takes time to master. Datamation, 42(9):40–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Gerber:1996:WIF

[Ger96b] Cheryl Gerber. Why is IBM first in services? IBM Global Services has become the no. 1 systems integrator by supporting multivendor systems and translating its inhouse development into experience for its customers. Datamation, 42(13):74–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Grill:1989:CSS

[GFR89] E. Grill, J. Flittner, and W. Rausch. A CAX success story (CAD/ CAM/ CAE). Datamation, 35(6):16–23, March 1989.

CODEN DTMNAT. ISSN 0011-6963.

Goldstein:1988:SDR

[GH88] M. Goldstein and J. Hagel. Systems discontinuity: roadblock to strategic change. Datamation, 34(20):34–37, 42, October 1988. CODEN DTMNAT. ISSN 0011-6963.

Gibson:1989:IPC

[Gib89]  M. L. Gibson.  Implement-

ing the promise (computer-aided software engineering). Datamation, 35(3):65–67, February 1989. CODEN DTMNAT. ISSN 0011-6963.

GibbonsPaul:1997:LS

[Gib97a] Lauren Gibbons Paul. AS/400 is a life saver. Datamation, 43 (9):90–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

GibbonsPaul:1997:RFT

[Gib97b] Lauren Gibbons Paul. The right formula for training train your employees too early and they forget everything. Train them too late and they hammer the help desk. Datamation, 43(9):96–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

GibbonsPaul:1998:WPO

[Gib98] Lauren Gibbons Paul. What price ownership? CheckFree finds total cost of ownership pays off in its efficient new datacenter. Datamation, 44(1):88–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Gill:1991:FBU

[Gil91] P. J. Gill. Finally, a binary UNIX PC standard. Datamation, 36 (24):59–60, December 1991. CO-DEN DTMNAT. ISSN 0011-6963.

Gill:1995:NIN

[Gil95] Philip J. Gill. Needed: Integrated network and system management. Datamation, 41(17):55, 57, 59, 61, 63, September 1995. CODEN DTMNAT. ISSN 0011-6963.

Gill:1996:UTH

[Gil96a] Philip J. Gill. UNIX takes the high road — while NT looks to be a great success in the low-end of the server market, more UNIXbased enterprise servers are being used as IS managers turn away from mainframes and legacy applications. Datamation, 42(12): 58–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Gill:1996:UVG

[Gil96b] Philip J. Gill. UNIX vendors get server fever — as NT moves in on the PC server market, commercial UNIX vendors move upscale. Datamation, 42(11):90–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Goldman:1995:TTD

[Gol95] S. J. Goldman. ThinkWrap — Technology is developing a life of its own beyond the scope of the typical user. Datamation, 41(13): 84–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Goslin:1988:ESA

[Gos88] R. Goslin. Eight steps to applications engineering. Datamation, 34(21):59–62, 66, November 1988. CODEN DTMNAT. ISSN 0011-6963.

Graham:1991:DHS

[Gra91] C. Graham. Decentralizing HR software. Datamation, 37(15): 71–73, August 1991. CODEN DTMNAT. ISSN 0011-6963.

Gregory:1988:DML

[Gre88] E. Gregory. Database machines: the least-cost route? Datamation, 34(21):85–86, 90, November 1988. CODEN DTMNAT. ISSN 0011-6963.

Greenfield:1996:DLD

[Gre96] Larry Greenfield. Don’t let data warehousing gotchas getcha — running a data warehouse isn’t a job for the faint of heart, but avoiding these 10 pitfalls will make it easier. Datamation, 42 (5):76–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Greenblatt:1997:TNP

[Gre97] Ellen C. Greenblatt. Top new products of 1997. Datamation, 43(2):60–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Garcia-Rose:1990:MD

[GRF90] L. Garcia-Rose and H. Fosdick. The maturation of DB2. Datamation, 36(6):75–82, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Grindley:1989:CTC

[Gri89] K. Grindley. Cutting through the CASE hype. Datamation, 35(7):88–89, 88–10, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Guterl:1992:TMP

[Gut92] F. V. Guterl. Twin mainframes power lufthansa’s reservations. Datamation, 38(20):95–96, October 1992. CODEN DTMNAT. ISSN 0011-6963.

Guterl:1994:IED

[Gut94] Frederick V. Guterl. IT Expands Despite Bargain Basement Budgets. Datamation, 40(16):46–48, August 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Guterl:1995:MBGa

[Gut95a] Fred Guterl. Mainframes Are Breaking Out of the Glass House. Datamation, 41(11):34–??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Guterl:1995:MBGb

[Gut95b] Fred Guterl. Mainframes Are Breaking Out of the Glass House. Datamation, 41(11):34–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Guterl:1996:HMY

[Gut96] Fred Guterl. How to manage your outsourcer — when you hire an outsourcer, you’re inviting a subcontractor to become an integral part of your business. to get the most value, you’ll need to manage the relationship. here’s how. Datamation, 42(5):79–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Haber:1989:BNO

[Hab89] L. Haber. Balancing network options (communications security). Datamation, 35(15):63–64, August 1989. CODEN DTMNAT. ISSN 0011-6963.

Haber:1994:GRE

[Hab94] Lynn Haber. Get ready for easier PC support (plug and play technology). Datamation, 40(21):49– 50, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Haber:1995:CSFb

[Hab95a] L. Haber. Client/server financials: take a hard look at soft services. Datamation, 41(1):64, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Haber:1995:CSFa

[Hab95b] Lynn Haber. Client/Server Financials: Take a Hard Look at Soft Services. Datamation, 41(1): 64–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Haber:1995:SIA

[Hab95c] Lynn Haber. In Search of the Ideal App Server. Datamation, 41(3):63–??, February 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Haber:1996:CTB

[Hab96] Lynn Haber. Computer telephony: Big help for customer support — if your company is trying to distinguish itself through improved customer service, you should probably look into the next generation of telephone response-computer telephony integration. Datamation, 42(8):71–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Haber:1997:BTS

[Hab97] Lynn Haber. Building truck stops on the Web. Datamation, 43(7):110–??, ???? 1997. CO-DEN DTMNAT. ISSN 0011-6963.

Hackathorn:1995:DWE

[Hac95] Richard Hackathorn. Data Warehousing Energizes Your Enterprise. Datamation, 41(2):38–40, 43, 45, February 01, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Halper:1988:PGR

[Hal88] M. Halper. Portables get raves on the road. Datamation, 34(23): 55–58, November 1988. CODEN DTMNAT. ISSN 0011-6963.

Halvorsen:1990:WLL

[Hal90] J.-M. Halvorsen. What to look for in leasing. Datamation, 36 (6):88–94, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Halper:1996:UDS

[Hal96] Mark Halper. Universal databases are still on the bleeding edge — if you can digitize it, you can retrieve it, claim the database leaders. but users say marrying video, audio, image, and other data types with traditional bits and bytes is not going to be a snap. Datamation, 42(8):48–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Hallows:1997:PP

[Hal97] Jolyon Hallows. The politics of projects. Datamation, 43(11): 105–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hamilton:1990:CLS

[Ham90a] D. Hamilton. Can leasing software save you money 2Q. Datamation, 36(10):47–50, May 1990. CODEN DTMNAT. ISSN 0011-6963.

Hamilton:1990:UAP

[Ham90b] D. Hamilton. A united approach to PC purchasing. Datamation, 36(21):59–60, 62, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Hamilton:1994:DLC

[Ham94a] Dennis Hamilton. Don’t Let Client/Server Performance Gotchas Getcha. Datamation, 40 (21):39–40, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Hamilton:1994:GHC

[Ham94b] Dennis Hamilton. Get a Handle on Client/Server Performance. Datamation, 40(19):45–??, October 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Hamilton:1994:GYD

[Ham94c] Dennis Hamilton. Give Your Database Administrators More Leverage. Datamation, 40(23):

51–??, December 01, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Hamilton:1994:STH

[Ham94d] Dennis Hamilton. Stop Throwing Hardware at Performance! Datamation, 40(19):43–??, October 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Hamilton:1996:BPD

[Ham96]  Dennis Hamilton.  Build a

plan for database performance — your company’s databases have evolved from mere data stores to profit centers. how are you going to make sure. Datamation, 42 (15):71–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Haskin:1995:VPS

[Has95] Denis W. Haskin. VisualWorks:

Powerful — but Smalltalk Underneath. Datamation, 41(19): 81–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Hayashi:1996:CAR

[Hay96a] Alden M. Hayashi. Is corporate America ready for e-commerce? Not quite. Cost, security issues, and lack of mature software are major concerns. Still, many IS managers are plowing full-steam ahead with e-commerce systems. Datamation, 42(16):54–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Hayashi:1996:SPW

[Hay96b] Alden M. Hayashi. Starting point: Why you should rethink reengineering. Datamation, 42 (18):5–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Hayashi:1996:WWV

[Hay96c] Alden M. Hayashi. When will VisualAge come of age? IBM’s application development product line including the VisualAge family, has been garnering rave reviews. so, why. Datamation, 42 (13):58–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Hayashi:1996:WIL

[Hay96d] Alden M. Hayashi. Will IBM lose the PC price war? IBM’s adherence to IBM-developed hardware and software architectures has taken a toll. Datamation, 42 (13):46–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Hayes:1996:TDA

[Hay96e] Linda Hayes. Testing distributed applications: Unraveling the Web — as Internet use continues to surge, companies have been furiously working on adding more and more functionality to their Web sites. not many companies, however, know how. Datamation, 42(13):108–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Hayashi:1997:SPIb

[Hay97a] Alden M. Hayashi. Squeezing profits from IT. Datamation, 43 (7):42–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayashi:1997:SPIa

[Hay97b] Alden M. Hayashi. Starting point: If you haven’t project yet. Datamation, 43(1):5–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayashi:1997:TITh

[Hay97c] Alden M. Hayashi. Thinking it through. Datamation, 43(11):42– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Hayashi:1997:TITf

[Hay97d] Alden M. Hayashi. Thinking it through: a mercenary mentality. Datamation, 43(9):36–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayashi:1997:TITg

[Hay97e] Alden M. Hayashi. Thinking it through: An IT rashomon. Datamation, 43(10):34–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayashi:1997:TITc

[Hay97f] Alden M. Hayashi. Thinking it through: Are you prepared to embrace chaos? Datamation, 43 (6):34–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayashi:1997:TITe

[Hay97g] Alden M. Hayashi. Thinking it through: Of IT and anesthesia. Datamation, 43(8):33–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayashi:1997:TITb

[Hay97h] Alden M. Hayashi. Thinking it through: The dangerous secret of the web. Datamation, 43(5):41– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Hayashi:1997:TITa

[Hay97i] Alden M. Hayashi. Thinking it through: The testosterone test. Datamation, 43(4):38–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayashi:1997:TITd

[Hay97j] Alden M. Hayashi. Thinking it through: Why your boss says “no”. Datamation, 43(7):33–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayes:1997:QQB

[Hay97k] Linda Hayes. Quality quest: Boost your test team’s value. Datamation, 43(5):46–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayes:1997:QQD

[Hay97l] Linda Hayes. Quality quest: Don’t let your business processes regress. Datamation, 43(7):34– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Hayes:1997:QQF

[Hay97m] Linda Hayes. Quality quest: Forget about “quality”. Datamation, 43(8):31–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayes:1997:QQP

[Hay97n] Linda Hayes. Quality quest: Process of perish. Datamation, 43 (11):29–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayes:1997:QQTc

[Hay97o] Linda Hayes. Quality quest: Testing: No easy way out. Datamation, 43(6):39–??, ???? 1997.

CODEN DTMNAT. ISSN 0011-6963.

Hayes:1997:QQH

[Hay97p] Linda Hayes. Quality quest:

The high cost of low investment. Datamation, 43(10):38–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayes:1997:QQTd

[Hay97q] Linda Hayes. Quality quest: The three faces of testing. Datamation, 43(9):32–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayes:1997:QQTb

[Hay97r] Linda Hayes. Quality quest:

The truth about automated test tools. Datamation, 43(4):45–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayes:1997:QQTa

[Hay97s] Linda Hayes. Quality quest: The tyranny of The Schedule. Datamation, 43(3):29–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayes:1997:YSC

[Hay97t] Linda Hayes. The Year 2000 and the S&L crisis. Datamation, 43 (1):75–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Hayashi:1998:TIT

[Hay98a] Alden M. Hayashi. Thinking it through. Datamation, 44(1):46– ??, ???? 1998. CODEN DTM-NAT. ISSN 0011-6963.

Hayes:1998:QQB

[Hay98b] Linda Hayes. Quality quest: The big lie. Datamation, 44(1):48–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Hughes:1990:SCU

[HC90] C. T. Hughes and J. D. Clark. The stages of CASE usage. Datamation, 36(3):41–44, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Hershey:1990:WEF

[HE90] G. L. Hershey and J. L. Eatman. Why IS execs feel left out of big decisions. Datamation, 36 (10):97–99, May 1990. CODEN DTMNAT. ISSN 0011-6963.

Healy:1989:IPA

[Hea89]  T. Healy.  Impressive per-

formance (australian data centres). Datamation, 35(19):72/9– 10, October 1989. CODEN DTMNAT. ISSN 0011-6963.

Healey:1990:DLE

[Hea90a] M. Healey. Diversity at the low end (PC/workstation market). Datamation, 36(7):100/7–8, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Healey:1990:PPW

[Hea90b] M. Healey. A power play for workstations. Datamation, 36(2): 83–85, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Hebditch:1989:RS

[Heb89] D. Hebditch. To Russia with software. Datamation, 35(23):

78/3–4, December 1989. CODEN DTMNAT. ISSN 0011-6963.

Hebditch:1990:OCO

[Heb90] D. Hebditch. Opening up CIM opportunities. Datamation, 36 (11):111–113, June 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Hecht:1996:HDM

[Hec96] Bradley Hecht. The help desk market: disarray and disrepair. Datamation, 42(18):57–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Hecht:1997:CRE

[Hec97] Bradley Hecht. Choose the right ERP software. Datamation, 43 (3):56–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Henry:1990:MBV

[Hen90] B. Henry. Measuring IS for business value. Datamation, 36 (7):89–91, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Henderson:1994:TQA

[Hen94] Tom Henderson. Tough Questions to Ask a Client/Server Integrator. Datamation, 40(16):59– 61, August 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Henderson:1995:TIW

[Hen95] Paul Henderson. ThinkWrap — If we don’t watch out, U.S.-made software is going to be about as reliable as the Chevy Vega of the ’70s. Datamation, 41(23):84–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Henning:1996:IYF

[Hen96] Jeffrey Henning. Intel in your face — with faster, cheaper microprocessors coming from Intel every three months, IS managers, who once turned to RISC workstations for. Datamation, 42(12): 44–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Herbst:1990:MOM

[Her90] K. Herbst. A more open market for supercomputers. Datamation, 36(18):123–125, September 1990. CODEN DTMNAT. ISSN 0011-6963.

Hindus:1992:WMS

[Hin92] L. A. Hindus. What MIS should know about videoconferencing. Datamation, 38(17):60–62, 64, August 1992. CODEN DTM-NAT. ISSN 0011-6963.

Hindus:1993:TFC

[Hin93] L. A. Hindus. Tardy fax choice can waste cash. Datamation, 39 (5):94, 96, 98, March 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Hirschfield:1996:SSS

[Hir96] Rebecca Hirschfield. Shared services save big money — you may soon be called on to support a new trend in human resource administration — shared service centers. or you may even be asked to join one. better know what it is. Datamation, 42(15):76–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Hodges:1988:NMC

[Hod88] P. Hodges. The new maturity of computer science. Datamation, 34(18):37–40, September 1988. CODEN DTMNAT. ISSN 0011-6963.

Hodges:1989:BGS

[Hod89a] P. Hodges. IS budget growth slides. Datamation, 35(7):18–22, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Hodges:1989:LGS

[Hod89b] P. Hodges. LAN growth surges.

Datamation, 35(8):32–34, 36, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Hodges:1989:MAP

[Hod89c] P. Hodges. Manufacturing automation’s problem. Datamation, 35(22):32–35, 38, November 1989. CODEN DTMNAT. ISSN 0011-6963.

Hodges:1989:RSD

[Hod89d] P. Hodges. A relational successor? (databases). Datamation, 35(21):47–48, 50, November 1989. CODEN DTMNAT. ISSN 0011-6963.

Hodges:1990:ADB

[Hod90] P. Hodges. The application decade begins. Datamation, 36 (2):22–25, 28–30, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Hoke:1996:LTC

[Hok96] Gordon E. J. Hoke. LAN-to-As/400: a trio of connections — both IBM and Microsoft have products to link Windows desktops to the AS/400, as do several third-parties, like. Datamation, 42(6):98–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Hackathorn:1994:HPC

[HS94] Richard D. Hackathorn and Mark Schlack. How To Pick Client/Server Middleware. Datamation, 40(14):52–54, 56, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Huggins:1990:FVL

[Hug90a] T. Huggins. 1993 forges vendor links. Datamation, 36(5):95–96, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Huggins:1990:HPF

[Hug90b] T. Huggins. 1993 holds promise but few guarantees. Datamation, 36(5):91–93, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Humphrey:1989:ISD

[Hum89] W. Humphrey. Improving the software development process. Datamation, 35(7):28–30, 52, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Hurwicz:1994:WHA

[Hur94a] M. Hurwicz. WANs: how to add branches without going bankrupt. Datamation, 40(6):42– 44, March 1994. CODEN DTM-NAT. ISSN 0011-6963.

Hurwicz:1994:CYW

[Hur94b] Michael Hurwicz. Check Your Wallet Before Buying Pentium.

Datamation, 40(10):71–72, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Hurwicz:1994:SEP

[Hur94c] Michael Hurwicz. In Sight: Enterprise PC Management. Datamation, 40(15):53–54, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Hurwicz:1994:HIU

[Hur94d] Mike Hurwicz. How To Integrate UNIX Servers and Net-Ware LANs. Datamation, 40 (23):57–??, December 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Hurwicz:1995:NEG

[Hur95a] Michael Hurwicz. NetWare 4.1 Eases Global Directory Tensions. Datamation, 41(1):53–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Hurwicz:1995:LNM

[Hur95b] Mike Hurwicz. Lower the Net Management Burden Now. Datamation, 41(18):53–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Hurwicz:1995:NNW

[Hur95c] Mike Hurwicz. NetWare and NT:

When Servers Collide. Datamation, 41(21):51–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Hurwicz:1996:BNW

[Hur96] Mike Hurwicz. Blue Notes: Is 4.0 worth the update hassle? not so long ago, Notes was pretty much the only game in town for groupware. now the Internet and the World Wide Web have some users questioning whether they should stick with Notes. here’s how Lotus is meeting the challenge. Datamation, 42(3):57–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Inmon:1997:MDW

[Inm97a] W. H. Inmon. Are multiple data warehouses too much of a good thing? Certain types of data warehouses go well together and certain types don’t. If you make the right choice, you can achieve a symbiosis that will give your organization a competitive advantage. Datamation, 43(4):94–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Inmon:1997:DYD

[Inm97b] W. H. Inmon. Does your datamart vendor care about your architecture? Your enterprise is hungry for the benefits of a datamart, but beware: Your datamart vendor might be trying to sell you a seven-course meal of Hershey bars. Datamation, 43 (3):105–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Jager:1996:HBY

[Jag96a] Peter De Jager. How big is your year 2000 mess? A systems inventory will tell you which systems have bad dates in them. an impact analysis will tell you how big a risk those systems are to your business and get you started on developing a plan to fix the problem. Datamation, 42(15): 109–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Jager:1996:TYT

[Jag96b] Peter De Jager. Take the Year 2000 test! Will your computer systems function when the new century begins? Test your systems now or die the death of a thousand cuts. Datamation, 42 (16):76–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Jager:1997:MYO

[Jag97] Peter De Jager. Manage your own expectations. Datamation, 43(1):69–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

James:1997:CS

[Jam97a] Geoffrey James. Cover story.

Datamation, 43(8):36–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

James:1997:IFH

[Jam97b] Geoffrey James. IT fiascos… and how to avoid them. Datamation, 43(11):84–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

James:1997:IHH

[Jam97c] Geoffrey James. IT helps HR become strategic. Datamation, 43 (4):110–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Jefferson:1997:CYT

[Jef97a] Steve Jefferson. Connect your telecommuters. Datamation, 43 (10):112–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Jefferson:1997:CPY

[Jef97b] Steve Jefferson. Cut and paste your way to a powerful prototype. Datamation, 43(10):104– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Jefferson:1997:FF

[Jef97c] Steve Jefferson. Filling up fiber. Datamation, 43(10):98–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Jefferson:1997:VS

[Jef97d] Steve Jefferson. Vital signs. Datamation, 43(10):54–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Jefferson:1997:WYS

[Jef97e] Steve Jefferson. When your system’s gushing data. Datamation, 43(9):56–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Jefferson:1998:SR

[Jef98] Steve Jefferson. Selling on the run. Datamation, 44(1):108–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Jenkins:1989:LRA

[Jen89a] A. Jenkins. The longer reach of AppleTalk. Datamation, 35(15): 59–60, August 1989. CODEN DTMNAT. ISSN 0011-6963.

Jenkins:1989:NPI

[Jen89b] A. Jenkins. A network puzzle (IBM systems network architecture). Datamation, 35(19):63–64, October 1989. CODEN DTM-NAT. ISSN 0011-6963.

Jenson:1991:TDI

[Jen91] T. Jenson. From telephone to database (interactive voice response systems). Datamation, 37 (25):46–48, December 1991. CO-DEN DTMNAT. ISSN 0011-6963.

Johnson:1989:FF

[Joh89] R. C. Johnson. That fuzzy feeling. Datamation, 35(14):39, 42– 43, July 1989. CODEN DTM-NAT. ISSN 0011-6963.

Johnson:1990:HFQ

[Joh90a] J. R. Johnson. Hallmark’s formula for quality (software quality assurance). Datamation, 36 (4):119–120, 122, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Johnston:1990:BTN

[Joh90b] M. Johnston. Betting on a transaction network (horse racing). Datamation, 36(9):99–101, May 1990. CODEN DTMNAT. ISSN 0011-6963.

Johnston:1990:WJI

[Joh90c] M. Johnston. Why Japan isn’t the hottest IT spot in Asia. Datamation, 36(12):148–153, 157– 161, 163, 165, 167, 170–171, 174, 176, June 1990. CODEN DTM-NAT. ISSN 0011-6963.

Jones:1997:YWR

[Jon97a] Capers Jones. Year 2000: What’s the real cost? U.S. organizations will spend USD 71 billion to make their applications Year 2000 compliant. And that figure doesn’t. . .. Datamation, 43 (3):88–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Jones:1997:OSF

[Jon97b] Dennis H. Jones. Overnight success: Four targets for success. Datamation, 43(10):132–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Jones:1997:OSC

[Jon97c] Dennis H. Jones. Overnight success: The 21st Century work force. Datamation, 43(11):126– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Jones:1998:OSB

[Jon98] Dennis H. Jones. Overnight success: Best wishes for 1998. Datamation, 44(1):140–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Joyce:1988:IHD

[Joy88a] E. J. Joyce. Innovative in-house development. Datamation, 34 (20):81–83, October 1988. CO-DEN DTMNAT. ISSN 0011-6963.

Joyce:1988:RSP

[Joy88b] E. J. Joyce. Reusable software: passage to productivity? Datamation, 34(18):97–102, September 1988. CODEN DTMNAT. ISSN 0011-6963.

Joyce:1989:DHP

[Joy89a] E. J. Joyce. Dialing for help. as the PC software field narrows, customer support is becoming more competitive as users demand better, faster, and more service than ever before. Datamation, 34(26):53–55, January 1989. CODEN DTMNAT. ISSN 0011-6963.

Joyce:1989:SCN

[Joy89b] E. J. Joyce. A system for the 21st century (NASA system development). Datamation, 35(10):51– 54, May 1989. CODEN DTM-NAT. ISSN 0011-6963.

Joyce:1989:EFS

[Joy89c] J. Joyce. Is error-free software achievable? Datamation, 35(4): 53, 56, February 1989. CODEN DTMNAT. ISSN 0011-6963.

Juan:1990:SSS

[Jua90] E. San Juan. Systematizing Spanish semantics (IBM’s Technological Centre for Languages). Datamation, 36(18):127, September 1990. CODEN DTM-NAT. ISSN 0011-6963.

Judge:1990:IDS

[Jud90a] P. Judge. IT delivers at securicor. Datamation, 36(7):100/1–3, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Judge:1990:UPE

[Jud90b] P. Judge. UNIX ports in europe.

Datamation, 36(3):78–80, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Kador:1997:ARQ

[Kad97] John Kador. Ask the right questions. Datamation, 43(6):59–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Kaufman:1996:LOC

[Kau96] Lar Kaufman. Linux opens corporate IS doors — most corporate IS shops turn up their noses at shareware, but Linux is good code for free. well, almost. Datamation, 42(7):72–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Kaye:1990:FEI

[Kay90] J. Kaye. Feeding Europe’s info appetite. Datamation, 36(1): 85–86, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Kay:1995:BCMc

[Kay95a] A. S. Kay. The business case for multimedia. Datamation, 41 (11):55–56, June 1995. CODEN DTMNAT. ISSN 0011-6963.

Kay:1995:BCMa

[Kay95b] Alan Kay. The Business Case for Multimedia. Datamation, 41(11): 55–??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Kay:1995:NMSa

[Kay95c] Alan S. Kay. Networked Multimedia Still a Challenge. Datamation, 41(11):56–??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Kay:1995:NMSb

[Kay95d] Alan S. Kay. Networked multimedia Still a Challenge. Datamation, 41(11):56–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Kay:1995:BCMb

[Kay95e] Alan S. Kay. The Business Case for Multimedia. Datamation, 41 (11):55–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Kay:1995:WWF

[Kay95f] Alan S. Kay. Where Warp Fits.

Datamation, 41(19):75–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Kay:1995:CCW

[Kay95g] E. Kay. Control the cost (windows95). Datamation, 41(16):38– 41, September 1995. CODEN DTMNAT. ISSN 0011-6963.

Kay:1995:CC

[Kay95h] Emily Kay. Control the Cost.

Datamation, 41(16):38–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Kay:1995:PPE

[Kay95i] Emily Kay. PowerBuilder: Power Enough for Complex Apps? Datamation, 41(5):49–??, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Kay:1995:SEI

[Kay95j] Emily Kay. Selling Enters the Information Age. Datamation, 41 (8):38–42, May 01, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Kay:1995:SYO

[Kay95k] Emily Kay. Should You Outsource Your Move to Windows95? Datamation, 41(20):55– 56, 59, November 1995. CODEN DTMNAT. ISSN 0011-6963.

Kay:1995:SYC

[Kay95l] Emily Kay. Surprise: Your Client Is Sinking Your Server. Datamation, 41(2):51–??, February 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Kay:1996:LOT

[Kay96a] Alan S. Kay. Land of opportunity and turmoil — chances are good that datamarts will pop up at your IS site. but what are they? for starters, they offer most of the benefits of data warehousing without a lot of the pain and agony. Datamation, 42(12):40– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Kay:1996:DSS

[Kay96b] Emily Kay. Desperately seeking SAP support — need help with SAP? you’re not alone. and there aren’t enough qualified consultants to go around. here’s how to avoid high rates and low levels of experience. Datamation, 42 (4):42–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Kay:1996:HCH

[Kay96c] Emily Kay. How to choose a help desk outsourcer — outsourcers, consultants, system integrators, even computer vendors are falling all over each other to offer you help desk outsourcing services. that makes it a dangerous place to shop. Datamation, 42(7):65–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Kay:1996:MYH

[Kay96d] Emily Kay. Make your help desk network-aware. Datamation, 42 (1):61–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Kay:1996:PPD

[Kay96e] Emily Kay. Pentium Pros on the desktop? not yet — if you’re thinking about putting PCs powered with Intel’s Pentium Pro on your users’ desktops, think about it a little more. Datamation, 42 (7):60–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Kay:1996:SYO

[Kay96f] Emily Kay. Should you outsource your Intranet? despite some hiccups, early adopters say Network Notes and other private data network services are helping control cost and complexity for remote sites and intraenterprise computing. Datamation, 42(3):37–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Kay:1997:HSI

[Kay97] Alan S. Kay. High speed Internet access: What’s the payoff? Datamation, 43(9):106–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Kay:1998:FP

[Kay98] Emily Kay. Flexed pricing. Datamation, 44(2):58–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Keen:1997:RAH

[Kee97a] Jack M. Keen. ROI: Aim high to get upper-level management approval. Datamation, 43(8):34–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Keen:1997:RSF

[Kee97b] Jack M. Keen. ROI: The six faces of business cases. Datamation, 43 (11):40–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Keen:1997:RTSb

[Kee97c] Jack M. Keen. ROI: Turn setbacks into successes. Datamation, 43(10):32–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Keen:1997:RTSa

[Kee97d] Jack M. Keen. ROI: Turn “soft” benefits into hard savings. Datamation, 43(9):38–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Keen:1998:RBB

[Kee98a] Jack M. Keen. ROI: Benefit’s biggest hits. Datamation, 44 (2):30–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Keen:1998:RFS

[Kee98b] Jack M. Keen. ROI: Fallow Sherlock’s lead to a better ROI. Datamation, 44(1):44–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Kelly:1989:TMS

[Kel89] J. Kelly. Three markets shape one industry (DP industry). Datamation, 35(12):7–29, June 1989. CODEN DTMNAT. ISSN 0011-6963.

Keller:1990:MMM

[Kel90a] W. J. Keller. Has the mini met its match? Datamation, 36(9):62– 66, May 1990. CODEN DTM-NAT. ISSN 0011-6963.

Kelly:1990:ASD

[Kel90b] J. Kelly. Automated storage draws interest. Datamation, 36 (4):81–82, 86, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Kelly:1990:ITS

[Kel90c] J. Kelly. Information technology sales soar to US$256B. Datamation, 36(12):22–28, 32–35, June 1990. CODEN DTMNAT. ISSN 0011-6963.

Kelly:1990:OWP

[Kel90d] J. Kelly. Outsourcing: who pulls the strings? (DP cost control). Datamation, 36(18):103– 104, 106, September 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Kerr:1988:CIR

[Ker88a] S. Kerr. Can IBM revive the RT PC? Datamation, 34(21):95–98, November 1988. CODEN DTM-NAT. ISSN 0011-6963.

Kerr:1988:DIR

[Ker88b] S. Kerr. Document interchange reigns at DOD. Datamation, 34 (23):81–84, November 1988. CO-DEN DTMNAT. ISSN 0011-6963.

Kerr:1988:TB

[Ker88c] S. Kerr. IS trims down (IS budgets). Datamation, 34(19):46–49, October 1988. CODEN DTM-NAT. ISSN 0011-6963.

Kerr:1988:PNM

[Ker88d] S. Kerr. The politics of network management. Datamation, 34 (18):50–54, September 1988. CO-DEN DTMNAT. ISSN 0011-6963.

Kerr:1989:CGD

[Ker89a] S. Kerr. The coming global directory for E-mail. Datamation, 35(22):105–108, November 1989. CODEN DTMNAT. ISSN 0011-6963.

Kerr:1989:IFP

[Ker89b] S. Kerr. IBM’s fixed price pitch (maintenance). Datamation, 35 (7):59–64, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Kerr:1989:INA

[Ker89c] S. Kerr. IBM’s NFS alternative. Datamation, 34(26):63–65, January 1989. CODEN DTMNAT. ISSN 0011-6963.

Kerr:1989:LLF

[Ker89d] S. Kerr. Legal laissez-faire (electronic data interchange). Datamation, 35(8):54, 56, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Kerr:1989:NCI

[Ker89e] S. Kerr. New CICS out of IBM. Datamation, 35(3):59–60, 62, February 1989. CODEN DTMNAT. ISSN 0011-6963.

Kerr:1989:NFC

[Ker89f] S. Kerr. The new IS force (corporate cultures). Datamation, 35 (15):18–23, August 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Kerr:1989:PCC

[Ker89g] S. Kerr. Playing in concert (corporate LANs). Datamation, 35 (11):55–57, June 1989. CODEN DTMNAT. ISSN 0011-6963.

Kerr:1989:SNM

[Ker89h] S. Kerr. A secret no more (security and encryption). Datamation, 35(13):53–55, July 1989. CODEN DTMNAT. ISSN 0011-6963.

Kerr:1989:TPE

[Ker89i] S. Kerr. Telecom pros evolve.

Datamation, 35(2):45, 48, January 1989. CODEN DTMNAT. ISSN 0011-6963.

Kerr:1989:TCP

[Ker89j] S. Kerr. A tide of change in the Pacific (telecommunications). Datamation, 35(6):55–62, March 1989. CODEN DTMNAT. ISSN 0011-6963.

Kerr:1989:TMG

[Ker89k] S. Kerr. Tl’s middle ground gets filled. Datamation, 35(5):61–62, 66, 68, March 1989. CODEN DTMNAT. ISSN 0011-6963.

Kerr:1989:WDL

[Ker89l] S. Kerr. What’s developing in large systems? (user spending survey). Datamation, 35(11):18– 26, June 1989. CODEN DTM-NAT. ISSN 0011-6963.

Kerr:1990:AWB

[Ker90a] S. Kerr. The applications wave behind ISDN. Datamation, 36 (3):64–66, February 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Kerr:1990:UAI

[Ker90b] S. Kerr. Using AI to improve security. Datamation, 36(3):57– 58, 60, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Kerr:1990:VNB

[Ker90c] S. Kerr. Virtual net battle yields user savings. Datamation, 36(2): 71–73, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Keyes:1989:SDU

[Key89] J. Keyes. Systems design with the user in mind. Datamation, 35(7):48–50, 55, 56, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Kight:1996:SWB

[Kig96] Betty Kight. The smart way to build a data warehouse. Datamation, 42(16):91–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Kim:1990:DOD

[Kim90a] Won Kim. Defining object databases anew. Datamation, 36(3): 33–34, 36, February 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Kim:1990:NDN

[Kim90b] Won Kim. A new database for new times. Datamation, 36(2):

35–36, 40, 42, January 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Kirkley:1989:NMJ

[Kir89] J. Kirkley. No magic, just hard work (DP management). Datamation, 35(10):67–73, May 1989. CODEN DTMNAT. ISSN 0011-6963.

Keene:1988:SSP

[KJH88] C. Keene, P. Jessel, and J. Hagel.

Solving the systems puzzle. 3 (systems integrators). Datamation, 34(21):119–125, November 1988. CODEN DTMNAT. ISSN 0011-6963.

Katz:1995:CRT

[KK95] Peter L. Katz and Cynthia J.

Katz. Choose the Right Training Strategy. Datamation, 41(23): 45–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Kniskern:1990:EVS

[Kni90] J. Kniskern. Engineering a visionary solution (image processing). Datamation, 36(8):90–91, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1996:MPI

[Kno96a] Anne Knowles. Microsoft plays Internet catch-up — Microsoft has announced a bevy of tools that, when released, will help users build, deploy, and manage MS Internet systems based on Microsoft apps. Datamation, 42 (6):75–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1996:SWJ

[Kno96b] Anne Knowles. Second-wave Java tools tuned to perform. Datamation, 42(18):82–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1996:CEB

[Kno96c] James H. Knowles. Competing effectively: Build an IT architecture on a business foundation. Datamation, 42(13):25–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1996:CEL

[Kno96d] James H. Knowles. Competing effectively: Learn to bungee jump. Datamation, 42(16):29–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1996:CEE

[Kno96e] Jim Knowles. Competing effectively: Explore data warehousing. Datamation, 42(18):30–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1996:CEM

[Kno96f] Jim Knowles. Competing effectively: Make the holes bigger. Datamation, 42(15):34–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1997:CWT

[Kno97a] Anne Knowles. Can we talk? Datamation, 43(9):76–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1997:CUB

[Kno97b] Anne Knowles. Copy this! A USD 3.2 billion outsourcing deal could have killed the career of Xerox CIO Patricia Wallington. But “team Xerox” is so real, she had the . . .. Datamation, 43 (6):42–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1997:DHB

[Kno97c] Anne Knowles. A database with a head for business. Datamation, 43(9):62–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1997:DYD

[Kno97d] Anne Knowles. Dump your dirty data for added profits. Datamation, 43(9):80–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1997:ECN

[Kno97e] Anne Knowles. Euro: Creation of a new currency. Datamation, 43(11):80–??, ???? 1997. CO-DEN DTMNAT. ISSN 0011-6963.

Knowles:1997:FWY

[Kno97f] Anne Knowles. Find out what your customers really want. Datamation, 43(2):68–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1997:GCP

[Kno97g] Anne Knowles. Get the complete picture. Datamation, 43(10):74– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Knowles:1997:WRM

[Kno97h] Anne Knowles. Where the rubber meets the road. Datamation, 43(8):49–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1997:CEA

[Kno97i] Jim Knowles. Competing effectively: Average doesn’t win customers. Datamation, 43(3):41– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Knowles:1997:CEB

[Kno97j] Jim Knowles. Competing effectively: Buy versus build in six words. Datamation, 43(1):31–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1997:CERa

[Kno97k] Jim Knowles. Competing effectively: Raising the Internet ante. Datamation, 43(4):35–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1997:CERb

[Kno97l] Jim Knowles. Competing effectively: Ready or not, here come net transactions. Datamation, 43 (5):38–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Knowles:1997:CES

[Kno97m] Jim Knowles. Competing effectively: Seeing through the browser war. Datamation, 43 (2):31–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Koland:1994:CSM

[Kol94] Cordell Koland. It’s a Cinch To Synch Mobile User’s Files.

Datamation, 40(20):67–??, October 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Korzeniowski:1993:MBA

[Kor93] P. Korzeniowski. Moving business abroad? take a VAN! Datamation, 39(3):80/13–15, February 1993. CODEN DTMNAT. ISSN 0011-6963.

Korzeniowski:1994:DSS

[Kor94] Paul Korzeniowski. Desperately Seeking Storage Solutions. Datamation, 40(16):62–64, August 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Krepchin:1993:FR

[Kre93a] I. Krepchin. A faster, roomier AS/400. Datamation, 39(9):89– 90, May 1993. CODEN DTM-NAT. ISSN 0011-6963.

Krepchin:1993:ISS

[Kre93b] I. Krepchin. Imaging: start small, think big. Datamation, 39 (13):47–48, July 1993. CODEN DTMNAT. ISSN 0011-6963.

Krepchin:1993:WCM

[Kre93c] I. Krepchin. Will cheaper mainframes slow downsizing? Datamation, 39(19):63–65, October 1993. CODEN DTMNAT. ISSN 0011-6963.

Kimball:1994:WDS

[KS94] Ralph Kimball and Kevin Strehlo. Why Decision Support Fails and How To Fix It. Datamation, 40(11):40–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Lamb:1990:EGA

[Lam90] J. Lamb. Europe’s grand IS alliances. Datamation, 36(3):81, 83, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Lazar:1998:DD

[Laz98] Gerald Lazar. Dash from the desktop. Datamation, 44(2):82– ??, ???? 1998. CODEN DTM-NAT. ISSN 0011-6963.

Lee:1995:BGP

[Lee95] The Lee. Behind Good Products Are Methods that Work. Datamation, 41(4):63–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Leibowitz:1990:UWA

[Lei90a] M. R. Leibowitz. UNIX workstations arrive! Datamation, 36(11): 24–27, 30, June 1990. CODEN DTMNAT. ISSN 0011-6963.

Leinfuss:1990:DER

[Lei90b] E. Leinfuss. Digital’s EDI remedy. Datamation, 36(19):63–73, October 1990. CODEN DTM-NAT. ISSN 0011-6963.

Leinfuss:1990:MTA

[Lei90c] E. Leinfuss. Making terminal access dynamic (IBM’s virtual telecommunications access method). Datamation, 36(4):99– 100, 102, 104, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Leinfuss:1990:RDF

[Lei90d] E. Leinfuss. R and D without the fee (business/academia relationships). Datamation, 36(9):93–95, May 1990. CODEN DTMNAT. ISSN 0011-6963.

Leinfuss:1990:UIS

[Lei90e] E. Leinfuss. USAA’s image of success (image processing). Datamation, 36(10):77–78, 80, 82, May 1990. CODEN DTM-NAT. ISSN 0011-6963.

Levine:1995:SFS

[Lev95] Ron Levine. Surprise! Fax Servers Smarten Up. Datamation, 41(9):63–64, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Libersat:1988:WSC

[Lib88] C. Libersat. Weaving seamless computing systems. Datamation, 34(17):59–66, September 1988. CODEN DTMNAT. ISSN 0011-6963.

Lichtenthal:1994:CBT

[Lic94] Sigfrido Lichtenthal. Connect business and technology. Datamation, 40(22):61–62, 64, November 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Lindsey:1991:CLS

[Lin91a] S. H. Lindsey. Color laptop saves power. Datamation, 37(25):43– 44, December 1991. CODEN DTMNAT. ISSN 0011-6963.

Lindsey:1991:DPR

[Lin91b] S. H. Lindsey. A database publishing recipe. Datamation, 37 (22):83–84, November 1991. CO-DEN DTMNAT. ISSN 0011-

Lindholm:1992:CD

[Lin92a] E. Lindholm. CAD on the desktop. Datamation, 38(1):68, 69, January 1992. CODEN DTM-NAT. ISSN 0011-6963.

Lindholm:1992:MOS

[Lin92b] E. Lindholm. Manufacturing open systems. Datamation, 38 (20):105–106, October 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Lindsey:1992:VGH

[Lin92c] S. H. Lindsey. VAX 6000 gets high marks. Datamation, 38(2): 55–56, January 1992. CODEN DTMNAT. ISSN 0011-6963.

Lindholm:1993:TCO

[Lin93] E. Lindholm. Two computers in one (DUET). Datamation, 39 (14):67–68, July 1993. CODEN DTMNAT. ISSN 0011-6963.

Lindholm:1994:BYL

[Lin94a] E. Lindholm. Button up your LAN security. Datamation, 40 (3):59–60, February 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Lindholm:1994:SMA

[Lin94b] E. Lindholm. System management across the enterprise. Datamation, 40(8):74–76, April 1994. CODEN DTMNAT. ISSN 0011-6963.

Lindholm:1994:DAC

[Lin94c] Elizabeth Lindholm. Data Access Without Chaos. Datamation, 40(15):59–??, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Lindholm:1995:PY

[Lin95a] Elizabeth Lindholm. Products of the Year. Datamation, 41(2):46– ??, February 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Linthicum:1995:BP

[Lin95b] David Linthicum. Backing into Portability. Datamation, 41(8): 54–??, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Linthicum:1995:RRT

[Lin95c] David S. Linthicum. The Rugged Road to Totally Transportable Apps. Datamation, 41(8):51–??, May 01, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Linthicum:1995:WUB

[Lin95d] David S. Linthicum. What UNIX Branding Means to You. Datamation, 41(13):53–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Lindholm:1996:FFF

[Lin96a] E. Lindholm. Framework — finding and fixing all the lurking date problems in your computer systems is a big, expensive, messy problem based on what appears to be a trivial technical glitch. Datamation, 42(1):7–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Lindholm:1996:TDP

[Lin96b] Elizabeth Lindholm. ThinkWrap — despite predictions of imminent demise, mainframes are proving to be very adaptive technology. Datamation, 42(8):102– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Linthicum:1996:ADT

[Lin96c] David S. Linthicum. App dev tackles the intranet — as client/ server embraces the intranet, tool vendors are recasting their products for the Web while developers have been busy learning new technologies, such as Java and ActiveX. Datamation, 42(15): 113–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Livingston:1992:OLB

[Liv92] D. Livingston. Outsourcing: look beyond the price tag. Datamation, 38(23):93, 96–97, November 1992. CODEN DTMNAT. ISSN 0011-6963.

Livingston:1995:TYH

[Liv95] Brian Livingston. 10 Things Your Help Desk Better Know. Datamation, 41(16):67–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Loh:1989:RCH

[LN89] M. Loh and R. R. Nelson. Reaping CASE harvests. Datamation, 35(13):31–32, 34, July 1989. CODEN DTMNAT. ISSN 0011-6963.

Lonsford:1990:USF

[Lon90] E. H. Lonsford. UNIX security: fact not fiction. Datamation, 36(4):44–46, 48, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Lemay:1996:YJS

[LP96] Laura Lemay and Charles Perkins. Yes, Java is secure. here’s why — secure transactions on the Internet? yup. here, today, now. just run Java-compiled bytecode on a Java Virtual Machine, and you can keep the cyberpunks off the Internet. Datamation, 42(5):47–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Lodahl:1989:ABT

[LR89] T. M. Lodahl and K. L. Redditt.

Aiming IS at business targets. Datamation, 35(4):93, 95, 97, 99– 100, February 1989. CODEN DTMNAT. ISSN 0011-6963.

Lederer:1989:PPS

[LS89] A. L. Lederer and V. Sethi. Pitfalls in planning (strategic information systems). Datamation, 35 (11):59–62, June 1989. CODEN DTMNAT. ISSN 0011-6963.

MacGregor:1995:HTR

[Mac95] Charlotte MacGregor. How Telecom Regulations Affect IS. Datamation, 41(18):78–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Madsen:1990:ANE

[Mad90] W. Madsen. African nations emphasizing security. Datamation, 36(9):104/2–8, May 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Mael:1995:WEB

[Mae95] Susan Mael. Want To Earn Big Money? Go West or Become C10. Datamation, 41(18):45–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Mael:1996:BDY

[Mae96] Susan Mael. Bigger by design — you’d think DATAMATION 100 companies just couldn’t get enough of each other last year — through acquisition, that is. mergers, solid. Datamation, 42 (12):32–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Majkiewicz:1989:TS

[Maj89] J. Majkiewicz. A test for SAA? Datamation, 35(20):37–38, 40, October 1989. CODEN DTM-NAT. ISSN 0011-6963.

Majkiewicz:1990:WDV

[Maj90] J. Majkiewicz. Will desktop video play in business? Datamation, 36(1):53–54, 56, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Mann:1991:IDM

[Man91] J. Mann. An image of document management. Datamation, 37 (23):81–82, November 1991. CO-DEN DTMNAT. ISSN 0011-6963.

Mantel:1992:AM

[Man92a] K. Mantel. An accounting masterpiece? Datamation, 38(21): 115–116, October 1992. CODEN DTMNAT. ISSN 0011-6963.

Mantel:1992:GFP

[Man92b] K. Mantel. Go figure with a pen (pen-based spreadsheet). Datamation, 38(10):91–92, May 1992.

CODEN DTMNAT. ISSN 0011-6963.

Martin:1988:CAS

[Mar88] J. Martin. Computing for art’s sake. Datamation, 34(24):86/9, 86/12–13, December 1988. CO-DEN DTMNAT. ISSN 0011-6963.

Martin:1989:ADI

[Mar89a] J. Martin. Adaptation drives IS integration (european single market). Datamation, 35(5):76/18, 20, March 1989. CODEN DTM-NAT. ISSN 0011-6963.

Martin:1989:LRD

[Mar89b] J. Martin. The long road to DB2 migration (relational databases). Datamation, 35(2):29, 32, January 1989. CODEN DTMNAT. ISSN 0011-6963.

Martin:1989:PPU

[Mar89c] J. Martin. Pursuing pirates (unauthorized software copying). Datamation, 35(15):41–42, August 1989. CODEN DTMNAT. ISSN 0011-6963.

Martin:1989:SF

[Mar89d] J. Martin. Synergy at fininvest. Datamation, 35(19):72/15– 16, October 1989. CODEN DTMNAT. ISSN 0011-6963.

Martin:1989:STP

[Mar89e] R. Martin. The standards test for portability (open systems). Datamation, 35(10):38–40, May 1989. CODEN DTMNAT. ISSN 0011-6963.

Martin:1990:SSG

[Mar90] J. Martin. Software that spans the globe. Datamation, 36 (21):96–95–96, 96–98, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Marion:1992:WIS

[Mar92] L. Marion. World IT sales stagger 4.1% to US$290 billion. Datamation, 38(13):12–17, 20, 22, 26– 27, 30–31, June 1992. CODEN DTMNAT. ISSN 0011-6963.

Marion:1995:WGT

[Mar95a] Larry Marion. Who’s Guarding the Till at the CyberMall? Datamation, 41(3):38–41, February 15, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Marsh:1995:TCSb

[Mar95b] V. Marsh. Test client/server apps before they bug you to death. Datamation, 41(14):32–35, August 1995. CODEN DTMNAT. ISSN 0011-6963.

Marsh:1995:TCSa

[Mar95c] Vivien Marsh. Test Client/ Server Apps Before They Bug You to Death. Datamation, 41 (14):32–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Marshall:1995:URM

[Mar95d] Ralph Marshall. USAir’s Ralph Marshall says you have to try different approaches to data warehousing before it flies. Datamation, 41(19):93–??, 1995. CO-DEN DTMNAT. ISSN 0011-

May:1990:JII

[May90] T. May. Justifying the image (image processing). Datamation, 36(8):82–84, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Maynard:1996:BMB

[May96] Jack Maynard. Break down manufacturing barriers — the face of manufacturing is changing rapidly, thanks to technologies that are allowing everyone in the process to have input. Datamation, 42(11):74–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Merlyn:1990:IOA

[MB90] V. Merlyn and G. Boone. The ins and outs of AD/cycle. Datamation, 36(5):59–64, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Merhoff:1994:OCM

[MB94] Eric Merhoff and Raj Bhela.

Overwhelmed by change? manage it! (client/server computing). Datamation, 40(18):49–50, 52, 54, September 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Mankin:1988:MTC

[MBGS88] D. Mankin, T. Bikson, B. Gutek, and C. Stasz. Managing technological change: the process is key. Datamation, 34(18):69–80, September 1988. CODEN DTM-NAT. ISSN 0011-6963.

McAfee:1989:VC

[McA89] J. McAfee. The virus cure. Datamation, 35(4):29–40, February 1989. CODEN DTMNAT. ISSN 0011-6963.

Miron:1988:MRC

[MCBH88] M. Miron, J. Cecil, K. Bradicich, and G. Hall. The myths and realities of competitive advantage (information technology). Datamation, 34(19):71–82, October 1988. CODEN DTMNAT. ISSN 0011-6963.

McCready:1988:TDF

[McC88] S. C. McCready. Top drug firms to digitize documentation. Datamation, 34(23):63–66, November 1988. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1989:PPM

[McC89a] T. McCusker. Project planning made easy. Datamation, 35(20): 49–50, October 1989. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1989:TWD

[McC89b] T. McCusker. Tracking the wild data base (DB2). Datamation, 35(13):3–8, July 1989. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1990:BVN

[McC90a] T. McCusker. A broader view of networking. Datamation, 36(7): 83–84, 86, April 1990. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1990:HDP

[McC90b] T. McCusker. How to decode 150 protocols. Datamation, 36(23): 47–48, December 1990. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1990:LMG

[McC90c] T. McCusker. LAN manager grows up. Datamation, 36(23): 53–54, December 1990. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1990:NPC

[McC90d] T. McCusker. Now playing: compact disks on LANs. Datamation, 36(20):55–56, October 1990. CO-DEN DTMNAT. ISSN 0011-6963.

McCusker:1990:UNH

[McC90e] T. McCusker. Users now hail expert systems upgrade. Datamation, 36(20):69–70, October 1990. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1990:WOS

[McC90f] T. McCusker. Where open systems are making headway. Datamation, 36(9):55–60, May 1990. CODEN DTMNAT. ISSN 0011-6963.

McClusker:1992:LCM

[McC92] T. McClusker. Low-cost messaging, the LAN fax way. Datamation, 38(1):46, 48, January 1992. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1993:CCM

[McC93a] T. McCusker. Credit card modems now do fax, too. Datamation, 39(16):65–67, August 1993. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1993:DWI

[McC93b] T. McCusker. The database way to integrate e-mail. Datamation, 39(2):61–62, 64, January 1993. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1993:ELF

[McC93c] T. McCusker. Easing the LAN fax burden. Datamation, 39(14): 73–74, 76, July 1993. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1993:HGE

[McC93d] T. McCusker. How to get from 80 to 100% in EDI. Datamation, 39(3):45–48, February 1993. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1993:LYL

[McC93e] T. McCusker. Link your LAN to your AS/400 (telephone company). Datamation, 39(19):85– 88, October 1993. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1993:NLP

[McC93f] T. McCusker. Network laser printers tackle high-end jobs. Datamation, 39(21):103–105, November 1993. CODEN DTM-NAT. ISSN 0011-6963.

McCusker:1993:WTE

[McC93g] T. McCusker. Workflow takes on the enterprise. Datamation, 39 (23):88, 90–92, December 1993. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1994:BLM

[McC94a] T. McCusker. Build links to multiple databases. Datamation, 40 (2):65–67, January 1994. CO-DEN DTMNAT. ISSN 0011-

McCusker:1994:HGM

[McC94b] T. McCusker. How to get more value from EDI. Datamation, 40 (9):56, 58, 60, May 1994. CO-DEN DTMNAT. ISSN 0011-6963.

McCusker:1994:MCB

[McC94c] T. McCusker. Maintenance contract bargains abound. Datamation, 40(12):64, June 1994. CO-DEN DTMNAT. ISSN 0011-6963.

McCusker:1994:MYL

[McC94d] Tom McCusker. Manage Your LAN Fax Software. Datamation, 40(14):70–73, 76–78, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1994:PCV

[McC94e] Tom McCusker. Pen computing: vertical reality. Datamation, 40(18):80, 82–85, September 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1994:NP

[McC94f] Tom McCusker. The Network Is the Printer. Datamation, 40(15): 55–??, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1995:VLS

[McC95a] Vance McCarthy. Are Virtual LANs Still Virtual Reality? Datamation, 41(12):45–46, 48– 49, July 1995. CODEN DTM-NAT. ISSN 0011-6963.

McCarthy:1995:AT

[McC95b] Vance McCarthy. ATM on Trial.

Datamation, 41(17):73–??, 1995.

CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1995:CEU

[McC95c] Vance McCarthy. Can Ethernet unclog your WAN? Datamation, 41(14):47, 49, 51, August 1995. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1995:DVS

[McC95d] Vance McCarthy. Desktop Videoconferencing Still a Rough Cut. Datamation, 41(9):51–52, 54–55, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1995:GNE

[McC95e] Vance McCarthy. Global Networking: Easier? Sure! Cheaper? Forget it! Datamation, 41(23):36–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

McCarthy:1995:GNL

[McC95f] Vance McCarthy. Going Network Live with Win95. Datamation, 41(16):63–65, September 1995. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1995:ILN

[McC95g] Vance McCarthy. IBM/Lotus: Notes rules. Datamation, 41(18): 75–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1995:LS

[McC95h] Vance McCarthy. Living with SNA. Datamation, 41(15):49–51, August 1995. CODEN DTM-NAT. ISSN 0011-6963.

McCarthy:1995:MIN

[McC95i] Vance McCarthy. Managing IP Nets: a Hitchhiker’s Guide. Datamation, 41(19):63–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1995:WPP

[McC95j] Vance McCarthy. WIN95: plan, plan, plan… then automate. Datamation, 41(20):45, 47–48, November 1995. CODEN DTM-NAT. ISSN 0011-6963.

McCusker:1995:USR

[McC95k] Tom McCusker. UNIX Systems Rule Now — but Make Way for SMP. Datamation, 41(10):57–??, June 01, 1995. CODEN DTM-NAT. ISSN 0011-6963.

McCarthy:1996:ABT

[McC96a] Vance McCarthy. ATM backbones tame Ethernet traffic — no longer just for IS sites with multimedia, ATM backbones can help you get that huge upsurge in Ethernet traffic under control. Datamation, 42(9):58–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1996:BFJ

[McC96b] Vance McCarthy. Building a firewall — just about everyone’s in a rush to get on the Internet. slow down! remember to look at network security products first before the world gets a chance to look at your secure data. Datamation, 42(10):74–??, ???? 1996. CODEN DTMNAT. ISSN 0011-

McCarthy:1996:CLY

[McC96c] Vance McCarthy. CTI lets you coddle customers at lower cost. Datamation, 42(18):46–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1996:CCP

[McC96d] Vance McCarthy. Cut costs! pipe voice and data together — if you’d like to cut line costs by adding voice to your data traffic, but you’re not ready for fullblown ATM, some new frame relay technologies may be the answer. Datamation, 42(2):47–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1996:EBB

[McC96e] Vance McCarthy. Exchange builds a backbone for the Enterprise. the much-delayed Microsoft Exchange can be a messaging backbone for your enterprise, and it can hook your users up to the Internet, too. Datamation, 42(6):64–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1996:GJD

[McC96f] Vance McCarthy. Gosling on Java — A DATAMATION exclusive interview with the creator of Java, Sun Microsystems’ James Gosling. Datamation, 42(5):30– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

McCarthy:1996:IAT

[McC96g] Vance McCarthy. The IP address time bomb — “the Internet’s backbone networks are clogged. the ISPs have a plan to open them up, but their solution may explode in your face”. Datamation, 42(8):54–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1996:JSY

[McC96h] Vance McCarthy. Jump start your I-Nets — hundreds of corporations are taking what they’ve learned from their Internet Web sites and are building intranets to make internal documents, files, and even important data more easily available to their own employees. here’s a simple guide to help you through the process, p us a look at a new Internet version of CA-Unicenter to help you manage it. Datamation, 42 (3):30–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1996:KMV

[McC96i] Vance McCarthy. Keep the Millennium virus off your net. Datamation, 42(1):55–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1996:NTA

[McC96j] Vance McCarthy. New tools authenticate remote users — do you really know who your remote users are? authentication tools tighten up remote security, but costs can be. Datamation, 42 (15):92–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1996:NNS

[McC96k] Vance McCarthy. Novell’s Net2000: Superglue for network apps? forget App-Ware, forget OpenDoc. Novell’s working on yet another new way to let you use your current object-oriented technologies to write networkready applications. Datamation, 42(4):38–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1996:PTK

[McC96l] Vance McCarthy. Performance tools kill the gremlins on your net — how’d you like to increase the performance and capacity of your network by at least 20%–without spending a dime on hardware? new network capacity-planning tools make it possible. Datamation, 42(15):88–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1996:SSR

[McC96m] Vance McCarthy. Seven steps to rolling out ISDN — ISDN is finally becoming a useful and affordable tool for IS managers who want to support users working from home or screaming to hook up to the internet. but it’s still not easy to install. Datamation, 42(11):86–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1996:TMP

[McC96n] Vance McCarthy. ThinkWrap — are Microsoft’s plans for the Internet real or not? Datamation, 42(3):68–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1996:TRI

[McC96o] Vance McCarthy. Tivoli to the rescue — IBM’s acquisition of Tivoli Systems could mark a watershed in the rush for more integrated, intelligent management systems. Datamation, 42(13):68– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

McCarthy:1996:UFL

[McC96p] Vance McCarthy. Use the fast lane on the Information Superhighway. Datamation, 42(18):88– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

McCool:1996:TNR

[McC96q] Rob McCool. ThinkWrap — Netscape’s Rob McCool: What’s wrong with HTTPD? Datamation, 42(7):94–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1997:BVL

[McC97a] Vance McCarthy. Bringing video on line: It’s more than smiling for the camera. Datamation, 43 (2):100–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1997:HSY

[McC97b] Vance McCarthy. How to switch on your intranet. Datamation, 43 (2):86–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1997:SIR

[McC97c] Vance McCarthy. Strike it rich!

Thanks to new datamining tools, companies are unearthing valuable information about their customers. The result? Closer customer relationships and a healthier bottom line. Datamation, 43 (2):44–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1997:UWM

[McC97d] Vance McCarthy. Use the Web to manage your network. Datamation, 43(1):128–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1997:WSH

[McC97e] Vance McCarthy. Web security: How much is enough? A group of Fortune 500 companies has developed a road map to help you find a balance between security and performance. Datamation, 43(1):112–??, ???? 1997. CO-DEN DTMNAT. ISSN 0011-6963.

McCartney:1997:SC

[McC97f] Laton McCartney. Safe computing. Datamation, 43(10):42–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

McMullen:1989:EUC

[McM89a] J. McMullen. End users create PC applications. Datamation, 35 (23):41–43, December 1989. CO-DEN DTMNAT. ISSN 0011-6963.

McMullen:1989:SC

[McM89b] J. McMullen. Spreadsheet consolidation. Datamation, 35(23): 35–36, December 1989. CODEN DTMNAT. ISSN 0011-6963.

McMullen:1990:ATG

[McM90a] J. McMullen. Ashton-Tate gets back to dBASEics. Datamation, 36(18):38–40, September 1990. CODEN DTMNAT. ISSN 0011-6963.

McMullen:1990:CVW

[McM90b] J. McMullen. A clear view for windows. Datamation, 36(2): 47–48, January 1990. CODEN DTMNAT. ISSN 0011-6963.

McMullen:1990:PBG

[McM90c] J. McMullen. Have PCs buried the graphics terminals market? Datamation, 36(11):69–70, 72, June 1990. CODEN DTMNAT. ISSN 0011-6963.

McMullen:1990:HGM

[McM90d] J. McMullen. How to get the most out of DOS. Datamation, 36(19):51–52, October 1990. CO-DEN DTMNAT. ISSN 0011-6963.

McMullen:1990:LDC

[McM90e] J. McMullen. The latest desktop contender (X-terminals). Datamation, 36(7):63–64, 66, April 1990. CODEN DTMNAT. ISSN 0011-6963.

McMullen:1990:MAN

[McM90f] J. McMullen. Microsoft in the age of networks. Datamation, 36 (9):36–39, May 1990. CODEN DTMNAT. ISSN 0011-6963.

McMullen:1990:NAS

[McM90g] J. McMullen. New allies: IS and service suppliers. Datamation, 36 (5):42–51, March 1990. CODEN DTMNAT. ISSN 0011-6963.

McMullen:1990:NBM

[McM90h] J. McMullen. The new breed of Mac databases. Datamation, 36 (23):75–76, December 1990. CO-DEN DTMNAT. ISSN 0011-6963.

McMullen:1990:TTU

[McM90i] J. McMullen. Tools of the trade: utility software comes of age. Datamation, 36(21):53–54, November 1990. CODEN DTM-NAT. ISSN 0011-6963.

McMullen:1990:WWN

[McM90j] J. McMullen. What’s wrong with network licensing. Datamation, 36(19):43–44, October 1990. CO-DEN DTMNAT. ISSN 0011-6963.

McMullen:1991:BSC

[McM91a] J. McMullen. Borland’s spreadsheet contender (quattro pro 2.0). Datamation, 36(24):34–36, 38, December 1991. CODEN DTMNAT. ISSN 0011-6963.

McMullen:1991:TEA

[McM91b] J. McMullen. ToolBook’s easy approach to Windows. Datamation, 36(24):47–48, December 1991. CODEN DTMNAT. ISSN 0011-6963.

McWilliams:1988:BNP

[McW88a] G. McWilliams. Banks, network providers eye EDI (electronic document interchange). Datamation, 34(23):77–78, November 1988. CODEN DTMNAT. ISSN 0011-6963.

McWilliams:1988:OOC

[McW88b] G. McWilliams. Oracle’s olympian challenge. Datamation, 34(23):31–45, November 1988. CODEN DTMNAT. ISSN 0011-6963.

McWilliams:1989:BDB

[McW89a] G. McWilliams. Back to the drawing board (digital and OLTP). Datamation, 35(13): 57, 59, 61, July 1989. CODEN DTMNAT. ISSN 0011-6963.

McWilliams:1989:ICE

[McW89b] G. McWilliams. Integrated computing environments. Datamation, 35(9):18–21, May 1989. CODEN DTMNAT. ISSN 0011-6963.

McWilliams:1989:SLR

[McW89c] G. McWilliams. Soft landings for RISC systems. Datamation, 35 (2):35–36, January 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Meese:1996:OHJ

[Mee96] Philip David Meese. The one hour Java applet — everything you need to know to build your first Java applet. using this tutorial, at the end of one or two hours you will have a working Java bar-chart applet and HTML page. Datamation, 42(5):51–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Mehler:1988:VSM

[Meh88] M. Mehler. VDTs soup up to meet the PC challenge. Datamation, 34(17):79–82, September 1988. CODEN DTMNAT. ISSN 0011-6963.

Mehta:1989:FGG

[Meh89] A. Mehta. The fifth generation goes soft. Datamation, 35(9): 80/5, 9, May 1989. CODEN DTMNAT. ISSN 0011-6963.

Melford:1995:SUE

[Mel95] Robert J. Melford. Secure UNIX for Enterprise Computing. Datamation, 41(4):55–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Meyer:1997:LL

[Mey97] N. Dean Meyer. Leadership on the line. Datamation, 43(5):119– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Mann:1990:LBM

[MF90] J. Mann and M. Freely. LANbased e-mail takes off. Datamation, 36(22):105–106, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Manville:1996:HYW

[MF96] Brook Manville and Nathaniel Foote. Harvest your workers’ knowledge — it’s a brave new world out there. call it postmodern reengineering, and to make your organization perform you’ll have to build systems that support knowledge — not data. Datamation, 42(13):78–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Miksanek:1997:EAS

[Mik97a] Chris Miksanek. Over the edge: Analog storage comes full circle.

Datamation, 43(2):136–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Miksanek:1997:ECP

[Mik97b] Chris Miksanek. Over the edge:

Calling the plays when you’re on the road. Datamation, 43(10): 128–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Miksanek:1997:EEE

[Mik97c] Chris Miksanek. Over the edge:

Extension 911: The effective help desk. Datamation, 43(11):122– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Miksanek:1997:EGG

[Mik97d] Chris Miksanek. Over the edge:

Gartner Group gambles on new research paradigm. Datamation, 43(4):128–??, ???? 1997. CO-DEN DTMNAT. ISSN 0011-6963.

Miksanek:1997:ESI

[Mik97e] Chris Miksanek. Over the edge:

In search of IT soft skills. Datamation, 43(7):118–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Miksanek:1997:EOD

[Mik97f] Chris Miksanek. Over the edge:

Outsourcing: Don’t get thrown out with the bathwater. Datamation, 43(8):82–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Miksanek:1997:ERR

[Mik97g] Chris Miksanek. Over the edge:

Reap the rewards of datamining.

Datamation, 43(5):126–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Miksanek:1997:EYW

[Mik97h] Chris Miksanek. Over the edge: So you wanna be a CIO. Datamation, 43(3):128–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Miksanek:1997:ETY

[Mik97i] Chris Miksanek. Over the edge: Tinkering with your IT staff. Datamation, 43(6):114–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Miksanek:1997:EWO

[Mik97j] Chris Miksanek. Over the edge: When the operation’s a success, but the patient dies. Datamation, 43(9):126–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Miksanek:1998:EEG

[Mik98a] Chris Miksanek. Over the edge: The executive guide to benchmarking. Datamation, 44(2): 100–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Miksanek:1998:EIS

[Mik98b] Chris Miksanek. Over the edge: The IT spin doctor. Datamation, 44(1):136–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Mitnick:1996:CEF

[Mit96] Kevin Mitnick. Cutting edge — former fugitive Kevin Mitnick confesses. Datamation, 42(2):6– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

McWilliams:1989:IRC

[MK89] G. McWilliams and S. Kerr. IBM redefines the competition. Datamation, 34(26):37–40, January 1989. CODEN DTMNAT. ISSN 0011-6963.

Moad:1990:HAF

[MK90a] J. Moad and S. Kerr. How allies fit with the new IBM. Datamation, 36(1):34–35, 38–39, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Moad:1990:HCH

[MK90b] J. Moad and S. Kerr. How customers help the new IBM. Datamation, 36(1):20–24, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Moad:1990:HLM

[MK90c] J. Moad and S. Kerr. How line managers run the new IBM. Datamation, 36(1):28–30, 32, January 1990. CODEN DTM-NAT. ISSN 0011-6963.

Moad:1988:CUJ

[Moa88a] J. Moad. Can unisys juggle open systems, too? Datamation, 34 (17):50–54, September 1988. CO-DEN DTMNAT. ISSN 0011-6963.

Moad:1988:CWL

[Moa88b] J. Moad. CSP: the weak link in SAA? (systems application architecture). Datamation, 34(23): 49–52, November 1988. CODEN DTMNAT. ISSN 0011-6963.

Moad:1988:PBU

[Moa88c] J. Moad. The plight of the Brazilian user (IS users). Datamation, 34(21):30–44, November 1988. CODEN DTMNAT. ISSN 0011-6963.

Moad:1988:PDC

[Moa88d] J. Moad. The push is on for data center automation. Datamation, 34(21):69–72, November 1988. CODEN DTMNAT. ISSN 0011-6963.

Moad:1989:AUJ

[Moa89a] J. Moad. Asking users to judge IS. Datamation, 35(21):93–94, 98, 100, November 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Moad:1989:CI

[Moa89b] J. Moad. Contracting with integrators. Datamation, 35(10):18– 22, May 1989. CODEN DTM-NAT. ISSN 0011-6963.

Moad:1989:CBS

[Moa89c] J. Moad. Cultural barriers slow reusability (programming). Datamation, 35(22):87–92, November 1989. CODEN DTM-NAT. ISSN 0011-6963.

Moad:1989:DDR

[Moa89d] J. Moad. The database dimension (RDA protocol). Datamation, 35(10):59–63, May 1989. CODEN DTMNAT. ISSN 0011-6963.

Moad:1989:DTA

[Moa89e] J. Moad. It’s DASD time again. Datamation, 35(8):50, 52, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Moad:1989:NCF

[Moa89f] J. Moad. Navigating crossfunctional IS waters (IS integration). Datamation, 35(5):73–75, March 1989. CODEN DTMNAT. ISSN 0011-6963.

Moad:1989:NSW

[Moa89g] J. Moad. Next stop, world markets (fujitsu). Datamation, 35 (15):28–32, August 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Moad:1989:GHB

[Moa89h] J. Moad. Out of the glass house (business training for IS professionals). Datamation, 35(15):71– 72, August 1989. CODEN DTM-NAT. ISSN 0011-6963.

Moad:1989:SWM

[Moa89i] J. Moad. The second wave (management of PC applications). Datamation, 35(3):14–20, February 1989. CODEN DTMNAT. ISSN 0011-6963.

Moad:1989:SSF

[Moa89j] J. Moad. Son of SAA (FAA for banks). Datamation, 35(11):39– 40, June 1989. CODEN DTM-NAT. ISSN 0011-6963.

Moad:1989:UGU

[Moa89k] J. Moad. Unix goes upscale.

Datamation, 35(4):61, 63, 65, 68, 70, February 1989. CODEN DTMNAT. ISSN 0011-6963.

Moad:1989:WWB

[Moa89l] J. Moad. Workstations win a big blue CASE. Datamation, 35(19): 37–38, 40, October 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Moad:1990:DPY

[Moa90a] J. Moad. Disaster-proof your data. Datamation, 36(21):87–88, 90, 93, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Moad:1990:LSH

[Moa90b] J. Moad. Large systems are hot! (buying trends). Datamation, 36 (10):24–29, 32, May 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Moad:1990:MCE

[Moa90c] J. Moad. Maintaining the competitive edge. Datamation, 36 (4):61–62, 64, 66, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Moad:1990:NS

[Moa90d] J. Moad. NCR’s 486 strategy. Datamation, 36(23):34–36, 38, December 1990. CODEN DTMNAT. ISSN 0011-6963.

Moad:1990:NAO

[Moa90e] J. Moad. The new agenda for open systems (standards). Datamation, 36(7):22–27, 30, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Moad:1990:SR

[Moa90f] J. Moad. The software revolution. Datamation, 36(4):22–25, 28–30, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Moad:1990:SPB

[Moa90g] J. Moad. The standards process breaks down. Datamation, 36(18):24–27, 30, 32, September 1990. CODEN DTMNAT. ISSN 0011-6963.

Moad:1990:WDD

[Moa90h] J. Moad. Which database for digital’s customers? Datamation, 36 (22):87–90, November 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Moad:1990:WCS

[Moa90i] J. Moad. Windows 3.0: the corporate standard. Datamation, 36(19):30–32, 36, October 1990. CODEN DTMNAT. ISSN 0011-6963.

Moad:1991:CIR

[Moa91a] J. Moad. Can IBM realize its S/390 vision? Datamation, 37 (19):30–34, October 1991. CO-DEN DTMNAT. ISSN 0011-6963.

Moad:1991:WHU

[Moa91b] J. Moad. Why are HP users angry? Datamation, 36(24):69–71, December 1991. CODEN DTM-NAT. ISSN 0011-6963.

Moad:1993:DRR

[Moa93a] J. Moad. Does reengineering really work? Datamation, 39(15): 22–24, 28, August 1993. CODEN DTMNAT. ISSN 0011-6963.

Moad:1993:HIQ

[Moa93b] J. Moad. How ISO 9000 quality programs affect IS. Datamation, 39(15):65–66, August 1993. CODEN DTMNAT. ISSN 0011-6963.

Moad:1993:HCT

[Moa93c] J. Moad. How to certify that training works. Datamation, 39 (17):67–68, September 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Moad:1993:IOD

[Moa93d] J. Moad. Inside an outsourcing deal. Datamation, 39(4):20–23, 26–27, February 1993. CODEN DTMNAT. ISSN 0011-6963.

Moad:1993:SC

[Moa93e] J. Moad. IS satisfies the customer. Datamation, 39(19):79, October 1993. CODEN DTM-NAT. ISSN 0011-6963.

Moad:1994:ART

[Moa94a] Jeff Moad. After Reengineering: Taking Care of Business. Datamation, 40(20):40–42, 44, October 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Moad:1994:LSS

[Moa94b] Jeff Moad. At last! standards that stick (standard IT architecture). Datamation, 40(19):60– 62, October 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Moad:1994:DYC

[Moa94c] Jeff Moad. Does Your CEO Get It? Datamation, 40(18):59–??, September 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Moad:1994:OTO

[Moa94d] Jeff Moad. One True Operating System Out, Diversity In. Datamation, 40(10):44–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Moad:1994:OHD

[Moa94e] Jeff Moad. Outsourcing the Help Desk: First Aid or Folly? Datamation, 40(24):44–46, December 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Moad:1994:PTM

[Moa94f] Jeff Moad. Psych Tests for MIS Staff: Is this Nuts? Datamation, 40(13):27–??, July 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Moad:1994:RRT

[Moa94g] Jeff Moad. Reengineering: Report From The Trenches. Datamation, 40(6):36–??, March 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Moad:1994:STC

[Moa94h] Jeff Moad. Shock Therapy Cures Fossilitis. Datamation, 40(11): 57–59, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Moad:1994:TCS

[Moa94i] Jeff Moad. The Training Crisis: School’s Out! Datamation, 40(15):46–48, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Moad:1994:TDS

[Moa94j] Jeff Moad. Toss Dispersed IS a Safety Net. Datamation, 40(14): 61–??, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Moad:1995:CRB

[Moa95a] J. Moad. Calculating the real benefit of training. Datamation, 41(7):45–47, April 1995. CODEN DTMNAT. ISSN 0011-6963.

Moad:1995:CHP

[Moa95b] Jeff Moad. Can High Performance Be Cloned? Should It Be? Datamation, 41(4):44–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Moad:1995:LCI

[Moa95c] Jeff Moad. Let Customers Have It Their Way. Datamation, 41(6): 34–??, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Moad:1995:OMT

[Moa95d] Jeff Moad. Oject Methods Tame Reengineering Madness. Datamation, 41(9):43–??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Moad:1995:OGL

[Moa95e] Jeff Moad. Outsourcing? Go Out on the Limb Together. Datamation, 41(2):58–59, 61, February 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Moad:1995:SSSa

[Moa95f] Jeff Moad. The Services Sector Soared in 1994. Datamation, 41 (10):65–??, June 01, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Moad:1995:SSSb

[Moa95g] Jeff Moad. The Services Sector Soared in 1994. Datamation, 41 (10):65–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Moad:1995:TFA

[Moa95h] Jeff Moad. Time for a fresh approach to ROI (IT infrastructure investment analysis). Datamation, 41(3):57–59, February 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Moad:1995:WCL

[Moa95i] Jeff Moad. When the Career Ladder Crumbles. Datamation, 41(1):44–??, January 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Moda:1990:WYS

[Mod90] J. Moda. Why you should be making IS allies. Datamation, 36 (9):26–34, May 1990. CODEN DTMNAT. ISSN 0011-6963.

McCusker:1994:DPD

[MS94] Tom McCusker and Paul Strauss.

Disaster Proofing? Don’t Forget the WAN. Datamation, 40(13): 48–??, July 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Mattison:1995:OLM

[MS95] Rob Mattison and Michael J.

Sipolt. An object lesson in management (object-oriented technology). Datamation, 41(12): 51, 53, 55, July 1995. CODEN DTMNAT. ISSN 0011-6963.

Munshi:1997:WSB

[Mun97] Renee Munshi. Web servers by the bundle. Datamation, 43(11): 90–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Musgrave:1988:UCI

[Mus88] B. Musgrave. Users colonize ISDN islands. Datamation, 34 (21):101–108, November 1988. CODEN DTMNAT. ISSN 0011-6963.

Neely:1995:WFC

[Nee95] Kris Neely. Where the AS/ 400 fits into client/server systems. Datamation, 41(9):57–59, May 15, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Newquist:1990:ERE

[New90] H. P. Newquist, III. Experts at retail (expert systems). Datamation, 36(7):53–54, 56, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Norton:1988:PNE

[Nor88] J. Norton. Private networks: the end of an era? Datamation, 34(21):111–117, November 1988. CODEN DTMNAT. ISSN 0011-6963.

OConnell:1988:MSD

[O’C88] D. R. O’Connell. A matter of semantics (DBMS). Datamation, 34(25):51–54, December 1988. CODEN DTMNAT. ISSN 0011-6963.

OConnor:1991:VSL

[O’C91] R. O’Connor. Vendors slash laptop prices. Datamation, 37(15):

64/4–5, August 1991. CODEN DTMNAT. ISSN 0011-6963.

Orfali:1995:OCS

[OH95] Robert Orfali and Dan Harkey.

Object Component Suites: The Whole Is Greater Than the Parts. Datamation, 41(3):44– ??, February 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Orfali:1994:CDC

[OHE94] Robert Orfali, Dan Harkey, and Jeri Edwards. Compound documents and component programming are powerful tools, but things could get messy if you try to snap Legos into Tinker Toys. For early experimentation, there’s not much choice: Microsoft’s OLE 2.0 has already come charging out to the gate with shrink-wrapped products. But next year, OpenDoc toolkits will become available, followed by shrink wrap the following year. It’s time to start mapping out your component strategy. Datamation, 40(22):38–??, November 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Ostroff:1990:SSP

[Ost90] L. Ostroff. Spinning systems in the Pacific (Rohm and Haas chemical manufacturing company). Datamation, 36(4):129– 131, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Ovans:1995:SYT

[Ova95] Andrea Ovans. Should You Take the Reengineering Risk?

Datamation, 41(17):38–39, 43– 44, September 1995. CODEN DTMNAT. ISSN 0011-6963.

Ovans:1996:IDR

[Ova96] Andrea Ovans. Improve ISdepartment relations! Everybody talks about aligning IS with business goals. Here’s how one pretty traditional IS organization actually did it. Datamation, 42 (16):66–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Owen:1989:WH

[Owe89] J. W. Owen. Working at home. Datamation, 35(13):63–65, July 1989. CODEN DTMNAT. ISSN 0011-6963.

Palmer:1992:MYC

[Pal92] S. D. Palmer. Mark your calendars with LAN software. Datamation, 38(23):77–78, 80, November 1992. CODEN DTM-NAT. ISSN 0011-6963.

Palmer:1993:PCC

[Pal93] S. D. Palmer. A plan that cured chaos. Datamation, 39(1):77–78, January 1993. CODEN DTM-NAT. ISSN 0011-6963.

Pantages:1990:NOJ

[Pan90a] A. Pantages. The new order at Johnson Wax. Datamation, 36 (6):103–108, March 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Pantages:1990:SRS

[Pan90b] A. Pantages. Salaries rise 5.6% in soft market. Datamation, 36(21): 42–44, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Pantages:1996:BIB

[Pan96a] Angie Pantages. Big iron is back — IS managers are choosing bigger systems to satisfy exponential increases in transaction volumes and in numbers of users on line. Datamation, 42(12):66–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Pantages:1996:ECD

[Pan96b] Angie Pantages. Enterprise computing drives IS spending — while the desktop market has cooled, IS managers are spending more on building a client/server enterprise infrastructure. Datamation, 42(9):66–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Pantages:1996:TSS

[Pan96c] Angie Pantages. THINKWRAP — a score of scorecards: Look back over a score of years and you’ll see — yep, a lot of things are different. and a lot of things are the same. Datamation, 42 (12):96–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Pantages:1997:NSS

[Pan97] Angie Pantages. I-nets spur IS spending. Datamation, 43(6): 107–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Parker:1992:CFS

[Par92] J. Parker. Channeling the flood of stored corporate data. Datamation, 38(19):93–94, 96, September 1992. CODEN DTM-NAT. ISSN 0011-6963.

Patch:1996:HCN

[Pat96] Kimberly Patch. Here come the net PCs — net PCs are no longer just a figment of Larry Ellison’s imagination. A handful of products are readying to ship this year. Datamation, 42(11):58–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Paul:1991:NMP

[Pau91] S. Paul. New 386 muscle for 286 PCs. Datamation, 37(20):83, October 1991. CODEN DTMNAT. ISSN 0011-6963.

Paul:1995:MSPa

[Pau95a] Steve Paul. More To Store and Plenty To Print. Datamation, 41 (10):66–??, June 01, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Paul:1995:MSPb

[Pau95b] Steve Paul. More to Store and Plenty To Print. Datamation, 41 (10):66–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Popolizio:1989:NSO

[PC89] J. J. Popolizio and W. S. Cappelli. New shells for old iron (expert systems). Datamation, 35 (8):41–43, 45, 47–48, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Pendse:1997:VOG

[Pen97] Nigel Pendse. Virtual OLAP: getting real. Datamation, 43(9): 115–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Peterson:1988:DCD

[Pet88] R. O. Peterson. Data center depopulation. Datamation, 34 (24):49–51, 53, 55, 57, December 1988. CODEN DTMNAT. ISSN 0011-6963.

Pickering:1994:CTM

[Pic94a] W. Pickering. Computer: take a memo. Datamation, 40(1):51–52, January 1994. CODEN DTM-NAT. ISSN 0011-6963.

Pickering:1994:OUR

[Pic94b] W. Pickering. Outfitting the ultimate road warrior. Datamation, 40(5):37–39, March 1994. CODEN DTMNAT. ISSN 0011-6963.

Pickering:1994:PWO

[Pic94c] W. Pickering. Project 4.0 works in the office. Datamation, 40(6): 64, March 1994. CODEN DTM-NAT. ISSN 0011-6963.

Pickering:1994:UGS

[Pic94d] W. Pickering. UPSs get smarter.

Datamation, 40(9):69–73, May 1994. CODEN DTMNAT. ISSN 0011-6963.

Pinella:1990:OCA

[Pin90] P. Pinella. Organizational computing arrives. Datamation, 36 (22):42–44, 46, 48, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Pinella:1991:CRT

[Pin91] P. Pinella. CD-ROM takes on dial-up data. Datamation, 36 (24):43–44, December 1991. CO-DEN DTMNAT. ISSN 0011-6963.

Pinella:1992:CIR

[Pin92] P. Pinella. Can IBM resurge in PCs? Datamation, 38(22):28–32, 34, 36, November 1992. CODEN DTMNAT. ISSN 0011-6963.

Purton:1989:EET

[Pur89] P. Purton. Europe’s electronic trading bloc. Datamation, 35(5): 76/13–14, March 1989. CODEN DTMNAT. ISSN 0011-6963.

Quinn:1995:TYC

[Qui95] Stephen R. Quinn. Test Your Client/Server Apps. Datamation, 41(23):60–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Radding:1992:QJH

[Rad92a] A. Radding. Quality is job hash 1. Datamation, 38(20):98–100, October 1992. CODEN DTM-NAT. ISSN 0011-6963.

Radding:1992:WNM

[Rad92b] A. Radding. When non-IS managers take control. Datamation, 38(14):55–56, 58, July 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Radding:1993:NYC

[Rad93] N. Radding. Now you can outsource remote LANs. Datamation, 39(14):36, 38, 42, July 1993. CODEN DTMNAT. ISSN 0011-6963.

Radding:1994:DCSb

[Rad94a] A. Radding. From dBASE to client/server: now you can get there. Datamation, 40(23):45–46, December 1994. CODEN DTM-NAT. ISSN 0011-6963.

Radding:1994:DCSa

[Rad94b] Alan Radding. From dBase to Client/Server: Now You Can Get There. Datamation, 40(23):45– ??, December 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Radding:1995:SDM

[Rad95] Alan Radding. Support Decision Makers with a Data Warehouse. Datamation, 41(5):53–56, March 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Radosevich:1996:CYM

[Rad96a] Lynda Radosevich. Can you measure Web ROI? well, sort of. it’s tough to measure what you can’t see, and Web projects are riddled with hidden costs. here’s some advice from IS managers that can help. Datamation, 42 (13):92–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Radosevich:1996:DTE

[Rad96b] Lynda Radosevich. Development tools for the Enterprise? Microsoft has great Windows app dev tools, but its Enterprise tools still have a long way to go. Datamation, 42(6):54–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Radosevich:1996:IPC

[Rad96c] Lynda Radosevich. Internet plumbing comes to groupware — confused as to whether an intranet or groupware package will enable your users to better collaborate? it’s no longer an either/or choice, thanks to nextgeneration groupware on tap. Datamation, 42(10):58–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Rashid:1989:COS

[Ras89] R. Rashid. A catalyst for open systems (mach kernel). Datamation, 35(10):32–33, May 1989. CODEN DTMNAT. ISSN 0011-6963.

Raynor:1995:GMT

[Ray95] D. A. Raynor. Get the most from training dollars (development team training). Datamation, 41(7):47–49, April 1995. CODEN DTMNAT. ISSN 0011-6963.

Realini:1995:TMC

[Rea95]  C. Realini.  ThinkWrap —

IS must change the way it approaches distributed system management problems. Datamation, 41(17):100–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Reck:1989:SGG

[Rec89] R. H. Reck. The shock of going global (management organization). Datamation, 35(15):67– 68, 70, August 1989. CODEN DTMNAT. ISSN 0011-6963.

Rhodewalt:1995:T

[Rho95a] B. Rhodewalt. ThinkWrap.

Datamation, 41(11):84–??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Rhodewalt:1995:TOT

[Rho95b] B. Rhodewalt. ThinkWrap — Original thought is a rare and valuable thing. Datamation, 41 (11):84–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Ricciuti:1990:APL

[Ric90a] M. Ricciuti. Apple plays up a low-cost Mac. Datamation, 36(22):101–102, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1990:TNS

[Ric90b] M. Ricciuti. AT and T’s network star (StarGROUP server). Datamation, 36(23):91–92, December 1990. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1990:IES

[Ric90c] M. Ricciuti. IBM entry spans midrange gap. Datamation, 36 (21):101–102, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1990:SOC

[Ric90d] M. Ricciuti. Sun’s office contender (SPARCstation SLC). Datamation, 36(19):111–112, October 1990. CODEN DTM-NAT. ISSN 0011-6963.

Ricciuti:1991:NTS

[Ric91] M. Ricciuti. NaviGate through system slowdowns. Datamation, 36(24):75–76, December 1991. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1992:CME

[Ric92a] M. Ricciuti. Connect manufacturing to the enterprise (software). Datamation, 38(2):42, 44, January 1992. CODEN DTM-NAT. ISSN 0011-6963.

Ricciuti:1992:EUS

[Ric92b] M. Ricciuti. Excel 4.0 ups spreadsheet ante. Datamation, 38(7):81–82, April 1992. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1992:HCH

[Ric92c] M. Ricciuti. Here come the HR client/server systems! Datamation, 38(14):37–38, 40, July 1992. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1992:PSG

[Ric92d] M. Ricciuti. Publishing software gets serious. Datamation, 38(24): 65–66, 68–69, December 1992. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1993:CTE

[Ric93a] M. Ricciuti. Compaq takes on the enterprise. Datamation, 39 (19):18–21, 24, October 1993. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1993:MMW

[Ric93b] M. Ricciuti. Making multimedia work. Datamation, 39(17):30–31, September 1993. CODEN DTM-NAT. ISSN 0011-6963.

Ricciuti:1993:P

[Ric93c] M. Ricciuti. The PC 50. Datamation, 39(23):36–38, December 1993. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1993:SPY

[Ric93d] M. Ricciuti. Spy proof your data! (LAN security). Datamation, 39(5):101–102, 104, 107, March 1993. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1994:BCS

[Ric94a] M. Ricciuti. The best in client/ server computing. Datamation, 40(5):26–30, 33–35, March 1994. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1994:MAW

[Ric94b] M. Ricciuti. Multidimensional analysis: winning the competitive game. Datamation, 40(4): 21–23, February 1994. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1994:OST

[Ric94c] M. Ricciuti. Outsourcing as a survival tactic (client/server). Datamation, 40(8):48, 50–52, April 1994. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1994:DYT

[Ric94d] Mike Ricciuti. DCE: Yesterday’s Technology for Tomorrow’s Apps? Datamation, 40(13):52– ??, July 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Ricciuti:1994:PMC

[Ric94e] Mike Ricciuti. Parallel Mainframe: Cheap Chips, Moderate Upgrade, Expensive Solution. Datamation, 40(11):70–??, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1994:MSI

[Ric94f] Mike Ricciuti. The Mainframe as Server: Is IBM Totally Bonkers — Or Brillant? Datamation, 40 (10):61–??, May 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Ricciuti:1994:USA

[Ric94g] Mike Ricciuti. UNIX Servers Approach PC Prices. Datamation, 40(14):65–??, July 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Riccuiti:1994:PMA

[Ric94h] M. Riccuiti. PC makers adapt to slim margins. Datamation, 40 (12):55, 58, June 1994. CODEN DTMNAT. ISSN 0011-6963.

Ridgway:1988:CSP

[Rid88] M. Ridgway. Curriculum shortfall (personnel). Datamation, 34 (25):77, December 1988. CO-DEN DTMNAT. ISSN 0011-6963.

Riley:1995:GLD

[Ril95a] W. D. Riley. Get the legal desktop (software licensing). Datamation, 41(8):56–57, 59, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Riley:1995:L

[Ril95b] W. D. Riley. LANScape. Datamation, 41(11):33–??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Riley:1995:LDY

[Ril95c] W. D. Riley. LANScape — Does your company own the files you put on the network? Maybe. Maybe not. Datamation, 41(16): 30–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Riley:1995:LNG

[Ril95d] W. D. Riley. LANScape — It’s no good having backup if it’s not in a safe place. Datamation, 41 (11):33–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Riley:1995:LNA

[Ril95e] W. D. Riley. LANScape — Network administrators can get away with a lot of things. Losing all the data on a file server ain’t one of’em. Datamation, 41(13): 28–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Riley:1995:LOW

[Ril95f] W. D. Riley. LANScape — Okay, so Win95 was a little late. Name one OS that hit the streets on schedule. Datamation, 41(23): 32–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Riley:1995:LSYa

[Ril95g] W. D. Riley. LANScape — Should you be wearing a suit? Datamation, 41(20):36–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Riley:1995:LSI

[Ril95h] W. D. Riley. LANScape — Sure the Internet is a great place to goof off But it’s also a great resource for the adults among us.

Datamation, 41(17):32–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Riley:1995:LSYb

[Ril95i] W. D. Riley. LANScape — Sure, your users want apps that are the equivalent of Ferraris; but sometimes, what’s best for everyone is a practical Chevy. Datamation, 41(21):36–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Riley:1995:LWH

[Ril95j] W. D. Riley. LANScape — Want to hear a really bad idea? Microsoft is eliminating concurrent licensing for part-time users. Datamation, 41(15):33–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Riley:1995:LWG

[Ril95k] W. D. Riley. LANScape — What goes around comes around, and none too soon, either. Datamation, 41(18):36–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Riley:1995:LWN

[Ril95l] W. D. Riley. LANScape — Windows NT may not be perfect, but it’s probably the pick of the litter at the moment. Datamation, 41 (19):40–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Riley:1996:LVP

[Ril96a] W. D. Riley. LANScape — a virus protection system make time to set up, but they’ll save you from wasting time later.

Datamation, 42(8):31–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:LBW

[Ril96b] W. D. Riley. LANSCAPE — behind the wall — before you put up your firewall, there are a few things you have to do. Datamation, 42(10):33–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:LDD

[Ril96c] W. D. Riley. LANSCAPE — dialing in — dialup is the ultimate frustration engine. Datamation, 42(11):25–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:LDY

[Ril96d] W. D. Riley. LANScape — do you know who’s fixing your computers? you had better! Datamation, 42(3):28–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:LET

[Ril96e] W. D. Riley. LANSCAPE — encrypt this!! — there are a thousand reasons to use encryption. unfortunately, the current state of technology does not lend itself to e-mail platforms. Datamation, 42(9):27–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:LNM

[Ril96f] W. D. Riley. LANSCAPE — no more cows: What we want from software vendors is a little bit of this and a little bit of that, not the whole cow. Datamation, 42 (12):31–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:LOC

[Ril96g] W. D. Riley. LANScape — offering choices is okay as long as the number of choices can be reduced, thereby creating a “standard”. Datamation, 42(5):27–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:LRC

[Ril96h] W. D. Riley. LANScape — remote control of PCs may be a network administrator’s dream, but it feels downright creepy to a lot of end users. Datamation, 42(2):22–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:LTM

[Ril96i] W. D. Riley. LANScape — too many networks means too many duplicate applications. take control. Datamation, 42(7):31–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:LUN

[Ril96j] W. D. Riley. LANScape — upgrades are a necessary part of our lives, and in every case, the same pitfalls await. Datamation, 42 (6):33–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:LWH

[Ril96k] W. D. Riley. LANScape — whatever happened to Apple Computer, anyway? Datamation, 42 (1):32–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:LIU

[Ril96l] W. D. Riley. LANScape it’s an inept upgrade policy that uses outdated software under the guise of saving money. Datamation, 42(4):27–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:NMA

[Ril96m] W. D. Riley. Network management: Anyone for rent-an-app? Datamation, 42(16):38–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:NMF

[Ril96n] W. D. Riley. Network management: Fire in the sky. Datamation, 42(15):43–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:NML

[Ril96o] W. D. Riley. Network management: Let’s chat. Datamation, 42 (18):36–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1996:NMB

[Ril96p] W. D. Riley. Network management: The best-laid plans. Datamation, 42(13):32–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Riley:1997:L

[Ril97a] W. D. Riley. LANScape. Datamation, 43(3):37–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Riley:1997:LAM

[Ril97b] W. D. Riley. LANscape: Audit more, suffer less. Datamation, 43 (2):33–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Riley:1997:LCH

[Ril97c] W. D. Riley. LANscape: Cheap help for pack rats. Datamation, 43(9):30–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Riley:1997:LGW

[Ril97d] W. D. Riley. LANscape: Get what you want. Datamation, 43 (5):45–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Riley:1997:LUR

[Ril97e] W. D. Riley. LANscape: I used to be really good at this. Datamation, 43(4):43–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Riley:1997:LWP

[Ril97f] W. D. Riley. LANscape: Want to play Internet cop? Datamation, 43(1):39–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Riley:1997:LWA

[Ril97g] W. D. Riley. LANscape: What about Bob? Datamation, 43 (7):31–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Rivard:1989:BQ

[RK89] E. Rivard and K. Kaiser. The benefit of quality IS. Datamation, 35(2):53–55, 58, January 1989. CODEN DTMNAT. ISSN 0011-6963.

Robinson:1996:TRS

[Rob96] R. Robinson. ThinkWrap RAD is not simply about replacing development tools. it’s about making fundamental changes in the system development process itself. Datamation, 42(4):80–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Roche:1997:HCW

[Roc97a] Eileen Roche. The hunt: Cast a wide net for talent. Datamation, 43(9):120–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Roche:1997:HFR

[Roc97b] Eileen Roche. The hunt: Find the right search firm. Datamation, 43(7):114–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Roche:1997:HHC

[Roc97c] Eileen Roche. The hunt: How to choose your successor. Datamation, 43(6):116–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Roche:1997:HMI

[Roc97d] Eileen Roche. The hunt: Money isn’t everything. Datamation, 43 (8):76–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Roche:1997:HSM

[Roc97e] Eileen Roche. The hunt: Staffing for the millennium. Datamation, 43(10):118–??, ???? 1997. CO-DEN DTMNAT. ISSN 0011-

Roche:1997:HHS

[Roc97f] Eileen Roche. The hunt: The hunt stops here: End personnel churn. Datamation, 43(11):112– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Roche:1998:HLE

[Roc98a] Eileen Roche. The hunt: Leading by example. Datamation, 44 (1):126–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Roche:1998:HSY

[Roc98b] Eileen Roche. The hunt: Send your staff home. Datamation, 44 (2):94–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Ross:1989:PCE

[Ros89] G. Ross. Planning for change (european single market). Datamation, 35(5):76/5–8, March 1989. CODEN DTMNAT. ISSN 0011-6963.

Rosenfeld:1996:HTD

[Ros96] Eric Rosenfeld. How to turn data into useful information. Datamation, 42(18):98–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Rothschild:1988:AIU

[Rot88a] K. Rothschild. The age of influence (user groups). Datamation, 34(25):18–21, 24, December 1988. CODEN DTMNAT. ISSN 0011-6963.

Rothstein:1988:RHE

[Rot88b] P. J. Rothstein. Up and running: how to ensure disaster recovery. Datamation, 34(20):86– 88, 92, 94, 96, October 1988. CODEN DTMNAT. ISSN 0011-6963.

Rothschild:1989:IUG

[Rot89] K. Rothschild. International user group directory. Datamation, 35(23):81–106, December 1989. CODEN DTMNAT. ISSN 0011-6963.

Rothschild:1991:TGD

[Rot91] K. Rothschild. Timeplex girds for data deluge. Datamation, 36 (24):57, December 1991. CO-DEN DTMNAT. ISSN 0011-6963.

Rothke:1997:WED

[Rot97] Ben Rothke. Want to encrypt data? Try Triple-DES. Datamation, 43(3):122–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Reck:1989:SCS

[RR89] R. H. Reck and V. P. Reck.

Steering IS committees straight. Datamation, 35(20):89–92, October 1989. CODEN DTMNAT. ISSN 0011-6963.

Ruber:1994:HRY

[Rub94] Peter Ruber. How Redundant Is Your RAID? Datamation, 40 (21):54–??, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Ruber:1995:CPW

[Rub95a] Peter Ruber. Color printing — who needs it? Datamation, 41 (20):73–74, November 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Ruber:1995:DHB

[Rub95b] Peter Ruber. Drive a hard bargain with network outsourcers. Datamation, 41(21):61, 63–64, November 1995. CODEN DTM-NAT. ISSN 0011-6963.

Ruber:1995:TRW

[Rub95c] Peter Ruber. Take the Risk Out of Wireless Communications. Datamation, 41(13):35–36, 39, July 1995. CODEN DTM-NAT. ISSN 0011-6963.

Ruhl:1988:HSP

[Ruh88]  J. Ruhl.  How to survive

post-implementation aftershock. Datamation, 34(24):77, 80, 82– 83, December 1988. CODEN DTMNAT. ISSN 0011-6963.

Runyan:1989:GS

[Run89a] L. Runyan. Global IS strategies. Datamation, 35(23):71–72, 78, December 1989. CODEN DTMNAT. ISSN 0011-6963.

Runyan:1989:HT

[Run89b] L. Runyan. Hot technologies for 1989. Datamation, 35(2):18– 22, 24, January 1989. CODEN DTMNAT. ISSN 0011-6963.

Runyan:1989:OOO

[Run89c] L. Runyan. The open opportunity (open systems). Datamation, 35(10):26–28, May 1989. CODEN DTMNAT. ISSN 0011-6963.

Ryan:1994:KHD

[Rya94] Alan Ryan. Kick the Hard Disk Habit: Add Hierarchical Storage. Datamation, 40(23):62–??, December 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Ryan:1995:AD

[Rya95] Alan J. Ryan. It’s all in the Disc.

Datamation, 41(3):61–??, February 15, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Rymer:1994:AAO

[Rym94] John Rymer. The AppWare Answer to Object Chaos. Datamation, 40(19):55–??, October 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Rypdal:1988:WAI

[Ryp88] J. Rypdal. What’s ahead for IBM’s on-line systems. Datamation, 34(23):69–70, 74–75, November 1988. CODEN DTM-NAT. ISSN 0011-6963.

Sarna:1995:TIM

[Sar95] D. Sarna. ThinkWrap — It may sell a lot of books to say that we’re all descending to the level of a computer, but it just ain’t so. Datamation, 41(15):80–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Schatz:1988:SLN

[Sch88a] W. Schatz. Suffolk law, new studies reinvigorate VDT debate (health aspects). Datamation, 34 (16):39–41, August 1988. CO-DEN DTMNAT. ISSN 0011-6963.

Schultz:1988:UAF

[Sch88b] B. Schultz. Users await fruitful DEC/ Apple yield. Datamation, 34(21):75–81, November 1988. CODEN DTMNAT. ISSN 0011-6963.

Schatz:1989:PPP

[Sch89a] W. Schatz. Programming is the problem parallel processing. Datamation, 35(9):57, 60– 61, May 1989. CODEN DTM-NAT. ISSN 0011-6963.

Schatz:1989:UCR

[Sch89b] W. Schatz. US computer research’s basic dilemma. Datamation, 35(23):44–47, December 1989. CODEN DTMNAT. ISSN 0011-6963.

Schatz:1989:WBS

[Sch89c] W. Schatz. Where’s the beef? (systems development at the department of agriculture). Datamation, 35(11):49–53, June 1989. CODEN DTMNAT. ISSN 0011-6963.

Schatz:1989:WWS

[Sch89d] W. Schatz. Who’s winning the supercomputer race? Datamation, 35(14):18–21, July 1989. CODEN DTMNAT. ISSN 0011-6963.

Schultz:1989:TLM

[Sch89e] B. Schultz. Take a letter, mr. postman (electronic mail). Datamation, 35(2):37, 40–42, January 1989. CODEN DTMNAT. ISSN 0011-6963.

Schatz:1990:WLG

[Sch90a] W. Schatz. Washington’s IS leadership gap. Datamation, 36(1): 81–83, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Schlack:1990:MGV

[Sch90b] M. Schlack. Macs get a VAX connection. Datamation, 36(22):69– 70, 72, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Schussel:1990:PRA

[Sch90c] G. Schussel. The promise and the reality of AD/cycle (CASE tool). Datamation, 36(18):69–70, 72, September 1990. CODEN DTMNAT. ISSN 0011-6963.

Schlack:1991:HKV

[Sch91] M. Schlack. How to keep viruses off your LAN. Datamation, 37 (20):87–88, 90, October 1991. CODEN DTMNAT. ISSN 0011-6963.

Schell:1992:HMM

[Sch92] E. H. Schell. How to make millions with database marketing. Datamation, 38(16):77–79, August 1992. CODEN DTMNAT. ISSN 0011-6963.

Schlack:1993:NII

[Sch93] M. Schlack. The new IT industry takes shape. Datamation, 39 (12):12–136, June 1993. CODEN DTMNAT. ISSN 0011-6963.

Schlack:1994:DYW

[Sch94] Mark Schlack. Do You Want The Phone Co. To Run Your LAN? Datamation, 40(24):38– ??, December 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Schlack:1995:FCC

[Sch95a] M. Schlack. Framework — Cut costs and extinguish that cigar at the same time. Datamation, 41 (18):6–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Schlack:1995:CSW

[Sch95b] Mark Schlack. Cover Story — Windows95 Migration Survival Guide. Datamation, 41(16):??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Schlack:1995:FFW

[Sch95c] Mark Schlack. Facing the Future with Windows95. Datamation, 41(16):43–44, September 1995. CODEN DTMNAT. ISSN 0011-6963.

Schlack:1995:HBNa

[Sch95d] Mark Schlack. Hello Bay Networks and Dial-Up LAN Servers. Datamation, 41(10):64–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Schlack:1995:HBNb

[Sch95e] Mark Schlack. Hello Bay Networks and Dial-Up LAN Servers. Datamation, 41(10):64–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Schlack:1995:KCS

[Sch95f] Mark Schlack. The Key to Client/Server OLTP. Datamation, 41(6):53–??, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Schreiber:1995:MD

[Sch95g] Rich Schreiber. Middleware Demystified. Datamation, 41(6): 41–??, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Schreiber:1995:CSD

[Sch95h] Richard Schreiber. Customer Service Demands New Architecture. Datamation, 41(23):48–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Schreiber:1995:GDA

[Sch95i] Richard Schreiber. Getting Distributed Apps To Communicate: Let’s Talk Middleware. Datamation, 41(12):57–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Schreiber:1995:GEA

[Sch95j] Richard Schreiber. Glue Enterprisewide Apps Together. Datamation, 41(15):41–43, August 1995. CODEN DTMNAT. ISSN 0011-6963.

Schreiber:1995:USS

[Sch95k] Richard Schreiber. Unify Security to Speed Development. Datamation, 41(21):78–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Schreiber:1995:WIO

[Sch95l] Richard Schreiber. Workflow imposes order on transaction processing. Datamation, 41(13): 57, 59–60, July 1995. CODEN DTMNAT. ISSN 0011-6963.

Schoenung:1997:ICN

[Sch97a] Michelle Schoenung. IT chokes on network constraints. Datamation, 43(7):71–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Schoenung:1997:SD

[Sch97b] Michelle Schoenung. Switching for dollars. Datamation, 43(5): 106–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Schwartz:1997:CS

[Sch97c] Karen D. Schwartz. Countdown to scalability. Datamation, 43 (9):48–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Schwartz:1997:MYB

[Sch97d] Karen D. Schwartz. Manage your bandwidth. Datamation, 43(10): 48–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Schwartz:1997:NST

[Sch97e] Karen D. Schwartz. Network speed takes off. Datamation, 43 (10):58–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Schwartz:1998:BD

[Sch98] Karen D. Schwartz. Benchmarking for dollars. Datamation, 44 (2):50–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Scotkin:1996:BUJ

[Sco96] Joel Scotkin. Business uses for Java — think client/ server computing and the Internet have changed your life? Java’s going to be twice as important-it will totally change the way you write applications and distribute them to clients. Datamation, 42(5):40– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Scrupski:1995:SV

[Scr95a]  S. Scrupski.  SerVices and

Virtues. Datamation, 41(11):29– ??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Scrupski:1995:SVB

[Scr95b] S. Scrupski. SerVices and Virtues — It’s becoming fashionable to outsource, so now IS can concentrate on doing it intelligently. Datamation, 41(11):29–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Sedacca:1990:LUA

[Sed90a] B. Sedacca. Lloyd’s ups the ante for claims processing. Datamation, 36(11):115–116, June 1990. CODEN DTMNAT. ISSN 0011-6963.

Sedacca:1990:SSH

[Sed90b] B. Sedacca. Systems in store for harrods. Datamation, 36(6):114– 116, March 1990. CODEN DTM-NAT. ISSN 0011-6963.

Semich:1990:CP

[Sem90] J. W. Semich. CASE power for OS/ 2. Datamation, 36(22):34– 37, 40, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Semich:1993:MTE

[Sem93] J. W. Semich. Multimedia tools are enterprise ready. Datamation, 39(20):92, 94, October 1993. CODEN DTMNAT. ISSN 0011-6963.

Semich:1994:HHQ

[Sem94a] J. W. Semich. Here’s how to quantify IT investment benefits. Datamation, 40(1):45–46, January 1994. CODEN DTMNAT. ISSN 0011-6963.

Semich:1994:BDJ

[Sem94b] J. William Semich. Big Development Jobs Need CASE. Datamation, 40(17):72–??, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Semich:1994:CYO

[Sem94c] J. William Semich. Can You Orchestrate Client/Server Computing? Datamation, 40(16):36–37, 40, 43, August 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Semich:1994:CSC

[Sem94d] J. William Semich. Client/Server Case: Oxymoron or Essential? Datamation, 40(17):67–??, September 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Semich:1994:HEM

[Sem94e] J. William Semich. Hermes: Enterprise Manager It Ain’t. Datamation, 40(13):31–??, July 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Semich:1994:IRI

[Sem94f] J. William Semich. Information Replaces Inventory at the Virtual Corp. Datamation, 40(14):37– 39, 42, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Semich:1994:LVM

[Sem94g] J. William Semich. The Long View from Microsoft: Component DBMSs. Datamation, 40 (15):40–??, August 01, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Semich:1994:YMN

[Sem94h] J. William Semich. You Mean NT Can’t Really Run Windows. Datamation, 40(10):67–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Semich:1995:BYJ

[Sem95a] J. William Semich. Bet Your Job on BackOffice? Datamation, 41 (20):38–42, November 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Semich:1995:CSUa

[Sem95b] J. William Semich. Client/Server Unchained: Finally, Hardware Independence. Datamation, 41 (11):40–??, June 15, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Semich:1995:CSUb

[Sem95c] J. William Semich. Client/Server Unchained: Finally, Hardware Independence. Datamation, 41 (11):40–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Semich:1995:CMB

[Sem95d] J. William Semich. C/S manufacturing: Build, buy, or reengineer? Datamation, 41(17):84–87, September 1995. CODEN DTM-NAT. ISSN 0011-6963.

Semich:1995:FSW

[Sem95e] J. William Semich. Four steps to working Web servers. Datamation, 41(4):42–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Semich:1995:ROSa

[Sem95f] J. William Semich. The Rush to Objects Speeds Up. Datamation, 41(10):47–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Semich:1995:ROSb

[Sem95g] J. William Semich. THE RUSH TO OBJECTS SPEEDS UP. Datamation, 41(10):47–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Semich:1995:WWW

[Sem95h] J. William Semich. The World Wide Web: Internet Boomtown? Datamation, 41(1):37–41, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Semich:1996:FLM

[Sem96a] J. W. Semich. Framework — look for more pages with in-depth analyses of technology choices for IS managers. Datamation, 42(4): 5–??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Semich:1996:FDM

[Sem96b] J. William Semich. FRAME-WORK — DATAMATION means business: Look to DATA-MATION to explain how new technologies can be used by your company to stay competitive.

Datamation, 42(12):5–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Semich:1996:FME

[Sem96c] J. William Semich. FRAME-WORK — Microsoft “embraces” Netscape — Netscape has a twoyear lead over Microsoft. does it matter? Datamation, 42(9):5– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Semich:1996:FSJ

[Sem96d] J. William Semich. Framework — Sun’s Java development language is a lot more than just another new toy for the Internet. Datamation, 42(5):5–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Semich:1996:FTW

[Sem96e] J. William Semich. FRAME-WORK — technology is the weapon — find a way to use data warehouse technologies to put your company a giant step ahead of the competition. Datamation, 42(10):5–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Semich:1996:FPU

[Sem96f] J. William Semich. FRAME-WORK — the promise of UNIX — who could possibly need 16 qbytes of memory? qbytes — you know, “quintillions of bytes.”. Datamation, 42(11):5–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Semich:1996:FWB

[Sem96g] J. William Semich. Framework — what’s business process management? Datamation, 42(8):6– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Semich:1996:SPD

[Sem96h] J. William Semich. Starting point — database technologies race to meet business demands. Datamation, 42(15):5–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Semich:1996:SPW

[Sem96i] J. William Semich. Starting point: What’s holding back ecommerce? Datamation, 42(16): 5–??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Serlin:1989:TEB

[Ser89] O. Serlin. Toward an equitable benchmark (DebitCredit). Datamation, 35(3):47, 50–51, 54, February 1989. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1995:PS

[SF95a] D. Sarna and G. Febish. Paradigm Shift. Datamation, 41 (11):24–??, June 15, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Sarna:1995:PSD

[SF95b] D. Sarna and G. Febish. Paradigm Shift — Does Windows do UNIX? That’s what the UNIX folks want to know. Datamation, 41(21):33–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Sarna:1995:PSGb

[SF95c] D. Sarna and G. Febish.

Paradigm Shift — George hosts DevDays and sees some more real-world, big-time Visual Basic app stuff. Datamation, 41(19): 37–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1995:PSGa

[SF95d] D. Sarna and G. Febish.

Paradigm Shift — George, who’s been out in the sun too long, begins to see a resemblance between OLE and Mickey Mouse. Datamation, 41(11):24–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1995:PSI

[SF95e] D. Sarna and G. Febish.

Paradigm Shift — If you think you don’t need T1 bandwidth to the desktop, you’re missing another looming paradigm shift. Datamation, 41(16):27–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1995:PSC

[SF95f] D. Sarna and G. Febish.

Paradigm Shift — Is the computing world moving toward fat servers and dumb terminals (er, clients) or the other way around? Only certain COMDEX visitors know for sure. Datamation, 41 (23):29–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Sarna:1995:PSO

[SF95g] D. Sarna and G. Febish.

Paradigm Shift — Our intrepid columnists tell you about their day job, during which they’re creating electronic City Halls for the Big Apple. Datamation, 41 (18):31–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Sarna:1995:PSS

[SF95h] D. Sarna and G. Febish. Paradigm Shift — Surfing the Net is one thing; developing for it is another. Here’s what you need to know to stay out of trouble. Datamation, 41(17):29–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1995:PSQ

[SF95i] D. Sarna and G. Febish. Paradigm Shift — The question is, Is chicken soup the cure for a lack of vision on IBM’s part? Datamation, 41(20):29–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1995:PSW

[SF95j] D. Sarna and G. Febish. Paradigm Shift — The world is awash in OLE objects. It’s a buyer’s market, but you should be wary. Datamation, 41(15): 29–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1995:VET

[SF95k] David E. Y. Sarna and George J. Febish. VB4: An Enterprise Tool Is Born. Datamation, 41(17):77– ??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Sharna:1995:PSC

[SF95l] D. Sharna and G. Febish. Paradigm Shift — Components should describe the business, not some part of a GUI. Datamation, 41(13):25–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSA

[SF96a] D. Sarna and G. Febish.

Paradigm shift — anyone can make predictions. we go back and see how ours came out last year before we do it again. Datamation, 42(1):29–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSE

[SF96b] D. Sarna and G. Febish.

Paradigm shift — evaluating Inet toolkits — if you’re taking the plunge into the Internet, here are some toolkit recommendations from our columnist duo. Datamation, 42(9):23–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSHa

[SF96c] D. Sarna and G. Febish.

Paradigm shift — how do programmers develop apps without code? magic. Datamation, 42 (7):28–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSIa

[SF96d] D. Sarna and G. Febish.

Paradigm shift — intranets on the runway — lately, fashion slaves have been mesmerized by the siren call of intranet development. Datamation, 42(11):21–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSIb

[SF96e] D. Sarna and G. Febish. Paradigm shift — invisible interfaces: About face is required reading for anyone building an interface for others to use. Datamation, 42(12):27–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSJ

[SF96f] D. Sarna and G. Febish. Paradigm shift — Java’s hot, but only part of the big picture. Datamation, 42(3):25–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSO

[SF96g] D. Sarna and G. Febish. Paradigm shift — our favorite pair of columnist’s pick their favorite object-oriented programming tools. Datamation, 42(4): 24–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSSa

[SF96h] D. Sarna and G. Febish. Paradigm shift — some say Microsoft ignored the Internet and is scrambling to catch up. well, you decide for yourself. Datamation, 42(5):24–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSB

[SF96i] D. Sarna and G. Febish. Paradigm shift — the business reality of VRML — VRML will have a profound impact on business. Datamation, 42(10):27–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSWa

[SF96j] D. Sarna and G. Febish.

Paradigm shift — what do you get when you combine sloppy object programmers with Internet connectivity? Datamation, 42 (2):17–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSWb

[SF96k] D. Sarna and Q. Febish.

Paradigm shift — while NT 4.0 will sport the Windows-95 GUI, as well as faster and networked OLE, Cairo it isn’t. Datamation, 42(6):27–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PST

[SF96l] David E. Y. Sarna and George J.

Febish. Paradigm shift — is there life after death? Datamation, 42 (13):27–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSIc

[SF96m] David E. Y. Sarna and George J.

Febish. Paradigm shift — the Internet ain’t no WAN, yet. Datamation, 42(15):36–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSF

[SF96n] David E. Y. Sarna and George J.

Febish. Paradigm shift: Feel the passion. Datamation, 42(16):32– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Sarna:1996:PSSb

[SF96o] David E. Y. Sarna and George J. Febish. Paradigm shift: Still out on a limb. Datamation, 42 (18):33–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1996:PSHb

[SF96p] I. D. Sarna and G. Febish. Paradigm shift — how Microsoft’s Internet strategy is different from all the Internet strategies. Datamation, 42(8): 25–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Semich:1996:JIT

[SF96q] J. William Semich and David Fisco. Java: Internet toy or enterprise tool — Java’s a whole lot more than just an animation tool for the Web. Sun’s renaming it the “Java Enterprise Platform”. better said, it’s the Java network-centric computing platform. Datamation, 42(5):28–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Semich:1996:RYB

[SF96r] J. William Semich and Mary Jo Foley. Run your business with Systems Management Tools! could you use a trouble-ticketing system to tell you when to replenish the inventory in your warehouse? A scheduling tool to monitor and refill bank ATMs? some systems management vendors are building the Business Process Management tools you’ll need to do that and more. Datamation, 42(8):32–??, ???? 1996.

CODEN DTMNAT. ISSN 0011-6963.

Sarna:1997:PSBa

[SF97a] David E. Y. Sarna and George J.

Febish. Paradigm shift: Back to our roots. Datamation, 43(3):25– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Sarna:1997:PSBb

[SF97b] David E. Y. Sarna and George J.

Febish. Paradigm shift: Bandwidth is (not yet) free. Datamation, 43(8):23–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1997:PSH

[SF97c] David E. Y. Sarna and George J.

Febish. Paradigm shift: Hanging on the wall. Datamation, 43 (1):33–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1997:PSI

[SF97d] David E. Y. Sarna and George J.

Febish. Paradigm shift: Integration now. Datamation, 43 (11):35–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1997:PSL

[SF97e] David E. Y. Sarna and George J.

Febish. Paradigm shift: Life without Microsoft. Datamation, 43(6):27–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1997:PSPb

[SF97f] David E. Y. Sarna and George J.

Febish. Paradigm shift: Past is Prolog? Datamation, 43(9):22– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Sarna:1997:PSR

[SF97g] David E. Y. Sarna and George J. Febish. Paradigm shift: Remote COM and the secret of life. Datamation, 43(10):25–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1997:PSPa

[SF97h] David E. Y. Sarna and George J. Febish. Paradigm shift: The OS pays off. Datamation, 43(4):27– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Sarna:1997:PSWa

[SF97i] David E. Y. Sarna and George J. Febish. Paradigm shift: The Web’s true power. Datamation, 43(2):35–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1997:PSS

[SF97j] David E. Y. Sarna and George J. Febish. Paradigm shift: To suite or not to suite. Datamation, 43 (5):29–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1997:PSWb

[SF97k] David E. Y. Sarna and George J. Febish. Paradigm shift: Whither Java. Datamation, 43(7):23–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Sarna:1998:PSA

[SF98] David E. Y. Sarna and George J. Febish. Paradigm shift: Again out on a limb. Datamation, 44 (1):31–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Sharp:1992:WFE

[Sha92] B. Sharp. Where to find expert help for downsizing. Datamation, 38(18):61–62, 64, September 1992. CODEN DTMNAT. ISSN 0011-6963.

Sharp:1993:BSC

[Sha93a] B. Sharp. Big screens come to color notebooks. Datamation, 39 (22):89–93, November 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Sharp:1993:ITI

[Sha93b] B. Sharp. Is it time to insource your financial apps? Datamation, 39(18):75–77, September 1993. CODEN DTMNAT. ISSN 0011-6963.

Sharp:1993:LEW

[Sha93c] B. Sharp. The low-end workstation race. Datamation, 39(2): 104–107, January 1993. CODEN DTMNAT. ISSN 0011-6963.

Sharp:1993:PWP

[Sha93d] B. Sharp. Preparing the way for Pentium (microprocessor). Datamation, 39(6):36–38, March 1993. CODEN DTMNAT. ISSN 0011-6963.

Sharp:1994:OUR

[Sha94a] Bill Sharp. It’s Official: UNIX Rules on the Server! Datamation, 40(22):50–??, November 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Sharp:1994:PNB

[Sha94b] Bill Sharp. Pick Notebooks for Business on the Road. Datamation, 40(16):66–71, August 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Sharp:1994:RNY

[Sha94c] Bill Sharp. Raid: Now You Can Take Your Pick. Datamation, 40 (22):81–??, November 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Sharp:1995:AAM

[Sha95] Bill Sharp. The Alpha Alternative to Mainframes. Datamation, 41(8):61–62, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Shamoon:1996:WBB

[Sha96] David N. Shamoon. Who’s the best at building Microsoft-based systems? of the 12,700 Microsoft Solution Provider firms, eight were recently recognized for their work in . . .. Datamation, 42 (15):139–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Sheldrick:1995:SBW

[She95] Dennis Sheldrick. Start becoming a Webmeister. Datamation, 41(4):40–??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1995:WNDb

[Sim95a] D. Simpson. Windows95 is no desktop management panacea. Datamation, 41(16):49–51, September 1995. CODEN DTM-NAT. ISSN 0011-6963.

Simpson:1995:DPP

[Sim95b] Dave Simpson. Downsizing: pull the plug slowly (mainframe to client/server migration). Datamation, 41(12):34–37, July 1995. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1995:USN

[Sim95c] David Simpson. A UNIX Server Is No Mainframe. Datamation, 41(23):40–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1995:CTS

[Sim95d] David Simpson. Can’t Tolerate Server Downtime? Cluster ’Em! Datamation, 41(15):45–47, August 1995. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1995:CCC

[Sim95e] David Simpson. Cut Costs with Client/Server Computing? Here’s How! Datamation, 41 (18):38–39, 41, October 1995. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1995:DCS

[Sim95f] David Simpson. Does CMOS Save Money? Datamation, 41 (20):69–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Simpson:1995:DHP

[Sim95g] David Simpson. Does HSM pay? be skeptical! Datamation, 41(14):53, 55–56, August 1995. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1995:BMM

[Sim95h] David Simpson. It’s a Buyers Market for Mainframe RAID. Datamation, 41(19):59–??, 1995. CODEN DTMNAT. ISSN 0011-

Simpson:1995:SNC

[Sim95i] David Simpson. Switch On the New Computer. Datamation, 41 (21):40–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Simpson:1995:WVN

[Sim95j] David Simpson. Win95 vs. NT Workstation. Datamation, 41 (21):69–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Simpson:1995:WNDa

[Sim95k] David Simpson. Windows95 Is No Desktop Management Panacea. Datamation, 41(16): 49–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:BYW

[Sim96a] David Simpson. Build your warehouse on MPP. Datamation, 42 (18):102–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:CNS

[Sim96b] David Simpson. Can NT scale? eager to deploy Windows NT in enterprise-level applications? there’s one problem: NT doesn’t always scale very well. Datamation, 42(6):78–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:CIY

[Sim96c] David Simpson. Cluster AS/ 400s? if you’re pushing the capacity of your AS/400s, or you need higher availability, you’re a prime candidate for clustering. Datamation, 42(4):47–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:CSB

[Sim96d] David Simpson. CMOS saves big bucks — next-generation CMOS mainframes are due by the end of the year. here’s how you can take advantage of them. Datamation, 42(13):44–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:CTP

[Sim96e] David Simpson. Compaq tries for the ’prise — Compaq is betting on a number of technologies to gain a foothold in the enterprise midrange server market. should you bet on them too? Datamation, 42(15):104–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:CSU

[Sim96f]  David Simpson.  COVER

STORY: UNIX. Datamation, 42 (11):??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:ERS

[Sim96g] David Simpson. The early returns are in: SNA Server 3.0 is a winner — Microsoft’s upcoming SNA Server 3.0 is garnering rave reviews from beta users. in fact, some sites are planning to switch from NetWare for SAA. Datamation, 42(15):126–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:GSP

[Sim96h] David Simpson. Go slow with the Pentium Pro — Intel’s Pentium Pro processor promises much faster speeds than its Pentium predecessor, but that doesn’t mean you should make the jump quickly. Datamation, 42(5):65– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Simpson:1996:HCN

[Sim96i] David Simpson. How to cross the NT, UNIX divide — more IS shops are finding themselves with a mix of UNIX and NT apps. here are some integration and migration options you should consider. Datamation, 42(10):68–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:ILW

[Sim96j] David Simpson. IBM launches OS/2 Warp Server 4 — IBM last month turned loose version 4 of OS/2 Warp server, with an impressive array of integrated system and network. Datamation, 42(7):55–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:TRY

[Sim96k] David Simpson. It’s time to reevaluate your tape options — tape prices are falling fast, so now’s a good time to take stock of your backup situation. tape arrays are coming on strong, and new giants in both the DLT and QIC arena will arrive by fall. Datamation, 42(8):58–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:LGB

[Sim96l] David Simpson. Let the games begin! more than 80 AS/400s are the work horses behind IBM’s Atlanta Olympics infrastructure, and they’re connected to all of IBM’s client/server platforms. Datamation, 42(13):117–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:MRI

[Sim96m] David Simpson. Midrange RAID-The innovators — thanks to evolving technology, RAID subsystems are providing enhanced management and faulttolerant features. DATAMA-TION picks the best of today’s RAID systems. Datamation, 42 (3):62–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:NWY

[Sim96n] David Simpson. NT workstation: Your next corporate desktop? Windows95 got the glory — and the hype — but it isn’t the corporate world’s only choice. here are some compelling reasons why NT workstation may be your next desktop standard. Datamation, 42(10):89–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:NBS

[Sim96o] David Simpson. NUMA busts the SMP bottleneck. Datamation, 42(16):82–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:PYW

[Sim96p] David Simpson. Put your AS/ 400 on the Web! OK. so it’s not the obvious choice. but the AS/ 400 may still have some advantages over other platforms as a Web server. more important, enabling software is finally becoming available. Datamation, 42 (12):91–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:RYD

[Sim96q] David Simpson. Replicate your AS/400 data! Emerging applications and IT requirements are driving a renewed surge toward AS/400-based data replication. Datamation, 42(16):137–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:SGM

[Sim96r] David Simpson. Say goodbye to MVS, say hello to OS/390 — OS/390 is bundled with more than 30 software packages and services. advantages for IS: reduced testing and lower licensing costs. Datamation, 42(9):44–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:TAS

[Sim96s] David Simpson. There are alternatives for AS/400 storage — when it comes to AS/400 storage, IBM rules the roost. but don’t ignore third parties if you need speed, lower costs or extra connectivity. Datamation, 42 (10):96–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:UNT

[Sim96t] David Simpson. UNIX? NT? time to decide! for enterpriselevel servers, solving the UNIX vs NT conundrum may be one of your biggest decisions over the next three to. Datamation, 42 (2):24–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:WNN

[Sim96u] David Simpson. Who needs a Network Computer? Network Computers are only blurry blips until they start shipping in volume. But a hybrid breed of machine is starting to show up on IS managers’ radar screens. Datamation, 42(16):96–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:WVN

[Sim96v] David Simpson. Win95 vs. NT workstation: You make the call. Datamation, 42(18):110–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:MCA

[Sim97a] David Simpson. Are mainframes cool again? Datamation, 43 (4):46–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:BIR

[Sim97b] David Simpson. Big Iron roars.

Datamation, 43(7):48–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:BIS

[Sim97c] David Simpson. Big iron shops flex the ’plex. Datamation, 43 (1):94–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:BB

[Sim97d] David Simpson. Bust the I/O bottleneck. Datamation, 43(3): 109–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:CYS

[Sim97e] David Simpson. Corral your storage management costs. Datamation, 43(4):88–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:CDM

[Sim97f] David Simpson. Cut desktop management costs! It’s tough to calculate desktop management costs, but one thing is clear: They’re steep. Here’s how some corporations are tackling the problem. Datamation, 43(1): 102–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:DSU

[Sim97g] David Simpson. Data sharing unites mainframes with open systems. Datamation, 43(2):113–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:DUC

[Sim97h] David Simpson. Do UNIX clusters pass muster? UNIX clusters are three steps ahead of NT clusters, but a few features are still missing. Datamation, 43(2):104– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Simpson:1997:NAF

[Sim97i] David Simpson. New applications fuel server sales. Datamation, 43(7):50–??, ???? 1997.

CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:NSD

[Sim97j] David Simpson. New storage for data downpour. Datamation, 43 (7):57–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:PMN

[Sim97k] David Simpson. PC management is no. 1 issue. Datamation, 43 (7):55–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:SS

[Sim97l] David Simpson. Shoot this server. Datamation, 43(6):86–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:UYT

[Sim97m] David Simpson. Untangle your tape storage costs. Datamation, 43(6):78–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:WTR

[Sim97n] David Simpson. Who’s taking to RISC and why? Mips and PowerPCs are no longer in the Windows NT arena. Yet some users continue to go for RISC architectures. Where is the payoff? Datamation, 43(5):70–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1997:WNS

[Sim97o] David Simpson. Will NCs save you a bundle? Network computers can indeed save big bucks, but not as much as vendors originally promised. Datamation, 43 (5):100–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Sivula:1989:CPC

[Siv89a] C. Sivula. Cooperative processing. 3. the client-server perspective. Datamation, 35(19):47–48, October 1989. CODEN DTM-NAT. ISSN 0011-6963.

Sivula:1989:GPM

[Siv89b] C. Sivula. Georgia-Pacific’s MRP II test. Datamation, 35(22):95– 101, November 1989. CODEN DTMNAT. ISSN 0011-6963.

Sivula:1989:NRE

[Siv89c] C. Sivula. A new rival to Ethernet and token ring? (ARCnet). Datamation, 35(23):53–58, December 1989. CODEN DTM-NAT. ISSN 0011-6963.

Sivula:1989:TBI

[Siv89d] C. Sivula. A test-bed for IBM’s strategy. Datamation, 35(20):67– 68, 70, October 1989. CODEN DTMNAT. ISSN 0011-6963.

Sivula:1990:CDC

[Siv90a] C. Sivula. A call for distributed computing (RPC). Datamation, 36(1):75–76, 78, 80, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Sivula:1990:DCD

[Siv90b] C. Sivula. Ducking the competition on the desktop (digital and the PC market). Datamation, 36 (5):36–39, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Sivula:1990:DIH

[Siv90c] C. Sivula. Duking it out on the high end (DEC’s new marketing strategies). Datamation, 36 (5):22–28, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Sivula:1990:GNL

[Siv90d] C. Sivula. GTE’s new line on customer service. Datamation, 36 (9):68–70, May 1990. CODEN DTMNAT. ISSN 0011-6963.

Sivula:1990:HOS

[Siv90e] C. Sivula. How open system suppliers differ (UNIX). Datamation, 36(10):57–58, 60, 62, May 1990. CODEN DTMNAT. ISSN 0011-6963.

Sivula:1990:EDI

[Siv90f] C. Sivula. The IS engine driving Infiniti. Datamation, 36(4):89– 90, 92, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Sivula:1990:PPF

[Siv90g] C. Sivula. The promise and problems of fractional T1. Datamation, 36(1):69–71, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Sivula:1990:WCC

[Siv90h] C. Sivula. What’s cookin’ in Campbell’s CIM lab. Datamation, 36(18):79–80, 82, September 1990. CODEN DTMNAT. ISSN 0011-6963.

Sivula:1990:WCP

[Siv90i] C. Sivula. The white-collar productivity push OA. Datamation, 36(2):52–54, 56, January 1990. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1994:DPD

[SM94] P. Strauss and T. McCusker. Disaster proofing? don’t forget the WAN. Datamation, 40(13):48– 50, July 1994. CODEN DTM-NAT. ISSN 0011-6963.

Snapp:1990:EAH

[Sna90a] C. D. Snapp. EDI aims high for global growth. Datamation, 36 (5):77–80, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Snapp:1990:SCN

[Sna90b] C. D. Snapp. Support for a complex network. Datamation, 36 (5):85–88, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Snell:1991:FNL

[Sne91a] N. Snell. Fast network lasers break new ground. Datamation, 37(19):79–80, October 1991. CO-DEN DTMNAT. ISSN 0011-6963.

Snell:1991:LS

[Sne91b] N. Snell. Leasing smart. Datamation, 37(23):73–74, November 1991. CODEN DTMNAT. ISSN 0011-6963.

Snell:1991:ZCP

[Sne91c] N. Snell. Zenith’s comeback portable. Datamation, 37(18): 28–30, September 1991. CODEN DTMNAT. ISSN 0011-6963.

Snell:1992:UWB

[Sne92a] B. Snell. Users want the bare fax. Datamation, 38(18):116– 119, September 1992. CODEN DTMNAT. ISSN 0011-6963.

Snell:1992:NN

[Sne92b] N. Snell. Are notebooks noteworthy? Datamation, 38(23):110– 112, November 1992. CODEN DTMNAT. ISSN 0011-6963.

Snell:1992:MBM

[Sne92c] N. Snell. Making the best mainframe deal. Datamation, 38(18): 73–74, 76, September 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Snell:1992:MGG

[Sne92d] N. Snell. Multimedia goes global.

Datamation, 38(10):88/8–9, May 1992. CODEN DTMNAT. ISSN 0011-6963.

Snell:1992:PPG

[Sne92e] N. Snell. PostScript printing gets a boost. Datamation, 38(19):89, September 1992. CODEN DTM-NAT. ISSN 0011-6963.

Snell:1992:PAA

[Sne92f] N. Snell. Print (almost) anything anywhere. Datamation, 38(19): 83–84, September 1992. CODEN DTMNAT. ISSN 0011-6963.

Snell:1992:TAR

[Sne92g] N. Snell. Top accountants rate accounting software. Datamation, 38(12):80–82, June 1992. CODEN DTMNAT. ISSN 0011-6963.

Snell:1993:FPF

[Sne93] N. Snell. Fast 486 PCs: features plus speed. Datamation, 39 (8):82–87, April 1993. CODEN DTMNAT. ISSN 0011-6963.

Snell:1994:CSCb

[Sne94a] N. Snell. Client/server chaos busters. Datamation, 40(20):45– 46, 50, October 1994. CODEN DTMNAT. ISSN 0011-6963.

Snell:1994:CSCa

[Sne94b] Ned Snell. Client/Server Chaos Busters. Datamation, 40(20):45– ??, October 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Snow:1998:BIP

[Sno98] C. Matthew Snow. Bridging intranet profit and value. Datamation, 44(1):20–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Spencer:1996:KYP

[Spe96] Will Spencer. Keep your password to yourself — here are some of the most useful software packages for UNIX security. Datamation, 42(9):76–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Stamps:1988:DCI

[Sta88] D. Stamps. Department computing. III. who’s in charge? Datamation, 34(19):85–92, October 1988. CODEN DTMNAT. ISSN 0011-6963.

Stamps:1989:SR

[Sta89a] D. Stamps. The AS/ 400’s surprise role. Datamation, 35(9): 49–50, 52, May 1989. CODEN DTMNAT. ISSN 0011-6963.

Stamps:1989:BTF

[Sta89b] D. Stamps. A bridge too far (networking). Datamation, 35 (21):83–84, 86, November 1989. CODEN DTMNAT. ISSN 0011-6963.

Stamps:1989:GI

[Sta89c] D. Stamps. The grooming of IS interns. Datamation, 35(19): 67–68, October 1989. CODEN DTMNAT. ISSN 0011-6963.

Stamps:1989:LUF

[Sta89d] D. Stamps. Loyalties under fire (control data corp. mainframe users). Datamation, 35(13):49– 50, July 1989. CODEN DTM-NAT. ISSN 0011-6963.

Stamps:1989:SDI

[Sta89e] D. Stamps. Spanning the DEC-IBM worlds. Datamation, 35(5): 45–48, March 1989. CODEN DTMNAT. ISSN 0011-6963.

Stamps:1989:TOL

[Sta89f] D. Stamps. Taking an objective look (object-oriented programming). Datamation, 35(10):45– 48, May 1989. CODEN DTM-NAT. ISSN 0011-6963.

Statland:1989:PPC

[Sta89g] N. Statland. Payoffs down the pike: a CASE study. Datamation, 35(7):32–33, 52, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Stamps:1990:FRF

[Sta90a] D. Stamps. Finding the right fixes for DB2. Datamation, 36 (19):89–90, 92, 96, October 1990. CODEN DTMNAT. ISSN 0011-6963.

Stamps:1990:HRS

[Sta90b] D. Stamps. Human resources: a strategic partner or IS burden? Datamation, 36(11):47–48, 50, June 1990. CODEN DTM-NAT. ISSN 0011-6963.

Stamps:1990:EIT

[Sta90c] D. Stamps. The IS eye on insider trading. Datamation, 36(8):35– 36, 38, 43, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Stamps:1990:SSM

[Sta90d] D. Stamps. Survival software for manufacturing. Datamation, 36 (23):81–82, December 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Staiti:1993:EJS

[Sta93] C. Staiti. Excel 5.0: not just a spreadsheet anymore. Datamation, 39(22):83, November 1993. CODEN DTMNAT. ISSN 0011-6963.

Staiti:1995:UGHb

[Sta95a] Chris Staiti. Users Give HP Top Rating. Datamation, 41(10):38– ??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Staiti:1995:UGHa

[Sta95b] Chris Staiti. Users Give HP Top Ratings. Datamation, 41(10): 38–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Stalters:1995:TIT

[Sta95c] R. Stalters. ThinkWrap — It takes teamwork to build a Competency Aligned Organization. Datamation, 41(20):94–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Stallman:1996:TAN

[Sta96] Richard Stallman. ThinkWrap — Americans now face the threat of two years in prison for indecent network postings; it would be helpful if they could access precise rules for avoiding imprisonment. Datamation, 42(5):98– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Stevens:1989:ODC

[Ste89] L. Stevens. The open data center (operations). Datamation, 35 (15):55–56, August 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Steyer:1990:WBO

[Ste90] H. C. Steyer, Jr. The world bank’s open network. Datamation, 36(4):109–110, 114, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Stevens:1992:MMM

[Ste92a] L. Stevens. Mastering E-mail management. Datamation, 38 (25):51–53, December 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Stevens:1992:NYP

[Ste92b] L. Stevens. Now your PC can ‘read’ your fax. Datamation, 38 (22):79–80, November 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Stevens:1994:YRD

[Ste94a] Larry Stevens. Are You Ready for the Desktop of the Future?? Datamation, 40(11):60–63, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Stevens:1994:CSG

[Ste94b] Larry Stevens. Clusters Stay Up and Grow Huge. Datamation, 40(17):39–40, 44, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Stevens:1995:CCb

[Ste95a] L. Stevens. To CD or not to CD? Datamation, 41(15):53–55, August 1995. CODEN DTM-NAT. ISSN 0011-6963.

Stevens:1995:SYS

[Ste95b] Larry Stevens. Safeguard Your Servers: Rack ’Em Up. Datamation, 41(2):63–??, February 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Stevens:1995:CCa

[Ste95c] Larry Stevens. To CD or Not To CD? Datamation, 41(15):53–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Stevens:1996:CTT

[Ste96a] Larry Stevens. Consider threetier client/server — just when client/server systems are needed to off-load some mainframe functions, they, too, are stretched to the limit. one solution is to move to a three-tiered architecture. Datamation, 42(4):61–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Stevens:1996:MCC

[Ste96b] Larry Stevens. MSDN cleans out code — MSDN 1 and 2 have been a boon to developers who want a repository of Microsoft products and documentation. Microsoft Developer. Datamation, 42(6):93–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1993:SMC

[Str93]  P. Strauss.  Secure E-mail

cheaply with software encryption. Datamation, 39(23):48, December 1993. CODEN DTM-NAT. ISSN 0011-6963.

Strassmann:1994:CSG

[Str94a] Paul Strassmann. CIOs Should Get Back to Basics. Datamation, 40(18):70–??, September 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1994:CCSb

[Str94b] P. Strauss. Cheaper client/ server networking for mainframes. Datamation, 40(21):69– 70, November 1994. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1994:ISG

[Str94c] Paul Strauss. An ISDN Survival Guide. Datamation, 40(20):59– ??, October 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1994:ABC

[Str94d] Paul Strauss. ATM Backbones: Complex? And How! Datamation, 40(14):44–??, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1994:BTH

[Str94e] Paul Strauss. Beyond Talking Heads: Videoconferencing Makes Money. Datamation, 40(19):38– 41, October 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1994:CCSa

[Str94f] Paul Strauss. Cheaper Client/ Server Networking for Mainframes. Datamation, 40(21):69– ??, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1994:GNS

[Str94g] Paul Strauss. Get NetWare Storage Under Control. Datamation, 40(13):36–??, July 01, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Strauss:1994:JWY

[Str94h] Paul Strauss. Just When You Thought IP Was Safe. Datamation, 40(21):71–??, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1994:MMM

[Str94i] Paul Strauss. Manage the Mess in Messaging. Datamation, 40 (24):65–??, December 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1994:NHA

[Str94j] Paul Strauss. NetWare: Higher Availability At Lower Cost. Datamation, 40(22):57–??, November 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Strauss:1994:PYN

[Str94k] Paul Strauss. PLAN Your 2001 Network Now. Datamation, 40 (15):32–35, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1994:SCY

[Str94l] Paul Strauss. Seize Control of Your 2001 Network. Datamation, 40(15):35–37, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1994:WCS

[Str94m] Paul Strauss. Welcome to Client-Server PBX Computing. Datamation, 40(11):49–50, 52, June 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1994:WYO

[Str94n] Paul Strauss. Write Your Own Wireless Apps. Datamation, 40 (18):76–??, September 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Strehlo:1994:SGM

[Str94o] Kevin Strehlo. Smartsourcing:

Getting the Most from Service Vendors. Datamation, 40(23):38– ??, December 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1995:LNM

[Str95a] Paul Strauss. At Last: Net Managers that Distribute the Load. Datamation, 41(3):50–51, 54, 56, February 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Strauss:1995:GTG

[Str95b] Paul Strauss. The Get Tough Guide to Telecom Contracts. Datamation, 41(5):38–39, 42–44, March 15, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Strauss:1995:TTY

[Str95c] Paul Strauss. Time To Terminate Your Trusted Token Ring? Datamation, 41(2):55–??, February 01, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Strehlo:1995:F

[Str95d] K. Strehlo. Framework. Datamation, 41(11):7–??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Strehlo:1995:FAY

[Str95e] K. Strehlo. Framework — After 13 years, Windows95 emerges from a cloud of barbecue potato chip powder. Datamation, 41 (16):7–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Strehlo:1995:FBP

[Str95f] K. Strehlo. Framework — Back-Office is probably no threat to existing OLTP setups. It’s something a lot more important. Datamation, 41(20):9–??, 1995. CODEN DTMNAT. ISSN 0011-

Strehlo:1995:FBS

[Str95g] K. Strehlo. Framework — Bestselling management books about IT never mention the IT wizardry that makes the magic happen. Datamation, 41(17):7–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Strehlo:1995:FNM

[Str95h] K. Strehlo. Framework — No matter where you go, IS faces the same challenges. Datamation, 41 (15):8–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Strehlo:1995:FST

[Str95i] K. Strehlo. Framework — Sometimes the truth hurts. But face it: Even the best OLTP data can create disaster in a data warehouse if it’s not cleaned up first. Datamation, 41(19):8–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Strehlo:1995:FSB

[Str95j] K. Strehlo. Framework — The switch has been thrown, and soon we’ll be seeing a new generation of computer architectures. Datamation, 41(21):9–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Strehlo:1995:FWS

[Str95k] K. Strehlo. Framework — We’re stepping up our coverage of international connectivity issues. Datamation, 41(23):9–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Strehlo:1995:FAD

[Str95l] K. Strehlo. Framework-Another day in the life of a reader. Datamation, 41(13):6–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Strehlo:1995:BBD

[Str95m] Kevin Strehlo. Business Benefits from Distributed Objects. Datamation, 41(18):71–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Strehlo:1995:WDW

[Str95n] Kevin Strehlo. Why Develop for Win95? OLE! Datamation, 41 (16):53–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

Strehlo:1996:FTY

[Str96a] K. Strehlo. Framework — two years ago, we promised a new and better magazine. we delivered. where do we go from here? Datamation, 42(3):7–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Strehlo:1996:QAO

[Str96b]  Kevin Strehlo.  1995 Quest

Awards — our second annual Quest Awards highlight IS shops that’ve made the most effective use of IT services. Datamation, 42(2):41–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Strehlo:1996:DWA

[Str96c] Kevin Strehlo. Data warehousing: Avoid planned obsolescence — you’ve got to do more than just make and maintain the data warehouse. you have to sustain it, nurture it, make it grow. here’s how. Datamation, 42(2):32–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Strassmann:1997:WBS

[Str97a]  Paul Strassmann.  Will big

spending on computers guarantee profitability? Maybe but maybe not. In Paul Strassmann’s new book The Squandered Computer, he offers evidence that the relationship between spending and profitability is no simple formula. Datamation, 43(2):75– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Strassmann:1997:FYI

[Str97b] Paul A. Strassmann. Forty years of IT history. Datamation, 43 (10):80–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Strom:1997:RME

[Str97c] David Strom. Road map to the extranet. Datamation, 43(11): 60–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Strom:1997:WWW

[Str97d] David Strom. Web wars: Which are the most valuable weapons? Datamation, 43(4):56–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Sudama:1995:GRD

[Sud95] Ram Sudama. Get Ready for Distributed Objects. Datamation, 41(18):67–??, 1995. CO-DEN DTMNAT. ISSN 0011-

Taber:1992:COT

[Tab92a] M. Taber. Can OfficeVision take off with Notes? Datamation, 38 (4):65–68, February 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Taber:1992:TDC

[Tab92b] M. Taber. Tracking the data center’s weak link. Datamation, 38 (14):45–46, 48, July 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Tate:1989:EEP

[Tat89a] P. Tate. Europe’s IS execs predict the future. Datamation, 35 (5):76/1–2, March 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Tate:1989:EVK

[Tat89b] P. Tate. Europe’s VM knight woos the blue dragon (IBM user group). Datamation, 35(11):64a– b, June 1989. CODEN DTM-NAT. ISSN 0011-6963.

Tate:1989:EMU

[Tat89c] P. Tate. With Europe’s market unity comes IS urgency. Datamation, 35(5):28–32, March 1989. CODEN DTMNAT. ISSN 0011-6963.

Taylor:1990:TMM

[Tay90] S. Taylor. T1 muxes meet the networking challenge. Datamation, 36(21):67–68, 70, November 1990. CODEN DTMNAT. ISSN 0011-6963.

Taylor:1996:PYU

[Tay96] Martin Taylor. Port your UNIX apps to OS/400 — think it’s a strange idea to port your UNIX apps to the AS/400? considering the cost is a lot lower, maybe you should be asking yourself why you’re not. Datamation, 42 (8):77–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Tibbetts:1990:WSW

[TB90] J. Tibbetts and B. Bernstein.

Why SAA will play on local area networks. Datamation, 36 (11):37–42, June 1990. CODEN DTMNAT. ISSN 0011-6963.

The:1991:MBM

[The91] L. The. MIS buys into mail order. Datamation, 37(25):27–29, December 1991. CODEN DTM-NAT. ISSN 0011-6963.

The:1992:AFW

[The92a] L. The. AST fields a Windows machine. Datamation, 38 (12):22–26, June 1992. CODEN DTMNAT. ISSN 0011-6963.

The:1992:PTH

[The92b] L. The. PCs tool up for hypertext. Datamation, 38(3):35–36, February 1992. CODEN DTM-NAT. ISSN 0011-6963.

The:1993:NLO

[The93a] L. The. New life for old general ledgers. Datamation, 39(17): 55–56, September 1993. CODEN DTMNAT. ISSN 0011-6963.

The:1993:NPT

[The93b] L. The. The new palmtops: tools or toys? Datamation, 39(13):35– 37, July 1993. CODEN DTM-NAT. ISSN 0011-6963.

The:1993:PFP

[The93c] L. The. Paper-free publishing gets real (electronic publishing). Datamation, 39(17):39–41, September 1993. CODEN DTM-NAT. ISSN 0011-6963.

The:1993:PHG

[The93d] L. The. Pen hardware: get practical. Datamation, 39(24):30–31, 34–35, December 1993. CODEN DTMNAT. ISSN 0011-6963.

The:1993:SME

[The93e] L. The. Scaling e-mail for the enterprise. Datamation, 39(8):20– 24, April 1993. CODEN DTM-NAT. ISSN 0011-6963.

The:1993:TYM

[The93f] L. The. Turn your e-mail into groupware. Datamation, 39(8): 28–30, 32, April 1993. CODEN DTMNAT. ISSN 0011-6963.

The:1993:PMR

[The93g] Lee The. Project management for the rest of us. Datamation, 39(1):41–42, 44, January 1993. CODEN DTMNAT. ISSN 0011-6963.

The:1994:APL

[The94a] Lee The. All Products Lead to the Help Desk. Datamation, 40 (21):73–??, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

The:1994:BIA

[The94b] Lee The. Big isn’t always better (databases). Datamation, 40 (13):66–71, July 1994. CODEN DTMNAT. ISSN 0011-6963.

The:1994:CFK

[The94c] Lee The. Cooking with forms (kentucky fried chicken: forms automation). Datamation, 40 (23):69–73, December 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

The:1994:FGS

[The94d] Lee The. E-Forms Get Smarter.

Datamation, 40(23):67–69, December 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

The:1994:GW

[The94e] Lee The. Getting into the Workflow. Datamation, 40(19):65– ??, October 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

The:1994:LMT

[The94f] Lee The. Look for More than Just a Little Support. Datamation, 40(21):75–??, November 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

The:1994:PTY

[The94g] Lee The. Pick Tools with — not for — Your Users. Datamation, 40(15):61–??, August 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

The:1994:SCO

[The94h] Lee The. Service Cinches Onsite Shootout. Datamation, 40(10):

77–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

The:1994:WMG

[The94i] Lee The. Workflow Makes the Grade. Datamation, 40(19):68– ??, October 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

The:1995:RHRc

[The95a] L. The. Retool human resources. Datamation, 41(11):66, 68–69, June 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:BGM

[The95b] Lee The. Bagging good manufacturing software. Datamation, 41 (17):87, 89–93, September 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:BWG

[The95c] Lee The. Beta Not Wait for Groupware. Datamation, 41 (13):69–74, July 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:BSA

[The95d] Lee The. Building suiteware apps (evaluation). Datamation, 41(6): 59–61, April 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:CFI

[The95e] Lee The. Can Frameworks Integrate System Management Apps? Datamation, 41(21):86– ??, 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:CUH

[The95f] Lee The. Check Under the Hood for OS Muscle. Datamation, 41 (23):66–??, 1995. CODEN DTM-NAT. ISSN 0011-6963.

The:1995:DWC

[The95g] Lee The. Draino for Workstation Clogs. Datamation, 41(15): 67–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:FSD

[The95h] Lee The. Financial Software Does More Than Just Crunch Numbers. Datamation, 41(2):67– 70, February 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:HTR

[The95i] Lee The. Hands-On Testing Reveals the Truth. Datamation, 41(2):69–??, February 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:HAMa

[The95j] Lee The. HR App Meets Critical Needs. Datamation, 41(11): 69–??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:HAMb

[The95k] Lee The. HR App Meets Critical Needs. Datamation, 41(11): 69–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:KYE

[The95l] Lee The. Keep Your Evaluation Open. Datamation, 41(8):67–??, May 01, 1995. CODEN DTM-NAT. ISSN 0011-6963.

The:1995:NGT

[The95m] Lee The. Need Groupware?

Think Functions, Not Products.

Datamation, 41(13):67–69, July 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:NYC

[The95n] Lee The. Now You Can Automate BPR. Datamation, 41(4): 61–63, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:OAT

[The95o] Lee The. OLAP Answers Tough Business Questions. Datamation, 41(8):65–??, May 01, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

The:1995:PDB

[The95p] Lee The. Patrolling database backups (associated grocers). Datamation, 41(21):89–93, November 1995. CODEN DTM-NAT. ISSN 0011-6963.

The:1995:PPS

[The95q] Lee The. Plug and play servers for DSS apps (USAir). Datamation, 41(19):93–98, October 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:RHRa

[The95r] Lee The. Retool Human Resources. Datamation, 41(11):66– ??, June 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:RHRb

[The95s] Lee The. Retool Human Resources. Datamation, 41(11):66– ??, 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:SS

[The95t] Lee The. Service Sells an OS.

Datamation, 41(23):68–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

The:1995:SSD

[The95u] Lee The. SMP Servers for DSS:

Go for the Package Buy. Datamation, 41(19):91–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

The:1995:SPS

[The95v] Lee The. Suite pricing seals the deal (Novell PerfectOffice review). Datamation, 41(6):62–65, April 01, 1995. CODEN DTM-NAT. ISSN 0011-6963.

The:1995:WTP

[The95w] Lee The. Workflow Tackles the Productivity Paradox. Datamation, 41(15):65–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

The:1996:AAH

[The96a] Lee The. Al automates a help desk — Konica’s help desk already was an award winner, but it needed to move up to the next level so it could take on products that were nearly impossible to support. The answer? A help desk with smarts of its own. Datamation, 42(2):54–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

The:1996:HHC

[The96b] Lee The. How to hire a consultant — IT research and consulting firms sell information sprinkled liberally with expertise, experience, and contacts. they (usually) know what technologies, vendors, and products are hot and which are not but how do you hire one? it doesn’t have to be a hit-or-miss proposition. Datamation, 42(4):66–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

The:1996:FPM

[The96c] Lee The. IS-Friendly project management — you can’t manage IS projects with generalpurpose project management software. here’s what you need to look for and why. Datamation, 42(7):79–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

The:1996:MYH

[The96d] Lee The. Morph your help desk into customer support — customer support and help desks do just about the same thing. now they’re coming together. the result? one package that can handle both needs, cutting your overall costs and improving your product. now you can make both your customers and your users happy. Datamation, 42(2):52–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

The:1996:TRB

[The96e] Lee The. Turn research into business decisions — it’s easy to get good IT information from research service. it’s not so easy to turn research into recommendations. here’s how one company does it. Datamation, 42(4):68–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Tillmann:1989:PRR

[Til89] G. Tillmann. Prototyping for the right results. Datamation, 35 (7):42, 45, April 1989. CODEN DTMNAT. ISSN 0011-6963.

Tittel:1996:LNC

[Tit96] Ed Tittel. Let NT connect your intranet. Datamation, 42(16):12– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Tittel:1996:UNY

[TMS96] Ed Tittel, Mary Madden, and David B. Smith. Use NT for your Web server — NT server integrates TCP/IP protocols and services to bring the flexibility and functionality of enterprise networking and Internet interoperability to the Windows environment. Datamation, 42(15): 131–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Tomlin:1988:EC

[Tom88] R. Tomlin. A European IS culture for the 1990s. Datamation, 34(17):48/14–20, September 1988. CODEN DTMNAT. ISSN 0011-6963.

Trautman:1989:CTP

[Tra89a] P. Trautman. Can they practice what they preach? (european single market and government DP). Datamation, 35(5): 76/19, 20, March 1989. CODEN DTMNAT. ISSN 0011-6963.

Trautman:1989:SSP

[Tra89b] P. Trautman. Systems solution to pollution. Datamation, 35(19): 72/1–2, October 1989. CODEN DTMNAT. ISSN 0011-6963.

Trautman:1990:ITT

[Tra90a] P. Trautman. Integrating text for TV databases. Datamation, 36(10):102–103, May 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Trautman:1990:MST

[Tra90b] P. Trautman. Midsized steps toward automation. Datamation, 36(8):73–74, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Travaglia:1995:TSC

[Tra95] Simon Travaglia. ThinkWrap — Status is claiming your body is racked with pain when you sit down at your machine. Datamation, 41(18):100–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Tucker:1996:NSS

[Tuc96a] Michael Jay Tucker. NFS servers soup up file I/O — NFS servers boost performance by speeding file operations, and the market has boomed as NFS has become the de facto file-sharing standard in PC networks. Datamation, 42 (13):121–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Tucker:1996:WDT

[Tuc96b] Michael Jay Tucker. Will DSM tools ever really work? Datamation, 42(18):115–??, ???? 1996.

CODEN DTMNAT. ISSN 0011-6963.

Tucker:1997:RDD

[Tuc97a] Michael J. Tucker. Raising dough from data. Datamation, 43(5): 50–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Tucker:1997:BYL

[Tuc97b] Michael Jay Tucker. Bridge your legacy systems to the Web. Datamation, 43(3):114–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Tucker:1997:CNH

[Tuc97c] Michael Jay Tucker. Can NT handle the big numbers? If your corporation has thousands of users and assorted platforms, NT must come under some kind of enterprise management. Some users believe NT can be managed by a framework right now, but most. Datamation, 43(11):54–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Tucker:1997:DPF

[Tuc97d] Michael Jay Tucker. The DSM payback: a framework for profit. Datamation, 43(3):69–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Tucker:1997:ENP

[Tuc97e] Michael Jay Tucker. EDI and the Net: a profitable partnering. Datamation, 43(4):62–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Tucker:1997:HRI

[Tuc97f] Michael Jay Tucker. The hunt for red ink. Datamation, 43(6):66– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Tucker:1997:JAB

[Tuc97g] Michael Jay Tucker. A juggling act between old and new. Datamation, 43(7):61–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Tucker:1997:LJ

[Tuc97h] Michael Jay Tucker. Lords of the jungle. Datamation, 43(9):70–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Tucker:1997:MYW

[Tuc97i] Michael Jay Tucker. Managing your Web-to-database performance. Datamation, 43(1):106– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Tucker:1997:MTT

[Tuc97j] Michael Jay Tucker. MQSeries tools take flight. Datamation, 43 (8):70–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Tucker:1997:NMT

[Tuc97k] Michael Jay Tucker. The new money: Transactions pour across the Web. Datamation, 43(4):73– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Tucker:1997:SSO

[Tuc97l] Michael Jay Tucker. Super solutions from ordinary software. Datamation, 43(7):90–??, ????

1997. CODEN DTMNAT. ISSN 0011-6963.

Tucker:1997:UP

[Tuc97m] Michael Jay Tucker. Unmask its power. Datamation, 43(10):92– ??, ???? 1997. CODEN DTM-NAT. ISSN 0011-6963.

Tucker:1997:WST

[Tuc97n] Michael Jay Tucker. Web servers tuned for e-commerce. Datamation, 43(2):92–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Tucker:1998:ESD

[Tuc98a] Michael Jay Tucker. Enterprise software development: Cooking up a sweet deal. Datamation, 44 (1):80–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Tucker:1998:SBN

[Tuc98b] Michael Jay Tucker. Strange bedfellows: NT, UNIX, and the integration option. Datamation, 44(2):42–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Ullman:1994:TPU

[Ull94] E. Ullman. Turn power users into power savers. Datamation, 40(9):53–55, May 1994. CODEN DTMNAT. ISSN 0011-6963.

Vacca:1994:SSS

[Vac94] John R. Vacca. Sign On to Streamlined Security. Datamation, 40(18):65–67, September 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Valigra:1996:XCE

[Val96] Lori Valigra. XTerminals: Cheap and easy — okay, so it’s not exactly on the cutting edge. but if you’re looking for cheap, secure, and easy to administer inter and. Datamation, 42(8):65– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

VanLoon:1990:WIC

[Van90] R. Van Loon. Working the information channels. Datamation, 36(9):85–88, May 1990. CODEN DTMNAT. ISSN 0011-6963.

Varhol:1994:TAW

[Var94a] Peter Varhol. The Truth About Windows on Alternative Platforms. Datamation, 40(13):43– ??, July 01, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Varhol:1994:WAH

[Var94b] Peter Varhol. What’s Ahead In Hot Chips. Datamation, 40(22): 73–??, November 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Varhol:1995:PLI

[Var95a] Peter Varhol. The P6: At Last, an Intel Heavy-Duty Architecture. Datamation, 41(17):70–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Varhol:1995:WWB

[Var95b] Peter Varhol. Warp’s Weakness: 32-Bit Windows Apps. Datamation, 41(6):47–??, April 01, 1995. CODEN DTMNAT. ISSN 0011-

Varhol:1995:TRO

[Var95c] Peter D. Varhol. Three Routes to OLAP. Datamation, 41(15): 57–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Varhol:1995:USD

[Var95d] Peter D. Varhol. Upsizing with SMP: Don’t Ignore the OS. Datamation, 41(1):59–??, January 1995. CODEN DTMNAT. ISSN 0011-6963.

Varhol:1996:FUB

[Var96a] Peter D. Varhol. The future of UNIX: 64 bits — so far, only Digital and SGI can give you the OS you need to get the main benefit of 64-bit UNIX RISC machines: gazillions of gigabytes of memory. Datamation, 42(11):26–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Varhol:1996:NRD

[Var96b] Peter D. Varhol. NT/RISC: a desktop combo for the future — computing managers looking for an alternative to UNIX on RISC systems may finally have found a good one in Windows NT. it can spread a single look and feel across an enterprise, but for desktop use more generalpurpose business apps need to be developed. Datamation, 42 (2):27–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Varney:1996:CSS

[Var96c] Sara E. Varney. Calm strategy solves year 2000 problem — IS managers at Merle Norman are spreading the word — there’s much to be done in getting 2 million lines of home-grown code converted in time for the year 2000. Datamation, 42(10):79–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Varney:1996:LYH

[Var96d] Sarah Varney. Link your help desk to the Web — linking your help desk to the Web can cut down on the number of support calls and help you provide better customer service. read about five companies that succeeded. Datamation, 42(10):64–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Varney:1996:AYS

[Var96e] Sarah E. Varney. Arm your salesforce with the Web. Datamation, 42(16):72–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Varney:1996:DMP

[Var96f] Sarah E. Varney. Database marketing predicts customer loyalty — it costs more to find new customers than it does to competitive keep the ones you already have. database marketing lets you predict customer loyalty and future buying patterns, as well as prevent customer turnover from churning away profits. Datamation, 42(15):50–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Varney:1996:DCI

[Var96g] Sarah E. Varney. Datamarts:

Coming to an IT mall near you! chances are good that datamarts will pop up at your IS site. but what are they? for starters, they offer most of the benefits of data warehousing without a lot of the pain and agony. Datamation, 42 (11):44–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Varney:1996:DLW

[Var96h] Sarah E. Varney. Datawebs! link the Web to your legacy data and apps — roll up your sleeves, it’s time to take a down-and-dirty look at the new Web-enabled tools that. Datamation, 42(7): 38–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Varney:1996:FYW

[Var96i] Sarah E. Varney. Framework — you will link your legacy app to the Web. Datamation, 42(7):5– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Varney:1996:IJC

[Var96j] Sarah E. Varney. IBM jumps into e-commerce — IBM is serious about selling new S/390 mainframes to host Web sites. for interactive its Big Iron story makes sense. Datamation, 42(13):52–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Varney:1996:WIL

[Var96k] Sarah E. Varney. Will intranets lay waste to groupware? Datamation, 42(18):72–??, ???? 1996.

CODEN DTMNAT. ISSN 0011-6963.

Vins:1995:BDO

[Vin95a] David Vins. Borland’s Delphi: The Object Is Control. Datamation, 41(9):67–??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Vins:1995:OMY

[Vin95b] David Vins. Obsydian: Model Your Way Through App Dev. Datamation, 41(19):85–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Vins:1995:VST

[Vin95c] David Vins. Visual Smalltalk: Team Objects for C/S Development. Datamation, 41(13):61–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Varhol:1996:WWW

[VM96a] Peter D. Varhol and Vance McCarthy. Who wins the Web server shootout? to tell the truth, there is no winner in the traditional sense. but DATAMATION’s evaluation of Web servers from Microsoft, Netscape, and Process Software highlights the factors to consider when making a choice. Datamation, 42(7):48–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Varney:1996:C

[VM96b] Sarah E. Varney and Vance Mc-Carthy. e-commerce. Datamation, 42(16):42–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Vigil:1996:MIS

[VM96c] Henry P. Vigil and Marianne Mueller. Making the Internet safe for e-commerce. Datamation, 42(16):64–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Vogel:1995:TNW

[Vog95] Peter Vogel. ThinkWrap — Now that Windows95 with Remote OLE is here, you really can make regular paper pushers productive. Datamation, 41(21):100–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Vogel:1996:KYB

[Vog96a] Peter Vogel. Know your business: Build a knowledgebase — a knowledgebase won’t build itself. here are five things to do — and three things not to do — to make sure it works. Datamation, 42 (13):84–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Vogel:1996:TIY

[Vog96b] Peter Vogel. ThinkWrap — if you think the maintenance group eats up too much of your department’s budget, think again. it’s the development group that’s killing you. Datamation, 42(6): 108–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Vogel:1996:TWD

[Vog96c] Peter Vogel. THINKWRAP — when in doubt… — the fastest way to achieve higher production is to eliminate tasks. Datamation, 42(11):96–??, ???? 1996.

CODEN DTMNAT. ISSN 0011-6963.

vanHoff:1996:WTC

[vSS96] Arthur van Hoff, Sami Shaio, and Orca Starbuck. What is this thing called Java — Java is a lot of things to a lot of people. here are a dozen of its most important features, according to the people. Datamation, 42(5):45– 46, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Waller:1989:CNP

[Wal89a] M. Waller. Cures for noise problems (power protection). Datamation, 35(15):47–48, 50, August 1989. CODEN DTMNAT. ISSN 0011-6963.

Waller:1989:PDC

[Wal89b] M. Waller. Planning data center design. Datamation, 35(21): 73–74, November 1989. CODEN DTMNAT. ISSN 0011-6963.

Wallace:1990:NLO

[Wal90a] S. Wallace. A new look for optical networks. Datamation, 36 (19):79–82, October 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Wallace:1990:RFI

[Wal90b] S. Wallace. Rx for faster image transmission. Datamation, 36 (18):95–96, 98, September 1990. CODEN DTMNAT. ISSN 0011-6963.

Warner:1997:BCB

[War97] Richard W. Warner. Blue Cross/ Blue Shield saves big bucks in Year 2000 project. Datamation, 43(1):54–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Wasson:1990:OFT

[Was90] R. E. Wasson. Organizing for future technologies. Datamation, 36(7):93–95, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Watterson:1996:DRE

[Wat96a] Karen Watterson. Database replication explained — there are almost as many replication schemes as there are databases that use the technology. our database consultant . . .. Datamation, 42(15):62–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Watterson:1996:IDD

[Wat96b] Karen Watterson. IBM does data — the technology pioneer is working double time to keep up in data warehousing and OLAP, says our database expert. Datamation, 42(13):64–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Watterson:1996:SSO

[Wat96c] Karen Watterson. SQL server 6.5: One for the Web — DATA-MATION first look: Microsoft’s SQL Server 6.5 is a solid database for data warehousing with great support for Internet applications. Datamation, 42(6):57– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Watterson:1997:PT

[Wat97]  Karen Watterson.  Parallel

tracks. Datamation, 43(5):92–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Watterson:1998:WIC

[Wat98] Karen Watterson. When it comes to choosing a database, the object is value. Datamation, 44 (1):100–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Weber:1989:CSF

[Web89] H. Weber. From CASE to software factories. Datamation, 35 (7):34–36, 52, April 1989. CO-DEN DTMNAT. ISSN 0011-6963.

Weston:1990:WMP

[Wes90] F. C. Weston, Jr. Why IS must be a partner in CIM. Datamation, 36(18):111–112, September 1990. CODEN DTMNAT. ISSN 0011-6963.

Williams:1990:NTC

[Wil90] D. Williams. New technology for coordinating work. Datamation, 36(10):92–96, May 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Wilent:1994:PDL

[Wil94] Steve Wilent. Parallel DBMS Lets UNIX Challenge Mainframes. Datamation, 40(17):45– ??, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Wildemuth:1997:ACN

[Wil97a] Scott Wildemuth. Aim carefully for I-net profits. Datamation, 43 (7):97–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Wildemuth:1997:BB

[Wil97b]  Scott Wildemuth.  Bankers’

beachhead. Datamation, 43(6): 91–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Wildemuth:1997:EW

[Wil97c] Scott Wildemuth. Employ the Web. Datamation, 43(8):67–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Wildemuth:1997:CC

[Wil97d] Scott Wildemuth. I-commerce is in the cards. Datamation, 43 (10):109–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Wildemuth:1997:SCS

[Wil97e] Scott Wildemuth. A simple case of supply and demand. Datamation, 43(11):99–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Wildemuth:1997:SWP

[Wil97f] Scott Wildemuth. Still waiting by the phone? Datamation, 43 (9):103–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Wildemuth:1998:WIP

[Wil98] Scott Wildemuth. Web investment pays off. Datamation, 44 (1):115–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Winkler:1990:BPM

[Win90] C. Winkler. Better project management in uncertain times. Datamation, 36(11):95–98, June 1990. CODEN DTMNAT. ISSN 0011-6963.

Winsberg:1995:LCD

[Win95] Paul Winsberg. Legacy Code: Don’t Bag It, Wrap It. Datamation, 41(9):36–??, May 15, 1995. CODEN DTMNAT. ISSN 0011-6963.

Williams:1990:SOD

[WN90] J. Williams and J. A. Nelson. Striking oil in decision support. Datamation, 36(6):83–87, March 1990. CODEN DTMNAT. ISSN 0011-6963.

Wood:1991:DPS

[Woo91] L. Wood. Dial 1-800 for PC support. Datamation, 37(20):69–70, October 1991. CODEN DTM-NAT. ISSN 0011-6963.

Wood:1992:DPG

[Woo92] L. Wood. Desktop publishing goes on line. Datamation, 38(25): 45–47, December 1992. CODEN DTMNAT. ISSN 0011-6963.

Wood:1993:BPB

[Woo93] L. Wood. Business profits from bulletin boards. Datamation, 39 (9):45–46, 48, May 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Worhach:1990:DIY

[Wor90] D. Worhach. A do-it-yourself guide to PC repair. Datamation, 36(20):39–40, October 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Watts:1994:CGW

[WP94] Michael Watts and Jeff Plewa. CIOs Go with what Works.

Datamation, 40(13):55–??, July 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Wray:1998:NBP

[Wra98] Nicholas Wray. Negotiate better phone rates. Datamation, 44 (2):79–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Yourdon:1997:MDM

[You97] Ed Yourdon. Making death march projects pay off. Datamation, 43(4):116–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Zawacki:1995:RBY

[ZNZA95] Robert Zawacki, Carol A. Norman, Paul A. Zawacki, and Paul D. Applegate. Reenergize Before You Reengineer. Datamation, 41(13):45–??, 1995. CO-DEN DTMNAT. ISSN 0011-6963.

Zawacki:1988:HCS

[ZSZ88] R. A. Zawacki, D. Kubert Scott, and P. A. Zawacki. How college students choose IS careers. Datamation, 34(18):40–48, September 1988. CODEN DTMNAT. ISSN 0011-6963.